### v1

# Chinese Academy of Sciences, Beijing 100190, China

Abstract
The primordial neutrino asymmetry leave profound imprints on the evolution history of the universe, which can be constrained by cosmological observations, including Big Bang Nucleosynthesis (BBN), Cosmic Microwave Background (CMB), and Large-Scale Structure (LSS). We present comprehensive analysis on implications and constraints of the primordial neutrino asymmetry $\xi_{\nu}$, based on a precise treatment of neutrino decoupling by solving the complete (anti)neutrino quantum kinetic equations in the Closed-Time-Path formalism. The effective number of neutrinos $N_{\rm eff}$ and (anti)neutrino spectral distortions are calculated, and we find that the non-instantaneous decoupling correction is given by $\delta N_{\rm eff}=0.0440+0.0102\,\xi_{\nu}^{2}$. Then we perform the state-of-the-art calculation for the abundance of light elements including (anti)neutrino spectral distortions, which indicate a positive asymmetry $0.032\leq\xi_{\nu}\leq 0.052$ from EMPRESS data. The implications of the neutrino asymmetry for the CMB and LSS are studied in detail, and we find that the Baryon Acoustic Oscillations (BAO) are also significantly affected by $\xi_{\nu}$ in addition to the sum of neutrino masses. A combined analysis with EMPRESS BBN, Planck CMB and BOSS BAO data yields a tighter constraint $\xi_{\nu}=0.024\pm 0.013$, which provides constraints on UV models capable of producing large asymmetries.
1 Introduction
[…]
A precise understanding of the neutrino decoupling process is essential for accurate constraints on neutrino properties. Notably, the partial reheating of neutrinos and antineutrinos due to the small overlap between neutrino decoupling and $e^{\pm}$ annihilations results in slightly non-thermal spectra and an increased total neutrino energy density, typically described by the effective number of relativistic species, $N_{\rm eff}$. The values of $N_{\rm eff}$ and the spectral distortions of neutrinos and antineutrinos are used as initial conditions to evaluate the impact of relic neutrinos in subsequent processes. For the Standard Model case, where asymmetries between neutrinos and antineutrinos are assumed to be negligible, significant progress has been made in detailing the neutrino decoupling process over recent decades Hannestad:1995rs ; Dolgov:1997mb ; Mangano:2001iu ; Mangano:2005cc ; Grohs:2015tfy ; deSalas:2016ztq ; Akita:2020szl ; Froustey:2019owm ; EscuderoAbenza:2020cmq ; Froustey:2020mcq ; Bennett:2020zkv . Including flavor oscillations, matter effects, finite-temperature QED corrections up to $\mathcal{O}(3)$, and full neutrino-electron and neutrino-neutrino collision terms, the current standard prediction yields $N_{\rm eff}(0)=3.0440\pm 0.0002$ Froustey:2020mcq ; Bennett:2020zkv .
On the other hand, inspired by the recent EMPRESS survey on primordial helium abundance empress and its potential to alleviate cosmological tensions Barenboim:2016lxv ; Yeung:2020zde ; Seto:2021tad ; Kumar:2022vee ; Yeung:2024krv , the scenario where asymmetries between neutrinos and antineutrinos are not negligible has recently attracted significant attention. Due to the sphaleron processes in the early universe Kuzmin:1985mm ; Khlebnikov:1988sr ; Harvey:1990qw ; Dreiner:1992vm , one might naively expect neutrino asymmetries, or lepton asymmetries, to be of the same order as the baryon asymmetry of the universe, which is strongly constrained by BBN and CMB observations: $\eta_{B}\equiv(n_{B}-n_{\bar{B}})/n_{\gamma}=(6.14\pm 0.04)\times 10^{-10}$ Planck:2018vyg , where $n_{B}$, $n_{\bar{B}}$, and $n_{\gamma}$ denote the number densities of baryons, antibaryons, and photons, respectively. However, substantial lepton asymmetries compared to the baryon asymmetry before the neutrino decoupling epoch can still be generated by various models, see $e.g.,$ Casas:1997gx ; Dolgov:1989us ; Bajc:1997ky ; Asaka:2005pn ; Asaka:2005an ; Pilaftsis:2003gt ; Borah:2022uos ; Kawasaki:2002hq ; Kawasaki:2022hvx .
The lepton asymmetries prior to the neutrino decoupling epoch, hereafter referred to as primordial neutrino asymmetries, have significant implications for the evolution of the universe. During the neutrino decoupling epoch, primordial neutrino asymmetries lead to an increased $N_{\rm eff}$ and specific spectral distortions of neutrinos and antineutrinos Grohs:2016cuu . Furthermore, if the asymmetries among different flavors are different, neutrinos and antineutrinos can undergo synchronous oscillations Bell:1998ds ; Pastor:2001iu ; Abazajian:2002qx ; Wong:2002fa , resulting in the redistribution of asymmetries among flavors Dolgov:2002ab ; Pastor:2008ti ; Mangano:2010ei ; Mangano:2011ip ; Castorina:2012md ; Barenboim:2016shh ; Johns:2016enc ; Froustey:2021azz . After neutrino decoupling, during the BBN epoch, primordial neutrino asymmetries primarily affect the final positron-to-neutron ratio and thus the final abundances of light elements. This effect mainly depends on the density distributions of electron neutrinos and antineutrinos, while corrections from the increased $N_{\rm eff}$ are subleading Sarkar:1995dd ; Iocco:2008va ; Pitrou:2018cgg ; Serpico:2005bc ; Mangano:2011ip ; Chu:2006ua ; Simha:2008mt ; Saviano:2013ktj ; Burns:2023sgx . Additionally, primordial neutrino asymmetries influence CMB anisotropies and LSS through increased $N_{\rm eff}$ and modified light element abundances, alongside the effects of massive neutrinos. Consequently, observations of BBN, CMB, and LSS can be utilized to constrain primordial neutrino asymmetries Simha:2008mt ; Oldengott:2017tzj ; Pitrou:2018cgg ; Matsumoto:2022tlr ; Burns:2022hkq ; Escudero:2022okz . In particular, recent studies from the EMPRESS survey indicate a preference for a positive asymmetry in electron neutrinos Matsumoto:2022tlr ; Burns:2022hkq ; Escudero:2022okz . Specifically, combining BBN and CMB observations yields $\xi_{\nu_{e}}=0.034\pm 0.014$ Escudero:2022okz , where $\xi_{\nu_{e}}=\mu_{\nu_{e}}/T_{\nu}$ represents the degeneracy parameter for electron neutrinos.
However, the current situation is less satisfactory compared to the Standard Model (SM) case due to persistent ambiguities regarding neutrino decoupling. These ambiguities particularly involve neutrino spectral distortions and the effects of flavor oscillations and QED corrections. As a result, the effects of primordial neutrino asymmetries on BBN and CMB, as well as the constraints derived from observational data, are generally evaluated using approximations. These approximations often include the instantaneous neutrino decoupling approximation and the assumption of massless neutrinos. To derive precise and accurate constraints on primordial neutrino asymmetries from both current and future observations, it is essential to rigorously analyze the implications of these asymmetries through a comprehensive treatment of the neutrino decoupling process. A recent study Froustey:2024mgf makes effort toward a rigorous analysis of primordial neutrino asymmetries. This study focuses on the implications of neutrino synchronous oscillations for BBN, while ignoring the correction from (anti)neutrino spectral distortions, etc, and furthermore, the implications for both the CMB and LSS are not addressed at all.
In this work, we focus on evaluating the neutrino decoupling process and its subsequent effects on BBN, CMB and LSS in the presence of primordial neutrino asymmetries. To accurately describe the evolution of neutrino density matrices during decoupling, including flavor oscillations and full collision effects, it is crucial to employ the complete set of neutrino quantum kinetic equations (QKEs). These equations can be derived using various approaches, such as perturbative expansions of the density matrix Sigl:1993ctk , relativistic generalizations of the Bogoliubov-Born-Green-Kirkwood-Yvon (BBGKY) equations Volpe:2013uxl ; Volpe:2015rla ; Froustey:2020mcq , and more generally from the Closed-Time Path (CTP) formalism of non-equilibrium quantum field theory Vlasenko:2013fja ; Blaschke:2016xxt . The main improvements in our study are as follows:
  • We re-derive the most general form of the QKEs for neutrinos and antineutrinos using the CTP formalism, incorporating primordial neutrino asymmetry.
  • We numerically solve the QKEs for the neutrino decoupling process in the presence of primordial neutrino asymmetry, including effects of flavor oscillations, matter interactions, finite-temperature QED (FTQED) corrections up to $\mathcal{O}(3)$ order, and complete neutrino-electron and neutrino-neutrino collision terms. This approach matches the accuracy used in current state-of-the-art predictions for the Standard Model (SM) case Froustey:2020mcq ; Bennett:2020zkv  [^1] Additional QED corrections, such as those to interaction vertices Cielo:2023bqp ; Jackson:2023zkl ; Drewes:2024wbw , are still debated for the SM case and are not included in this study..
  • We use the resulting density matrices to study the impacts of primordial neutrino asymmetry on the BBN, CMB, and LSS. The effects of neutrino spectral distortions are also highlighted by comparing these results with those obtained using the Fermi-Dirac distribution for neutrinos.
  • With the observation that the Baryon Acoustic Oscillations (BAO) are also affected by the primordial neutrino asymmetry, we derive the constraints on the primordial neutrino asymmetry by combining the traditional BBN and CMB datasets with the BAO observations for the first time.
Figure 1: A schematic illustration outlining the structure of this paper. Section 2 presents the derivation of the QKEs for neutrinos within the CTP formalism. These QKEs are subsequently applied to study neutrino decoupling in the presence of primordial neutrino asymmetry $\xi_{\nu}$ in Section 3. Based on such precise treatment of neutrino decoupling, Sections 4 and 5 explore the effects of primordial neutrino asymmetry on BBN, CMB and LSS. For each section, the primary input and output physical quantities that depend on $\xi_{\nu}$ are labelled in each box. Finally, in Section 6, cosmological constraints on $\xi_{\nu}$ are obtained by the MCMC analyses combining the theoretical predictions and observational data from BBN, CMB, and LSS.
The key physical results of this study are:
  • We numerically evaluate $N_{\rm eff}$ and the (anti)neutrino spectral distortions, incorporating primordial neutrino asymmetries with state-of-the-art accuracy. We find that $N_{\rm eff}$ can be expressed as $N_{\rm eff}=N_{\rm eff}^{\rm SM}+\delta N_{\rm eff}^{\xi_{\nu},\rm ther}+0.0102\,\xi_{\nu}^{2}$, where $N_{\rm eff}^{\rm SM}=3.0440\pm 0.002$ represents the SM prediction with comparable precision Froustey:2020mcq ; Bennett:2020zkv . The term $\delta N_{\rm eff}^{\xi_{\nu},\rm ther}$ accounts for the contribution from primordial neutrino asymmetry under the instantaneous decoupling approximation, and is given by $\delta N_{\rm eff}^{\xi_{\nu},\rm ther}\simeq 3\left(\frac{30}{7\pi^{2}}\xi_{\nu}^{2}+\frac{15}{7\pi^{4}}\xi_{\nu}^{4}\right)$, while the final term represents the additional correction from non-instantaneous decoupling effects in the presence with non-zero primordial neutrino asymmetry.
  • We provide state-of-the-art predictions for light element abundances produced by BBN, through which the EMPRESS measurement on helium abundance $Y_{P}|_{\rm EMPRESS}=0.2370^{+0.0034}_{-0.0033}$ immediately indicate a positive primordial neutrino asymmetry $0.032\leq\xi_{\nu}\leq 0.052$ with fixed baryon abundance $\omega_{b}$. We also find that the (anti)neutrino spectral distortions are important for precise predictions for BBN, especially for the abundance of Helium-4 $Y_{P}$.
  • For the constraints on primordial neutrino asymmetry, combing EMPRESS BBN and Planck CMB data yields $\xi_{\nu}=0.028\pm 0.013$, while incorporating BOSS BAO data in the analysis yields $\xi_{\nu}=0.024\pm 0.013$. Compared to previous result obtained in ref $\xi_{\nu}=0.034\pm 0.014$, our results indicate a $\sim 18\%$ and $\sim 29\%$ smaller central value and a slight reduction in the uncertainty, respectively.
  • The impacts of treating the sum of neutrino mass $\sum m_{\nu}$ and the primordial neutrino asymmetry $\xi_{\nu}$ as free parameters on the resulting constraints for each of them are also discussed. Finally, we show how the constraints on $\xi_{\nu}$ can be converted into the constraints on UV model parameters with a specific Q-ball decay model capable of producing large primordial neutrino asymmetries.
This paper is organized as follows. In Section 2, we re-derive the QKEs for Dirac and Majorana neutrinos using the CTP formalism. In Section 3, we present the QKEs in the context of the early universe and provides corresponding numerical results considering primordial neutrino asymmetries. The impacts of primordial neutrino asymmetries on the BBN epoch is discussed in Section 4, while their effects on both CMB and LSS are analyzed in Section 5. In Section 6, we present the results from a joint MCMC analysis of neutrino primordial asymmetries and neutrino masses, utilizing datasets from BBN, CMB and BAO observations. Finally, we conclude in Section 7. To guide the readers, a schematic illustration of this work are presented in Fig. 1. Throughout this manuscript, we use the natural units where $\hbar=c=k_{B}=1$.
2 Quantum Kinetic Equations for Dirac and Majorana Neutrinos
In this section, we derive the quantum kinetic equations (QKEs) for both Dirac and Majorana neutrinos using the closed-time-path (CTP) formalism of non-equilibrium quantum field theory. The QKEs generalize the traditional Boltzmann equation by governing the evolution of density matrices and incorporating both the phase-space distribution functions for the eigenstates and the quantum coherence between them. These equations have also been derived through various approaches and approximations in the literature (see $e.g.,$  Sigl:1993ctk ; Volpe:2013uxl ; Volpe:2015rla ; Froustey:2020mcq ).
2.1 The Closed-Time-Path formalism
The Closed-Time-Path (CTP) formalism, also known as the in-in formalism, provides a general framework for non-equilibrium quantum field theory (QFT). This approach is necessary because, in non-equilibrium situations, the final state of the system is not known a priori. The CTP formalism employs a closed time contour to handle such cases. Specifically, the path integral in this formalism is defined over a closed time path that starts at an initial time $t_{0}$, extends to infinity on the positive branch, and then returns to $t_{0}$ on the negative branch. This closed time path allows for the consistent treatment of quantum fields in non-equilibrium scenarios by incorporating both forward and backward evolution in time. Accordingly, depending on the location of the time arguments on the $\pm$ branches, there are four kinds of real-time Green functions, which we denote as
$\left(S_{ab}\right)^{>}(x,y)$ $\equiv\langle\nu_{a}(x_{-})\bar{\nu}_{b}(y_{+})\rangle~{},$ (1a)
$\left(S_{ab}\right)^{<}(x,y)$ $\equiv\langle\bar{\nu}_{b}(y_{-})\nu_{a}(x_{+})\rangle~{},$ (1b)
$\left(S_{ab}\right)^{\rm T}(x,y)$ $\equiv\langle\mathcal{T}\left(\nu_{a}(x_{+})\bar{\nu}_{b}(y_{+})\right)\rangle~{},$ (1c)
$\left(S_{ab}\right)^{\rm\bar{T}}(x,y)$ $\equiv\langle\bar{\mathcal{T}}\left(\nu_{a}(x_{-})\bar{\nu}_{b}(y_{-})\right)\rangle~{},$ (1d)
where $a$ and $b$ denote flavor indices of the neutrinos, while spinor indices are suppressed. The operators $\mathcal{T}$ and $\bar{\mathcal{T}}$ represent the usual time-ordering and reversed time-ordering operations, respectively. It is evident that only two of these four Green functions are independent. For example, by explicitly writing out the time-ordering operators, we obtain
$\left(S_{ab}\right)^{\rm T}(x,y)$ $=\theta(x^{0}-y^{0})\left(S_{ab}\right)^{>}(x,y)-\theta(y^{0}-x^{0})\left(S_{ab}\right)^{<}(x,y)~{},$ (2a)
$\left(S_{ab}\right)^{\rm\bar{T}}(x,y)$ $=\theta(y^{0}-x^{0})\left(S_{ab}\right)^{>}(x,y)-\theta(x^{0}-y^{0})\left(S_{ab}\right)^{<}(x,y)~{}.$ (2b)
In practice, it is also convenient to introduce the other two dependent propagators to describe the non-equilibrium systems of neutrinos, namely the statistical and spectral functions, defined as
$S^{F}_{ab}(x,y)$ $\equiv\frac{1}{2}\langle[\nu_{a}(x),\bar{\nu}_{b}(y)]\rangle=\frac{1}{2}\left(\left(S_{ab}\right)^{>}(x,y)-\left(S_{ab}\right)^{<}(x,y)\right)\,,$ (3a)
$S^{\mathcal{A}}_{ab}(x,y)$ $\equiv\frac{i}{2}\langle\{\nu_{a}(x),\bar{\nu}_{b}(y)\}\rangle=\frac{i}{2}\left(\left(S_{ab}\right)^{>}(x,y)+\left(S_{ab}\right)^{<}(x,y)\right)\,.$ (3b)
The advantage of this definition is that the statistical and spectral functions have straightforward physical interpretations Berges:2004yj ; Blaschke:2016xxt ; Vlasenko:2013fja . Specifically, the spectral function provides information about the spectrum of the theory, while the statistical function encodes information about the occupation numbers and quantum coherence of the available states, $i.e.,$ information about the density matrices. Additionally, we note that we do not specify whether the neutrinos are Dirac or Majorana particles here; the distinction between Dirac and Majorana neutrinos will be discussed later.
Therefore, the starting point for the QKEs is the equation of motion for the Green functions $S_{ab}(x,y)$, which can be expressed through the Dyson-Schwinger equation:
$\left(i\not{\partial}^{x}-m\right)\tilde{S}(x,y)=\mathbb{1}\,i\delta^{4}(x-y)+i\int d^{4}z\,\tilde{\Sigma}(x,z)\tilde{S}(z,y)\,.$ (4)
Here, $\tilde{S}$ denotes the matrix of Green functions in the Closed-Time-Path (CTP) space, and is defined as
$\tilde{S}=\begin{pmatrix}\left(S_{ab}\right)^{T}&\left(S_{ab}\right)^{<}\\ \left(S_{ab}\right)^{>}&-\left(S_{ab}\right)^{\bar{T}}\end{pmatrix}\,.$ (5)
In this context, $m\equiv\mathbb{1}_{2\times 2}\,m_{ab}$, where $m_{ab}$ is the neutrino mass matrix. The identity matrix preceding the Dirac delta function $\delta^{4}(x-y)$ is both an identity matrix in the CTP space and in the flavor space, i.e., $\mathbb{1}\,i\delta^{4}(x-y)=\mathbb{1}_{2n_{f}\times 2n_{f}}\,i\delta^{4}(x-y)$, with $n_{f}$ representing the number of neutrino flavors. Additionally, $\tilde{\Sigma}(x,y)$ is the neutrino proper self-energy, which is generally a functional of the Green functions and can be decomposed into a local singular term and a matrix in CTP space:
$\tilde{\Sigma}(x,y)=-\mathbb{1}_{2\times 2}i\Sigma(x)\delta^{4}(x-y)+\tilde{\Pi}\,,\quad\tilde{\Pi}=\begin{pmatrix}\left(\Pi_{ab}\right)^{T}&\left(\Pi_{ab}\right)^{<}\\ \left(\Pi_{ab}\right)^{>}&-\left(\Pi_{ab}\right)^{\bar{T}}\end{pmatrix}\,.$ (6)
In principle, $\Sigma(x)$ arises from the forward scattering potentials contributed by one-loop diagrams, while $\Pi^{\pm}(x,y)$ represents contributions from non-forward scatterings and annihilations that occur in at least two loops.
In the following, we suppress the flavor and spinor indices of the Green functions for simplicity. Expanding Eq. (4), the Dyson-Schwinger equation for $S^{\gtrless}$ can be written as
$\left(i\not{\partial}^{x}-m-\Sigma(x)\right)S^{\gtrless}(x,y)=i\int d^{4}z\left(\Pi^{H}(x,z)S^{\gtrless}(z,y)+\Pi^{\gtrless}(x,z)S^{H}(z,y)\right)\pm i\,\mathcal{C}_{\rm coll}(x,y)\,,$ (7)
which are known as the Kadanoff–Baym (KB) equations. Here, $S^{H}\equiv(S^{T}-S^{\bar{T}})/2$ and similarly for $\Pi^{H}$, while the collision term $\mathcal{C}_{\rm coll}$ is defined as
$\mathcal{C}_{\rm coll}(x,y)\equiv\frac{1}{2}\int d^{4}z\left(\Pi^{>}(x,z)S^{<}(z,y)+\Pi^{<}(x,z)S^{>}(z,y)\right)\,,$ (8)
[…]
Using the definition in Eq. (3), we can derive the equations for the statistical function $S^{F}$ and the spectral function $S^{\mathcal{A}}$:
$\left(i\not{\partial}^{x}-m-\Sigma(x)\right)S^{F}(x,y)$ $=i\int d^{4}z\left(\Pi^{H}(x,z)S^{F}(z,y)+\Pi^{F}(x,z)S^{H}(z,y)\right)-i\,\mathcal{C}_{\rm coll}(x,y)\,,$ (9a)
$\left(i\not{\partial}^{x}-m-\Sigma(x)\right)S^{\mathcal{A}}(x,y)$ $=i\int d^{4}z\left(\Pi^{H}(x,z)S^{\mathcal{A}}(z,y)+\Pi^{\mathcal{A}}(x,z)S^{H}(z,y)\right)\,.$ (9b)
We note that Eq. (9a) is equivalent to Eq. (26) for the statistical function given in Ref. Vlasenko:2013fja . As will be showed, our formulation is more convenient for the Wigner transform.
The purpose of the Wigner transform is to separate the evolution of the Green functions on the microscopic scale with respect to $r\equiv x-y$ from the evolution on the macroscopic scale with respect to $X\equiv(x+y)/2$, which is our primary focus. The (inverse) Wigner transform for any two-point (or self-energy) function $F(x,y)$ is given by
$F(k,X)$ $=\int d^{4}r\,e^{ik\cdot r}\,F\left(X+\frac{r}{2},X-\frac{r}{2}\right)\,,$ (10a)
$F(x,y)$ $=\int\frac{d^{4}k}{(2\pi)^{4}}\,e^{-ik\cdot(x-y)}\,F\left(k,\frac{x+y}{2}\right)\,.$ (10b)
To transform the equations for statistical and spectral functions, we also need the Wigner transform of the generalized convolution term like $\int d^{4}z\,F(x,z)\,G(z,y)$. The result is given by
$\int d^{4}z\,F(x,z)\,G(z,y)=\int\frac{d^{4}k}{(2\pi)^{4}}\,e^{-ik\cdot(x-y)}\,e^{-i\Diamond}\left\{F(k,X)\right\}\left\{G(k,X)\right\}\,,$ (11)
where the right-hand side is known as the Moyal product, and the diamond operator $\Diamond$ is defined as
$\Diamond\left\{F(k,X)\right\}\left\{G(k,X)\right\}=\frac{1}{2}\left(\frac{\partial F}{\partial X^{\mu}}\frac{\partial G}{\partial k_{\mu}}-\frac{\partial F}{\partial k_{\mu}}\frac{\partial G}{\partial X^{\mu}}\right)\,.$ (12)
The Wigner-transformed equations for the statistical and spectral functions, given by Eq. (9), can then be straightforwardly obtained as
[…]
$\quad+i\,e^{-i\Diamond}\left\{\Pi^{F}(k,X)\right\}\left\{S^{H}(k,X)\right\}-i\,\mathcal{C}_{\rm coll}(k,X)\,,$ (13b)
[…]
$\quad+i\,e^{-i\Diamond}\left\{\Pi^{\mathcal{A}}(k,X)\right\}\left\{S^{H}(k,X)\right\}\,,$
with the collision term
$\mathcal{C}_{\rm coll}(k,X)=\frac{1}{2}\left(e^{-i\Diamond}\left\{\Pi^{>}(k,X)\right\}\left\{S^{<}(k,X)\right\}-e^{-i\Diamond}\left\{\Pi^{<}(k,X)\right\}\left\{S^{>}(k,X)\right\}\right)\,.$ (14)
2.2 Gradient expansion of the kinetic equation
So far, the equations for statistical and spectral functions (13) involve infinite gradient expansions and are thus valid to all orders. In practice, we follow the procedure outlined in Ref. Blaschke:2016xxt and truncate these infinite series at the leading non-trivial order in a small parameter $\epsilon$. To achieve this, we first need to specify the power counting in the regime of interest, as follows:
  • The variation of physical quantities with respect to the macroscopic coordinate $X$ is small compared to its intrinsic de Broglie frequency, so each derivative $\partial_{X}$ carries one power of $\epsilon$.
  • Compared to the neutrino energy, the masses and interaction potentials are small, so terms involving $m$ or $\Sigma$ carry one power of $\epsilon$.
  • Contributions to the self-energy $\Pi$ appear at least at the 2-loop order in the Feynman diagram expansion, so terms involving $\Pi$ carry two power of $\epsilon$.
  • The variation of the two-point functions from equilibrium is small, so $S^{H}=(S^{T}-S^{\bar{T}})/2$ carries one power of $\epsilon$.
In conclusion, our power counting reads
$\frac{\partial_{X},m,\Sigma,S^{H}}{E_{\nu}}=O(\epsilon)\,,\quad\frac{\tilde{\Pi}^{\pm}}{E_{\nu}}=O(\epsilon^{2})\,,$ (15)
where $E_{\nu}$ is the neutrino energy. Additionally, we note that the contribution from the $\Pi^{H}$ term can be absorbed into the $\Sigma$ term at the considered order, so we will omit the relevant terms in the following equations.
Next, we expand the equations for statistical and spectral functions (13) order by order to derive the QKEs for neutrinos and antineutrinos. To do this, we first note that the kinetic equation (13b) involves the sixteen spinor components (scalar, pseudoscalar, vector, axial-vector, tensor) of the statistical function and the self-energies, which can be decomposed as follows:
$S^{F}$ $=\left[F_{S}+\left(F_{V}^{R}\right)^{\mu}\gamma_{\mu}-\frac{i}{4}\left(F_{T}^{L}\right)^{\mu\nu}\sigma_{\mu\nu}\right]P_{L}+\left[F_{S}^{\dagger}+\left(F_{V}^{L}\right)^{\mu}\gamma_{\mu}+\frac{i}{4}\left(F_{T}^{R}\right)^{\mu\nu}\sigma_{\mu\nu}\right]P_{R}\,,$ (16a)
$\Pi^{F}$ $=\left[\Pi_{S}+\Pi_{R}^{\mu}\gamma_{\mu}-\frac{i}{4}\left(\Pi_{T}^{L}\right)^{\mu\nu}\sigma_{\mu\nu}\right]P_{L}+\left[\Pi_{S}^{\dagger}+\Pi_{L}^{\mu}\gamma_{\mu}+\frac{i}{4}\left(\Pi_{T}^{R}\right)^{\mu\nu}\sigma_{\mu\nu}\right]P_{R}\,,$ (16b)
where $P_{L,R}\equiv\frac{1\mp\gamma_{5}}{2}$ and $\sigma_{\mu\nu}\equiv\frac{i}{2}[\gamma_{\mu},\gamma_{\nu}]$. Due to the hermiticity conditions, we have $F_{V}^{L\dagger}=F_{V}^{L}$, $F_{V}^{R\dagger}=F_{V}^{R}$, and $F_{T}^{L\dagger}=F_{T}^{R}$, with similar conditions applying to the components of $\Pi^{F}$. An analogous decomposition for the forward scattering potential $\Sigma(x)$ is also permissible.
Therefore, the kinetic equation (13b) encompasses both the QKEs and the algebraic constraints for various spinor components. The basic strategy to obtain the QKEs is as follows:
  1. 1.
    Solve Eq. (13) at $\mathcal{O}(\epsilon^{0})$ to obtain the leading order solutions for the statistical and spectral functions, and establish the relations between the remaining spinor components of the statistical function and the density matrices for Dirac and Majorana neutrinos.
  2. 2.
    Beyond $\mathcal{O}(\epsilon^{0})$, all spinor components of $S^{F}$ receive additional corrections due to neutrino masses and interactions. Thus, solve the kinetic equation (13b) at $\mathcal{O}(\epsilon^{1})$ to find the constraint relations among these small components.
  3. 3.
    Expand the kinetic equation (13b) to $\mathcal{O}(\epsilon^{2})$ and reorganize it using the constraint relations obtained at $\mathcal{O}(\epsilon^{1})$.
  4. 4.
    Finally, extract the QKEs for Dirac and Majorana neutrinos, incorporating the relations between the spinor components of the statistical function and the density matrices.
The detailed derivation of these steps is presented in the following sub-subsections.
2.2.1 Kinetic equation to $\mathcal{O}(\epsilon^{0})$
[…]
$\left(F_{V}^{L,R}\right)^{\mu}(k,x)$ $=\hat{\kappa}^{\mu}(k)\,F_{L,R}(k,x)\,,$ (17a)
$\left(F_{T}^{L}\right)_{\mu\nu}(k,x)$ $=e^{-i\varphi(k)}(\hat{\kappa}(k)\wedge\hat{x}^{-}(k))_{\mu\nu}\,\Phi(k,x)\,,$ (17b)
$\left(F_{T}^{R}\right)_{\mu\nu}(k,x)$ $=e^{i\varphi(k)}(\hat{\kappa}(k)\wedge\hat{x}^{+}(k))_{\mu\nu}\,\Phi^{\dagger}(k,x)\,.$ (17c)
Here, we have introduced a set of basis vectors used to express the Lorentz tensors and components of the two-point functions. The basis consists of two light-like vectors $\hat{\kappa}^{\mu}(k)=({\rm sign}(k^{0}),\hat{k})$ and $\hat{\kappa}^{\prime\mu}(k)=({\rm sign}(k^{0}),-\hat{k})$, and two transverse four-vectors $\hat{x}_{1,2}(k)$ which satisfy the relations $\hat{\kappa}\cdot\hat{\kappa}=\hat{\kappa}^{\prime}\cdot\hat{\kappa}^{\prime}=0$, $\hat{\kappa}\cdot\hat{\kappa}^{\prime}=2$, $\hat{\kappa}\cdot\hat{x}_{i}=\hat{\kappa}^{\prime}\cdot\hat{x}_{i}=0$, and $\hat{x}_{i}\cdot\hat{x}_{j}=-\delta_{ij}$. Additionally, we define $\hat{x}^{\pm}\equiv\hat{x}_{1}\pm i\hat{x}_{2}$, such that $\hat{x}^{+}\cdot\hat{x}^{-}=-2$.
In addition, $F_{L,R}(k,x)$ and $\Phi(k,x)$ can be organized into a $2n_{f}\times 2n_{f}$ matrix:
$\hat{F}=\left(\begin{array}[]{cc}F_{L}&\Phi\\ \Phi^{\dagger}&F_{R}\end{array}\right)~{}.$ (18)
The advantage of this organization is that it can be directly related to the conventional density matrices. At $\mathcal{O}(\epsilon^{0})$, the density matrices for neutrinos and antineutrinos can be obtained via integrals over positive and negative frequencies. For Dirac neutrinos, we have
$-2\int_{0}^{\infty}\frac{dk^{0}}{2\pi}\hat{F}(k,x)$ $=F(\vec{k},x)-\frac{1}{2}\mathbb{1}\,,$ (19a)
$-2\int_{-\infty}^{0}\frac{dk^{0}}{2\pi}\hat{F}(k,x)$ $=\bar{F}(-\vec{k},x)-\frac{1}{2}\mathbb{1}\,.$ (19b)
The density matrices $F$ and $\bar{F}$ are also $2n_{f}\times 2n_{f}$ matrices:
$F(\vec{p},x)=\left(\begin{array}[]{cc}\varrho_{LL}&\varrho_{LR}\\ \varrho_{RL}&\varrho_{RR}\end{array}\right)\,;\qquad\qquad\bar{F}(\vec{p},x)=\left(\begin{array}[]{cc}\bar{\varrho}_{RR}&\bar{\varrho}_{RL}\\ \bar{\varrho}_{LR}&\bar{\varrho}_{LL}\end{array}\right)\,,$ (20)
where $n_{f}$ is the total number of neutrino flavors, and each block is a square $n_{f}\times n_{f}$ matrix with $\varrho_{hh^{\prime}}=\varrho_{hh^{\prime}}^{ij}$, where $i,\,j$ denote the flavors and $h,\,h^{\prime}\in\{L,R\}$ denote the chirality. Among the various components, $\varrho_{LL}$ and $\bar{\varrho}_{RR}$ represent the density matrices for active left-handed neutrinos and right-handed antineutrinos, respectively. Conversely, $\varrho_{RR}$ and $\bar{\varrho}_{LL}$ correspond to sterile right-handed neutrinos and left-handed antineutrinos. The remaining density matrices account for the quantum coherence between different chiralities of neutrinos and antineutrinos across all flavors. Note that there is no single-particle density matrix for quantum coherence between Dirac neutrinos and antineutrinos, as they are fundamentally distinct particles.
For Majorana neutrinos, the antiparticle is the neutrino itself, and the density matrices for neutrinos and antineutrinos are related by $\varrho_{hh^{\prime}}^{ij}=\bar{\varrho}_{hh^{\prime}}^{ji}$ through a phase change. Therefore, only the density matrices for active neutrinos are necessary, which can be consolidated into a single $2n_{f}\times 2n_{f}$ matrix using the definitions $\varrho\equiv\varrho_{LL}$, $\bar{\varrho}\equiv\bar{\varrho}_{RR}=\varrho_{RR}^{T}$, and $\phi\equiv\varrho_{LR}$:
${\cal F}=\left(\begin{array}[]{cc}\varrho&\phi\\ \phi^{\dagger}&\bar{\varrho}^{T}\end{array}\right)~{},$ (21)
where $\varrho$ and $\bar{\varrho}$ denote the density matrices for active neutrinos and antineutrinos, respectively, and $\phi$ describes the quantum spin coherence between neutrinos and antineutrinos. Similar to the Dirac case, the density matrix for Majorana neutrinos, ${\cal F}$, can be extracted from the two-point functions via
$-2\int_{0}^{\infty}\frac{dk^{0}}{2\pi}\hat{F}(k,x)=\mathcal{F}(\vec{k},x)-\frac{1}{2}\mathbb{1}\,.$ (22)
Beyond $\mathcal{O}(\epsilon^{0})$, the equations (19) and (22) serve as the definitions of density matrices for Dirac and Majorana neutrinos, respectively, and are used to extract the QKEs from the evolution equation of $\hat{F}$. Additionally, to obtain the QKEs at $\mathcal{O}(\epsilon^{2})$, only the $\mathcal{O}(\epsilon^{0})$ expression for the vector component of the spectral function is required, which is given by $\hat{\rho}(k)=2i\pi|\vec{k}|\delta(k^{2})\text{sgn}(k^{0})$.
2.2.2 Kinetic equation to $\mathcal{O}(\epsilon^{1})$
Beyond $\mathcal{O}(\epsilon^{0})$, we still consider $F_{L,R}(k,x)$ and $\Phi(k,x)$ as the four independent spinor components of $S^{F}(k,x)$, which can be isolated using the projections:
$F_{L,R}(k,x)$ $\equiv\frac{1}{4}\mathrm{Tr}\left(\gamma_{\mu}P_{L,R}F^{(\nu)}(k,x)\right)\hat{\kappa}^{\prime\mu}(k)\,,$ (23a)
$\Phi^{(\dagger)}(k,x)$ $\equiv\mp\frac{i}{16}\mathrm{Tr}\left(\sigma_{\mu\nu}P_{L/R}F^{(\nu)}(k,x)\right)(\hat{\kappa}^{\prime}(k)\wedge\hat{x}^{\pm}(k))^{\mu\nu}e^{\pm i\varphi(k)}\,,$ (23b)
where the upper (lower) signs and indices refer to $\Phi$ ($\Phi^{\dagger}$). The dispersion relations for the functions $F_{L,R}(k,x)$ and $\Phi(k,x)$, as well as for the other spinor components of $S^{F}(k,x)$, are modified by neutrino masses and interactions at $\mathcal{O}(\epsilon^{1})$. Therefore, the general statistical function $S^{F}(k,x)$ can be expressed as
$S^{F}$ $=\left[F_{R}\hat{\kappa}^{\mu}\gamma_{\mu}-\frac{i}{4}e^{-i\varphi(k)}(\hat{\kappa}(k)\wedge\hat{x}^{-}(k))_{\mu\nu}\Phi\sigma_{\mu\nu}\right]P_{L}$ (24)
$+\left[F_{L}\hat{\kappa}^{\mu}\gamma_{\mu}+\frac{i}{4}e^{i\varphi(k)}(\hat{\kappa}(k)\wedge\hat{x}^{+}(k))_{\mu\nu}\Phi^{\dagger}\sigma_{\mu\nu}\right]P_{R}$
$+\left[\Delta_{S}+(\Delta_{R})^{\mu}\gamma_{\mu}-\frac{i}{4}(\Delta_{T}^{L})^{\mu\nu}\sigma_{\mu\nu}\right]P_{L}$
$+\left[\Delta_{S}^{\dagger}+(\Delta_{L})^{\mu}\gamma_{\mu}+\frac{i}{4}(\Delta_{T}^{R})^{\mu\nu}\sigma_{\mu\nu}\right]P_{R}\,,$
where $\Delta$ represents the $\mathcal{O}(\epsilon)$ small components.
Expand the kinetic equation for $S^{F}(k,x)$ to $O(\epsilon)$, we obtain
[…]
Substituting the general statistical function $S^{F}(k,x)$, including the $O(\epsilon)$ corrections, into the kinetic equation and decomposing into scalar, vector, and tensor components, we derive the dispersion relations for $F_{L,R}(k,x)$ and $\Phi(k,x)$ at $O(\epsilon)$:
$\left(k\cdot\hat{\kappa}\right)F_{R}-\frac{1}{2}\left\{\Sigma_{L}^{\kappa},F_{R}\right\}$ $=O\left(\epsilon^{2}\right),$ (26)
$\left(k\cdot\hat{\kappa}\right)F_{L}-\frac{1}{2}\left\{\Sigma_{R}^{\kappa},F_{L}\right\}$ $=O\left(\epsilon^{2}\right),$ (27)
$\left(k\cdot\hat{\kappa}\right)P_{+}^{ij}F_{T}^{j}-\frac{1}{2}\left(\Sigma_{R}^{\kappa}P_{+}^{ij}F_{T}^{j}+P_{+}^{ij}F_{T}^{j}\Sigma_{L}^{\kappa}\right)$ $=O\left(\epsilon^{2}\right),$ (28)
$\left(k\cdot\hat{\kappa}\right)P_{-}^{ij}F_{T}^{j}-\frac{1}{2}\left(\Sigma_{L}^{\kappa}P_{-}^{ij}F_{T}^{j}+P_{-}^{ij}F_{T}^{j}\Sigma_{R}^{\kappa}\right)$ $=O\left(\epsilon^{2}\right),$ (29)
and the expressions for the small components
$\Delta_{L/R}^{\kappa}$ $=O\left(\epsilon^{2}\right)\ \ \ \ \,,\ \ \ \ \ \Delta_{T}^{i}=O\left(\epsilon^{2}\right),$ (30)
$\Delta_{S}$ $=\frac{1}{2\left|\vec{k}\right|}\left(m^{\dagger}F_{R}+F_{L}m^{\dagger}\right)+\frac{P_{+}^{ij}}{2\left|\vec{k}\right|}\left(i\partial^{i}F_{T}^{j}-\left(\Sigma_{R}^{i}F_{T}^{j}-F_{T}^{j}\Sigma_{L}^{i}\right)\right)$ (31)
$\Delta_{T}$ $=-\frac{1}{2\left|\vec{k}\right|}\left(m^{\dagger}F_{R}-F_{L}m^{\dagger}\right)+\frac{P_{+}^{ij}}{2\left|\vec{k}\right|}\left(\Sigma_{R}^{i}F_{T}^{j}+F_{T}^{j}\Sigma_{L}^{i}\right),$ (32)
$\Delta_{L}^{i}$ $=\frac{1}{2\left|\vec{k}\right|}\left(m^{\dagger}P_{-}^{ij}F_{T}^{j}+P_{+}^{ij}F_{T}^{j}m\right)+\frac{1}{\left|\vec{k}\right|}\left(\frac{1}{2}\epsilon^{ij}\partial^{j}F_{L}-\left(P_{-}^{ij}\Sigma_{R}^{j}F_{L}+F_{L}P_{+}^{ij}\Sigma_{R}^{j}\right)\right),$ (33)
$\Delta_{R}^{i}$ $=-\frac{1}{2\left|\vec{k}\right|}\left(mP_{+}^{ij}F_{T}^{j}+P_{-}^{ij}F_{T}^{j}m^{\dagger}\right)-\frac{1}{\left|\vec{k}\right|}\left(\frac{1}{2}\epsilon^{ij}\partial^{j}F_{R}+\left(P_{+}^{ij}\Sigma_{L}^{j}F_{R}+F_{R}P_{-}^{ij}\Sigma_{L}^{j}\right)\right).$ (34)
2.2.3 Kinetic equation to $\mathcal{O}(\epsilon^{2})$
At $O(\epsilon^{2})$, the kinetic equation $S^{F}(k,x)$ becomes
[…]
Inserting the expressions for the small components obtained at $\mathcal{O}(\epsilon)$ and reorganizing the kinetic equations, we can derive the kinetic equation and the dispersion relation for $\hat{F}$ as
$\partial^{\kappa}\hat{F}+\frac{1}{2|\vec{k}|}\left\{\Sigma^{i},\partial^{i}\hat{F}\right\}+\frac{1}{2}\left\{\frac{\partial\Sigma^{\kappa}}{\partial x^{\mu}},\frac{\partial\hat{F}}{\partial k_{\mu}}\right\}$ $=-i\left[H,\hat{F}\right]+\hat{C}~{},$ (36a)
$\left\{\hat{\kappa}(k)\cdot k-\Sigma^{\kappa}\ ,\ \hat{F}\right\}$ $=0~{},$ (36b)
where we have defined $\partial^{\kappa}\equiv\hat{\kappa}(k)\cdot\partial$ and $\partial^{i}\equiv\hat{x}^{i}(k)\cdot\partial$, and $H$ is the Hamiltonian-like operator given by
$H=\left(\begin{array}[]{cc}H_{R}&H_{LR}\\ H_{LR}^{\dagger}&H_{L}\end{array}\right)~{},\qquad$ (37)
with
$H_{R}$ $=\Sigma_{R}^{\kappa}+\frac{1}{2|\vec{k}|}\left(m^{\dagger}m-\epsilon^{ij}\partial^{i}\Sigma_{R}^{j}+4\Sigma_{R}^{+}\Sigma_{R}^{-}\right)\,,$ (38a)
$H_{L}$ $=\Sigma_{L}^{\kappa}+\frac{1}{2|\vec{k}|}\left(mm^{\dagger}+\epsilon^{ij}\partial^{i}\Sigma_{L}^{j}+4\Sigma_{L}^{-}\Sigma_{L}^{+}\right)\,,$ (38b)
$H_{LR}$ $=-\frac{1}{|\vec{k}|}\left(\Sigma_{R}^{+}\,m^{\dagger}-m^{\dagger}\,\Sigma_{L}^{-}\right)~{},$ (38c)
where $\Sigma_{L,R}^{\pm}\equiv(1/2)\,e^{\pm i\varphi}\,(x_{1}\pm ix_{2})_{\mu}\,\Sigma^{\mu}_{L,R}$ and $\epsilon^{ij}$ is the two-dimensional Levi-Civita symbol. Finally, the collision term $\hat{C}$ is given by
$\hat{C}=-\frac{1}{2}\left\{\hat{\Pi}^{>},{\hat{S}}^{<}\right\}+\frac{1}{2}\left\{\hat{\Pi}^{<},{\hat{S}}^{>}\right\}\,,\$ (39)
with
[…]
and the $2n_{f}\times 2n_{f}$ self energy matrices $\hat{\Pi}^{\pm}(k)$ are given by
$\hat{\Pi}^{\pm}(k,x)$ $=\left(\begin{array}[]{cc}\Pi_{R}^{\kappa\pm}&2P_{T}^{\pm}\\ 2P_{T}^{\pm\dagger}&\Pi_{L}^{\kappa\pm}\end{array}\right)~{}.$ (43)
where the components are obtained by the projections
$\Pi_{L,R}^{\kappa}(k,x)$ $=\frac{1}{2}\hat{\kappa}_{\mu}\,\mathrm{Tr}\left[\tilde{\Pi}(k,x)\,\gamma^{\mu}P_{L,R}\right]\,,$ (44a)
$P_{T}(k,x)$ $=\frac{ie^{i\varphi}}{16}\,(\hat{\kappa}\wedge\hat{x}^{+})^{\mu\nu}\,\,\mathrm{Tr}\left[\tilde{\Pi}(k,x)\,\sigma_{\mu\nu}P_{R}\right]~{}.$ (44b)
2.3 QKEs for Dirac and Majorana neutrinos
Finally, integrating Eq. (36a) over positive and negative frequencies using Eq. (19), we obtain the quantum kinetic equations (QKEs) for Dirac neutrinos:
$\partial^{\kappa}F+\frac{1}{2|\vec{k}|}\left\{\Sigma^{i},\partial^{i}F\right\}-\frac{1}{2}\left\{\frac{\partial\Sigma^{\kappa}}{\partial\vec{x}},\frac{\partial F}{\partial\vec{k}}\right\}$ $=-i[H,F]+{\cal C}\,,$ (45a)
$\partial^{\kappa}\bar{F}-\frac{1}{2|\vec{k}|}\left\{\Sigma^{i},\partial^{i}\bar{F}\right\}+\frac{1}{2}\left\{\frac{\partial\Sigma^{\kappa}}{\partial\vec{x}},\frac{\partial\bar{F}}{\partial\vec{k}}\right\}$ $=-i[\bar{H},\bar{F}]+\bar{\cal C}\,.$ (45b)
The terms on the left-hand side (LHS) of the QKEs generalize the classical space-time derivative, drift, and force terms. The first term on the right-hand side (RHS) of the QKEs describes neutrino flavor oscillations, which are induced by neutrino masses and forward scattering between neutrinos and background particles. This corresponds to the generalization of the well-known Mikheyev-Smirnov-Wolfenstein (MSW) effect. The second term on the RHS represents the collision term with a non-trivial matrix structure:
${\cal C}$ $=\frac{1}{2}\left\{\Pi^{>},F\right\}-\frac{1}{2}\left\{\Pi^{<},\mathbb{1}-F\right\}\,,$ (46a)
$\bar{\cal C}$ $=\frac{1}{2}\left\{\bar{\Pi}^{>},\bar{F}\right\}-\frac{1}{2}\left\{\bar{\Pi}^{<},\mathbb{1}-\bar{F}\right\}\,,$ (46b)
where
[…]
On the other hand, the QKEs for Majorana neutrinos can be derived by integrating Eq. (36a) over positive frequencies using Eq. (22), yielding:
$\partial^{\kappa}{\cal F}+\frac{1}{2|\vec{k}|}\left\{\Sigma^{i},\partial^{i}{\cal F}\right\}-\frac{1}{2}\left\{\frac{\partial\Sigma^{\kappa}}{\partial\vec{x}},\frac{\partial{\cal F}}{\partial\vec{k}}\right\}$ $=-i[H,{\cal F}]+{\cal C}_{M}\,,$ (49)
with
${\cal C}_{M}$ $=\frac{1}{2}\left\{\Pi^{>},{\cal F}\right\}-\frac{1}{2}\left\{\Pi^{<},\mathbb{1}-{\cal F}\right\}\,.$ (50)
Explicitly, the potentials induced by forward scattering $\Sigma_{R,L}$ and the collision terms for Dirac and Majorana neutrinos should be calculated from the 1-loop and 2-loop self-energy diagrams, depending on the interaction terms for the neutrinos. For active neutrinos in the early universe, we will present the resulting forward potentials and collision terms in the next section.
3 Neutrino Decoupling in the Presence of Primordial Neutrino Asymmetry
3.1 QKEs in the early universe
In the early universe, both the Standard Model (SM) thermal bath and the neutrino system can be approximated as homogeneous and isotropic. Under these conditions, the QKEs for neutrinos can be significantly simplified. First, the drift and force terms on the left-hand side are eliminated, and a Hubble term is introduced to account for the expansion of the Universe. Second, the density matrix for neutrinos is considered to depend solely on the absolute value of the momentum. Finally, the coherence matrices between left-handed and right-handed neutrinos, namely $\varrho_{LR,RL}$ and $\bar{\varrho}_{LR,RL}$ for Dirac neutrinos, and $\phi$ for Majorana neutrinos, are reduced to zero, leading to corresponding collision terms also being zero. Consequently, the non-zero density matrices are $\varrho_{LL,RR}$ and $\bar{\varrho}_{LL,RR}$ for Dirac neutrinos, and $\varrho$ and $\bar{\varrho}$ for Majorana neutrinos.
For active neutrinos, the QKEs for $\varrho_{LL}$ and $\bar{\varrho}_{RR}$ in the case of Dirac neutrinos, are identical with those for $\varrho$ and $\bar{\varrho}$ for Majorana neutrinos. Therefore, in the subsequent analysis, we will not distinguish between Dirac and Majorana neutrinos and will use $\varrho$ and $\bar{\varrho}$ to represent the density matrices for neutrinos and antineutrinos, respectively. In the flavor basis, the QKEs can be written as:
$i\left[\frac{\partial}{\partial t}-Hp\frac{\partial}{\partial p}\right]\varrho$ $=\Big{[}H_{\nu},\varrho\Big{]}+i\mathcal{I}\,,$ (51)
$i\left[\frac{\partial}{\partial t}-Hp\frac{\partial}{\partial p}\right]\bar{\varrho}$ $=-\Big{[}\bar{H}_{\nu},\bar{\varrho}\Big{]}+i\bar{\mathcal{I}}\,,$ (52)
where $H\equiv\dot{a}/a$ is the Hubble rate, and $H_{\nu}$ and $\bar{H}_{\nu}$ are the Hamiltonian terms for neutrinos and antineutrinos, defined as:
$H_{\nu}$ $=U\frac{\mathbb{M}^{2}}{2p}U^{\dagger}+\sqrt{2}G_{F}\left(\mathbb{N}_{l}+\mathbb{N}_{\nu}\right)-2\sqrt{2}G_{F}p\left(\frac{\mathbb{E}_{l}+\mathbb{P}_{l}}{m_{W}^{2}}+\frac{4}{3}\frac{\mathbb{E}_{\nu}}{m_{Z}^{2}}\right)\,,$ (53)
$\bar{H}_{\nu}$ $=U\frac{\mathbb{M}^{2}}{2p}U^{\dagger}-\sqrt{2}G_{F}\left(\mathbb{N}_{l}+\mathbb{N}_{\nu}\right)-2\sqrt{2}G_{F}p\left(\frac{\mathbb{E}_{l}+\mathbb{P}_{l}}{m_{W}^{2}}+\frac{4}{3}\frac{\mathbb{E}_{\nu}}{m_{Z}^{2}}\right)\,.$ (54)
Here, $\mathbb{M}$ represents the neutrino mass matrix in the mass basis, and $U$ is the PMNS transformation matrix that relates the flavor eigenstates to the mass eigenstates. Using the standard parametrization of the PMNS matrix Giunti:2007ry ; Gariazzo:2019gyi , we have:
$\mathbb{M}=\begin{pmatrix}m_{1}^{2}&0&0\\ 0&m_{2}^{2}&0\\ 0&0&m_{3}^{2}\end{pmatrix}\,,\quad U=\begin{pmatrix}c_{12}c_{13}&s_{12}c_{13}&s_{13}\\ -s_{12}c_{23}-c_{12}s_{23}s_{13}&c_{12}c_{23}-s_{12}s_{23}s_{13}&s_{23}c_{13}\\ s_{12}s_{23}-c_{12}c_{23}s_{13}&-c_{12}s_{23}-s_{12}c_{23}s_{13}&c_{23}c_{13}\end{pmatrix}\,,$ (55)
where $c_{ij}=\cos{\theta_{ij}}$, $s_{ij}=\sin{\theta_{ij}}$, and $\theta_{ij}$ are the mixing angles. We neglect the potential CP-violating phase in this work.
The remaining terms in the Hamiltonian account for the matter potentials induced by coherent scatterings between neutrinos and other plasma particles. Here, $m_{W}$ and $m_{Z}$ denote the masses of the $W$ and $Z$ bosons, respectively, and $G_{F}$ is the Fermi constant. The second term in the Hamiltonian reflects the net number densities for charged leptons ($\mathbb{N}_{l}\equiv\mathrm{diag}(n_{e}-n_{\bar{e}},n_{\mu}-n_{\bar{\mu}},0)$) and neutrinos ($\mathbb{N}_{\nu}\equiv n_{\nu}-n_{\bar{\nu}}$). The third term accounts for the energy densities of charged leptons and neutrinos ($\mathbb{E}_{l}\equiv\mathrm{diag}(\rho_{e}+\rho_{\bar{e}},\rho_{\mu}+\rho_{\bar{\mu}},0)$, $\mathbb{P}_{l}\equiv\mathrm{diag}(P_{e}+P_{\bar{e}},P_{\mu}+P_{\bar{\mu}},0)$, $\mathbb{E}_{\nu}\equiv\rho_{\nu}+\rho_{\bar{\nu}}$). In practice, the net number densities for charged leptons are often negligible due to the small baryon asymmetry. However, the net number densities for neutrinos can lead to significant effects, especially if primordial neutrino asymmetries are present. This can result in pronounced neutrino flavor oscillations, including the well-known "synchronized oscillation" when this term dominates the Hamiltonian Bell:1998ds ; Pastor:2001iu ; Abazajian:2002qx ; Wong:2002fa ; Johns:2016enc ; Froustey:2021azz ; Froustey:2024mgf .
Finally, the collision terms $\mathcal{I}$ and $\bar{\mathcal{I}}$ account for contributions from neutrino-electron/positron scattering ($\nu e\leftrightarrow\nu e$), annihilation ($\nu\bar{\nu}\leftrightarrow e^{+}e^{-}$), and neutrino-neutrino self-interactions ($\nu\nu\leftrightarrow\nu\nu$, $\nu\bar{\nu}\leftrightarrow\nu\bar{\nu}$). These terms can be expressed as
$\mathcal{I}\equiv\mathcal{I}_{\rm sc}+\mathcal{I}_{\rm ann}+\mathcal{I}_{\nu\nu}$ (56)
with a similar decomposition for the collision term of antineutrinos $\bar{\mathcal{I}}$. The collision terms are generally functions of the neutrino density matrices, and their explicit forms are detailed in Appendix A. Contributions from other interaction processes, such as $\mu^{\pm}$ scattering and annihilation, are neglected due to their minimal impact at the temperature of neutrino decoupling.
In practice, comoving coordinates and corresponding comoving variables are employed to eliminate the Hubble expansion term in the QKEs. The comoving temperature is defined as $T_{\rm CM}\propto a^{-1}$ Grohs:2015tfy , with $T_{\rm CM}=T_{\nu}=T_{\gamma}$ at the initial time before neutrino decoupling when all species are tightly coupled. Using the comoving temperature, the reduced scale factor, the comoving momentum, and the dimensionless photon temperature are expressed as Mangano:2005cc ; Froustey:2020mcq ; Bennett:2020zkv
$x=m_{e}/T_{\rm CM}\,,\qquad y=p/T_{\rm CM}\,,\qquad z=T_{\gamma}/T_{\rm CM}\,.$ (57)
By definition, the dimensionless photon temperature $z$ equals 1 at high temperatures and increases beyond 1 later due to the annihilation of electrons and positrons into photons. For the convenience of the QKEs, we also define the dimensionless energy density and pressure as $\tilde{\rho}\equiv(x/m_{e})^{4}\rho$ and $\tilde{P}\equiv(x/m_{e})^{4}P$, respectively.
The comoving QKEs for neutrinos and antineutrinos then take the form as:
$\frac{{\rm d}\varrho(x,y)}{{\rm d}x}$ $=$ $\sqrt{\frac{3m^{2}_{\rm Pl}}{8\pi\tilde{\rho}_{\rm tot}}}\left\{-i\frac{x^{2}}{m_{e}^{3}}\left[\tilde{H}_{\nu},\varrho\right]+\frac{m_{e}^{3}}{x^{4}}\tilde{\mathcal{I}}(\varrho,\bar{\varrho})\right\}\,,$
$\frac{{\rm d}\bar{\varrho}(x,y)}{{\rm d}x}$ $=$ $\sqrt{\frac{3m^{2}_{\rm Pl}}{8\pi\rho}}\left\{-i\frac{x^{2}}{m_{e}^{3}}\left[\tilde{H}_{\bar{\nu}},\varrho\right]+\frac{m_{e}^{3}}{x^{4}}\tilde{\bar{\mathcal{I}}}(\bar{\varrho},\varrho)\right\}\,,$
with the comoving Hamiltonian terms
$\tilde{H}_{\nu}$ $=U\frac{\mathbb{M}^{2}}{2y}U^{\dagger}+\frac{\sqrt{2}G_{\rm F}m_{e}^{4}}{x^{4}}\,\mathbb{N}_{\nu}-\frac{2\sqrt{2}G_{\rm F}ym_{e}^{6}}{x^{6}}\left(\frac{\mathbb{E}_{\ell}+\mathbb{P}_{\ell}}{m_{W}^{2}}+\frac{4}{3}\,\frac{\mathbb{E}_{\nu}}{m_{Z}^{2}}\right)\,,$ (58)
$\tilde{H}_{\bar{\nu}}$ $=U\frac{\mathbb{M}^{2}}{2y}U^{\dagger}-\frac{\sqrt{2}G_{\rm F}m_{e}^{4}}{x^{4}}\,\mathbb{N}_{\nu}-\frac{2\sqrt{2}G_{\rm F}ym_{e}^{6}}{x^{6}}\left(\frac{\mathbb{E}_{\ell}+\mathbb{P}_{\ell}}{m_{W}^{2}}+\frac{4}{3}\,\frac{\mathbb{E}_{\nu}}{m_{Z}^{2}}\right)\,.$
Additionally, $\tilde{\mathcal{I}}$ and $\tilde{\bar{\mathcal{I}}}$ represent the comoving collision terms for neutrinos and antineutrinos, respectively. The computation of these collision terms is the most complex and time-consuming aspect of solving the QKEs. Fortunately, the nine-dimensional collision integrals can be simplified to two-dimensional integrals by exploiting the homogeneous and isotropic conditions and the specific forms of the scattering amplitudes Hannestad:1995rs ; Dolgov:1997mb ; Grohs:2015tfy ; Blaschke:2016xxt . The explicit forms of the reduced collision terms are provided in Appendix A.
Furthermore, as noted in the discussion following the QKEs (Eq. (53)), the chemical potential of electrons and the $\mathbb{N}_{l}$ term in the Hamiltonian are negligible due to ue to the very low baryon-to-photon ratio $\eta$ and the efficient electromagnetic interactions. Consequently, the Hamiltonian is predominantly influenced by the net number densities of neutrinos when primordial lepton asymmetries are sufficiently large and vary among different flavors. This results in synchronized flavor oscillations of neutrinos and antineutrinos Bell:1998ds ; Pastor:2001iu ; Abazajian:2002qx ; Wong:2002fa . Consequently, neutrino asymmetries tend to equilibrate among different flavors prior to the epoch of BBN Froustey:2021azz ; Froustey:2024mgf . Since this work focuses on the implications of neutrino asymmetries for BBN, CMB, and LSS, we will concentrate on flavor-equilibrated primordial asymmetries, i.e., the initial primordial asymmetries are set to $\xi_{\nu_{e}}=\xi_{\nu_{\mu}}=\xi_{\nu_{\tau}}=\xi_{n}$, where $\xi_{\nu_{\alpha}}\equiv\mu_{\nu_{\alpha}}/T_{\rm CM}$ is the degeneracy parameter for neutrinos and antineutrinos of flavor $\alpha$. By adopting this initial condition, we bypass the intricate mapping between arbitrary primordial neutrino asymmetries and those after flavor equilibration, leaving the investigation of this mapping for future work. Recent study on such mapping and the implications of distinct neutrino asymmetries among flavors for BBN can be found in  Froustey:2024mgf .
On the other hand, it is well-known that the annihilation of electrons and positrons shortly after neutrino decoupling partially heats neutrinos, leading to a slightly increased $N_{\rm eff}$. To accurately determine $N_{\rm eff}$, finite-temperature QED (FTQED) corrections must be considered Fornengo:1997wa ; Mangano:2001iu ; Bennett:2019ewm . In this work, we address three aspects of FTQED corrections: (1) Corrections to the total energy pressure densities, which primarily affect the Hubble rate; (2) Corrections to the electron mass, which primarily influence the collision terms in the QKEs; (3) Corrections to the continuity equation of the Universe, typically expressed as a differential equation for the dimensionless photon temperature $z$,
$\frac{\mathrm{d}z}{\mathrm{d}x}=\frac{rJ_{2}(r)+G_{1}(r)-\frac{1}{4\pi^{2}z^{3}}\int_{0}^{\infty}\mathrm{d}y\,y^{3}\sum_{\alpha}\left(\frac{\mathrm{d}\varrho_{\alpha\alpha}}{\mathrm{d}x}+\frac{\mathrm{d}\bar{\varrho}_{\alpha\alpha}}{\mathrm{d}x}\right)}{\left[r^{2}J_{2}(r)+J_{4}(r)\right]+G_{2}(r)+\frac{2\pi^{2}}{15}}\,,$ (59)
where $r=x/z$, $\frac{\mathrm{d}\varrho_{\alpha\alpha}}{\mathrm{d}x}$, and $\frac{\mathrm{d}\bar{\varrho}_{\alpha\alpha}}{\mathrm{d}x}$ are the diagonal components of the QKEs for neutrinos and antineutrinos. The explicit forms of the remaining $J_{i}(r)$ and $G_{i}(r)$ functions appearing in the equation, as well as other FTQED corrections, can be found in Appendix B. In practice, this equation should be solved in conjunction with the QKEs, and we consider corrections up to order $\mathcal{O}(e^{3})$ while neglecting logarithmic terms in the FTQED corrections as their effects are minimal Bennett:2019ewm ; Bennett:2020zkv ; further details are provided in Appendix B. Other FTQED corrections, such as those to the interaction vertices, are omitted in this work.
Furthermore, we restrict our analysis to the normal mass hierarchy of neutrinos and neglect any possible CP-violating phase in the transformation matrix between the neutrino flavor and mass bases. Based on the most recent values from the Particle Data Group pdg , we adopt the following neutrino mass differences and mixing angles: $\Delta m_{21}^{2}=7.53\times 10^{-5}\,\text{eV}^{2}$, $\Delta m_{31}^{2}=2.53\times 10^{-3}\,\text{eV}^{2}$, $\sin^{2}\theta_{12}=0.307$, $\sin^{2}\theta_{23}=0.545$, and $\sin^{2}\theta_{13}=0.0218$. The results for neutrino decoupling and BBN are independent of the minimal neutrino mass, provided $m_{1}\ll 0.1\,\text{MeV}$. On the other hand, the resulting CMB anisotropies and LSS of the Universe are sensitive to the choice of $m_{1}$, which will be discussed in more detail in Sections 5 and 6.
3.2 Numerical methods
To solve the QKEs for neutrinos and antineutrinos numerically, we have extended the public code FortEPiaNO Gariazzo:2019gyi ; Bennett:2020zkv , initially designed for solving the QKEs in the Standard Model case, to accommodate primordial neutrino asymmetries. This subsection outlines the key numerical methods employed in the code.
For the momentum grid, the comoving momentum $y$ is discretized on a linear grid with $y_{\rm min}=0.01$, $y_{\rm max}=20$, and a total of $N_{y}=40$ points. The integrals are computed using the Newton-Cotes (NC) formula of order 1. The initial conditions are specified with an initial reduced scale factor $x=0.005$, corresponding to a comoving temperature $T_{\rm CM,in}\simeq 102.2\text{ MeV}$. The initial common temperature for all species is computed by conserving total entropy from an even higher temperature, yielding $z_{\rm in}=2.81\times 10^{-6}$. The initial density matrices are given by
$\varrho(x_{\rm in},y)={\rm diag}\{f_{\nu}^{({\rm in})}(y),f_{\nu}^{({\rm in})}(y),f_{\nu}^{({\rm in})}(y)\},\quad\bar{\varrho}(x_{\rm in},y)={\rm diag}\{f_{\bar{\nu}}^{({\rm in})}(y),f_{\bar{\nu}}^{({\rm in})}(y),f_{\bar{\nu}}^{({\rm in})}(y)\}.$ (60)
where
$f_{\nu}^{({\rm in})}(y)\equiv\frac{1}{e^{(y-\xi_{\nu})/z_{\rm in}}+1}\quad\text{and}\quad f_{\bar{\nu}}^{({\rm in})}(y)\equiv\frac{1}{e^{(y+\xi_{\nu})/z_{\rm in}}+1}\,.$ (61)
Finally, the QKEs are solved together with the energy conservation equation for $z$ using the DLSODA method from the ODEPACK hindmarsh1982odepack Fortran library. The absolute and relative error tolerances are set to $10^{-7}$, ensuring that the numerical errors of the results remain below $10^{-5}$.
3.3 Numerical results of neutrino decoupling
We present the numerical results of neutrino decoupling in this subsection. First, we consider the final energy density of neutrinos, which is typically parameterized by the effective number of relativistic species $N_{\rm eff}$, defined as
$N_{\rm eff}\equiv\frac{87}{8}\left(\frac{11}{4}\right)^{4/3}\frac{\rho_{\nu}}{\rho_{\gamma}}\,,\quad\rho_{\nu}=\sum_{\alpha}\int\frac{d^{3}p}{(2\pi)^{3}}\,p\,(\varrho_{\nu_{\alpha}}+\bar{\varrho}_{\nu_{\alpha}})$ (62)
where $\rho_{\nu}$ represents the total neutrino energy density obtained by solving the QKEs and $\rho_{\gamma}$ is the photon energy density. For the SM in the instantaneous decoupling limit, $N_{\rm eff}$ equals the number of neutrino flavors. However, since the temperature of neutrino decoupling is close to that of electron-positron annihilation, neutrinos are also heated through weak interactions with electrons and positrons, resulting in a slightly larger $N_{\rm eff}$. The most accurate calculation for the SM $N_{\rm eff}$ to date, which includes neutrino oscillations, matter effects, FTQED corrections up to $\mathcal{O}(3)$ order, and neutrino-electron and neutrino-neutrino collision terms, yields $N_{\rm eff}^{\rm SM}=3.0440\pm 0.0002$ Froustey:2020mcq ; Bennett:2020zkv .
Figure 2: Left panel: Results of the neutrino non-instantaneous decoupling correction $\delta N_{\rm eff}^{\rm dec}=N_{\rm eff}-N_{\rm eff}^{\rm inst}$ as a function of the neutrino degeneracy parameter $\xi_{\nu}$. The solid and dashed curves represent the results of the numerical solution and the fitting function Eq. (63), respectively. Right panel: Relative differences in the number densities of electron neutrinos and antineutrinos compared to the SM predictions in the instantaneous decoupling limit, $n_{\nu_{e},\bar{\nu}_{e}}\simeq 56\,\text{cm}^{-3}$.
In the presence of non-zero primordial neutrino asymmetries, $N_{\rm eff}$ receives additional contributions from the neutrino chemical potentials. In the instantaneous decoupling limit, the comoving distribution of neutrinos and antineutrinos of flavor $\alpha$ can be approximated as $f_{\nu_{\alpha},\bar{\nu}_{\alpha}}=\frac{1}{e^{y\,\pm\,\xi_{\nu_{\alpha}}}+1}$. Therefore, the total $N_{\rm eff}$ can be estimated by
$N_{\rm eff}^{\rm inst}\simeq 3+\sum\limits_{\alpha}\left(\frac{30}{7\pi^{2}}\xi_{\nu_{\alpha}}^{2}+\frac{15}{7\pi^{4}}\xi_{\nu_{\alpha}}^{4}\right)=3+3\left(\frac{30}{7\pi^{2}}\xi_{\nu}^{2}+\frac{15}{7\pi^{4}}\xi_{\nu}^{4}\right),$ (63)
where we have used the condition $\xi_{\nu_{e}}=\xi_{\nu_{\mu}}=\xi_{\nu_{\tau}}=\xi_{\nu}$ in the second equality.
[…]
$N_{\rm eff}(\xi_{\nu})=N_{\rm eff}^{\rm SM}+3\left(\frac{30}{7\pi^{2}}\xi_{\nu}^{2}+\frac{15}{7\pi^{4}}\xi_{\nu}^{4}\right)+0.0102\,\xi_{\nu}^{2},$ (64)
where $N_{\rm eff}^{\rm dec,SM}=3.0440\pm 0.0002$.
[…]
Figure 3: Density matrices of neutrinos and antineutrinos in the flavor basis (left panel) and mass basis (right panel), normalized to the Fermi-Dirac distribution with $\xi_{\nu}=0$. The gray curves represent the density matrices for neutrinos and antineutrinos with $\xi_{\nu}=0$, while the red and green curves represent those for neutrinos and antineutrinos with non-zero $\xi_{\nu}$, respectively. For each case, the different flavor eigenstates ($\nu_{e}$, $\nu_{\mu}$, and $\nu_{\tau}$) and mass eigenstates ($\nu_{1}$, $\nu_{2}$, and $\nu_{3}$) are indicated by solid, dashed, and dash-dotted curves, respectively. Note that the density matrices for $\nu_{\mu}$ and $\nu_{\tau}$ nearly overlap. The Fermi-Dirac distributions with corresponding $\xi_{\nu}$ are also indicated by the gray dotted curves.
Next we present the resulting density matrices for neutrinos and antineutrinos after decoupling. The density matrices as a function of the comoving momentum $y\equiv p/T_{\rm CM}$ in the flavor basis and the mass basis for several choices of $\xi_{\nu}$ are shown in the left and right panels of Fig. 3 respectively, normalized to the Fermi-Dirac distribution with $\xi_{\nu}=0$. The density matrices of neutrinos and antineutrinos with $\xi_{\nu}=0$ coincide with each other and are represented by the gray curves. Our results also reproduce the previous density matrices obtained in the literature. For each choice of In the case of non-zero $\xi_{\nu}$, the density matrices of neutrinos and antineutrinos no longer coincide and are represented by red and green curves, respectively. For each case we use the solid, dashed and dash-dotted curve to indicate the three flavor eigenstates ($\nu_{e}$, $\nu_{\mu}$ and $\nu_{\tau}$) and the three mass eigenstates ($\nu_{1}$, $\nu_{2}$ and $\nu_{3}$), respectively. For each choice of the $\xi_{\nu}$, the corresponding thermal equilibrium distributions for neutrinos and antineutrinos $f_{\rm FD}^{\nu,\bar{\nu}}=1/(e^{y\,\pm\,\xi_{\nu}}+1)$ are also plotted by the dotted curves. Clearly, the density distributions of neutrinos and antineutrinos for the given $\xi_{\nu}$ coincide with the corresponding Fermi-Dirac distributions when $y$ is small, while the spectra are distorted at large $y$. This is due to the fact that the weak interaction rate between neutrinos and electrons/positrons is proportional to the neutrino momentum, so that the neutrino with smaller momentum decouples earlier, preserving the thermal Fermi-Dirac distribution, while the neutrino with larger momentum decouples later, resulting in a larger spectral distortion with respect to the Fermi-Dirac distribution. Regarding the density distributions for different flavors, we see that electron neutrinos possess larger spectral distortions due to the additional charge-current weak interaction, while the spectral distortions of muon neutrinos almost coincide with those of tauon neutrinos, since they have the same weak interactions. Instead, the spectral distortions of the three neutrino mass eigenstates are uniformly separated, as shown in the right panel.
Figure 4: The spectral distortions of neutrinos and antineutrinos in the flavor basis (top left) and mass basis (lower left) with respect to the Fermi-Dirac distribution with corresponding $\xi_{\nu}$. The meaning of the curves is the same as in the previous plot Fig. 3, while the spectral distortions of tauon neutrino in the left panel are not shown for brevity. The zoomed version of the spectral distortions in the flavor and mass basis are shown in the right panel.
Considering that the density distributions of neutrinos and antineutrinos coincide with the corresponding thermal Fermi-Dirac distributions for small momentum, it is natural to define the spectral distortions as $\delta\varrho_{\nu,\bar{\nu}}\equiv\delta\varrho_{\nu,\bar{\nu}}/f_{FD}(\pm\xi)-1$, where the plus sign corresponds to neutrinos and the minus sign to antineutrinos. The resulting spectral distortions as a function of the momentum $y$ in the flavor and mass basis are presented in the upper left and upper right panels of Fig. 4, respectively. The meaning of the curves is the same as in the previous plot Fig. 3, while the spectral distortions of the tauon neutrinos are not shown in the left panels, since they almost coincide with the spectral distortions of the muon neutrinos. We also plot the zoomed version of the spectral distortions in the lower panels. Interestingly, the spectral distortions for neutrinos and antineutrinos with non-zero $\xi_{\nu}$ differ from each other and lie on different sides of the spectral distortions with $\xi_{\nu}=0$. The antineutrinos with non-zero $\xi_{\nu}$ possess a relatively larger spectral distortion than that with $\xi_{\nu}=0$, while the neutrinos possess a relatively smaller spectral distortion. In addition, the larger the $\xi_{\nu}$, the larger the discrepancy between the spectral distortions of neutrinos and antineutrinos, clearly seen from the zoomed plots. Such properties are both valid in the flavor basis and in the mass basis. Finally, the above conclusions about neutrino density matrices also hold for negative $\xi_{\nu}$, except that the density matrices for neutrinos and antineutrinos with $-\xi_{\nu}$ are exchanged comparing with those for $\xi_{\nu}$. In other words, the density distributions and spectral distortions for antineutrinos in Fig. 3 and Fig. 4 are also the density distributions and spectral distortions for neutrinos with the opposite $\xi_{\nu}$.
As a brief conclusion of this section, we have presented the most accurate results of neutrino decoupling in the presence of primordial neutrino asymmetries and discussed the resulting $N_{\rm eff}$ and the spectral distortions of neutrinos and antineutrinos. In the next sections, we will study the effects of primordial neutrino asymmetries on the processes of BBN, CMB and structure formation, while emphasizing the corrections induced by the spectral distortions of neutrinos and antineutrinos.
4 Impacts of the Primordial Neutrino Asymmetry on BBN
Soon after neutrino decoupling, there is an epoch of primordial nucleosynthesis or the so-called Big Bang Nucleosynthesis (BBN), where the light elements are produced after protons and neutrons decouple from the SM thermal bath. Since the decoupling of protons depends on electron neutrinos, the resulting abundances of light elements after BBN depends on the properties of electron neutrinos and antineutrinos. BBN is affected by primordial neutrino asymmetries in three ways as follow.
  1. 1.
    The leading effect is the change in the weak interaction rate at the onset of BBN due to the chemical potential of the electron neutrinos, since the conversion between proton and neutron is related to the electron neutrino through the weak interaction, so that the difference of the chemical potentials of proton and neutron is decided by the chemical potential of electron neutrino, $i.e.,$ $\mu_{n}-\mu_{p}=\mu_{e}-\mu_{\nu_{e}}\simeq-\mu_{\nu_{e}}$. Therefore, the resulting neutron-to-proton ratio and the subsequent BBN processes are modified according to the absolute value of the electron neutrino chemical potential and its sign.
  2. 2.
    In addition, the expansion rate of the universe during BBN is changed by the additional neutrino energy density contributed by the chemical potentials. As we have seen in the last section, the neutrino energy density after neutrino decoupling is usually represented by $N_{\rm eff}$, and we have $N_{\rm eff}\simeq 3.044+3\left(\frac{30}{7\pi^{2}}\xi_{\nu}^{2}+\frac{15}{7\pi^{4}}\xi_{\nu}^{4}\right)$. The corresponding correction in BBN therefore depends on the absolute asymmetries of all three neutrino flavors.
  3. 3.
    Beyond the approximation of the thermal Fermi-Dirac distributions for neutrinos and antineutrinos, the spectral distortions resulting from the non-instantaneous decoupling of the neutrinos will also induce a correction of the weak interaction rate and thus affect the BBN. Such corrections are usually neglected in previous studies, especially in the presence of primordial neutrino asymmetries.
In the following, we first review the general treatment of the weak interaction rate and introduce the new correction induced by the neutrino spectral distortions, then we discuss the resulting light element abundances from the BBN.
4.1 The weak correction from neutrino spectral distortions
At the onset of the BBN, the conversion between neutron ($n$) and proton ($p$) is dominated by following interaction reactions,
$n+\nu_{e}$ $\leftrightarrow p+e^{-},$ (65)
$n$ $\leftrightarrow p+e^{-}+\bar{\nu}_{e},$ (66)
$n+e^{+}$ $\leftrightarrow p+\bar{\nu}_{e}.$ (67)
In the remainder of this section, we will omit the subscript for the electron-neutrino and the superscript for the electron/positron as long as there is no ambiguity. Due to the forceful coupling between photons and electrons and positrons, it is appropriate to assume that all particles expect for neutrinos are in local thermal equilibrium. Therefore, the neutron and proton number densities obey the Boltzmann equations
$\dot{n}_{n}+3Hn_{n}$ $=-n_{n}\Gamma_{n\to p}+n_{p}\Gamma_{p\to n},$ (68)
$\dot{n}_{p}+3Hn_{p}$ $=-n_{p}\Gamma_{p\to n}+n_{n}\Gamma_{n\to p},$ (69)
where the reaction rates $n_{n}\Gamma_{n\to p}$ and $n_{p}\Gamma_{p\to n}$ incorporate all relevant collision terms
$\Gamma_{n\to p}$ $=$ $\Gamma_{n+\nu\to p+e}+\Gamma_{n\to p+e+\bar{\nu}}+\Gamma_{n+e\to p+\bar{\nu}},$ (70)
$\Gamma_{p\to n}$ $=$ $\Gamma_{p+e\to n+\nu}+\Gamma_{p+e+\bar{\nu}\to n}+\Gamma_{p+\bar{\nu}\to n+e}.$ (71)
And the reaction rate for a specific interaction reaction with the nucleon $a$ in the initial state and nucleon $b$ in the final state takes the form as
$n_{a}\Gamma_{a\to b}=$ $\int\mathrm{d}\pi_{a}\mathrm{d}\pi_{b}\mathrm{d}\pi_{e}\mathrm{d}\pi_{\nu}\delta^{(4)}\left({p}_{a}-p_{b}+\alpha_{e}p_{e}+\alpha_{\nu}p_{\nu}\right)\left|M\right|^{2}_{a\to b}{}$ (72)
×$\times f_{a}(E_{a})f_{b}(-E_{b})f_{e}(\alpha_{e}{E}_{e})\tilde{f}_{\nu}(E_{\nu}),$
where $\mathrm{d}\pi_{i}=\frac{1}{(2\pi)^{3}}\frac{\mathrm{d}^{3}\mathbf{p}_{i}}{2E_{i}}$, $\alpha_{e}(\alpha_{\nu})=1$ if the electron (neutrino) is in the initial state and $\alpha_{e}(\alpha_{\nu})=-1$ if the electron (neutrino) is in the final state. In addition, $f_{i}(E)\equiv\frac{1}{e^{E/T_{\gamma}}+1}$ for $i=a,b,e$ is the Fermi-Dirac distribution with vanishing chemical potential, and we have used the relation $1-f_{i}(E)=f_{i}(-E)$ to incorporate the corresponding Pauli blocking factors. Lastly, $\tilde{f}_{\nu}(E_{\nu})$ equals $\varrho_{\nu}(E_{\nu})\,(\varrho_{\bar{\nu}}(E_{\nu}))$ for the neutrino (antineutrino) in the initial state and equals $1-\varrho_{\nu}(E_{\nu})\,(1-\varrho_{\bar{\nu}}(E_{\nu}))$ for the neutrino (antineutrino) in the initial state.
[…]
$\bar{\Gamma}_{n\to p}$ $=K\int_{0}^{\infty}p^{2}\mathrm{d}p[\chi_{+}(E)+\chi_{+}(-E)]\,,$ (73)
$\bar{\Gamma}_{p\to n}$ $=K\int_{0}^{\infty}p^{2}\mathrm{d}p[\chi_{-}(E)+\chi_{-}(-E)]\,,$ (74)
with $E=\sqrt{p^{2}+m_{e}^{2}}$ is the energy of the electron and $K=1/(\tau_{n}\lambda_{0}m_{e}^{5})$, where $\tau_{n}$ is the life time of the neutron and $\lambda_{0}$ the Born coefficient. The $\chi^{\pm}$ function is defined as
$\chi_{\pm}(E)\equiv(E_{\nu}^{\mp})^{2}f_{e}(-E)f_{\nu}^{\pm}(E^{\mp}_{\nu})\,,\,\,E^{\mp}_{\nu}\equiv E\mp\Delta,$ (75)
where $\Delta\equiv m_{n}-m_{p}$ is the energy gap between neutron and proton. The first term in Eq. (73) corresponds to the $n\to p+e$ processes Eq. (65) or Eq. (66) depending on whether $E-\Delta$ is positive or negative, and the second term corresponds to the $n+e\to p$ process Eq. (67) and similarly for the terms in Eq. (74). Note that here $E^{\mp}_{\nu}$ is equal to the energy of the neutrino $E_{\nu}$ for the processes where the neutrino is in the initial state, or its opposite $-E_{\nu}$ for the processes where the neutrino is in the final state. And the term $f_{\nu}^{\pm}(E^{\mp}_{\nu})$ in the function $\chi^{\pm}$ gives the corresponding distribution or the Pauli-blocking factor by the relation $1-f_{\nu}^{+}(E_{\nu})=f_{\nu}^{-}(-E_{\nu})$.
Now we introduce the corrections induced by the spectral distortions of neutrinos and antineutrinos. In the infinite nucleon mass limit, the corrections to the Born rates can be written as
$\delta\bar{\Gamma}_{n\to p}^{\rm SD}$ $=K\int_{0}^{\infty}p^{2}\mathrm{d}p[\delta\chi_{+}(E)+\delta\chi_{+}(-E)]\,,$ (76)
$\delta\bar{\Gamma}_{p\to n}^{\rm SD}$ $=K\int_{0}^{\infty}p^{2}\mathrm{d}p[\delta\chi_{-}(E)+\delta\chi_{-}(-E)]\,,$ (77)
with
$\delta\chi_{\pm}(E)\equiv$ $(E_{\nu}^{\mp})^{2}f_{e}(-E){\rm sgn}(E^{\mp}_{\nu}){}$ (78)
×$\times\left(\frac{1+\Theta(\pm E^{\mp}_{\nu})\delta\varrho_{\nu}+\Theta(\mp E^{\mp}_{\nu})\delta\varrho_{\bar{\nu}}}{e^{\left|E^{\mp}_{\nu}\right|/T_{\rm CM}\mp{\rm sgn}(E^{\mp}_{\nu})\xi_{\nu}}+1}-\frac{1}{e^{\left|E^{\mp}_{\nu}\right|/T_{\nu}\mp{\rm sgn}(E^{\mp}_{\nu})\xi_{\nu}}+1}\right)\,,$
where $\Theta(x)$ is the Heaviside step function, $\delta\varrho_{\nu}$ and $\delta\varrho_{\bar{\nu}}$ are the spectral distortions of electron neutrinos and antineutrinos. For a particular process, the various sgn functions and the $\Theta$ functions together represent the nature and the position of the neutrino particle: Whether it is a neutrino or an antineutrino, and whether it is in the initial state or the final state. We also note that $T_{\rm CM}$ and $T_{\nu}$ has slight difference after neutrino decoupling. If we neglect such difference and set $\xi_{\nu}=0$, so that $\delta\varrho_{\nu}=\delta\varrho_{\bar{\nu}}$, then Eq. (78) coincides with Eq. (29) in Ref. Froustey:2019owm , where they considered the corrections to the weak rates from neutrino spectral distortions in the SM.
Finally, we briefly review the other corrections to the weak interaction rates that we have considered in this work. First are the zero temperature radiative corrections, including the so-called pure radiative corrections corresponding to the emission and absorption of a virtual photon in the interaction and the so-called bremsstrahlung correction due to the emission of a real photon. Second are the finite temperature radiative corrections due to the finite temperature corrections to the proton-electron interactions and the electron energy. Third and fourth are the corrections from the finite nucleon mass and the weak magnetism. Combining all these corrections, the total weak reaction rates are given by
$\Gamma_{n\to p}=\Gamma^{\rm RC0}_{n\to p}+\delta\Gamma^{\rm SD,RC0}_{n\to p}+\Gamma^{T}_{n\to p}+\delta\Gamma^{\rm BS}_{n\to p}+\delta\Gamma^{\rm FM}_{n\to p}\,,$ (79)
where the superscript ’RC0’, ’SD’, ’T’, ’BS’ and ’FM’ represent the zero temperature radiative correction, the spectral distortion, the finite temperature radiative correction, the bremsstrahlung and the finite nucleon mass correction, respectively. And the form for the $p\to n$ rates is similar. Note that the effects of different corrections may be coupled. For example, the effects of the zero temperature radiative correction on the corrections induced by the spectral distortions are also considered in this work. Specifically, the Born rates with the zero temperature radiative correction take the form of
${\Gamma}^{{\rm RC}0}_{n\to p}$ $=K\int_{0}^{\infty}p^{2}\mathrm{d}p[\mathcal{F}_{+}(E)R(E,|\Delta-E|)\chi_{+}(E)+\mathcal{F}_{+}(-E)R(E,\Delta+E)\chi_{+}(-E)]\,,$ (80)
${\Gamma}^{{\rm RC}0}_{p\to n}$ $=K\int_{0}^{\infty}p^{2}\mathrm{d}p[\mathcal{F}_{-}(E)R(E,\Delta+E)\chi_{-}(E)+\mathcal{F}_{-}(-E)R(E,|E-\Delta|)\chi_{-}(-E)]\,.$ (81)
The explicit expressions for the Fermi function $F_{\pm}$ and the radiative correction factor $R(E,k_{\rm max})$ can be found in $e.g.,$   Ref. Pitrou:2018cgg and are not repeated here for brevity. We see that the zero temperature radiative correction only modifies the coefficients of the $\chi_{\pm}$ function. The $\delta\Gamma^{\rm RC0,SD}_{n\to p}$ and $\delta\Gamma^{\rm RC0,SD}_{p\to n}$ can also be obtained simply by the replacement of $\chi_{\pm}\to\delta\chi_{\pm}$. In principle, the other corrections are also affected by the spectral distortions of neutrinos and antineutrinos. However, such effects must be negligible since the corrections themselves are already sub-leading. Therefore, we neglect these interferences in this work.
4.2 Results of BBN abundances
In this subsection we present the resulting light element abundances incorporating all the corrections discussed in the last subsection. To do this, we have added the aforementioned correction induced by the neutrino spectral distortions in the public BBN code PRIMATPitrou:2018cgg and solved the abundances of the light elements. For the computation we have used the latest values of other physical and cosmological parameters, $e.g.,$ the neutron lifetime is taken as $\tau_{n}=879.4s$, the axial coupling of the nucleons is $g_{A}=1.2756$, and the baryon density is $\omega_{b}h^{2}=0.024$. For convenience, we define the mass fraction of isotope $i$ as $X_{i}\equiv A_{i}\frac{n_{i}}{n_{b}}$, where $A_{i}$ and $n_{i}$ are the baryon number and the number density of isotope $i$ and $n_{b}$ is the total number density of baryon. Customarily, we also denote $Y_{P}\equiv X_{{}^{4}{\rm He}}$ and $i/\rm H\equiv n_{i}/n_{{}^{1}\rm H}$.
Figure 5: Light element abundances as a function of the neutrino degeneracy parameter $\xi_{\nu}$, normalized to the results for $\xi_{\nu}=0$. The ${}^{3}{\rm He}$ stands for ${}^{3}{\rm He}+\rm T$ and ${}^{7}{\rm Li}$ stands for ${}^{7}{\rm Li}+\rm^{7}Be$ due to the slow radioactive decays. All the optional weak rate corrections in PRIMAT (see Ref. Pitrou:2018cgg ) and the new correction from neutrino spectral distortions are included in the calculation. The $68\%$ constraints on the helium abundance $Y\rm_{P}$ from the CMB (${Y\rm_{P}}=0.242\pm 0.012$) and the recent EMPRESS experiment (${Y\rm_{P}}=0.2370^{+0.0034}_{-0.0033}$) are indicated by the light and dark blue shaded regions, respectively.
We show in Fig. 5 the resulting light element abundances, including all the aforementioned weak rate corrections in (79) and the QED corrections up to $\mathcal{O}(e^{3})$ for plasma thermodynamics and certain nuclear rates, as a function of the neutrino degeneracy parameter $\xi_{\nu}$. The abundance of ${}^{3}{\rm He}$ stands for ${}^{3}{\rm He}+\rm T$ and ${}^{7}{\rm Li}$ for ${}^{7}{\rm Li}+\rm^{7}Be$ to account for the slow radioactive decays. Since the abundances for different elements can vary by several orders of magnitude, we normalize the element abundances by their corresponding abundance with $\xi_{\nu}=0$. Therefore all curves in Fig. 5 intersect at $\xi_{\nu}=0$. We see that the abundances of both $\rm D,\,^{3}{\rm He},\,^{4}{\rm He}$ and ${}^{7}{\rm Li}$ decrease with increasing $\xi_{\nu}$ due to the fact that the larger the $\xi_{\nu}$, the smaller the number density ratio between neutron and proton. In addition, the variation of the ${}^{4}{\rm He}$ abundance is the most sensitive to the variation of $\xi_{\nu}$, while the ${}^{3}{\rm He}$ abundance is the least sensitive. We also show in Fig. 5 the constraints on the helium abundance from obtained by CMB and EMPRESS experiments,
${Y\rm_{P}}|_{\rm CMB}=0.242\pm 0.012,\,\,{Y\rm_{P}}|_{\rm EMPRESS}=0.2370^{+0.0034}_{-0.0033},$ (82)
by the light and dark blue shaded regions. From our results, the constraints on ${Y\rm_{P}}$ within the $68\%$ confidence interval can be easily translated into constraints on the neutrino degeneracy parameter $\xi_{\nu}$ as
${\rm CMB}:\,\,-0.026\leq$ $\,\xi_{\nu}\leq 0.069\,,$ (83)
EMPRESS:  ${\rm EMPRESS}:\,\,0.027\leq$ $\,\xi_{\nu}\leq 0.054\,.$ (84)
Clearly, the EMPRESS experiment prefer a nonzero positive $\xi_{\nu}$ while the CMB experiment is still compatible with the SM case $\xi_{\nu}=0$.
Figure 6: The relative difference of the light element abundances compared to the results in the neutrino instantaneous decoupling limit as a function of the neutrino degeneracy parameter $\xi_{\nu}$. The solid and dotted curves correspond to the results with and without the weak corrections induced by the neutrino spectral distortions, respectively. The meaning of the color for the curves is the same as in the plot Fig. 5, except that the negative results of ${}^{7}{\rm Li}$ are shown here for simplicity.
BBN framework ${Y\rm_{P}}$ ${\rm D}/{\rm H}\times 10^{5}$ ${{}^{3}{\rm He}}/{\rm H}\times 10^{5}$ ${{}^{7}{\rm Li}}/{\rm H}\times 10^{10}$
Inst.dec, $\xi_{\nu}=-0.1$ 0.27359 2.58214 1.06094 5.87974
Inst.dec, $\xi_{\nu}=0$ 0.24711 2.42833 1.03799 5.53146
Inst.dec, $\xi_{\nu}=0.1$ 0.22260 2.30372 1.01897 5.14551
No-SD, $\xi_{\nu}=-0.1$ 0.27375 2.59246 1.06239 5.85555
No-SD, $\xi_{\nu}=0$ 0.24712 2.43725 1.03926 5.5077
No-SD, $\xi_{\nu}=0.1$ 0.2225 2.31162 1.02011 5.12186
Actual-SD, $\xi_{\nu}=-0.1$ 0.27375 2.59247 1.06239 5.85558
Actual-SD, $\xi_{\nu}=0$ 0.24721 2.43769 1.03933 5.50905
Actual-SD, $\xi_{\nu}=0.1$ 0.22266 2.31244 1.02024 5.12449
Table 1: Light element abundances including all weak rate corrections in (79), for various implementations of neutrino decoupling and degeneracy parameters. ${}^{3}{\rm He}$ stands for $(^{3}{\rm He}+\mathrm{T})$ and ${}^{7}{\rm Li}$ stands for $(^{7}{\rm Li}+\rm Be)$ to account for slow radioactive decays.
BBN framework $\delta{Y\rm_{P}}\,(\%)$ $\delta{\rm D}/{\rm H}\,(\%)$ $\delta{{}^{3}{\rm He}}/{\rm H}\,(\%)$ $\delta{{}^{7}{\rm Li}}/{\rm H}\,(\%)$
No-SD, $\xi_{\nu}=-0.1$ 0.058 0.4 0.137 -0.411
No-SD, $\xi_{\nu}=0$ 0.007 0.367 0.122 -0.43
No-SD, $\xi_{\nu}=0.1$0.048 0.343 0.112 -0.46
Actual-SD, $\xi_{\nu}=-0.1$ 0.059 0.4 0.137 -0.411
Actual-SD, $\xi_{\nu}=0$ 0.042 0.385 0.129 -0.405
Actual-SD, $\xi_{\nu}=0.1$ 0.026 0.379 0.125 -0.408
Table 2: Relative difference of the light element abundances for the case with and without the corrections induced by the neutrino spectral distortions compared to the results in the neutrino instantaneous decoupling limit. ${}^{3}{\rm He}$ stands for $(^{3}{\rm He}+\mathrm{T})$ and ${}^{7}{\rm Li}$ stands for $(^{7}{\rm Li}+\rm Be)$ to account for slow radioactive decays.
It is also interesting to discuss the effects of the corrections induced by the neutrino spectral distortions on BBN. Focusing on the corrections induced by different treatments of the neutrino decoupling process, we consider the following three implementations: (1) The neutrino instantaneous decoupling limit, $i.e.,$ the $N_{\rm eff}^{\rm inst}$ as discussed in the last section and the thermal Fermi-Dirac distributions for neutrinos and antineutrinos are used in the calculation of the BBN; (2) The no-SD case, where the actual $N_{\rm eff}$ and neutrino temperatures are used, including the non-instantaneous decoupling effect, but considering only the thermal Fermi-Dirac distributions for neutrinos and antineutrinos. This is the most common treatment for BBN in the studies to date; (3) The actual SD case, using the actual $N_{\rm eff}$ with the non-instantaneous decoupling effect and the actual density distributions with the spectral distortions for neutrinos and antineutrinos, corresponding to the results shown in Fig. 5. Using the BBN results in the instantaneous decoupling limit as the benchmark, we show the relative differences of the results for the no-SD and the actual-SD cases with the dotted and solid curves in Fig. 6, respectively. Note that we have plotted the opposite of the relative differences for ${}^{7}{\rm Li}$ in Fig. 6 as it is negative due to the particular nuclear reaction rates used in PRIMAT. We also report the explicit values of the abundances of helium-4, deuterium, helium-3 and lithium-7 for the degeneracy parameters $\xi_{\nu}=0,\pm 1$ in table 1, and the associated relative variations compared to the baseline abundance in table 2. We also note that our results with $\xi_{\nu}=0$ are consistent with the results reported in Froustey:2020mcq , where they also considered the effects of neutrino spectral distortions to BBN within the SM case.
As we can see in Fig. 6 and the tables, the actual-SD case and the no-SD case yield almost the same light element abundances for $\xi_{\nu}=-0.1$. And the variation of the BBN abundances from the results in the instantaneous decoupling limit for the actual-SD case show little dependence on $\xi_{\nu}$, while the traditional treatment without the spectral distortions predicts a stronger dependence for both elements. In particular, the no-SD case predicts negative variations for the helium abundance with positive $\xi_{\nu}$ compared to the result in the instantaneous decoupling limit. However, the more precise results with the spectral distortions show that the variations of the helium abundance are positive, showing that the inclusion of the neutrino spectral distortions is necessary for a precise and accurate prediction of the BBN abundances.
Finally, we also provide a set of fitting functions for the light element abundances as
$Y_{\mathrm{P}}\left(\xi_{\nu}\right)$ $\simeq\left.Y_{\mathrm{P}}\right|_{\xi_{\nu}=0}\times e^{-1.033\xi-0.131\xi^{2}},\,\,$ (85)
$\mathrm{D/H}\left(\xi_{\nu}\right)$ $\simeq\left.\mathrm{D/H}\right|_{\xi_{\nu}=0}\times e^{-0.57\,\xi_{\nu}(1-\xi_{\nu})}\,,$ (86)
${}^{3}\mathrm{He/H}\left(\xi_{\nu}\right)$ $\simeq\left.{}^{3}\mathrm{He/H}\right|_{\xi_{\nu}=0}\times e^{-0.19\,\xi_{\nu}(1-\xi_{\nu})},\,\,$ (87)
${}^{7}\mathrm{Li/H}\left(\xi_{\nu}\right)$ $\simeq\left.{}^{7}\mathrm{Li/H}\right|_{\xi_{\nu}=0}\times e^{-0.63\,\xi_{\nu}(1+\xi_{\nu})},$ (88)
where the subscript $\xi_{\nu}=0$ labels the results obtained in the SM case $\xi_{\nu}=0$ using the actual density matrices, see the explicit values in Table 1. Such fitting functions benefit for further studies, and the fractional errors are smaller than $0.1\%$ with the range $-0.1\leq\xi_{\nu}\leq 0.1$.  [^3] The coefficients of our fitting functions are slightly different from the results in Pitrou:2018cgg ; Froustey:2024mgf , which we think is due to the different corrections to the weak interaction rates we considered.
5 Impacts of the Primordial Neutrino Asymmetry on CMB and LSS
In this section we focus on the impacts of primordial neutrino asymmetries on the evolution of the late universe, especially the cosmic microwave background (CMB) and the large scale structure (LSS). In general, the primordial neutrino asymmetry influence CMB and LSS in three ways.
  1. 1.
    The expansion rate of the Universe is changed by the increased $N_{\rm eff}$ due to the primordial neutrino asymmetries. Particularly, the time of matter-radiation equality is delayed by the additional radiation density, leading to an enhanced early integrated Sachs-Wolfe effect for CMB angular power spectrum.
  2. 2.
    As we have seen in Fig. 5, the abundance of helium is significantly modified by the electron neutrino asymmetry, which will in turn alter the evolution of the late Universe, such as the tail of the CMB angular power spectrum by the amount of diffusion damping and the baryon acoustic oscillation (BAO).
  3. 3.
    In addition, in order to incorporate the effect of neutrino oscillation in the process of neutrino decoupling, the mass of neutrinos must be considered. The neutrino masses will alter the evolution of the late universe by such as, affect the matter-radiation equality, even in the SM case with the thermal Fermi-Dirac distributions. For the SM case, the experiment on CMB and BAO can put a relative stringent bound on the sum of neutrino masses $\sum m_{\nu}<0.07\,{\rm eV}$ DESI:2024mwx , we will return to the constraint on neutrino masses in the next section. Here we emphasize that the neutrino Boltzmann hierarchy for massive neutrinos will be furthermore modified by the neutrino degeneracy parameter and the spectral distortions.
In the following we first review the impact of massive neutrinos on the universe and present the Boltzmann hierarchies of electrons, photons and neutrinos. Then we discuss the effects of the primordial neutrino asymmetries on the CMB temperature anisotropies and the linear matter power spectrum computed with the Einstein-Boltzmann code CLASS Lesgourgues:2011re ; Lesgourgues:2011rh .
5.1 Effects of the primordial neutrino asymmetries in structure formation
To describe the flavor oscillations in the neutrino decoupling process, the mass of the neutrinos must be taken into account, which in turn has significant implications for the process of structure formation, since massive neutrinos are the only species in the SM that scale as radiation in the early universe and as matter in the late universe. In the early universe, neutrinos and antineutrinos remain relativistic after decoupling as long as their effective temperatures are much larger than their masses. In this case, the neutrino and antineutrino energy densities contribute to the total radiation energy density, parameterised by the aforementioned $N_{\rm eff}$. Since the neutrino effective temperatures decrease as $a^{-1}$ with the expansion of the universe, neutrinos and antineutrinos become non-relativistic when their effective temperatures fall below their masses, and contribute to matter after that. The non-relativistic transitions and the free-steaming effects of massive neutrinos then leave significant imprints on the CMB anisotropies and the LSS. As mentioned above, the situation is modified in the presence of the primordial neutrino asymmetries. In addition to the indirect change induced by the modified Hubble rate and the helium abundance, we highlight several other modifications due to the primordial neutrino asymmetries in the following.
5.1.1 Effects on neutrino abundance
First, the total relic abundance of neutrinos and antineutrinos are directly increased due to the primordial neutrino asymmetries. At present, the total neutrino relic abundance can be written as
$\Omega_{\nu}h^{2}$ $=\frac{\sum\limits_{i}T_{\rm CM,0}^{4}\int_{0}^{\infty}\mathrm{d}yy^{2}\sqrt{y^{2}+m_{i}}\left(\varrho_{i}(y)+\bar{\varrho}_{i}(y)\right)/(2\pi^{2})}{\rho_{\rm crit}}$ (89)
$\overset{\delta\varrho=0}{\simeq}\frac{\sum\limits_{i}m_{i}}{93.12\,{\rm eV}}\frac{{\rm Li}_{3}(-\exp(-\xi_{\nu}))+{\rm Li}_{3}(-\exp(\xi_{\nu}))}{3\zeta(3)/2},$
where $y=p/T_{\rm CM}$ is the comoving momentum defined in section 3, $T_{\rm CM,0}$ is the current comoving temperature, $\rho_{\rm crit}$ is the critical energy density, $\varrho_{i}(y)$ and $\bar{\varrho}_{i}(y)$ are the respective comoving density distributions for neutrinos and antineutrinos with $i=1,2,3$, corresponding to the three mass eigenstates with mass $m_{i}$. Obviously, the present neutrino abundance is also modified by the spectral distortions of neutrinos and antineutrinos, and the integrals over the neutrino distributions cannot be carried out analytically in this case. The effect of the primordial neutrino asymmetries will be clearer by neglecting the spectral distortions and performing the integral analytically, as seen in the second row of (89). The first factor in the second row $\frac{\sum\limits_{i}m_{i}}{93.12\,{\rm eV}}$ is the usual neutrino relic abundance in the SM case, while the second factor represents the effect of neutrino asymmetries with ${\rm Li}_{j}(x)$ the polylogarithms of order $j$ and $\zeta(x)$ the Riemann zeta function. For the SM case one finds ${\rm Li}_{3}(-1)=3\zeta(3)/4$, so the second factor becomes 1. For non-zero $\xi_{\nu}$, the second factor depends only on the absolute value of $\xi_{\nu}$, and the larger the $\xi_{\nu}$, the larger the second factor and the larger the neutrino abundance. However, we note that such corrections from primordial neutrino asymmetries are quite small. For example, the resulting relative difference of $\Omega_{\nu}h^{2}$ for $\xi_{\nu}=0.2$ is only $1.54\%$. Similarly, the corrections to the neutrino abundance induced by the spectral distortions of neutrinos and antineutrinos are also negligible.
5.1.2 Effects on the neutrino Boltzmann hierarchies
In order to obtain the accurate CMB spectral power spectrum and the matter power spectrum, we need to solve a coupled set of the Boltzmann-Einstein equations for the perturbations of the metric, photon, baryon, cold dark matter and massive neutrinos. Here we briefly review how neutrino asymmetries and spectral distortions affect the Boltzmann-Einstein equations. Following Ref. Ma:1995ey , the perturbed distribution functions for the three mass eigenstate neutrinos can be written as
$\varrho_{i}\left(\vec{x},\vec{p},\tau\right)=\varrho_{i}^{0}(y)\left[1+\Psi_{i}\left(\vec{x},y,\hat{n},\tau\right)\right]\,,$ (90)
[…]
$ds^{2}=a^{2}(\tau)\{-d\tau^{2}+(\delta_{ij}+h_{ij})dx^{i}dx^{j}\}.$ (91)
The perturbation of the metric $h_{ij}$ can be decomposed as $h_{ij}=h\delta_{ij}+h_{ij}^{\parallel}+h_{ij}^{\perp}+h_{ij}^{T}$, where $h\equiv h_{ii}$ is the trace part of $h_{ij}$ , and $h_{ij}^{\parallel}$, $h_{ij}^{\perp}$ and $h_{ij}^{T}$ are the traceless parts, satisfying
$\epsilon_{ijk}\partial_{j}\partial_{l}h_{lk}^{\parallel}=0,\quad\partial_{i}\partial_{j}h_{ij}^{\perp}=0,\quad\partial_{i}h_{ij}^{T}=0.$ (92)
By definition, $h_{ij}^{\parallel}$ and $h_{ij}^{\perp}$ can be written in terms of a scalar filed $\mu$ and a divergenceless vector $\vec{A}$,
$h_{ij}^{\parallel}$ $=\left(\partial_{i}\partial_{j}-\frac{1}{3}\delta_{ij}\nabla^{2}\right)\mu$ (93)
$h_{ij}^{\perp}$ $=\partial_{i}A_{j}+\partial_{j}A_{i},\quad\partial_{i}A_{i}=0.$
Therefore, the scalar modes of the metric perturbations are characterised by the two scalar fields $h$ and $\mu$, while the vector and the tensor modes are characterised by $\vec{A}$ and $h_{ij}^{T}$, respectively.
In this work we will only consider the scalar modes of $h_{ij}$, which can be Fourier transformed as
$h_{ij}(\vec{x},\tau)=\int d^{3}k~{}e^{i\vec{k}\cdot\vec{x}}\left\{\frac{{k}_{i}{k}_{j}}{k^{2}}h(\vec{k},\tau)+6\eta(\vec{k},\tau)\left(\frac{{k}_{i}{k}_{j}}{k^{2}}-\frac{1}{3}\delta_{ij}\right)\right\}\,,$ (94)
where $h(\vec{k},\tau)$ and $\eta(\vec{k},\tau)$ are the trace part and the traceless part of the scalar modes in the Fourier space which are related to the scalar fields $h$ and $\mu$ in the real space. Since neutrinos are collisionless after decoupling, the evolution of $\Psi_{i}$ obeys the first-order collsionless Boltzmann equation in the Fourier-space
$\frac{\partial\Psi_{i}}{\partial\tau}+i\frac{q}{\epsilon}(\vec{k}\cdot\hat{n})\Psi_{i}+\frac{d\ln\varrho_{i}^{0}}{d\ln q}\left[\dot{\eta}-\frac{\dot{h}+6\dot{\eta}}{2}(\hat{k}\cdot\hat{n})^{2}\right]=0\,,$ (95)
where $\epsilon\equiv aE=\sqrt{q^{2}+a^{2}m^{2}_{\nu_{i}}}$ is the comoving energy. Since Eq. (95) is independent of the azimuth angle, we can expand $\Psi$ in a Legendre series
$\Psi_{i}(\vec{k},\hat{n},q,\tau)=\sum_{l=0}^{\infty}(-i)^{l}(2l+1)\Psi_{i,l}(k,q,\tau)P_{l}(\hat{k}\cdot\hat{n})\,.$ (96)
Inserting Eq. (96) into Eq. (95) and using the orthonormality of the Legendre polynomials, we can obtain the Boltzmann hierarchies for $\Psi_{i,l}$:
$\dot{\Psi}_{i,0}$ $=-\frac{qk}{\epsilon}\Psi_{1}+\frac{\dot{h}}{6}\frac{{d}\ln\varrho_{i}^{0}}{{d}\ln q}\,,$ (97)
$\dot{\Psi}_{i,1}$ $=\frac{qk}{3\epsilon}\left(\Psi_{0}-2\Psi_{2}\right)\,,$
$\dot{\Psi}_{i,2}$ $=\frac{qk}{5\epsilon}\left(2\Psi_{1}-3\Psi_{3}\right)-\left(\frac{\dot{h}}{15}+\frac{2\dot{\eta}}{5}\right)\frac{{d}\ln\varrho_{i}^{0}}{{d}\ln q}\,,$
$\dot{\Psi}_{i,l\geq 3}$ $=\frac{qk}{(2l+1)\epsilon}\left[l\Psi_{l-1}-(l+1)\Psi_{l+1}\right]\,.$
Note that the effect of neutrino asymmetries and spectral distortions is manifested by the neutrino density distribution term $\frac{{d}\ln\varrho_{i}^{0}}{{d}\ln q}$. Since the density matrices for neutrinos and antineutrinos with different flavors are different in the presence of primordial neutrino asymmetries, the Boltzmann hierarchies for the corresponding perturbations $\Psi_{i}$ and $\bar{\Psi}_{i}$ are in general different and should be treated separately. With the moments $\Psi_{i,l}$, the perturbations in physical quantities, such as the energy density, pressure, energy flux and shear stress, can be computed straightforwardly by integrating different moments. For our purpose, the fractional density perturbation for the neutrinos with family number $i$ can be expressed as
[…]
where $\rho_{i,0}(t)$ is the zeroth-order homogeneous background energy density for neutrinos. Eventually, the effects of neutrino masses and asymmetries at different length scales can be reflected in the evolution of the Fourier modes $\delta_{i,k}$, which we obtain via the CLASS code.
5.1.3 Effects on the BAO
Following Ma:1995ey , the evolution of the perturbations for photon obey the Boltzmann hierarchy,
$\dot{\delta}_{\gamma}$ $=$ $-{4\over 3}\theta_{\gamma}-{2\over 3}\dot{h}\,,$
$\dot{\theta}_{\gamma}$ $=$ $k^{2}\left(\frac{1}{4}\delta_{\gamma}-\sigma_{\gamma}\right)+an_{e}\sigma_{T}(\theta_{b}-\theta_{\gamma})\,,$
[…]
$\dot{F}_{\gamma\,l}$ $=$ ${k\over 2l+1}\left[lF_{\gamma\,(l-1)}-(l+1)F_{\gamma\,(l+1)}\right]-an_{e}\sigma_{T}F_{\gamma\,l}\,,\quad l\geq 3\,,$
[…]
where $\delta_{\gamma}$ is the density fluctuation of photon, $\theta_{\gamma}$ is the divergence of photon velocity, and ${F}_{\gamma\,l}$ is the density perturbation of photon. In addition, $n_{e}$ is the electron number density and $\sigma_{T}$ is the Thomson scattering cross section. The Boltzmann equations for baryon are
$\dot{\delta}_{b}$ $=$ $-\theta_{b}-{1\over 2}\dot{h}\,,$
$\dot{\theta}_{b}$ $=$ $-{\dot{a}\over a}\theta_{b}+c_{s}^{2}k^{2}\delta_{b}+{4\bar{\rho}_{\gamma}\over 3\bar{\rho}_{b}}an_{e}\sigma_{T}(\theta_{\gamma}-\theta_{b})\,,$ (100)
where $\delta_{b}$ and $\theta_{b}$ is the baryon density fluctuation and the divergence of baryon fluid velocity, and $c_{s}^{2}$ is the square of the baryon sound speed.
Due to the electromagnetic interactions between photons and electrons, they evolve differently before and after recombination. Before recombination, photons and baryons are tightly coupled so that they can be treated as a single photon-baryon fluid, called as the "tight-coupling approximation". During such "tight-coupling epoch", the perturbations of baryon and photon undergo harmonic motion with very slowly decaying amplitude (Peebles:1970ag, ; Sunyaev:1970bma, ). After recombination, photons release their hold on the baryons, so that the baryon acoustic oscillations (BAO) become frozen-in, and the the baryon perturbations begin to grow as $\propto a$. The BAO feature then gets transferred to the gravitational potential and the matter fluctuations, leaving imprints in the galaxy power spectrum that depends on the sound horizon at decoupling and the angular diameter distance to the surface of photon-electron last-scattering.
In the presence of primordial neutrino asymmetries, the matter-radiation equality will be directly modified by the additional contribution to $N_{\rm eff}$, leading to the modification of the BAO properties. In addition, the effects of primordial neutrino asymmetries on the evolution of neutrino perturbations are also transported to photons and baryons through metric perturbations encoded in $h$ and $\eta$, even though neutrinos are collisionless. As a consequence, the primordial neutrino asymmetries have significant implications for both the CMB angular power spectrum and the BAO in the matter power spectrum, as we will see later.
5.2 Results of the CMB anisotropies and the matter power spectrum
Figure 7: Left panel: The CMB angular power spectrum $C_{l}^{\rm TT}$ within the SM case for different total neutrino masses, compared to the results with massless neutrinos. Right panel: The linear matter power spectrum $P(k)$ within the SM case for different total neutrino masses, compared to the results with massless neutrinos.
We present in this subsection the numerical results of the CMB and LSS obtained via the CLASS code. Since both observable are affected by the neutrino masses and asymmetries at the same time, we first discuss the effect of neutrino masses on the CMB and LSS within the SM. We present the ratio of the CMB angular power spectrum between the results with massive neutrinos and massless neutrinos $C_{l}^{\rm TT}/C_{l}^{\rm TT}(\sum m_{\nu}=0)$ and the similar ratio for the matter power spectrum $P_{k}/P_{k}(\sum m_{\nu}=0)$ for several choices of total neutrino masses in the left and right panels of Fig. 7, respectively. Note that the normal order of the neutrino mass hierarchy is considered in this work, so we adjust the $\mu_{\nu_{1}}$ for each choice of total mass, while $\mu_{\nu_{2}}$ and $\mu_{\nu_{3}}$ are obtained by the fixed mass differences between neutrinos. Focusing on the left panel, the ratio of the $C_{l}^{\rm TT}$ between massive and massless neutrinos can be divided into three stages. At large angular scales (lower multipoles ${\ell\lesssim 20}$), massive neutrinos induce a small suppression which comes from the late time integrated Sachs-Wolfe effect. On the intermediate scales (${20\lesssim\ell\lesssim 500}$), dominated by the imprints of the acoustic oscillations in the baryon-photon fluid, the first peak of $C_{l}^{\rm TT}$ is affected by the modified evolution of the gravitational potential due to the correlation between massive neutrinos and metric perturbations. As we can see, the height of the first peak given by the early integrated Sachs-Wolfe effect is increased for larger neutrino masses compared to the case with massless neutrinos. The acoustic peaks at small scales (${\ell\gtrsim 500}$) are also affected by the massive neutrinos. Similar with the first peak, the height of the acoustic peaks is increased for larger neutrno masses.
Similarly, the resulting matter power spectrum with massive neutrinos can also be divided to three stages. At large scales corresponding to $k\lesssim 10^{-3}\,h/{\rm Mpc}$, neutrinos do not enter the horizon, and the small suppression of $P(k)$ is induced by the modification of the matter-radiation equality from the non-relativistic transition of massive neutrinos. At intermediate scales corresponding to $10^{-3}\,h/{\rm Mpc}\lesssim k\lesssim 1\,h/{\rm Mpc}$, neutrinos gradually enter the horizon, so that the free streaming of these neutrinos results in an additional suppression of the matter power spectrum. Therefore, $P_{k}/P_{k}(\sum m_{\nu}=0)$ decreases with increasing wavenumber $k$. And at small scales ($k\gtrsim 1\,h/{\rm Mpc}$), the majority of neutrinos stream freely, leading to an almost constant suppression of the power spectrum. The larger the total neutrino mass, the greater the suppression of the matter power spectrum. It is also worth noting that the matter power spectrum is more sensitive to the mass of the neutrinos, as can be seen from the plots. Therefore, it is more efficient to constrain the neutrino masses with the LSS observations, as we will see in the next section.
Figure 8: Left panel: The CMB angular power spectrum $C_{l}^{\rm TT}$ for different neutrino degeneracy parameters $\xi_{\nu}$, compared to the results with $\xi_{\nu}=0$. Right panel: The linear matter power spectrum $P(k)$ for different neutrino degeneracy parameters $\xi_{\nu}$, compared to the results with $\xi_{\nu}=0$. The total neutrino mass is fixed to $\sum m_{\nu}=0.1{\rm eV}$ for both panels.
Figure 9: The sound horizon at the baryon drag epoch of the early universe $r_{d}$ as a function of the primordial neutrino asymmetry $\xi_{\nu}$.
We now present the results for non-zero $\xi_{\nu}$. To isolate the effects of neutrino asymmetries from the total neutrino mass, the left panel of Fig. 8 shows the ratio of the CMB angular power spectrum for $\xi_{\nu}\neq 0$ to $\xi_{\nu}=0$, with $\sum m_{\nu_{i}}=0.1\,\text{eV}$. A similar ratio for the matter power spectrum is shown in the right panel. As seen in the left panel, the effects of neutrino asymmetries on the CMB angular power spectrum manifest primarily at small scales ($\ell\gtrsim 400$). At $\ell\lesssim 400$, $C_{l}^{\rm TT}$ for non-zero $\xi_{\nu}$ matches the results for $\xi_{\nu}=0$. However, as $\ell$ increases, deviations due to $\xi_{\nu}$ become apparent. These deviations depend on both the magnitude and sign of $\xi_{\nu}$: positive $\xi_{\nu}$ enhances the small-scale power, while negative $\xi_{\nu}$ leads to suppression. This behavior is largely driven by changes in helium abundance, which alters the CMB spectrum via diffusion damping. In addition to overall enhancement or suppression, the acoustic peaks are also affected by $\xi_{\nu}$, as larger $\xi_{\nu}$ values modify the oscillations at the last scattering surface. The amplitude of these modifications grows with the absolute value of $\xi_{\nu}$, with negative $\xi_{\nu}$ producing more pronounced effects.
For the linear matter power spectrum in the right panel of Fig. 8, we observe that at large scales, the result for non-zero $\xi_{\nu}$ is identical to that for $\xi_{\nu}=0$. At small scales, however, the free-streaming of neutrinos is enhanced with larger $\xi_{\nu}$, resulting in a slightly stronger suppression of the matter power spectrum, although this suppression is insensitive to the sign of $\xi_{\nu}$. The more intriguing effect of non-zero $\xi_{\nu}$ is the oscillations of the matter power spectrum on the scales $10^{-1}\,h/{\rm Mpc}\gtrsim k\gtrsim 10^{0}\,h/{\rm Mpc}$, which represents the effect of the BAO. Similar to the acoustic oscillations in the CMB angular power spectrum, we find that the amplitudes of the BAO in the matter power spectrum are also larger with a larger absolute value of $\xi_{\nu}$, while the negative $\xi_{\nu}$ leads to a larger effect than that for the positive $\xi_{\nu}$. In addition to the imprints of BAO on the matter power spectrum, we show in Fig. 9 the sound horizon at the baryon drag epoch of the early universe $r_{d}$ as a function of the primordial neutrino asymmetry $\xi_{\nu}$. Due to the additional contribution to $N_{\rm eff}$ for non-zero $\xi_{\nu}$ according to Eq. 63, the larger the $\left|\xi_{\nu}\right|$, the smaller the $r_{d}$, resulting in the significant imprints on the matter power spectrum. These results imply that BAO experiments can also provide powerful constraints on the primordial neutrino asymmetry, as we will in Section 6, which is often overlooked in previous studies. Note also that these conclusions are drawn with a fixed total neutrino mass $\sum m_{\nu_{i}}=0.1\,\text{eV}$. However, they remain valid for different total neutrino masses, although the resulting ratios $C_{l}^{\rm TT}/C_{l}^{\rm TT,\,\xi_{\nu}=0}$ and $P(k)/P^{\xi_{\nu}=0}(k)$ may vary slightly at intermediate scales due to the non-relativistic transition of massive neutrinos.
Figure 10: The CMB angular power spectrum $C_{l}^{\rm TT}$ (left panel) and the linear matter power spectrum $P(k)$ (right panel) for different neutrino degeneracy parameters $\xi_{\nu}$, compared to the corresponding results with $\xi_{\nu}=0$. The total neutrino mass is fixed to $\sum m_{\nu}=0.1{\rm eV}$ for both panels. The solid curves are the same with Fig. 5, corresponding to the results including the spectral distortions of the neutrinos and antineutrinos, while the dashed curves correspond to the results without the spectral distortions.
Finally, we briefly discuss the effects of spectral distortions of the neutrinos and antineutrinos on the CMB and LSS. Similar to Fig. 8, we present the resulting CMB angular power spectrum $C_{l}^{\rm TT}$ and linear matter power spectrum $P(k)$ with non-zero $\xi_{\nu}$, neglecting the neutrino spectral distortions, as the dashed curves in the left and right panels of Fig. 10. The solid curves in Fig. 10 are the same as those in Fig. 8, corresponding to the results including the spectral distortions of the neutrinos and antineutrinos. Comparing the resulting CMB angular power spectrum $C_{l}^{\rm TT}$ with and without the neutrino spectral distortions, we see that the primary effect of including the spectral distortions is a slight suppression of the first three acoustic peaks, with an almost negligible effect on the remaining peaks in the small scales. For the linear power spectrum, we find that the effect of spectral distortions on the BAO is similar with that on the CMB angular power spectrum. Here the first two acoustic peaks are slightly enhanced while the effects on the remaining peaks are negligible. In addition, we find that the calculations without the neutrino spectral distortions cannot capture the right behavior of the matter power spectrum at small scales, where the results without the spectral distortions predict a near constant suppression at the scales $k\gtrsim 1\,h/{\rm Mpc}$, while the preciser results with the spectral distortions show that the matter power spectrum receives larger suppression at these scales. However, the discussion here refers only to the linear matter power spectrum, and we expect that a comprehensive study of the effects of neutrino spectral distortions on the LSS in the nonlinear regime can be done in the future with the N-body simulations.
6 Combined Analysis of Primordial Neutrino Asymmetries
In previous sections, we have shown that the primordial neutrino asymmetry significantly influences the neutrino decoupling process, leaving significant observable imprints on the Universe, such as the abundance of light elements produced during the BBN, the CMB anisotropies, and the BAO imprints in the LSS. These cosmological observations can then be used to constrain the primordial neutrino asymmetries. In this section, we consolidate the results of the previous sections and discuss the constraints on the primordial neutrino asymmetry derived from contemporary observations of the BBN, CMB and BAO. In addition, we briefly explore the corresponding constraints on UV model parameters using a specific model that generates primordial neutrino asymmetries in the early universe.
6.1 Observational Constraints on Primordial Neutrino Asymmetry
Bounds for the model $\mathbf{\Lambda}$CDM + $\mathbf{\xi_{\nu}}$ + $\mathbf{m_{\nu_{1}}}$
Data Sets $\qquad\,\,\,{\bf\xi_{\nu}}\,\,\,\qquad$ $\omega_{b}$ $\omega_{cdm}$ ${\sum m_{\nu}\,\,(\mathrm{eV})}$
BBN $0.039\pm 0.013$ $2.155\pm 0.016$
BBN + CMB $0.028\pm 0.012$ $2.189\pm 0.012$ $0.1225\pm 0.0013$ 0.414
BBN + CMB + BAO $0.024\pm 0.012$ $2.200\pm 0.011$ $0.1205\pm 0.0009$ 0.131
Bounds from the datasets BBN + CMB + BAO
Models $\qquad\,\,\,{\bf\xi_{\nu}}\,\,\,\qquad$ $\omega_{b}$ $\omega_{cdm}$ ${\sum m_{\nu}\,\,(\mathrm{eV})}$
$\mathbf{\Lambda}$CDM + $\mathbf{m_{\nu_{1}}}$ $2.211\pm 0.010$ $0.1203\pm 0.0009$ 0.136
$\mathbf{\Lambda}$CDM + $\mathbf{\xi_{\nu}}$ $0.023\pm 0.012$ $2.200\pm 0.011$ $0.1207\pm 0.0009$
$\mathbf{\Lambda}$CDM + $\mathbf{\xi_{\nu}}$ + $\mathbf{m_{\nu_{1}}}$ $0.024\pm 0.012$ $2.200\pm 0.011$ $0.1205\pm 0.0009$ 0.131
Table 3: Summary of constraints on the primordial lepton asymmetry $\xi_{\nu}$, the baryon abundance $\omega_{b}$, the cold dark matter abundance $\omega_{cdm}$ and the total neutrino mass $\sum m_{\nu}$ from various datasets and cosmological models.
In this work we utilize the following datasets for BBN, CMB, and BAO in our analysis:
BBN
The EMPRESS measurement of helium abundance, ${Y\rm_{P}}=0.2370^{+0.0034}_{-0.0033}$ empress , and the deuterium abundance adopted by the PDG group, $10^{5}\times{\rm D/H}=2.547\pm 0.025$ pdg , which is based on the results of Cooke:2017cwo .
CMB
The Planck 2018 low-$l$ and high-$l$ TT, TE, and EE spectra, along with the reconstructed CMB lensing spectrum Planck:2018vyg ; Planck:2018lbu ; Planck:2019nip .
BAO
The BOSS DR12 BOSS:2016wmc with its full covariance matrix for BAO, as well as the 6dFGS Beutler:2011hx and MGS of SDSS Ross:2014qpa .
We note that the helium abundance from the BBN data set primarily constrains the primordial neutrino asymmetries, while the deuterium abundance is more sensitive to the baryon density $\omega_{b}$. The BAO dataset is mainly used to constrain the total neutrino mass $\sum m_{\nu}$, but as we will discuss, it also places additional constraints on primordial neutrino asymmetries.
To extract the constraints from these datasets, we conduct MCMC analyses using the Montepython engine Audren:2012wb ; Brinckmann:2018cvx . For analysis, the $N_{\rm eff}$ is adopted as Eq. (63), derived in Section 3. The helium and deuterium abundances predicted by BBN are computed using our fitting formulas in Equation (85). As shown in Fig. 10, the effects of neutrino spectral distortions on the CMB angular power spectrum and the BAO in the matter power spectrum are small. Therefore, for simplicity, the (anti)neutrino spectral distortions are neglected in the MCMC analysis. The parameter set for the CMB and BAO datasets includes the six independent parameters of the $\Lambda$CDM model, along with the neutrino degeneracy parameter $\xi_{\nu}$ and the mass of the lightest neutrino for the normal hierarchy $m_{\nu_{1}}$, i.e., $\{\omega_{b},\omega_{cdm},\theta_{s},A_{s},n_{s},\tau_{\mathrm{reio}},\xi_{\nu},m_{\nu_{1}}\}$, where $\theta_{s}$ represents the sound horizon, $A_{s}$ is the amplitude of the primordial curvature perturbations at the pivot scale $k=0.05\,\mathrm{Mpc}^{-1}$, $n_{s}$ is the scalar spectral index, and $\tau_{\mathrm{reio}}$ is the optical depth to reionization. On the other hand, we only consider $\{\omega_{b},\xi_{\nu}\}$ for the BBN dataset, as the predicted light element abundances from BBN are largely independent of other cosmological parameters within the relevant range.
Figure 11: 2D and 1D marginalised posterior distributions for the baryon abundance $\omega_{b}$, the neutrino degeneracy parameter $\xi_{\nu}$, the cold dark matter abundance $\omega_{cdm}$, the mass of the lightest neutrino for the normal hierarchy $m_{\nu_{1}}$, and scalar spectral index $n_{s}$. The red, blue, and green regions represent constraints from BBN, CMB + BAO, and their combination, respectively.
Figure 12: 1D marginalised distributions for different parameter sets derived from the BBN + CMB + BAO datasets.
With the aforementioned choices of parameters and datasets, we modified the MontePython sampler and CLASS codes to perform the MCMC analysis. The Gelman-Rubin criterion for all chains converged to $R-1<0.01$. The resulting $1\,\sigma$ cosmological constraints are summarized in Table 3, where the $95\%$ confidence limits on the total neutrino mas are also presented from the constraints on the mass of the lightest neutrino for the normal hierarchy $m_{\nu_{1}}$. For the same model $\Lambda$CDM + ${\xi_{\nu}}$ + ${m_{\nu_{1}}}$, we present the 1D and 2D marginalized posterior distributions of key parameters that show significant variations across different datasets in Fig. 11, including the baryon abundance $\omega_{b}$, the neutrino degeneracy parameter $\xi_{\nu}$, the cold dark matter abundance $\omega_{cdm}$, the sound horizon $\theta_{s}$, and the mass of the lightest neutrino for the normal hierarchy $m_{\nu_{1}}$.
Concerning the constraints on the primordial neutrino asymmetries, we see that the recent EMPRESS data in the BBN dataset points to a positive primordial neutrino asymmetry $\xi_{\nu}=0.039\pm 0.013$, which is different from zero with a $3.0\,\sigma$ significance. Combining the BBN and CMB data, the resulting constraints on $\xi_{\nu}$ are significantly reduced to $\xi_{\nu}=0.028\pm 0.012$, different from zero with a $2.3\,\sigma$ significance. Compared to the constraints $\xi_{\nu}=0.034\pm 0.014$ reported in previous study with the same BBN and CMB datasets Escudero:2022okz , our result indicate a $\sim 18\%$ smaller central value and a slight reduction in the uncertainty. Incorporating BOSS BAO observations in the MCMC analysis, the primordial neutrino asymmetry constraint further reduce to $\xi_{\nu}=0.024\pm 0.013$, indicating a $\sim 29\%$ reduction in the central value compared to previous results, while the significance of a positive neutrino asymmetry is reduced to $2.0\,\sigma$. Our study demonstrate the additional constraining power of the BAO experiments on the primordial neutrino asymmetry, which are not aware in previous studies.
We find that the constraints on other parameters also vary significantly with different datasets, especially for $\omega_{b}$, $\omega_{cmb}$, $n_{s}$ and $m_{\nu_{1}}$, as shown in Fig. 11. For the constraints on $\omega_{b}$, the difference is largely due to the fact that we chose the nuclear reaction rates in PRIMAT in the process of BBN, which predicted a deuterium abundance lower than the current constraints from PDG, thus leading to significantly lower constraints on $\omega_{b}$ Pitrou:2018cgg ; Burns:2022hkq ; Escudero:2022okz . While the constraints on the primordial neutrino asymmetry $\xi_{\nu}$ do not strongly depend on the choice of nuclear reaction rates, as discussed in  Escudero:2022okz , a detailed study with different nuclear reaction rates will be performed in our companion paper. The inclusion of the BOSS BAO data significantly changes the corresponding constraints on $\omega_{cmb}$, $n_{s}$ and $m_{\nu_{1}}$ due to its sensitivity to the evolution of the late universe. Especially, the resulting $95\%$ upper limit for the total mass of neutrinos are greatly tightened from $0.414$ eV by BBN + CMB to $0.131$ eV by BBN + CMB + BAO, since the BAO data help to break the CMB degeneracy. Recently, the BAO observations of DESI have put an even stronger bound on neutrino masses $\sum m_{\nu_{i}}\lesssim 0.07\,eV$ DESI:2024mwx . The relevant studies on primordial neutrino asymmetry and neutrino masses with DESI BAO data will also be discussed in our companion paper.
Finally we briefly discuss the effects of different parameter sets on the resulting constraints, especially the interplay between the primordial neutrino asymmetry and the total mass of neutrinos. Using the data set of BBN + CMB + BAO, we present in Fig. 12 the resulting 1d marginalised distributions for three choices of the parameter set, first is the set of $\Lambda CDM+m_{\nu_{1}}$, second is the $\Lambda CDM+\xi_{\nu}$, while the third is $\Lambda CDM+m_{\nu_{1}}+\xi_{\nu}$. As we can see, while the different choices of the parameter set lead to quite modifications to the constraints on the six parameters of $\Lambda CDM$, especially for $\omega_{b}$ and $n_{s}$, the resulting constraints on the additional parameters are almost the same for different parameter sets. The constraints on the total mass of neutrinos are only slightly tightened while including primordial neutrino asymmetry $\xi_{\nu}$ as a free parameter, while the constraints on $\xi_{\nu}$ are almost the mass as the inclusion of free $\sum m_{\nu}$. Therefore, we conclude that the constraints on neutrino masses and the primordial neutrino asymmetries can indeed be analysed separately if one do not concern the constraints on the other parameters in $\Lambda CDM$, otherwise the parameter sets with both the neutrino masses and the primordial neutrino asymmetries must be considered.
6.2 Constraints on UV model parameters
With the constraints on the primordial neutrino asymmetries, we discuss in this subsection the corresponding constraints on the UV model parameters with a specific model to generate the primordial neutrino asymmetries. Here we consider a Q-ball decay model Kawasaki:2022hvx , where Q-balls are produced in the Affleck-Dine mechanism Kusenko:1997si ; Enqvist:1997si ; Kasuya:1999wu , which utilizes flat directions in the minimal supersymmetric standard model (MSSM). When some flat direction with lepton number has a large field value during inflation, it generates lepton number by field dynamics after inflation. At the same time, Q-balls are produced and the generated lepton number is all confined inside the Q-balls. Thus, such Q-balls have a large lepton number and are called L-balls. Since the sphaleron effect is inactive inside the L-balls, the lepton asymmetry produced by the Affleck-Dine mechanism are kept without being converted to a baryon asymmetry Kawasaki:2002hq ; Gelmini:2020ekg . The L-balls decay into neutrinos below the electroweak scale and realize a large lepton asymmetry in the universe.
Following the results of Kawasaki:2022hvx , the primordial lepton asymmetry $\eta_{L}$ is defined as the ratio between the difference of the number density of neutrinos over antineutrinos and the photon number density,
$\eta_{L}\equiv\sum_{\alpha}\frac{n_{\nu_{\alpha}}-n_{\bar{\nu}_{\alpha}}}{s}\simeq 0.105\,\xi_{\nu},$ (101)
where $s=\frac{2\pi^{2}}{45}g_{\star}T_{\gamma}^{3}$ is the entropy density and $g_{\star}=10.75$ is the relativistic degree of freedom at the neutrino decoupling epoch. In addition, the resulting $\eta_{L}$ in the Q-ball decay model can be expressed as
$\eta_{L}=\frac{3T_{\mathrm{D}}}{4m_{3/2}},$ (102)
where $m_{3/2}$ is the the gravitino mass, $T_{\mathrm{D}}$ is the decay temperature of the L-balls and can be expressed with the other UV model parameters with
$T_{\mathrm{D}}$ $\simeq 2.69~{}\mathrm{MeV}\,$
×$\hskip 15.0pt\times\left(\frac{g_{*}}{10.75}\right)^{-1/4}\left(\frac{\beta}{6\times 10^{-4}}\right)^{-5/8}\left(\frac{m_{3/2}}{0.5~{}\mathrm{GeV}}\right)^{5/2}\left(\frac{M_{F}}{5\times 10^{6}~{}\mathrm{GeV}}\right)^{-2}\left(\frac{N_{\ell}}{3}\right)^{1/2}\left(\frac{\zeta}{2.5}\right)^{1/2},$ (103)
where $g_{*}(T_{\mathrm{D}})$ is the relativistic number of degrees of freedom at $T_{\mathrm{D}}$, $\beta\simeq 6\times 10^{-4}$ is a dimensionless constant, $M_{F}$ is the SUSY breaking scale, $N_{\ell}$ is the number of decay channels. Therefore, once we have obtained the constraints on the lepton asymmetry from cosmological observations, they can be directly translated into constraints on the model parameters of the UV model.
According to our study, the constraints on $\eta_{L}$ from the BBN + CMB + BAO datasets are $\eta_{L}=0.0025\pm 0.0013$. Therefore, the constraint on the L-balls decay temperature is
$T_{\mathrm{D}}$ $=\frac{4}{3}\eta_{L}m_{3/2}\simeq 3.3~{}\mathrm{MeV}\,\left(\frac{m_{3/2}}{0.5~{}\mathrm{GeV}}\right)\left(\frac{\eta_{L}}{5\times 10^{-3}}\right)$
[…]
and conversely,
$m_{3/2}$ $=\frac{3T_{\mathrm{D}}}{4T\eta_{L}}\simeq 0.5~{}\mathrm{GeV}\,\left(\frac{T_{\mathrm{D}}}{3.3~{}\mathrm{MeV}}\right)\left(\frac{5\times 10^{-3}}{\eta_{L}}\right)$
[…]
These constraints can also be converted into the constraints on other UV model parameters, such as the SUSY breaking scale, as seen in the preliminary study of Kawasaki:2022hvx .
7 Conclusion
The presence of primordial neutrino asymmetries before the neutrino decoupling epoch has significant implications for both the neutrino decoupling process and subsequent cosmological phenomena, and can therefore be constrained by observations of BBN, CMB and LSS. However, the effects of primordial neutrino asymmetries on neutrino decoupling—particularly concerning the resulting $N_{\rm eff}$ and the spectral distortions of (anti)neutrinos, including the impacts of flavor oscillations and FTQED corrections—remain ambiguous. These ambiguities affect the interpretation of subsequent processes. Consequently, studies of primordial neutrino asymmetries in relation to BBN and the CMB often employ approximations, such as using thermal Fermi-Dirac distributions for (anti)neutrinos, without thorough scrutiny.
In this work, we aim to address these ambiguities by numerically solving the neutrino decoupling process and analyzing the implications of the resulting (anti)neutrino density matrices for BBN, CMB, and LSS. We begin by re-deriving the quantum kinetic equations (QKEs) for neutrinos and antineutrinos using the Closed Time Path (CTP) formalism for two-point functions. Focusing on primordial neutrino asymmetries with complete flavor equilibration, where $\xi_{\nu_{e}}=\xi_{\nu_{\mu}}=\xi_{\nu_{\tau}}=\xi_{\nu}$, we numerically solve the QKEs in the early universe using the public code FortEPiaNO Gariazzo:2019gyi ; Bennett:2020zkv , including the effects of neutrino oscillations, matter effects, FTQED corrections up to $\mathcal{O}(3)$, and comprehensive neutrino-electron and neutrino-neutrino collision terms. The final $N_{\rm eff}$ is then evaluated based on the resulting density matrices, and can be expressed as $N_{\rm eff}\simeq N_{\rm eff}^{\rm SM}+3\left(\frac{30}{7\pi^{2}}\xi_{\nu}^{2}+\frac{15}{7\pi^{4}}\xi_{\nu}^{4}\right)+0.0102\xi_{\nu}^{2}$, where $N_{\rm eff}^{\rm SM}=3.0440\pm 0.0002$, while the second and last term represent the corrections from nonzero $\xi_{\nu}$ in the instantaneous decoupling limit and from the non-instantaneous effect, respectively. We also confirm that the spectral distortions of neutrinos and antineutrinos with non-zero $\xi_{\nu}$ differ due to distinct Pauli-blocking effects, in agreement with the priliminary study in Grohs:2016cuu , where neutrino spectral distortions were evaluated while neglecting flavor oscillations and FTQED corrections.
Using the density matrices for neutrinos and antineutrinos, we further investigate the implications of primordial neutrino asymmetries for BBN, CMB and LSS. For BBN, we introduce a novel correction to the weak interaction rates arising from the distinct spectral distortions of neutrinos and antineutrinos. The light element abundances are then computed with the public BBN code PRIMAT Pitrou:2018cgg , incorporating the other additional corrections on weak rates. According to the the EMPRESS measurement on the helium abundance $Y_{P}|_{\rm EMPRESS}=0.2370^{+0.0034}_{-0.0033}$, our results immediately indicate a positive primordial neutrino asymmetry $0.032\leq\xi_{\nu}\leq 0.052$ with fixed baryon abundance $\omega_{b}$. The effects of the new proposed correction are also examined, and we find that the spectral distortions for neutrinos and antineutrinos are important for precise predictions for BBN, especially for the abundance of Helium-4. Compared with the results in the instantaneous limit, the traditional treatment with Fermi-Dirac distributions predict a negative relative difference for a positive primordial neutrino asymmetry, while our precise treatment with spectral distortions show that the relative difference is still positive. Additionally, we have provided semi-analytic expressions for these abundances as functions of $\xi_{\nu}$, which will be useful for future studies. For the CMB and LSS, a novel finding of our study is that primordial neutrino asymmetries also affect the amplitudes of the acoustic oscillations, both for the CMB and for baryon acoustic oscillations (BAO). This shows that BAO observations could serve as a tool for constraining primordial neutrino asymmetries.
Based on our systematic understanding of primordial asymmetries, we perform a joint MCMC analysis including the neutrino degeneracy parameter $\xi_{\nu}$ and the total neutrino mass $\sum m_{\nu}$, using data from EMPRESS BBN, Planck CMB, and BOSS BAO observations. Combining the EMPRESS BBN and Planck CMB datasets, the primordial neutrino asymmetry constraint yields $\xi_{\nu}=0.028\pm 0.013$, indicating a $\sim 18\%$ smaller central value and a slight reduction in the uncertainty compared to previous results $\xi_{\nu}=0.034\pm 0.014$ obtained in Escudero:2022okz . Incorporating BOSS BAO observations in the MCMC analysis, the primordial neutrino asymmetry constraint further reduce to $\xi_{\nu}=0.024\pm 0.013$, indicating a $\sim 29\%$ reduction in the central value compared to previous results and a $2.0\,\sigma$ significance of a positive neutrino asymmetry. We also discuss the impacts of treating the sum of neutrino mass $\sum m_{\nu}$ and the primordial neutrino asymmetry $\xi_{\nu}$ as free parameters on the resulting constraints for each of them. Finally, we also show how the constraints on the neutrino asymmetries can be translated into constraints on the UV model parameters with a specific model in which the large neutrino asymmetries are generated by the decay of Q-balls Kawasaki:2002hq ; Gelmini:2020ekg .
Our work provides the framework to obtain the implications of new physics beyond the SM and $\Lambda$CDM model for neutrino decoupling and subsequent processes, and the corresponding cosmological constraints from present and future BBN, CMB and LSS observations. Thus, our work facilitates a more comprehensive analysis for primordial neutrino asymmetry and other models with new physics, such as precise consideration of the neutrino synchronous oscillations Froustey:2021azz ; Froustey:2024mgf or neutrino non-standard interactions Du:2021idh ; Du:2023upj , based on more and more precise cosmological observations in the future.
Acknowledgements.
This work is supported by the National Science Foundation of China under Grants No. 12347105, No. 12375099 and No. 12047503, and the National Key Research and Development Program of China Grant No. 2020YFC2201501, No. 2021YFA0718304. The authors gratefully acknowledge the use of publicly available codes FortEPiaNO Gariazzo:2019gyi ; Bennett:2020zkv , PRIMAT Pitrou:2018cgg , CLASS Lesgourgues:2011re ; Lesgourgues:2011rh and MontePythonAudren:2012wb ; Brinckmann:2018cvx .
Appendix A Collision Terms for Neutrinos and Antineutrinos
A.1 Complete expressions for the collision terms
In this work, we consider the interaction processes of neutrino-electron/positron scattering ($\nu{}e\leftrightarrow\nu{}e$) and annihilation ($\nu\bar{\nu}\leftrightarrow e^{+}e^{-}$), plus neutrino-neutrino self-interactions ($\nu\nu\leftrightarrow\nu\nu$, $\nu\bar{\nu}\leftrightarrow\nu\bar{\nu}$). Therefore, the total collision term for neutrinos $\mathcal{I}$ can be decomposed as
$\mathcal{I}\equiv\mathcal{I}_{\rm sc}+\mathcal{I}_{\rm ann}+\mathcal{I}_{\nu\nu}$ (106)
where $\mathcal{I}_{\rm sc}$ and $\mathcal{I}_{\rm ann}$ represent the collision terms for neutrino-electron/positron scattering and annihilation respectively, while $\mathcal{I}_{\nu\nu}$ represents the collision terms for neutrino self-interactions. The collision term for antineutrinos $\bar{\mathcal{I}}$ also admits a similar decomposition. And the contributions from other interaction processes such as $\mu^{\pm}$ scattering and annihilation are neglected, as their contributions are highly suppressed at the temperature of neutrino decoupling.
In the following we present the explicit form of the collision terms, which are also presented in various studies (see $e.g.,$  Blaschke:2016xxt ). The collision terms for the neutrino-electron/positron scattering processes ($\nu{}e\leftrightarrow\nu{}e$) and the annihilation processes ($\nu\bar{\nu}\leftrightarrow e^{+}e^{-}$) take the form as
$\mathcal{I}_{\rm sc}=$ $\frac{1}{2}\frac{2^{5}G_{F}^{2}}{2p_{1}}\int{\mathrm{d}\pi_{2}\,\mathrm{d}\pi_{3}\,\mathrm{d}\pi_{4}\,(2\pi)^{4}\delta^{(4)}(p_{1}+p_{2}-p_{3}-p_{4})}$
$\Big{[}4(p_{1}\cdot p_{2})(p_{3}\cdot p_{4})\left(F_{\mathrm{sc}}^{LL}(\nu^{(1)},e^{(2)},\nu^{(3)},e^{(4)})+F_{\mathrm{sc}}^{RR}(\nu^{(1)},\bar{e}^{(2)},\nu^{(3)},\bar{e}^{(4)})\right)$
$+4(p_{1}\cdot p_{4})(p_{2}\cdot p_{3})\left(F_{\mathrm{sc}}^{RR}(\nu^{(1)},e^{(2)},\nu^{(3)},e^{(4)})+F_{\mathrm{sc}}^{LL}(\nu^{(1)},\bar{e}^{(2)},\nu^{(3)},\bar{e}^{(4)})\right)$
$-2(p_{1}\cdot p_{3})m_{e}^{2}\left(F_{\mathrm{sc}}^{LR}(\nu^{(1)},e^{(2)},\nu^{(3)},e^{(4)})+F_{\mathrm{sc}}^{LR}(\nu^{(1)},\bar{e}^{(2)},\nu^{(3)},\bar{e}^{(4)})+\{L\leftrightarrow R\}\right)\Big{]}\,,$ (107)
$\mathcal{I}_{\rm ann}=$ $\frac{1}{2}\frac{2^{5}G_{F}^{2}}{2p_{1}}\int{\mathrm{d}\pi_{2}\,\mathrm{d}\pi_{3}\,\mathrm{d}\pi_{4}\,(2\pi)^{4}\delta^{(4)}(p_{1}+p_{2}-p_{3}-p_{4})}$
$\Big{[}4(p_{1}\cdot p_{4})(p_{2}\cdot p_{3})F_{\mathrm{ann}}^{LL}(\nu^{(1)},\bar{\nu}^{(2)},e^{(3)},\bar{e}^{(4)})$
$+4(p_{1}\cdot p_{3})(p_{2}\cdot p_{4})F_{\mathrm{ann}}^{RR}(\nu^{(1)},\bar{\nu}^{(2)},e^{(3)},\bar{e}^{(4)})$
$+2(p_{1}\cdot p_{2})m_{e}^{2}\left(F_{\mathrm{ann}}^{LR}(\nu^{(1)},\bar{\nu}^{(2)},e^{(3)},\bar{e}^{(4)})+F_{\mathrm{ann}}^{RL}(\nu^{(1)},\bar{\nu}^{(2)},e^{(3)},\bar{e}^{(4)})\right)\Big{]}\,.$ (108)
where $\mathrm{d}\pi_{i}=\frac{\\ dd^{3}p_{i}}{(2\pi)^{3}2E_{i}}$ with $E_{i}=p_{i}$ if the particle $i$ is an (anti)neutrino or $E_{i}=\sqrt{p_{i}^{2}+m_{e}^{2}}$ if the particle $i$ is an electron/positron. The statistical factors incorporating the density distributions for all participating particles are
$F_{\mathrm{sc}}^{AB}(\nu^{(1)},e^{(2)},\nu^{(3)},e^{(4)})$ $=f_{e}^{(4)}(1-f_{e}^{(2)})\left[G^{A}\varrho^{(3)}G^{B}(1-\varrho^{(1)})+(1-\varrho^{(1)})G^{B}\varrho^{(3)}G^{A}\right]$
$-f_{e}^{(2)}(1-f_{e}^{(4)})\left[\varrho^{(1)}G^{B}(1-\varrho^{(3)})G^{A}+G^{A}(1-\varrho^{(3)})G^{B}\varrho^{(1)}\right]\,,$ (109)
$F_{\mathrm{ann}}^{AB}(\nu^{(1)},\bar{\nu}^{(2)},e^{(3)},\bar{e}^{(4)})$ $=f_{e}^{(3)}\bar{f}_{e}^{(4)}\left[G^{A}(1-\bar{\varrho}^{(2)})G^{B}(1-\varrho^{(1)})+(1-\varrho^{(1)})G^{B}(1-\bar{\varrho}^{(2)})G^{A}\right]$
$-(1-f_{e}^{(3)})(1-\bar{f}_{e}^{(4)})\left[G^{A}\bar{\varrho}^{(2)}G^{B}\varrho^{(1)}+\varrho^{(1)}G^{B}\bar{\varrho}^{(2)}G^{A}\right]\,,$ (110)
where $a,b\in\{L,R\}$ denote the charalitics for the electrons/positrons involved in the processes, $f(\bar{f})_{e}^{(i)}=1/(\exp{p_{i}/T_{\gamma}}+1)$ are the distribution functions for electrons/positrons, and $\varrho^{(i)},\bar{\varrho}^{(i)}$ are the corresponding density matrices for neutrinos and antineutrinos with momentum $p_{i}$. The interaction matrices used in the collision terms are
$G^{L}=\text{diag}(g_{L},\tilde{g}_{L},\tilde{g}_{L})\,,\qquad G^{R}=\text{diag}(g_{R},g_{R},g_{R})\,,$ (111)
where $g_{L}=\sin^{2}\theta_{W}+1/2$, $\tilde{g}_{L}=\sin^{2}\theta_{W}-1/2$, $g_{R}=\sin^{2}\theta_{W}$, and $\theta_{W}$ is the weak mixing angle.
The collision term for the neutrino self-interactions, including the scattering between neutrinos and (anti)neutrinos ($\nu\nu\leftrightarrow\nu\nu$, $\nu\bar{\nu}\leftrightarrow\nu\bar{\nu}$) and the annihilation between a pair of neutrinos and antineutrinos ($\nu\bar{\nu}\leftrightarrow\nu\bar{\nu}$), can be written as
$\mathcal{I}_{\nu\nu}=$ $\frac{1}{2}\frac{2^{5}G_{F}^{2}}{2p_{1}}\int{\mathrm{d}\pi_{2}\,\mathrm{d}\pi_{3}\,\mathrm{d}\pi_{4}\,(2\pi)^{4}\delta^{(4)}(p_{1}+p_{2}-p_{3}-p_{4})}$ (112)
$\Big{[}(p_{1}\cdot p_{2})(p_{3}\cdot p_{4})F_{\mathrm{sc}}(\nu^{(1)},\nu^{(2)},\nu^{(3)},\nu^{(4)})$
$+(p_{1}\cdot p_{4})(p_{2}\cdot p_{3})\left(F_{\mathrm{sc}}(\nu^{(1)},\bar{\nu}^{(2)},\nu^{(3)},\bar{\nu}^{(4)})+F_{\mathrm{ann}}(\nu^{(1)},\bar{\nu}^{(2)},\nu^{(3)},\bar{\nu}^{(4)})\right)\Big{]}\,,$
with the statistical factors for scattering and annihilation processes:
$F_{\mathrm{sc}}(\nu^{(1)},\nu^{(2)},\nu^{(3)},\nu^{(4)})=\\ \left[\varrho^{(4)}(\mathbb{1}-\varrho^{(2)})+\mathrm{Tr}(\cdots)\right]\varrho^{(3)}(\mathbb{1}-\varrho^{(1)})+(\mathbb{1}-\varrho^{(1)})\varrho^{(3)}\left[(\mathbb{1}-\varrho^{(2)})\varrho^{(4)}+\mathrm{Tr}(\cdots)\right]\\ -\left[(\mathbb{1}-\varrho^{(4)})\varrho^{(2)}+\mathrm{Tr}(\cdots)\right](\mathbb{1}-\varrho^{(3)})\varrho^{(1)}-\varrho^{(1)}(\mathbb{1}-\varrho^{(3)})\left[\varrho^{(2)}(\mathbb{1}-\varrho^{(4)})+\mathrm{Tr}(\cdots)\right]\,,$ (113)
$F_{\mathrm{sc}}(\nu^{(1)},\bar{\nu}^{(2)},\nu^{(3)},\bar{\nu}^{(4)})=\\ \left[(\mathbb{1}-\bar{\varrho}^{(2)})\bar{\varrho}^{(4)}+\mathrm{Tr}(\cdots)\right]\varrho^{(3)}(\mathbb{1}-\varrho^{(1)})+(\mathbb{1}-\varrho^{(1)})\varrho^{(3)}\left[\bar{\varrho}^{(4)}(\mathbb{1}-\bar{\varrho}^{(2)})+\mathrm{Tr}(\cdots)\right]\\ -\left[\bar{\varrho}^{(2)}(\mathbb{1}-\bar{\varrho}^{(4)})+\mathrm{Tr}(\cdots)\right](\mathbb{1}-\varrho^{(3)})\varrho^{(1)}-\varrho^{(1)}(\mathbb{1}-\varrho^{(3)})\left[(\mathbb{1}-\bar{\varrho}^{(4)})\bar{\varrho}^{(2)}+\mathrm{Tr}(\cdots)\right]\,,$ (114)
$F_{\mathrm{ann}}(\nu^{(1)},\bar{\nu}^{(2)},\nu^{(3)},\bar{\nu}^{(4)})=\\ \left[\varrho^{(3)}\bar{\varrho}^{(4)}+\mathrm{Tr}(\cdots)\right](\mathbb{1}-\bar{\varrho}^{(2)})(\mathbb{1}-\varrho^{(1)})+(\mathbb{1}-\varrho^{(1)})(\mathbb{1}-\bar{\varrho}^{(2)})\left[\bar{\varrho}^{(4)}\varrho^{(3)}+\mathrm{Tr}(\cdots)\right]\\ -\left[(\mathbb{1}-\varrho^{(3)})(\mathbb{1}-\bar{\varrho}^{(4)})+\mathrm{Tr}(\cdots)\right]\bar{\varrho}^{(2)}\varrho^{(1)}-\varrho^{(1)}\bar{\varrho}^{(2)}\left[(\mathbb{1}-\bar{\varrho}^{(4)})(\mathbb{1}-\varrho^{(3)})+\mathrm{Tr}(\cdots)\right]\,,$ (115)
where, as before, $\varrho^{(i)},\bar{\varrho}^{(i)}$ are the corresponding density matrices for neutrinos and antineutrinos with momentum $p_{i}$, and $\mathrm{Tr}(\cdots)$ means the trace of the term in front of it.
Finally, the collision terms for the antineutrinos ¯¯$\bar{\mathcal{I}}\equiv\bar{\mathcal{I}}_{\rm sc}+\bar{\mathcal{I}}_{\rm ann}+\bar{\mathcal{I}}_{\nu\nu}$ can be obtained simply by the transformations $\varrho\leftrightarrow\bar{\varrho}$ and $L\leftrightarrow R$.
A.2 Reduced expressions for the comoving collision terms
The comoving collision terms $\tilde{\mathcal{I}}$ in (3.1) are just the above collision terms expressed with the comoving momentum and comoving energies differing by a total factor. In addition, for computational convenience, the nine-dimensional integrals in the collision terms are usually simplified to two-dimensional integrals, using the homogeneous and isotropic condition and the special form of the scattering amplitudes Hannestad:1995rs ; Dolgov:1997mb ; Grohs:2015tfy ; Blaschke:2016xxt . As a result, the collision term can be simplified as
$\tilde{\mathcal{I}}$ $=$ $\frac{G_{F}^{2}}{(2\pi)^{3}y^{2}}\left(\tilde{\mathcal{I}}_{\rm sc}+\tilde{\mathcal{I}}_{\rm ann}+\tilde{\mathcal{I}}_{\nu\nu}\right)\,,$ (116)
~$\tilde{\mathcal{I}}_{\rm sc}$ $=$ $\int{\rm d}y_{2}{\rm d}y_{3}\frac{y_{2}}{E_{2}}$
$\left\{\left(\Pi_{2}^{s}(y,y_{4})+\Pi_{2}^{s}(y,y_{2})\right)\left[F_{\mathrm{sc}}^{LL}(\nu^{(1)},e^{(2)},\nu^{(3)},e^{(4)})+F_{\mathrm{sc}}^{RR}(\nu^{(1)},\bar{e}^{(2)},\nu^{(3)},\bar{e}^{(4)})\right]\right.$
$\left.-2(x^{2}+\delta m_{e}^{2})\Pi_{1}^{s}(y,y_{3})\left[F_{\mathrm{sc}}^{LR}(\nu^{(1)},e^{(2)},\nu^{(3)},e^{(4)})+F_{\mathrm{sc}}^{LR}(\nu^{(1)},\bar{e}^{(2)},\nu^{(3)},\bar{e}^{(4)})\right]\right\}\,,$
~$\tilde{\mathcal{I}}_{\rm ann}$ $=$ $\int{\rm d}y_{2}{\rm d}y_{3}\frac{y_{3}}{E_{3}}$
$\left\{\Pi_{2}^{a}(y,y_{4})F_{\rm ann}^{LL}(\nu^{(1)},\bar{\nu}^{(2)},e^{(3)},\bar{e}^{(4)})+\Pi_{2}^{a}(y,y_{3})F_{\rm ann}^{RR}(\nu^{(1)},\bar{\nu}^{(2)},e^{(3)},\bar{e}^{(4)})\right.$
$\left.+(x^{2}+\delta m_{e}^{2})\Pi_{1}^{a}(y,y_{2})\left[F_{\rm ann}^{RL}(\nu^{(1)},\bar{\nu}^{(2)},e^{(3)},\bar{e}^{(4)})+F_{\rm ann}^{LR}(\nu^{(1)},\bar{\nu}^{(2)},e^{(3)},\bar{e}^{(4)})\right]\right\}\,,$
~$\tilde{\mathcal{I}}_{\nu\nu}$ $=$ $\frac{1}{4}\int{\rm d}y_{2}{\rm d}y_{3}\,$
$\left\{\Pi_{2}^{\nu}(y,y_{2})F_{\nu\nu}(\nu^{(1)},\bar{\nu}^{(2)},\nu^{(3)},\bar{\nu}^{(4)})+\Pi_{2}^{\nu}(y,y_{4})F_{\nu\bar{\nu}}(\nu^{(1)},\bar{\nu}^{(2)},\nu^{(3)},\bar{\nu}^{(4)})\right\}\,,$
where $E^{2}_{i}=\sqrt{x^{2}+y_{i}^{2}+\delta m_{e}^{2}}$ is the comoving energy for electrons/positrons taking into account the mass contribution induced by FTQED corrections (see Appendix B for more details) and the collision kernels in the collision terms are
$\Pi_{1}^{s}(y,y_{3})$ $=$ $y\,y_{3}\,D_{1}+D_{2}(y,y_{3},y_{2},y_{4}),$ (120)
$\Pi_{1}^{a}(y,y_{2})$ $=$ $y\,y_{2}\,D_{1}-D_{2}(y,y_{2},y_{3},y_{4}),$ (121)
$\Pi_{2}^{s}(y,y_{2})/2$ $=$ $y\,E_{2}\,y_{3}\,E_{4}\,D_{1}+D_{3}-y\,E_{2}D_{2}(y_{3},y_{4},y,y_{2})$
$-y_{3}\,E_{4}D_{2}(y,y_{2},y_{3},y_{4}),$
$\Pi_{2}^{s}(y,y_{4})/2$ $=$ $y\,E_{2}\,y_{3}\,E_{4}\,D_{1}+D_{3}+E_{2}\,y_{3}D_{2}(y,y_{4},y_{2},y_{3})$
$+y\,E_{4}D_{2}(y_{2},y_{3},y,y_{4}),$
$\Pi_{2}^{a}(y,y_{3})/2$ $=$ $y\,y_{2}\,E_{3}\,E_{4}\,D_{1}+D_{3}+y\,E_{3}D_{2}(y_{2},y_{4},y,y_{3})$
$+y_{2}\,E_{4}D_{2}(y,y_{3},y_{2},y_{4}),$
$\Pi_{2}^{a}(y,y_{4})/2$ $=$ $y\,y_{2}\,E_{3}\,E_{4}\,D_{1}+D_{3}+y_{2}\,E_{3}D_{2}(y,y_{4},y_{2},y_{3})$
$+y\,E_{4}D_{2}(y_{2},y_{3},y,y_{4}),$
$\Pi_{2}^{\nu}(y,y_{2})/2$ $=$ $y\,y_{2}\,y_{3}\,y_{4}\,D_{1}+D_{3}-y\,y_{2}D_{2}(y_{3},y_{4},y,y_{2})-y_{3}\,y_{4}D_{2}(y,y_{2},y_{3},y_{4}),$ (126)
$\Pi_{2}^{\nu}(y,y_{4})/2$ $=$ $y\,y_{2}\,y_{3}\,y_{4}\,D_{1}+D_{3}+y_{2}\,y_{3}D_{2}(y,y_{4},y_{2},y_{3})+y\,y_{4}D_{2}(y_{2},y_{3},y,y_{4}),$ (127)
where the functions $D_{i}$ are defined as Dolgov:1997mb :
$D_{1}(a,b,c,d)$ $=$ $\frac{16}{\pi}\int_{0}^{\infty}\frac{{\rm d}\lambda}{\lambda^{2}}\prod_{i=a,b,c,d}\sin(\lambda i)\,,$ (128)
$D_{2}(a,b,c,d)$ $=$ $-\frac{16}{\pi}\int_{0}^{\infty}\frac{{\rm d}\lambda}{\lambda^{4}}\prod_{i=a,b}\Big{[}\lambda i\cos(\lambda i)-\sin(\lambda i)\Big{]}\prod_{j=c,d}\sin(\lambda j)\,,$ (129)
$D_{3}(a,b,c,d)$ $=$ $\frac{16}{\pi}\int_{0}^{\infty}\frac{\mathrm{d}\lambda}{\lambda^{6}}\prod_{i=a,b,c,d}\Big{[}\lambda i\cos(\lambda i)-\sin(\lambda i)\Big{]}\,.$ (130)
Again, the comoving collision terms for the antineutrinos ¯~¯~$\tilde{\bar{\mathcal{I}}}\equiv\tilde{\bar{\mathcal{I}}}_{\rm sc}+\tilde{\bar{\mathcal{I}}}_{\rm ann}+\tilde{\bar{\mathcal{I}}}_{\nu\nu}$ can be obtained simply by the transformations $\varrho\leftrightarrow\bar{\varrho}$ and $L\leftrightarrow R$.
Appendix B Finite Temperature QED Corrections
The finite temperature QED (FTQED) corrections on neutrino decoupling must be taking into account to obtain accurate density matrices and the resulting $N_{\rm eff}$ Fornengo:1997wa ; Mangano:2001iu ; Bennett:2019ewm . In this work we consider the following three aspects of the FTQED corrections.
B.1 FTQED correction to the continuity equation
The continuity equation of the Universe with the FTQED corrections is usually expressed as a differential equation for the dimensionless photon temperature $z$ and are solved combing with the QKEs in practice,
$\frac{\mathrm{d}z}{\mathrm{d}x}=\cfrac{rJ_{2}(r)+G_{1}(r)-\cfrac{1}{4\pi^{2}z^{3}}{\displaystyle\int_{0}^{\infty}\mathrm{d}y\,y^{3}\sum_{\alpha}\left(\cfrac{\mathrm{d}\varrho_{\alpha\alpha}}{\mathrm{d}x}+\cfrac{\mathrm{d}\bar{\varrho}_{\alpha\alpha}}{\mathrm{d}x}\right)}}{\Big{[}r^{2}J_{2}(r)+J_{4}(r)\Big{]}+G_{2}(r)+\cfrac{2\pi^{2}}{15}}\,,$ (131)
where $r=x/z$, $\frac{\mathrm{d}\varrho_{\alpha\alpha}}{\mathrm{d}x}$ and $\frac{\mathrm{d}\bar{\varrho}_{\alpha\alpha}}{\mathrm{d}x}$ are the diagonal components of the QKEs for neutrinos and antineutrinos. The remaining functions appearing in the equation are defined as
$J_{a}(r)$ $=$ $\frac{1}{\pi^{2}}\int_{0}^{\infty}{\rm d}u\,u^{a}\frac{\exp(\sqrt{u^{2}+r^{2}})}{\Big{[}\exp(\sqrt{u^{2}+r^{2}})+1\Big{]}^{2}}\,,$ (132)
$K_{a}(r)$ $=$ $\frac{1}{\pi^{2}}\int_{0}^{\infty}{\rm d}u\,\frac{u^{a}}{\sqrt{u^{2}+r^{2}}}\,\frac{1}{\exp(\sqrt{u^{2}+r^{2}})+1}\,.$ (133)
and $G_{1,2}$ functions are expanded as a series of powers of the electron charge $e^{2}=4\pi\alpha$ with $\alpha$ the fine structure constant,
$G_{1,2}(x,z)$ $=$ $G_{1,2}^{(2)}(x/z)+G_{1,2}^{(2+\ln)}(x,z)+G_{1,2}^{(3)}(x/z)+\ldots\,.$ (134)
At the $\mathcal{O}(e^{2})$ order we have
$G_{1}^{(2)}(r)$ $=$ $2\pi\alpha\left[\frac{1}{r}\left(\frac{K_{2}}{3}+2K_{2}^{2}-\frac{J_{2}}{6}-K_{2}J_{2}\right)+G_{a}\right]\,,$ (135)
$G_{2}^{(2)}(r)$ $=$ $-8\pi\alpha\left(\frac{K_{2}}{6}+\frac{J_{2}}{6}-\frac{1}{2}K_{2}^{2}+K_{2}J_{2}\right)+2\pi\alpha rG_{a}\,,$ (136)
$G_{a}(r)$ $=$ $\frac{K_{2}^{\prime}}{6}-K_{2}K_{2}^{\prime}+\frac{J_{2}^{\prime}}{6}+K_{2}^{\prime}J_{2}+K_{2}J_{2}^{\prime}\,,$ (137)
where the prime denotes derivative with respect to $r$ and the explicit dependence on $r$ for the $G$ functions are dropped for simplicity. An additional log term at the $\mathcal{O}(e^{2})$ order is
$G_{1}^{(2+\ln)}(x,z)$ $=$ $\frac{e^{2}x}{16\pi^{4}z^{3}}\int\!\!\!\int_{0}^{\infty}{\rm d}y\,{\rm d}k\,\frac{y\,k}{E_{y}E_{k}}\ln\left|\frac{y+k}{y-k}\right|\Bigg{\{}-x\bigg{[}z\Big{(}\partial_{x}\mathcal{N}_{y}\partial_{z}\mathcal{N}_{k}+\mathcal{N}_{y}\partial_{x}\partial_{z}\mathcal{N}_{k}\Big{)}-\mathcal{N}_{y}\partial_{x}\mathcal{N}_{k}\bigg{]}$ (138)
$\qquad\qquad\qquad\qquad-\mathcal{N}_{y}\mathcal{N}_{k}-z\mathcal{N}_{y}\partial_{z}\mathcal{N}_{k}+\frac{x^{2}(E_{y}^{2}+E_{k}^{2})}{2E_{y}^{2}E_{k}^{2}}\Big{(}2z\mathcal{N}_{y}\partial_{z}\mathcal{N}_{k}-\mathcal{N}_{y}\mathcal{N}_{k}\Big{)}\Bigg{\}}\,,$
$G_{2}^{(2+\ln)}(x,z)$ $=$ $\frac{e^{2}x^{2}}{16\pi^{4}z^{2}}\int\!\!\!\int_{0}^{\infty}{\rm d}y\,{\rm d}k\,\frac{y\,k}{E_{y}E_{k}}\ln\left|\frac{y+k}{y-k}\right|\partial_{z}\Big{(}\mathcal{N}_{y}\partial_{z}\mathcal{N}_{k}\Big{)}\,,$ (139)
where
$\mathcal{N}_{p}$ $=$ $\frac{2}{e^{E_{p}/z}+1}\,,$ (140)
$\partial_{x}\mathcal{N}_{p}$ $=$ $-\frac{x\,e^{E_{p}/z}\,\mathcal{N}_{p}^{2}}{2z\,E_{p}}\,,$ (141)
$\partial_{z}\mathcal{N}_{p}$ $=$ $\frac{e^{E_{p}/z}\,E_{p}\,\mathcal{N}_{p}^{2}}{2z^{2}}\,,$ (142)
$\partial_{x}\partial_{z}\mathcal{N}_{p}$ $=$ $\frac{x\,e^{E_{p}/z}\,\mathcal{N}_{p}^{2}}{2z^{3}}\left(1-e^{E_{p}/z}\,\mathcal{N}_{p}+\frac{z}{E_{p}}\right)\,.$ (143)
At the $\mathcal{O}(e^{3})$ order we have
$G_{1}^{(3)}(r)$ $=$ $\frac{e^{3}}{4\pi}G_{b}\left\{\frac{1}{r}\Big{(}2J_{2}-4K_{2}\Big{)}-2J_{2}^{\prime}-r^{2}J_{0}^{\prime}-r\Big{(}2K_{0}+J_{0}\Big{)}\right.$ (144)
$\left.-G_{c}\Big{[}r(K_{0}-J_{0})+K_{2}^{\prime}\Big{]}\right\}\,,$
$G_{2}^{(3)}(r)$ $=$ $\frac{e^{3}}{4\pi}G_{b}\left[G_{c}\Big{(}2J_{2}+r^{2}J_{0}\Big{)}-\frac{2}{r}J_{4}^{\prime}-r\Big{(}3J_{2}^{\prime}+r^{2}J_{0}^{\prime}\Big{)}\right]\,,$ (145)
with
$G_{b}(r)$ $=$ $\sqrt{K_{2}+r^{2}\frac{K_{0}}{2}}\,,$ (146)
$G_{c}(r)$ $=$ $\frac{2J_{2}+r^{2}J_{0}}{2\Big{(}2K_{2}+r^{2}K_{0}\Big{)}}\,.$ (147)
B.2 FTQED corrections to the total energies and the pressures
The total pressure and energy density of the universe are also modified by the FTQED corrections:
$P_{\rm tot}$ $=$ $\sum_{i=\gamma,{\nu_{i}},e}P_{i}+\delta P(x,z)\,,$ (148)
$\rho_{\rm tot}$ $=$ $\sum_{i=\gamma,{\nu_{i}},e}\rho_{i}+\delta\rho(x,z)\,,$ (149)
where $\delta P$ and $\delta\rho$ can also be expanded as a series of powers of the electron charge, so that we have Bennett:2019ewm
$\delta P(x,z)$ $=$ $\delta P^{(2)}(x/z)+\delta P^{(2+\ln)}(x,z)+\delta P^{(3)}(x/z)+\ldots\,,$ (150)
$\delta P^{(2)}(r)$ $=$ $-e^{2}z^{4}\,K_{2}\left(\frac{1}{6}+\frac{K_{2}}{2}\right)\,,$ (151)
$\delta P^{(2+\ln)}(x,z)$ $=$ $\frac{e^{2}x^{2}}{16\pi^{4}}\int\!\!\!\int_{0}^{\infty}{\rm d}y\,{\rm d}k\,\frac{y\,k}{E_{y}E_{k}}\ln\left|\frac{y+k}{y-k}\right|\mathcal{N}_{y}\,\mathcal{N}_{k}\,,$ (152)
$\delta P^{(3)}(r)$ $=$ $\frac{2e^{3}z^{4}}{3\pi}\left(K_{2}+\frac{r^{2}}{2}K_{0}\right)^{3/2}\,,$ (153)
for the pressure, while
$\delta\rho(x,z)$ $=$ $\delta\rho^{(2)}(x/z)+\delta\rho^{(2+\ln)}(x,z)+\delta\rho^{(3)}(x/z)+\ldots\,,$ (154)
$\delta\rho^{(2)}(r)$ $=$ $e^{2}z^{4}\left(\frac{K_{2}^{2}}{2}-\frac{K_{2}+J_{2}}{6}-K_{2}J_{2}\right)\,,$ (155)
$\delta\rho^{(2+\ln)}(x,z)$ $=$ $\frac{e^{2}x^{2}}{16\pi^{4}}\int\!\!\!\int_{0}^{\infty}{\rm d}y\,{\rm d}k\,\frac{y\,k}{E_{y}E_{k}}\ln\left|\frac{y+k}{y-k}\right|\mathcal{N}_{y}\Big{(}2z\partial_{z}\mathcal{N}_{k}-\mathcal{N}_{k}\Big{)}\,,$ (156)
$\delta\rho^{(3)}(r)$ $=$ $\frac{e^{3}z^{4}}{\pi}\left(K_{2}+\frac{r^{2}}{2}K_{0}\right)^{1/2}\left(J_{2}+\frac{r^{2}}{2}J_{0}\right)\,.$ (157)
for the energy densities.
B.3 FTQED correction to the electron mass
In addition, the mass of electron/positrons are also modified by the FTQED correction, which is mainly used in the collision terms. In the comoving coordinates, the additional FTQED contribution to the electron mass can be written as Fornengo:1997wa ; Mangano:2001iu ; Bennett:2019ewm :
$\delta m_{e}^{2}(x,y,z)=\frac{2\pi\alpha z^{2}}{3}+\frac{4\alpha}{\pi}\int_{0}^{\infty}{\rm d}k\,\frac{k^{2}}{E_{k}}\frac{1}{e^{E_{k}/z}+1}-\frac{x^{2}\alpha}{\pi y}\int_{0}^{\infty}{\rm d}k\,\frac{k}{E_{k}}\log\left|\frac{y+k}{y-k}\right|\mathcal{N}_{k}\,,$ (158)
so that the comoving electron mass in the collision terms is replaced using $x^{2}\rightarrow x^{2}+\delta m_{e}^{2}$. Finally, we note that the the log term that depends on $y$ in the FTQED corrections to the continuity equation, the total energy densities and the pressure, as long as the electron mass are ignored in the computation, as it is shown that their contribution is negligible Froustey:2020mcq ; Bennett:2020zkv .

### v2

@@
 The implications of the neutrino asymmetry for the CMB and LSS are studied in detail, and we find that the Baryon Acoustic Oscillations (BAO) are also significantly affected by $\xi_{\nu}$ in addition to the sum of neutrino masses.
-A combined analysis with EMPRESS BBN, Planck CMB and BOSS BAO data yields a tighter constraint $\xi_{\nu}=0.024\pm 0.013$, which provides constraints on UV models capable of producing large asymmetries.
+A combined analysis with EMPRESS BBN, Planck CMB and BOSS BAO data yields a tighter constraint $\xi_{\nu}=0.024\pm 0.012$, which provides constraints on UV models capable of producing large asymmetries.
 1 Introduction
@@
 For the Standard Model case, where asymmetries between neutrinos and antineutrinos are assumed to be negligible, significant progress has been made in detailing the neutrino decoupling process over recent decades Hannestad:1995rs ; Dolgov:1997mb ; Mangano:2001iu ; Mangano:2005cc ; Grohs:2015tfy ; deSalas:2016ztq ; Akita:2020szl ; Froustey:2019owm ; EscuderoAbenza:2020cmq ; Froustey:2020mcq ; Bennett:2020zkv .
-Including flavor oscillations, matter effects, finite-temperature QED corrections up to $\mathcal{O}(3)$, and full neutrino-electron and neutrino-neutrino collision terms, the current standard prediction yields $N_{\rm eff}(0)=3.0440\pm 0.0002$ Froustey:2020mcq ; Bennett:2020zkv .
+Including flavor oscillations, matter effects, Finite-Temperature QED (FTQED) corrections up to $\mathcal{O}(3)$, and full neutrino-electron and neutrino-neutrino collision terms, the current standard prediction yields $N_{\rm eff}^{\rm SM}=3.0440\pm 0.0002$ Froustey:2020mcq ; Bennett:2020zkv ; Akita:2020szl .
 On the other hand, inspired by the recent EMPRESS survey on primordial helium abundance empress and its potential to alleviate cosmological tensions Barenboim:2016lxv ; Yeung:2020zde ; Seto:2021tad ; Kumar:2022vee ; Yeung:2024krv , the scenario where asymmetries between neutrinos and antineutrinos are not negligible has recently attracted significant attention.
-Due to the sphaleron processes in the early universe Kuzmin:1985mm ; Khlebnikov:1988sr ; Harvey:1990qw ; Dreiner:1992vm , one might naively expect neutrino asymmetries, or lepton asymmetries, to be of the same order as the baryon asymmetry of the universe, which is strongly constrained by BBN and CMB observations: $\eta_{B}\equiv(n_{B}-n_{\bar{B}})/n_{\gamma}=(6.14\pm 0.04)\times 10^{-10}$ Planck:2018vyg , where $n_{B}$, $n_{\bar{B}}$, and $n_{\gamma}$ denote the number densities of baryons, antibaryons, and photons, respectively.
+Due to the sphaleron processes in the early universe Kuzmin:1985mm ; Khlebnikov:1988sr ; Harvey:1990qw ; Dreiner:1992vm , one might naively expect neutrino asymmetries, or lepton asymmetries, to be of the same order as the baryon asymmetry of the universe, which is strongly constrained by BBN and CMB observations: $\eta_{B}\equiv(n_{B}-n_{\bar{B}})/n_{\gamma}=(6.14\pm 0.04)\times 10^{-10}$ Planck:2018vyg , where $n_{B}$, $n_{\bar{B}}$, and $n_{\gamma}$ denote the number densities of baryons, antibaryons and photons, respectively.
 However, substantial lepton asymmetries compared to the baryon asymmetry before the neutrino decoupling epoch can still be generated by various models, see $e.g.,$ Casas:1997gx ; Dolgov:1989us ; Bajc:1997ky ; Asaka:2005pn ; Asaka:2005an ; Pilaftsis:2003gt ; Borah:2022uos ; Kawasaki:2002hq ; Kawasaki:2022hvx .
-The lepton asymmetries prior to the neutrino decoupling epoch, hereafter referred to as primordial neutrino asymmetries, have significant implications for the evolution of the universe.
-During the neutrino decoupling epoch, primordial neutrino asymmetries lead to an increased $N_{\rm eff}$ and specific spectral distortions of neutrinos and antineutrinos Grohs:2016cuu .
-Furthermore, if the asymmetries among different flavors are different, neutrinos and antineutrinos can undergo synchronous oscillations Bell:1998ds ; Pastor:2001iu ; Abazajian:2002qx ; Wong:2002fa , resulting in the redistribution of asymmetries among flavors Dolgov:2002ab ; Pastor:2008ti ; Mangano:2010ei ; Mangano:2011ip ; Castorina:2012md ; Barenboim:2016shh ; Johns:2016enc ; Froustey:2021azz .
-After neutrino decoupling, during the BBN epoch, primordial neutrino asymmetries primarily affect the final positron-to-neutron ratio and thus the final abundances of light elements.
-This effect mainly depends on the density distributions of electron neutrinos and antineutrinos, while corrections from the increased $N_{\rm eff}$ are subleading Sarkar:1995dd ; Iocco:2008va ; Pitrou:2018cgg ; Serpico:2005bc ; Mangano:2011ip ; Chu:2006ua ; Simha:2008mt ; Saviano:2013ktj ; Burns:2023sgx .
-Additionally, primordial neutrino asymmetries influence CMB anisotropies and LSS through increased $N_{\rm eff}$ and modified light element abundances, alongside the effects of massive neutrinos.
-Consequently, observations of BBN, CMB, and LSS can be utilized to constrain primordial neutrino asymmetries Simha:2008mt ; Oldengott:2017tzj ; Pitrou:2018cgg ; Matsumoto:2022tlr ; Burns:2022hkq ; Escudero:2022okz .
-In particular, recent studies from the EMPRESS survey indicate a preference for a positive asymmetry in electron neutrinos Matsumoto:2022tlr ; Burns:2022hkq ; Escudero:2022okz .
-Specifically, combining BBN and CMB observations yields $\xi_{\nu_{e}}=0.034\pm 0.014$ Escudero:2022okz , where $\xi_{\nu_{e}}=\mu_{\nu_{e}}/T_{\nu}$ represents the degeneracy parameter for electron neutrinos.
-However, the current situation is less satisfactory compared to the Standard Model (SM) case due to persistent ambiguities regarding neutrino decoupling.
-These ambiguities particularly involve neutrino spectral distortions and the effects of flavor oscillations and QED corrections.
-As a result, the effects of primordial neutrino asymmetries on BBN and CMB, as well as the constraints derived from observational data, are generally evaluated using approximations.
-These approximations often include the instantaneous neutrino decoupling approximation and the assumption of massless neutrinos.
-To derive precise and accurate constraints on primordial neutrino asymmetries from both current and future observations, it is essential to rigorously analyze the implications of these asymmetries through a comprehensive treatment of the neutrino decoupling process.
-A recent study Froustey:2024mgf makes effort toward a rigorous analysis of primordial neutrino asymmetries.
-This study focuses on the implications of neutrino synchronous oscillations for BBN, while ignoring the correction from (anti)neutrino spectral distortions, etc, and furthermore, the implications for both the CMB and LSS are not addressed at all.
-In this work, we focus on evaluating the neutrino decoupling process and its subsequent effects on BBN, CMB and LSS in the presence of primordial neutrino asymmetries.
-To accurately describe the evolution of neutrino density matrices during decoupling, including flavor oscillations and full collision effects, it is crucial to employ the complete set of neutrino quantum kinetic equations (QKEs).
-These equations can be derived using various approaches, such as perturbative expansions of the density matrix Sigl:1993ctk , relativistic generalizations of the Bogoliubov-Born-Green-Kirkwood-Yvon (BBGKY) equations Volpe:2013uxl ; Volpe:2015rla ; Froustey:2020mcq , and more generally from the Closed-Time Path (CTP) formalism of non-equilibrium quantum field theory Vlasenko:2013fja ; Blaschke:2016xxt .
-The main improvements in our study are as follows:
+The lepton asymmetries present before the neutrino decoupling epoch, henceforth referred to as primordial neutrino asymmetries, have significant implications for the universe’s evolution.
+During the neutrino decoupling epoch, these asymmetries contribute to an increased $N_{\rm eff}$ and induce specific spectral distortions in neutrinos and antineutrinos Grohs:2016cuu .
+If the asymmetries vary between different neutrino flavors, neutrinos and antineutrinos can undergo synchronous oscillations Bell:1998ds ; Pastor:2001iu ; Abazajian:2002qx ; Wong:2002fa , leading to a redistribution of asymmetries among the flavors Dolgov:2002ab ; Pastor:2008ti ; Mangano:2010ei ; Mangano:2011ip ; Castorina:2012md ; Barenboim:2016shh ; Johns:2016enc ; Froustey:2021azz .
+After neutrino decoupling, during the Big Bang Nucleosynthesis (BBN) epoch, primordial neutrino asymmetries primarily affect the final proton-to-neutron ratio and the resulting light element abundances.
+This effect depends largely on the distribution of electron neutrinos and antineutrinos, while corrections from the increased $N_{\rm eff}$ are secondary Sarkar:1995dd ; Iocco:2008va ; Pitrou:2018cgg ; Serpico:2005bc ; Mangano:2011ip ; Chu:2006ua ; Simha:2008mt ; Saviano:2013ktj ; Burns:2023sgx .
+Additionally, these asymmetries influence cosmic microwave background (CMB) anisotropies and large-scale structure (LSS) through the increased $N_{\rm eff}$ and modified light element abundances, along with the effects of massive neutrinos.
+Consequently, observations of BBN, CMB, and LSS can constrain primordial neutrino asymmetries Simha:2008mt ; Oldengott:2017tzj ; Pitrou:2018cgg ; Matsumoto:2022tlr ; Burns:2022hkq ; Escudero:2022okz .
+Notably, recent findings from the EMPRESS survey suggest a preference for a positive asymmetry in electron neutrinos Matsumoto:2022tlr ; Burns:2022hkq ; Escudero:2022okzx.
+Specifically, combining BBN and CMB observations yields $\xi_{\nu_{e}}=0.034\pm 0.014$ Escudero:2022okz , where $\xi_{\nu_{e}}=\mu_{\nu_{e}}/T_{\nu}$ is the degeneracy parameter for electron neutrinos.
+However, the current situation is less satisfactory compared to the Standard Model (SM) scenario due to persistent uncertainties concerning neutrino decoupling.
+These uncertainties involve neutrino spectral distortions, flavor oscillations, and finite-temperature quantum electrodynamics (FTQED) corrections.
+Consequently, the effects of primordial neutrino asymmetries on BBN and CMB, as well as the constraints derived from observational data, are generally evaluated using approximations, such as the instantaneous neutrino decoupling approximation and the assumption of massless neutrinos.
+To derive precise constraints on primordial neutrino asymmetries from both current and future observations, it is essential to rigorously analyze these asymmetries within a comprehensive framework of neutrino decoupling.
+A recent study Froustey:2024mgf makes an effort towards a rigorous analysis of primordial neutrino asymmetries.
+This study focuses on the implications of neutrino synchronous oscillations on the neutrino flavor equilibration and BBN process, while the corresponding constraints on the primordial neutrino asymmetries are obtained by combining the BBN measurements and the constraints on the parameter set ($N_{\rm eff},Y_{P},\omega_{b}$) from CMB + BAO experiments.
+However, the resulting (anti)neutrino spectral distortions and their implications for the BBN and CMB are not discussed in Froustey:2024mgf .
+Furthermore, the implications of primordial neutrino asymmetries for both the CMB angular power spectrum and the BAO are not addressed in Froustey:2024mgf , which according to our results are actually important for the accurate determination of the cosmological constraints.
+In this work, we focus on evaluating the neutrino decoupling process and its subsequent effects on BBN, CMB, and LSS in the presence of primordial neutrino asymmetries.
+To accurately describe the evolution of neutrino density matrices during decoupling, including flavor oscillations and full collision effects, we employ the complete set of neutrino quantum kinetic equations (QKEs).
+These equations are derived using several approaches, such as perturbative expansions of the density matrix Sigl:1993ctk , relativistic generalizations of the Bogoliubov-Born-Green-Kirkwood-Yvon (BBGKY) equations Volpe:2013uxl ; Volpe:2015rla ; Froustey:2020mcq , and the Closed-Time Path (CTP) formalism from non-equilibrium quantum field theory Vlasenko:2013fja ; Blaschke:2016xxt ; Kainulainen:2023ocv ; Kainulainen:2024fdg .
+The main improvements of our study are as follows:
   • We re-derive the most general form of the QKEs for neutrinos and antineutrinos using the CTP formalism, incorporating primordial neutrino asymmetry.
-  • We numerically solve the QKEs for the neutrino decoupling process in the presence of primordial neutrino asymmetry, including effects of flavor oscillations, matter interactions, finite-temperature QED (FTQED) corrections up to $\mathcal{O}(3)$ order, and complete neutrino-electron and neutrino-neutrino collision terms.
-This approach matches the accuracy used in current state-of-the-art predictions for the Standard Model (SM) case Froustey:2020mcq ; Bennett:2020zkv  [^1] Additional QED corrections, such as those to interaction vertices Cielo:2023bqp ; Jackson:2023zkl ; Drewes:2024wbw , are still debated for the SM case and are not included in this study..
-  • We use the resulting density matrices to study the impacts of primordial neutrino asymmetry on the BBN, CMB, and LSS.
-The effects of neutrino spectral distortions are also highlighted by comparing these results with those obtained using the Fermi-Dirac distribution for neutrinos.
-  • With the observation that the Baryon Acoustic Oscillations (BAO) are also affected by the primordial neutrino asymmetry, we derive the constraints on the primordial neutrino asymmetry by combining the traditional BBN and CMB datasets with the BAO observations for the first time.
-Figure 1: A schematic illustration outlining the structure of this paper.
-Section 2 presents the derivation of the QKEs for neutrinos within the CTP formalism.
-These QKEs are subsequently applied to study neutrino decoupling in the presence of primordial neutrino asymmetry $\xi_{\nu}$ in Section 3.
-Based on such precise treatment of neutrino decoupling, Sections 4 and 5 explore the effects of primordial neutrino asymmetry on BBN, CMB and LSS.
-For each section, the primary input and output physical quantities that depend on $\xi_{\nu}$ are labelled in each box.
-Finally, in Section 6, cosmological constraints on $\xi_{\nu}$ are obtained by the MCMC analyses combining the theoretical predictions and observational data from BBN, CMB, and LSS.
+  • We numerically solve the QKEs for the neutrino decoupling process, accounting for flavor oscillations, matter interactions, FTQED corrections up to $\mathcal{O}(3)$, and complete neutrino-electron and neutrino-neutrino collision terms, matching the accuracy of state-of-the-art Standard Model (SM) predictions Froustey:2020mcq ; Bennett:2020zkv  [^1] Additional QED corrections, such as those to interaction vertices Cielo:2023bqp ; Jackson:2023zkl ; Drewes:2024wbw , remain under debate for the SM case and are not included in this study..
+  • We use the resulting density matrices to study the impacts of primordial neutrino asymmetry on BBN, CMB, and LSS, with a particular emphasis on (anti)neutrino spectral distortions by comparing these results with those obtained using the Fermi-Dirac distribution for neutrinos.
+  • Recognizing that Baryon Acoustic Oscillations (BAO) are also influenced by primordial neutrino asymmetry, we derive constraints on primordial neutrino asymmetry by combining traditional BBN and CMB datasets with BAO observations for the first time.
+Figure 1: A schematic illustration of the paper’s structure.
+Section 2 derives the QKEs for neutrinos using the CTP formalism.
+These QKEs are then applied to study neutrino decoupling in the presence of primordial neutrino asymmetry $\xi_{\nu}$ in Section 3.
+Based on this precise treatment of neutrino decoupling, Sections 4 and 5 explore the effects of primordial neutrino asymmetry on BBN, the CMB, and LSS.
+For each section, the primary input and output physical quantities dependent on $\xi_{\nu}$ are labeled in each box.
+Finally, in Section 6, we obtain cosmological constraints on $\xi_{\nu}$ through Markov Chain Monte Carlo (MCMC) analyses, combining theoretical predictions with observational data from BBN, CMB, and LSS.
 The key physical results of this study are:
   • We numerically evaluate $N_{\rm eff}$ and the (anti)neutrino spectral distortions, incorporating primordial neutrino asymmetries with state-of-the-art accuracy.
-We find that $N_{\rm eff}$ can be expressed as $N_{\rm eff}=N_{\rm eff}^{\rm SM}+\delta N_{\rm eff}^{\xi_{\nu},\rm ther}+0.0102\,\xi_{\nu}^{2}$, where $N_{\rm eff}^{\rm SM}=3.0440\pm 0.002$ represents the SM prediction with comparable precision Froustey:2020mcq ; Bennett:2020zkv .
-The term $\delta N_{\rm eff}^{\xi_{\nu},\rm ther}$ accounts for the contribution from primordial neutrino asymmetry under the instantaneous decoupling approximation, and is given by $\delta N_{\rm eff}^{\xi_{\nu},\rm ther}\simeq 3\left(\frac{30}{7\pi^{2}}\xi_{\nu}^{2}+\frac{15}{7\pi^{4}}\xi_{\nu}^{4}\right)$, while the final term represents the additional correction from non-instantaneous decoupling effects in the presence with non-zero primordial neutrino asymmetry.
-  • We provide state-of-the-art predictions for light element abundances produced by BBN, through which the EMPRESS measurement on helium abundance $Y_{P}|_{\rm EMPRESS}=0.2370^{+0.0034}_{-0.0033}$ immediately indicate a positive primordial neutrino asymmetry $0.032\leq\xi_{\nu}\leq 0.052$ with fixed baryon abundance $\omega_{b}$.
-We also find that the (anti)neutrino spectral distortions are important for precise predictions for BBN, especially for the abundance of Helium-4 $Y_{P}$.
-  • For the constraints on primordial neutrino asymmetry, combing EMPRESS BBN and Planck CMB data yields $\xi_{\nu}=0.028\pm 0.013$, while incorporating BOSS BAO data in the analysis yields $\xi_{\nu}=0.024\pm 0.013$.
-Compared to previous result obtained in ref $\xi_{\nu}=0.034\pm 0.014$, our results indicate a $\sim 18\%$ and $\sim 29\%$ smaller central value and a slight reduction in the uncertainty, respectively.
-  • The impacts of treating the sum of neutrino mass $\sum m_{\nu}$ and the primordial neutrino asymmetry $\xi_{\nu}$ as free parameters on the resulting constraints for each of them are also discussed.
-Finally, we show how the constraints on $\xi_{\nu}$ can be converted into the constraints on UV model parameters with a specific Q-ball decay model capable of producing large primordial neutrino asymmetries.
+For primordial neutrino asymmetries with complete flavor equilibration before neutrino decoupling, $i.e.,$ $\xi_{\nu_{e}}=\xi_{\nu_{\mu}}=\xi_{\nu_{\tau}}=\xi_{\nu}$, we find that $N_{\rm eff}$ can be expressed as $N_{\rm eff}=N_{\rm eff}^{\rm SM}+\delta N_{\rm eff}^{\xi_{\nu},\rm ther}+0.0102\,\xi_{\nu}^{2}$, where $N_{\rm eff}^{\rm SM}=3.0440\pm 0.002$ represents the SM prediction Froustey:2020mcq ; Bennett:2020zkv .
+The term $\delta N_{\rm eff}^{\xi_{\nu},\rm ther}$ accounts for the contribution from primordial neutrino asymmetry under the instantaneous decoupling approximation and is given by $\delta N_{\rm eff}^{\xi_{\nu},\rm ther}\simeq 3\left(\frac{30}{7\pi^{2}}\xi_{\nu}^{2}+\frac{15}{7\pi^{4}}\xi_{\nu}^{4}\right)$.
+The final term represents the additional correction from non-instantaneous decoupling effects in the presence of non-zero primordial neutrino asymmetry.
+  • We provide state-of-the-art predictions for the abundances of light elements produced by BBN.
+Using the EMPRESS measurement of helium abundance, $Y_{P}|_{\rm EMPRESS}=0.2370^{+0.0034}_{-0.0033}$, we immediately indicate a positive primordial neutrino asymmetry, $0.032\leq\xi_{\nu}\leq 0.052$, with fixed baryon abundance $\omega_{b}$.
+We also find that (anti)neutrino spectral distortions are crucial for precise BBN predictions, especially for the abundance of Helium-4, $Y_{P}$.
+  • For constraints on primordial neutrino asymmetry, combining EMPRESS BBN and Planck CMB data yields $\xi_{\nu}=0.028\pm 0.012$, while incorporating BOSS BAO data results in $\xi_{\nu}=0.024\pm 0.012$.
+Compared to the previous result obtained in Ref.
+Escudero:2022okz , $\xi_{\nu}=0.034\pm 0.014$, our results show an approximately 18% and 29% reduction in the central value and a slight decrease in uncertainty, respectively.
+  • We also discuss the impacts of treating both the sum of neutrino masses $\sum m_{\nu}$ and the primordial neutrino asymmetry $\xi_{\nu}$ as free parameters on the resulting constraints for each.
+Finally, we demonstrate how the constraints on $\xi_{\nu}$ can be converted into constraints on ultraviolet (UV) model parameters, using a specific Q-ball decay model capable of generating large primordial neutrino asymmetries.
 This paper is organized as follows.
 In Section 2, we re-derive the QKEs for Dirac and Majorana neutrinos using the CTP formalism.
-In Section 3, we present the QKEs in the context of the early universe and provides corresponding numerical results considering primordial neutrino asymmetries.
-The impacts of primordial neutrino asymmetries on the BBN epoch is discussed in Section 4, while their effects on both CMB and LSS are analyzed in Section 5.
-In Section 6, we present the results from a joint MCMC analysis of neutrino primordial asymmetries and neutrino masses, utilizing datasets from BBN, CMB and BAO observations.
+Section 3 presents the QKEs in the context of the early universe and provides corresponding numerical results, considering primordial neutrino asymmetries.
+The impact of primordial neutrino asymmetries on the BBN epoch is discussed in Section 4, while their effects on both the CMB and LSS are analyzed in Section 5.
+In Section 6, we present results from a joint MCMC analysis of neutrino primordial asymmetries and neutrino masses, utilizing datasets from BBN, CMB, and BAO observations.
 Finally, we conclude in Section 7.
-To guide the readers, a schematic illustration of this work are presented in Fig.
+To guide the reader, a schematic illustration of this work is presented in Fig.
 1.
-Throughout this manuscript, we use the natural units where $\hbar=c=k_{B}=1$.
+Throughout this manuscript, we use natural units where $\hbar=c=k_{B}=1$.
 2 Quantum Kinetic Equations for Dirac and Majorana Neutrinos
@@
 3.1 QKEs in the early universe
-In the early universe, both the Standard Model (SM) thermal bath and the neutrino system can be approximated as homogeneous and isotropic.
+In the early universe, both the SM thermal bath and the neutrino system can be approximated as homogeneous and isotropic.
 Under these conditions, the QKEs for neutrinos can be significantly simplified.
 First, the drift and force terms on the left-hand side are eliminated, and a Hubble term is introduced to account for the expansion of the Universe.
-Second, the density matrix for neutrinos is considered to depend solely on the absolute value of the momentum.
-Finally, the coherence matrices between left-handed and right-handed neutrinos, namely $\varrho_{LR,RL}$ and $\bar{\varrho}_{LR,RL}$ for Dirac neutrinos, and $\phi$ for Majorana neutrinos, are reduced to zero, leading to corresponding collision terms also being zero.
+Second, the neutrino density matrix is assumed to depend solely on the absolute value of the momentum.
+Finally, the coherence matrices between left-handed and right-handed neutrinos, denoted as $\varrho_{LR,RL}$ and $\bar{\varrho}_{LR,RL}$ for Dirac neutrinos, and $\phi$ for Majorana neutrinos, vanish, leading to zero collision terms for these components.
 Consequently, the non-zero density matrices are $\varrho_{LL,RR}$ and $\bar{\varrho}_{LL,RR}$ for Dirac neutrinos, and $\varrho$ and $\bar{\varrho}$ for Majorana neutrinos.
-For active neutrinos, the QKEs for $\varrho_{LL}$ and $\bar{\varrho}_{RR}$ in the case of Dirac neutrinos, are identical with those for $\varrho$ and $\bar{\varrho}$ for Majorana neutrinos.
-Therefore, in the subsequent analysis, we will not distinguish between Dirac and Majorana neutrinos and will use $\varrho$ and $\bar{\varrho}$ to represent the density matrices for neutrinos and antineutrinos, respectively.
-In the flavor basis, the QKEs can be written as:
+For active neutrinos, the QKEs for $\varrho_{LL}$ and $\bar{\varrho}_{RR}$ in Dirac neutrinos are identical to those for $\varrho$ and $\bar{\varrho}$ in Majorana neutrinos.
+Therefore, in the subsequent analysis, we do not distinguish between Dirac and Majorana neutrinos and use $\varrho$ and $\bar{\varrho}$ to represent the density matrices for neutrinos and antineutrinos, respectively.
+In the flavor basis, the QKEs are written as:
 $i\left[\frac{\partial}{\partial t}-Hp\frac{\partial}{\partial p}\right]\varrho$ $=\Big{[}H_{\nu},\varrho\Big{]}+i\mathcal{I}\,,$ (51)
@@
 $\bar{H}_{\nu}$ $=U\frac{\mathbb{M}^{2}}{2p}U^{\dagger}-\sqrt{2}G_{F}\left(\mathbb{N}_{l}+\mathbb{N}_{\nu}\right)-2\sqrt{2}G_{F}p\left(\frac{\mathbb{E}_{l}+\mathbb{P}_{l}}{m_{W}^{2}}+\frac{4}{3}\frac{\mathbb{E}_{\nu}}{m_{Z}^{2}}\right)\,.$ (54)
-Here, $\mathbb{M}$ represents the neutrino mass matrix in the mass basis, and $U$ is the PMNS transformation matrix that relates the flavor eigenstates to the mass eigenstates.
-Using the standard parametrization of the PMNS matrix Giunti:2007ry ; Gariazzo:2019gyi , we have:
+Here, $\mathbb{M}$ represents the neutrino mass matrix in the mass basis, and $U$ is the PMNS matrix, which relates flavor eigenstates to mass eigenstates.
+Using the standard parameterization of the PMNS matrix Giunti:2007ry ; Gariazzo:2019gyi , we have:
 $\mathbb{M}=\begin{pmatrix}m_{1}^{2}&0&0\\ 0&m_{2}^{2}&0\\ 0&0&m_{3}^{2}\end{pmatrix}\,,\quad U=\begin{pmatrix}c_{12}c_{13}&s_{12}c_{13}&s_{13}\\ -s_{12}c_{23}-c_{12}s_{23}s_{13}&c_{12}c_{23}-s_{12}s_{23}s_{13}&s_{23}c_{13}\\ s_{12}s_{23}-c_{12}c_{23}s_{13}&-c_{12}s_{23}-s_{12}c_{23}s_{13}&c_{23}c_{13}\end{pmatrix}\,,$ (55)
-where $c_{ij}=\cos{\theta_{ij}}$, $s_{ij}=\sin{\theta_{ij}}$, and $\theta_{ij}$ are the mixing angles. We neglect the potential CP-violating phase in this work.
-The remaining terms in the Hamiltonian account for the matter potentials induced by coherent scatterings between neutrinos and other plasma particles.
-Here, $m_{W}$ and $m_{Z}$ denote the masses of the $W$ and $Z$ bosons, respectively, and $G_{F}$ is the Fermi constant.
-The second term in the Hamiltonian reflects the net number densities for charged leptons ($\mathbb{N}_{l}\equiv\mathrm{diag}(n_{e}-n_{\bar{e}},n_{\mu}-n_{\bar{\mu}},0)$) and neutrinos ($\mathbb{N}_{\nu}\equiv n_{\nu}-n_{\bar{\nu}}$).
+where $c_{ij}=\cos{\theta_{ij}}$, $s_{ij}=\sin{\theta_{ij}}$, and $\theta_{ij}$ are the mixing angles. We neglect the CP-violating phase in this work.
+The remaining terms in the Hamiltonian represent matter potentials induced by coherent scattering between neutrinos and other plasma particles.
+Here, $m_{W}$ and $m_{Z}$ are the masses of the $W$ and $Z$ bosons, respectively, and $G_{F}$ is the Fermi constant.
+The second term in the Hamiltonian reflects the net number densities of charged leptons ($\mathbb{N}_{l}\equiv\mathrm{diag}(n_{e}-n_{\bar{e}},n_{\mu}-n_{\bar{\mu}},0)$) and neutrinos ($\mathbb{N}_{\nu}\equiv n_{\nu}-n_{\bar{\nu}}$).
 The third term accounts for the energy densities of charged leptons and neutrinos ($\mathbb{E}_{l}\equiv\mathrm{diag}(\rho_{e}+\rho_{\bar{e}},\rho_{\mu}+\rho_{\bar{\mu}},0)$, $\mathbb{P}_{l}\equiv\mathrm{diag}(P_{e}+P_{\bar{e}},P_{\mu}+P_{\bar{\mu}},0)$, $\mathbb{E}_{\nu}\equiv\rho_{\nu}+\rho_{\bar{\nu}}$).
-In practice, the net number densities for charged leptons are often negligible due to the small baryon asymmetry.
-However, the net number densities for neutrinos can lead to significant effects, especially if primordial neutrino asymmetries are present.
-This can result in pronounced neutrino flavor oscillations, including the well-known "synchronized oscillation" when this term dominates the Hamiltonian Bell:1998ds ; Pastor:2001iu ; Abazajian:2002qx ; Wong:2002fa ; Johns:2016enc ; Froustey:2021azz ; Froustey:2024mgf .
+In practice, the net number densities of charged leptons are often negligible due to the small baryon asymmetry.
+However, the net number densities of neutrinos can significantly affect neutrino flavor oscillations, especially in the presence of primordial neutrino asymmetries.
+This can result in enhanced neutrino flavor transitions, including the well-known "synchronized oscillations" when this term dominates the Hamiltonian Bell:1998ds ; Pastor:2001iu ; Abazajian:2002qx ; Wong:2002fa ; Johns:2016enc ; Froustey:2021azz ; Froustey:2024mgf .
 Finally, the collision terms $\mathcal{I}$ and $\bar{\mathcal{I}}$ account for contributions from neutrino-electron/positron scattering ($\nu e\leftrightarrow\nu e$), annihilation ($\nu\bar{\nu}\leftrightarrow e^{+}e^{-}$), and neutrino-neutrino self-interactions ($\nu\nu\leftrightarrow\nu\nu$, $\nu\bar{\nu}\leftrightarrow\nu\bar{\nu}$).
-These terms can be expressed as
+These terms are decomposed as
 $\mathcal{I}\equiv\mathcal{I}_{\rm sc}+\mathcal{I}_{\rm ann}+\mathcal{I}_{\nu\nu}$ (56)
-with a similar decomposition for the collision term of antineutrinos $\bar{\mathcal{I}}$.
-The collision terms are generally functions of the neutrino density matrices, and their explicit forms are detailed in Appendix A.
-Contributions from other interaction processes, such as $\mu^{\pm}$ scattering and annihilation, are neglected due to their minimal impact at the temperature of neutrino decoupling.
-In practice, comoving coordinates and corresponding comoving variables are employed to eliminate the Hubble expansion term in the QKEs.
-The comoving temperature is defined as $T_{\rm CM}\propto a^{-1}$ Grohs:2015tfy , with $T_{\rm CM}=T_{\nu}=T_{\gamma}$ at the initial time before neutrino decoupling when all species are tightly coupled.
-Using the comoving temperature, the reduced scale factor, the comoving momentum, and the dimensionless photon temperature are expressed as Mangano:2005cc ; Froustey:2020mcq ; Bennett:2020zkv
+with a similar decomposition for the antineutrino collision term $\bar{\mathcal{I}}$.
+The collision terms depend on the neutrino density matrices, and their explicit forms are detailed in Appendix A.
+Contributions from other processes, such as $\mu^{\pm}$ scattering and annihilation, are neglected due to their minimal impact at the temperatures relevant to neutrino decoupling.
+In practice, comoving coordinates and their corresponding variables are employed to eliminate the Hubble expansion term in the QKEs.
+The comoving temperature is defined as $T_{\rm CM}\propto a^{-1}$ Grohs:2015tfy , with $T_{\rm CM}=T_{\nu}=T_{\gamma}$ at the initial time before neutrino decoupling, when all species are tightly coupled.
+Using the comoving temperature, the reduced scale factor, comoving momentum, and dimensionless photon temperature are expressed as Mangano:2005cc ; Froustey:2020mcq ; Bennett:2020zkv :
 $x=m_{e}/T_{\rm CM}\,,\qquad y=p/T_{\rm CM}\,,\qquad z=T_{\gamma}/T_{\rm CM}\,.$ (57)
-By definition, the dimensionless photon temperature $z$ equals 1 at high temperatures and increases beyond 1 later due to the annihilation of electrons and positrons into photons.
-For the convenience of the QKEs, we also define the dimensionless energy density and pressure as $\tilde{\rho}\equiv(x/m_{e})^{4}\rho$ and $\tilde{P}\equiv(x/m_{e})^{4}P$, respectively.
-The comoving QKEs for neutrinos and antineutrinos then take the form as:
+By definition, the dimensionless photon temperature $z$ equals 1 at high temperatures and increases beyond 1 due to the annihilation of electrons and positrons into photons.
+For convenience in the QKEs, we also define the dimensionless energy density and pressure as $\tilde{\rho}\equiv(x/m_{e})^{4}\rho$ and $\tilde{P}\equiv(x/m_{e})^{4}P$, respectively.
+The comoving QKEs for neutrinos and antineutrinos are then written as:
 $\frac{{\rm d}\varrho(x,y)}{{\rm d}x}$ $=$ $\sqrt{\frac{3m^{2}_{\rm Pl}}{8\pi\tilde{\rho}_{\rm tot}}}\left\{-i\frac{x^{2}}{m_{e}^{3}}\left[\tilde{H}_{\nu},\varrho\right]+\frac{m_{e}^{3}}{x^{4}}\tilde{\mathcal{I}}(\varrho,\bar{\varrho})\right\}\,,$
-$\frac{{\rm d}\bar{\varrho}(x,y)}{{\rm d}x}$ $=$ $\sqrt{\frac{3m^{2}_{\rm Pl}}{8\pi\rho}}\left\{-i\frac{x^{2}}{m_{e}^{3}}\left[\tilde{H}_{\bar{\nu}},\varrho\right]+\frac{m_{e}^{3}}{x^{4}}\tilde{\bar{\mathcal{I}}}(\bar{\varrho},\varrho)\right\}\,,$
-with the comoving Hamiltonian terms
+$\frac{{\rm d}\bar{\varrho}(x,y)}{{\rm d}x}$ $=$ $\sqrt{\frac{3m^{2}_{\rm Pl}}{8\pi\tilde{\rho}_{\rm tot}}}\left\{-i\frac{x^{2}}{m_{e}^{3}}\left[\tilde{H}_{\bar{\nu}},\bar{\varrho}\right]+\frac{m_{e}^{3}}{x^{4}}\tilde{\bar{\mathcal{I}}}(\bar{\varrho},\varrho)\right\}\,,$
+with the comoving Hamiltonian terms:
 $\tilde{H}_{\nu}$ $=U\frac{\mathbb{M}^{2}}{2y}U^{\dagger}+\frac{\sqrt{2}G_{\rm F}m_{e}^{4}}{x^{4}}\,\mathbb{N}_{\nu}-\frac{2\sqrt{2}G_{\rm F}ym_{e}^{6}}{x^{6}}\left(\frac{\mathbb{E}_{\ell}+\mathbb{P}_{\ell}}{m_{W}^{2}}+\frac{4}{3}\,\frac{\mathbb{E}_{\nu}}{m_{Z}^{2}}\right)\,,$ (58)
@@
 Additionally, $\tilde{\mathcal{I}}$ and $\tilde{\bar{\mathcal{I}}}$ represent the comoving collision terms for neutrinos and antineutrinos, respectively.
-The computation of these collision terms is the most complex and time-consuming aspect of solving the QKEs.
-Fortunately, the nine-dimensional collision integrals can be simplified to two-dimensional integrals by exploiting the homogeneous and isotropic conditions and the specific forms of the scattering amplitudes Hannestad:1995rs ; Dolgov:1997mb ; Grohs:2015tfy ; Blaschke:2016xxt .
+Calculating these collision terms is the most complex and time-consuming aspect of solving the QKEs.
+Fortunately, the nine-dimensional collision integrals can be reduced to two-dimensional integrals by exploiting the homogeneous and isotropic conditions and the specific forms of the scattering amplitudes Hannestad:1995rs ; Dolgov:1997mb ; Grohs:2015tfy ; Blaschke:2016xxt .
 The explicit forms of the reduced collision terms are provided in Appendix A.
 Furthermore, as noted in the discussion following the QKEs (Eq.
-(53)), the chemical potential of electrons and the $\mathbb{N}_{l}$ term in the Hamiltonian are negligible due to ue to the very low baryon-to-photon ratio $\eta$ and the efficient electromagnetic interactions.
+(53)), the chemical potential of electrons and the $\mathbb{N}_{l}$ term in the Hamiltonian are negligible due to the very low baryon-to-photon ratio $\eta$ and the efficient electromagnetic interactions.
 Consequently, the Hamiltonian is predominantly influenced by the net number densities of neutrinos when primordial lepton asymmetries are sufficiently large and vary among different flavors.
 This results in synchronized flavor oscillations of neutrinos and antineutrinos Bell:1998ds ; Pastor:2001iu ; Abazajian:2002qx ; Wong:2002fa .
-Consequently, neutrino asymmetries tend to equilibrate among different flavors prior to the epoch of BBN Froustey:2021azz ; Froustey:2024mgf .
-Since this work focuses on the implications of neutrino asymmetries for BBN, CMB, and LSS, we will concentrate on flavor-equilibrated primordial asymmetries, i.e., the initial primordial asymmetries are set to $\xi_{\nu_{e}}=\xi_{\nu_{\mu}}=\xi_{\nu_{\tau}}=\xi_{n}$, where $\xi_{\nu_{\alpha}}\equiv\mu_{\nu_{\alpha}}/T_{\rm CM}$ is the degeneracy parameter for neutrinos and antineutrinos of flavor $\alpha$.
+As a result, neutrino asymmetries tend to equilibrate among different flavors before the epoch of BBN Froustey:2021azz ; Froustey:2024mgf .
+Since this work focuses on the implications of neutrino asymmetries for BBN, CMB, and LSS, we concentrate on primordial neutrino asymmetries with complete flavor equilibration before neutrino decoupling.
+Specifically, the initial primordial asymmetries are set to $\xi_{\nu_{e}}=\xi_{\nu_{\mu}}=\xi_{\nu_{\tau}}=\xi_{\nu}$, where $\xi_{\nu_{\alpha}}\equiv\mu_{\nu_{\alpha}}/T_{\rm CM}$ represents the degeneracy parameter for neutrinos and antineutrinos of flavor $\alpha$.
 By adopting this initial condition, we bypass the intricate mapping between arbitrary primordial neutrino asymmetries and those after flavor equilibration, leaving the investigation of this mapping for future work.
-Recent study on such mapping and the implications of distinct neutrino asymmetries among flavors for BBN can be found in  Froustey:2024mgf .
-On the other hand, it is well-known that the annihilation of electrons and positrons shortly after neutrino decoupling partially heats neutrinos, leading to a slightly increased $N_{\rm eff}$.
-To accurately determine $N_{\rm eff}$, finite-temperature QED (FTQED) corrections must be considered Fornengo:1997wa ; Mangano:2001iu ; Bennett:2019ewm .
-In this work, we address three aspects of FTQED corrections: (1) Corrections to the total energy pressure densities, which primarily affect the Hubble rate; (2) Corrections to the electron mass, which primarily influence the collision terms in the QKEs; (3) Corrections to the continuity equation of the Universe, typically expressed as a differential equation for the dimensionless photon temperature $z$,
+Recent studies on this mapping and the implications of distinct neutrino asymmetries among flavors for BBN can be found in Froustey:2024mgf .
+On the other hand, it is well known that the annihilation of electrons and positrons shortly after neutrino decoupling partially heats neutrinos, leading to a slightly increased $N_{\rm eff}$.
+To accurately determine $N_{\rm eff}$, finite-temperature QED (FTQED) corrections must be taken into account Fornengo:1997wa ; Mangano:2001iu ; Bennett:2019ewm .
+In this work, we address three aspects of FTQED corrections: (1) corrections to the total energy and pressure densities, which primarily affect the Hubble rate; (2) corrections to the electron mass, which mainly influence the collision terms in the QKEs; and (3) corrections to the continuity equation of the Universe, typically expressed as a differential equation for the dimensionless photon temperature $z$,
 $\frac{\mathrm{d}z}{\mathrm{d}x}=\frac{rJ_{2}(r)+G_{1}(r)-\frac{1}{4\pi^{2}z^{3}}\int_{0}^{\infty}\mathrm{d}y\,y^{3}\sum_{\alpha}\left(\frac{\mathrm{d}\varrho_{\alpha\alpha}}{\mathrm{d}x}+\frac{\mathrm{d}\bar{\varrho}_{\alpha\alpha}}{\mathrm{d}x}\right)}{\left[r^{2}J_{2}(r)+J_{4}(r)\right]+G_{2}(r)+\frac{2\pi^{2}}{15}}\,,$ (59)
-where $r=x/z$, $\frac{\mathrm{d}\varrho_{\alpha\alpha}}{\mathrm{d}x}$, and $\frac{\mathrm{d}\bar{\varrho}_{\alpha\alpha}}{\mathrm{d}x}$ are the diagonal components of the QKEs for neutrinos and antineutrinos.
-The explicit forms of the remaining $J_{i}(r)$ and $G_{i}(r)$ functions appearing in the equation, as well as other FTQED corrections, can be found in Appendix B.
-In practice, this equation should be solved in conjunction with the QKEs, and we consider corrections up to order $\mathcal{O}(e^{3})$ while neglecting logarithmic terms in the FTQED corrections as their effects are minimal Bennett:2019ewm ; Bennett:2020zkv ; further details are provided in Appendix B.
+where $r=x/z$, and $\frac{\mathrm{d}\varrho_{\alpha\alpha}}{\mathrm{d}x}$ and $\frac{\mathrm{d}\bar{\varrho}_{\alpha\alpha}}{\mathrm{d}x}$ are the diagonal components of the QKEs for neutrinos and antineutrinos.
+The explicit forms of the remaining $J_{i}(r)$ and $G_{i}(r)$ functions in the equation, as well as other FTQED corrections, can be found in Appendix B.
+In practice, this equation should be solved in conjunction with the QKEs, and we consider corrections up to order $\mathcal{O}(e^{3})$ while neglecting logarithmic terms in the FTQED corrections, as their effects are minimal Bennett:2019ewm ; Bennett:2020zkv .
+Further details are provided in Appendix B.
 Other FTQED corrections, such as those to the interaction vertices, are omitted in this work.
@@
 The results for neutrino decoupling and BBN are independent of the minimal neutrino mass, provided $m_{1}\ll 0.1\,\text{MeV}$.
-On the other hand, the resulting CMB anisotropies and LSS of the Universe are sensitive to the choice of $m_{1}$, which will be discussed in more detail in Sections 5 and 6.
+However, the resulting CMB anisotropies and LSS of the Universe are sensitive to the choice of $m_{1}$, which will be discussed in more detail in Sections 5 and 6.
 3.2 Numerical Methods
-To solve the QKEs for neutrinos and antineutrinos numerically, we have extended the public code FortEPiaNO Gariazzo:2019gyi ; Bennett:2020zkv , initially designed for solving the QKEs in the Standard Model case, to accommodate primordial neutrino asymmetries.
-This subsection outlines the key numerical methods employed in the code.
+To numerically solve the QKEs for neutrinos and antineutrinos, we have extended the public code FortEPiaNO Gariazzo:2019gyi ; Bennett:2020zkv , originally developed for solving QKEs in the Standard Model, to account for primordial neutrino asymmetries.
+This subsection outlines the primary numerical methods employed in the code.
 For the momentum grid, the comoving momentum $y$ is discretized on a linear grid with $y_{\rm min}=0.01$, $y_{\rm max}=20$, and a total of $N_{y}=40$ points.
-The integrals are computed using the Newton-Cotes (NC) formula of order 1.
-The initial conditions are specified with an initial reduced scale factor $x=0.005$, corresponding to a comoving temperature $T_{\rm CM,in}\simeq 102.2\text{ MeV}$.
-The initial common temperature for all species is computed by conserving total entropy from an even higher temperature, yielding $z_{\rm in}=2.81\times 10^{-6}$.
-The initial density matrices are given by
-$\varrho(x_{\rm in},y)={\rm diag}\{f_{\nu}^{({\rm in})}(y),f_{\nu}^{({\rm in})}(y),f_{\nu}^{({\rm in})}(y)\},\quad\bar{\varrho}(x_{\rm in},y)={\rm diag}\{f_{\bar{\nu}}^{({\rm in})}(y),f_{\bar{\nu}}^{({\rm in})}(y),f_{\bar{\nu}}^{({\rm in})}(y)\}.$ (60)
+The integrals are evaluated using the first-order Newton-Cotes (NC) formula.
+The initial conditions are set at an initial reduced scale factor of $x=0.005$, corresponding to a comoving temperature of $T_{\rm CM,in}\simeq 102.2~{}\text{MeV}$.
+The initial common temperature for all species is calculated by conserving total entropy from a higher temperature, resulting in $z_{\rm in}=2.81\times 10^{-6}$.
+The initial density matrices are specified as:
+$\varrho(x_{\rm in},y)={\rm diag}\{f_{\nu}^{({\rm in})}(y),f_{\nu}^{({\rm in})}(y),f_{\nu}^{({\rm in})}(y)\},\quad\bar{\varrho}(x_{\rm in},y)={\rm diag}\{f_{\bar{\nu}}^{({\rm in})}(y),f_{\bar{\nu}}^{({\rm in})}(y),f_{\bar{\nu}}^{({\rm in})}(y)\},$ (60)
 where
-$f_{\nu}^{({\rm in})}(y)\equiv\frac{1}{e^{(y-\xi_{\nu})/z_{\rm in}}+1}\quad\text{and}\quad f_{\bar{\nu}}^{({\rm in})}(y)\equiv\frac{1}{e^{(y+\xi_{\nu})/z_{\rm in}}+1}\,.$ (61)
-Finally, the QKEs are solved together with the energy conservation equation for $z$ using the DLSODA method from the ODEPACK hindmarsh1982odepack Fortran library.
-The absolute and relative error tolerances are set to $10^{-7}$, ensuring that the numerical errors of the results remain below $10^{-5}$.
+$f_{\nu}^{({\rm in})}(y)\equiv\frac{1}{e^{(y-\xi_{\nu})/z_{\rm in}}+1},\quad f_{\bar{\nu}}^{({\rm in})}(y)\equiv\frac{1}{e^{(y+\xi_{\nu})/z_{\rm in}}+1}.$ (61)
+The QKEs are solved in conjunction with the energy conservation equation for $z$ using the DLSODA solver from the ODEPACK hindmarsh1982odepack Fortran library.
+Absolute and relative error tolerances are set to $10^{-7}$, ensuring numerical errors remain below $10^{-5}$.
 3.3 Numerical Results of Neutrino Decoupling
-We present the numerical results of neutrino decoupling in this subsection.
-First, we consider the final energy density of neutrinos, which is typically parameterized by the effective number of relativistic species $N_{\rm eff}$, defined as
-$N_{\rm eff}\equiv\frac{87}{8}\left(\frac{11}{4}\right)^{4/3}\frac{\rho_{\nu}}{\rho_{\gamma}}\,,\quad\rho_{\nu}=\sum_{\alpha}\int\frac{d^{3}p}{(2\pi)^{3}}\,p\,(\varrho_{\nu_{\alpha}}+\bar{\varrho}_{\nu_{\alpha}})$ (62)
-where $\rho_{\nu}$ represents the total neutrino energy density obtained by solving the QKEs and $\rho_{\gamma}$ is the photon energy density.
+In this subsection, we present the numerical results for neutrino decoupling.
+First, we focus on the final energy density of neutrinos, typically parameterized by the effective number of relativistic species, $N_{\rm eff}$, defined as
+$N_{\rm eff}\equiv\frac{7}{8}\left(\frac{11}{4}\right)^{4/3}\frac{\rho_{\nu}}{\rho_{\gamma}}\,,\quad\rho_{\nu}=\sum_{\alpha}\int\frac{d^{3}p}{(2\pi)^{3}}\,p\,(\varrho_{\nu_{\alpha}}+\bar{\varrho}_{\nu_{\alpha}}),$ (62)
+where $\rho_{\nu}$ is the total neutrino energy density, obtained by solving the QKEs, and $\rho_{\gamma}$ is the photon energy density.
 For the SM in the instantaneous decoupling limit, $N_{\rm eff}$ equals the number of neutrino flavors.
-However, since the temperature of neutrino decoupling is close to that of electron-positron annihilation, neutrinos are also heated through weak interactions with electrons and positrons, resulting in a slightly larger $N_{\rm eff}$.
-The most accurate calculation for the SM $N_{\rm eff}$ to date, which includes neutrino oscillations, matter effects, FTQED corrections up to $\mathcal{O}(3)$ order, and neutrino-electron and neutrino-neutrino collision terms, yields $N_{\rm eff}^{\rm SM}=3.0440\pm 0.0002$ Froustey:2020mcq ; Bennett:2020zkv .
-Figure 2: Left panel: Results of the neutrino non-instantaneous decoupling correction $\delta N_{\rm eff}^{\rm dec}=N_{\rm eff}-N_{\rm eff}^{\rm inst}$ as a function of the neutrino degeneracy parameter $\xi_{\nu}$.
-The solid and dashed curves represent the results of the numerical solution and the fitting function Eq.
+However, since the temperature of neutrino decoupling is near that of electron-positron annihilation, neutrinos are also heated through weak interactions with electrons and positrons, leading to a slightly increased $N_{\rm eff}$.
+The most precise calculation for $N_{\rm eff}$ in the SM, including neutrino oscillations, matter effects, FTQED corrections up to $\mathcal{O}(3)$, and neutrino-electron and neutrino-neutrino collision terms, gives $N_{\rm eff}^{\rm SM}=3.0440\pm 0.0002$ Froustey:2020mcq ; Bennett:2020zkv .
+Figure 2: Left panel: Neutrino non-instantaneous decoupling correction $\delta N_{\rm eff}^{\rm dec}=N_{\rm eff}-N_{\rm eff}^{\rm inst}$ as a function of the neutrino degeneracy parameter $\xi_{\nu}$.
+Solid and dashed curves represent the numerical solution and the fitting function of Eq.
 (63), respectively.
-Right panel: Relative differences in the number densities of electron neutrinos and antineutrinos compared to the SM predictions in the instantaneous decoupling limit, $n_{\nu_{e},\bar{\nu}_{e}}\simeq 56\,\text{cm}^{-3}$.
+Right panel: Relative differences in the number densities of electron neutrinos and antineutrinos compared to SM predictions in the instantaneous decoupling limit, $n_{\nu_{e},\bar{\nu}_{e}}\simeq 56\,\text{cm}^{-3}$.
 In the presence of non-zero primordial neutrino asymmetries, $N_{\rm eff}$ receives additional contributions from the neutrino chemical potentials.
-In the instantaneous decoupling limit, the comoving distribution of neutrinos and antineutrinos of flavor $\alpha$ can be approximated as $f_{\nu_{\alpha},\bar{\nu}_{\alpha}}=\frac{1}{e^{y\,\pm\,\xi_{\nu_{\alpha}}}+1}$.
+In the instantaneous decoupling limit, the comoving distribution of neutrinos and antineutrinos of flavor $\alpha$ is approximated as $f_{\nu_{\alpha},\bar{\nu}_{\alpha}}=\frac{1}{e^{y\,\pm\,\xi_{\nu_{\alpha}}}+1}$.
 Therefore, the total $N_{\rm eff}$ can be estimated by
-$N_{\rm eff}^{\rm inst}\simeq 3+\sum\limits_{\alpha}\left(\frac{30}{7\pi^{2}}\xi_{\nu_{\alpha}}^{2}+\frac{15}{7\pi^{4}}\xi_{\nu_{\alpha}}^{4}\right)=3+3\left(\frac{30}{7\pi^{2}}\xi_{\nu}^{2}+\frac{15}{7\pi^{4}}\xi_{\nu}^{4}\right),$ (63)
-where we have used the condition $\xi_{\nu_{e}}=\xi_{\nu_{\mu}}=\xi_{\nu_{\tau}}=\xi_{\nu}$ in the second equality.
+$N_{\rm eff}^{\rm inst}\simeq 3+\sum_{\alpha}\left(\frac{30}{7\pi^{2}}\xi_{\nu_{\alpha}}^{2}+\frac{15}{7\pi^{4}}\xi_{\nu_{\alpha}}^{4}\right)=3+3\left(\frac{30}{7\pi^{2}}\xi_{\nu}^{2}+\frac{15}{7\pi^{4}}\xi_{\nu}^{4}\right),$ (63)
 […]
@@
 The Fermi-Dirac distributions with corresponding $\xi_{\nu}$ are also indicated by the gray dotted curves.
-Next we present the resulting density matrices for neutrinos and antineutrinos after decoupling.
-The density matrices as a function of the comoving momentum $y\equiv p/T_{\rm CM}$ in the flavor basis and the mass basis for several choices of $\xi_{\nu}$ are shown in the left and right panels of Fig.
-3 respectively, normalized to the Fermi-Dirac distribution with $\xi_{\nu}=0$.
-The density matrices of neutrinos and antineutrinos with $\xi_{\nu}=0$ coincide with each other and are represented by the gray curves.
-Our results also reproduce the previous density matrices obtained in the literature.
-For each choice of In the case of non-zero $\xi_{\nu}$, the density matrices of neutrinos and antineutrinos no longer coincide and are represented by red and green curves, respectively.
-For each case we use the solid, dashed and dash-dotted curve to indicate the three flavor eigenstates ($\nu_{e}$, $\nu_{\mu}$ and $\nu_{\tau}$) and the three mass eigenstates ($\nu_{1}$, $\nu_{2}$ and $\nu_{3}$), respectively.
-For each choice of the $\xi_{\nu}$, the corresponding thermal equilibrium distributions for neutrinos and antineutrinos $f_{\rm FD}^{\nu,\bar{\nu}}=1/(e^{y\,\pm\,\xi_{\nu}}+1)$ are also plotted by the dotted curves.
-Clearly, the density distributions of neutrinos and antineutrinos for the given $\xi_{\nu}$ coincide with the corresponding Fermi-Dirac distributions when $y$ is small, while the spectra are distorted at large $y$.
-This is due to the fact that the weak interaction rate between neutrinos and electrons/positrons is proportional to the neutrino momentum, so that the neutrino with smaller momentum decouples earlier, preserving the thermal Fermi-Dirac distribution, while the neutrino with larger momentum decouples later, resulting in a larger spectral distortion with respect to the Fermi-Dirac distribution.
-Regarding the density distributions for different flavors, we see that electron neutrinos possess larger spectral distortions due to the additional charge-current weak interaction, while the spectral distortions of muon neutrinos almost coincide with those of tauon neutrinos, since they have the same weak interactions.
-Instead, the spectral distortions of the three neutrino mass eigenstates are uniformly separated, as shown in the right panel.
-Figure 4: The spectral distortions of neutrinos and antineutrinos in the flavor basis (top left) and mass basis (lower left) with respect to the Fermi-Dirac distribution with corresponding $\xi_{\nu}$.
-The meaning of the curves is the same as in the previous plot Fig.
-3, while the spectral distortions of tauon neutrino in the left panel are not shown for brevity.
-The zoomed version of the spectral distortions in the flavor and mass basis are shown in the right panel.
-Considering that the density distributions of neutrinos and antineutrinos coincide with the corresponding thermal Fermi-Dirac distributions for small momentum, it is natural to define the spectral distortions as $\delta\varrho_{\nu,\bar{\nu}}\equiv\delta\varrho_{\nu,\bar{\nu}}/f_{FD}(\pm\xi)-1$, where the plus sign corresponds to neutrinos and the minus sign to antineutrinos.
-The resulting spectral distortions as a function of the momentum $y$ in the flavor and mass basis are presented in the upper left and upper right panels of Fig.
+We now present the resulting density matrices for neutrinos and antineutrinos after decoupling.
+The density matrices as a function of the comoving momentum $y\equiv p/T_{\rm CM}$ in both the flavor and mass bases for several values of $\xi_{\nu}$ are shown in the left and right panels of Fig.
+3, respectively, and are normalized to the Fermi-Dirac distribution for $\xi_{\nu}=0$.
+The density matrices for neutrinos and antineutrinos with $\xi_{\nu}=0$ coincide and are represented by the gray curves, reproducing results from previous studies.
+For non-zero $\xi_{\nu}$, the density matrices of neutrinos and antineutrinos no longer coincide and are represented by red and green curves, respectively.
+In each case, the solid, dashed, and dash-dotted curves represent the three flavor eigenstates ($\nu_{e}$, $\nu_{\mu}$, and $\nu_{\tau}$) and the three mass eigenstates ($\nu_{1}$, $\nu_{2}$, and $\nu_{3}$), respectively.
+The corresponding thermal equilibrium distributions for neutrinos and antineutrinos, $f_{\rm FD}^{\nu,\bar{\nu}}=1/(e^{y\pm\xi_{\nu}}+1)$, are also shown by the dotted curves.
+The density distributions of neutrinos and antineutrinos for a given $\xi_{\nu}$ coincide with the Fermi-Dirac distributions at small $y$, but deviate at larger $y$.
+This occurs because the weak interaction rate between neutrinos and electrons/positrons is proportional to the neutrino momentum, causing neutrinos with smaller momenta to decouple earlier and maintain the thermal Fermi-Dirac distribution, while neutrinos with larger momenta decouple later, leading to greater spectral distortions relative to the Fermi-Dirac distribution.
+Regarding different flavors, electron neutrinos exhibit larger spectral distortions due to additional charged-current weak interactions, while the spectral distortions of muon and tau neutrinos nearly coincide, as they experience the same weak interactions.
+In contrast, the spectral distortions of the three neutrino mass eigenstates are uniformly separated, as shown in the right panel.
+Figure 4: Spectral distortions of neutrinos and antineutrinos in the flavor basis (left panel) and mass basis (right panel), relative to the Fermi-Dirac distribution with the corresponding $\xi_{\nu}$.
+The curves follow the same convention as in Fig.
+3, but the spectral distortions of tau neutrinos in the left panel are omitted for brevity.
+Since the density distributions of neutrinos and antineutrinos coincide with the thermal Fermi-Dirac distributions at small momenta, it is natural to define the spectral distortions as
+$\delta\varrho_{\nu,\bar{\nu}}\equiv\frac{\delta\varrho_{\nu,\bar{\nu}}}{f_{FD}(\pm\xi)}-1\,,$ (65)
+where the plus sign corresponds to neutrinos and the minus sign to antineutrinos.
+The resulting spectral distortions as a function of momentum $y$ in the flavor and mass bases are shown in the left and right panels of Fig.
 4, respectively.
-The meaning of the curves is the same as in the previous plot Fig.
-3, while the spectral distortions of the tauon neutrinos are not shown in the left panels, since they almost coincide with the spectral distortions of the muon neutrinos.
-We also plot the zoomed version of the spectral distortions in the lower panels.
-Interestingly, the spectral distortions for neutrinos and antineutrinos with non-zero $\xi_{\nu}$ differ from each other and lie on different sides of the spectral distortions with $\xi_{\nu}=0$.
-The antineutrinos with non-zero $\xi_{\nu}$ possess a relatively larger spectral distortion than that with $\xi_{\nu}=0$, while the neutrinos possess a relatively smaller spectral distortion.
-In addition, the larger the $\xi_{\nu}$, the larger the discrepancy between the spectral distortions of neutrinos and antineutrinos, clearly seen from the zoomed plots.
-Such properties are both valid in the flavor basis and in the mass basis.
-Finally, the above conclusions about neutrino density matrices also hold for negative $\xi_{\nu}$, except that the density matrices for neutrinos and antineutrinos with $-\xi_{\nu}$ are exchanged comparing with those for $\xi_{\nu}$.
-In other words, the density distributions and spectral distortions for antineutrinos in Fig.
-3 and Fig.
-4 are also the density distributions and spectral distortions for neutrinos with the opposite $\xi_{\nu}$.
-As a brief conclusion of this section, we have presented the most accurate results of neutrino decoupling in the presence of primordial neutrino asymmetries and discussed the resulting $N_{\rm eff}$ and the spectral distortions of neutrinos and antineutrinos.
-In the next sections, we will study the effects of primordial neutrino asymmetries on the processes of BBN, CMB and structure formation, while emphasizing the corrections induced by the spectral distortions of neutrinos and antineutrinos.
+The curves follow the same convention as in Fig.
+3, with tau neutrino spectral distortions omitted in the left panels, as they nearly coincide with those of muon neutrinos.
+Notably, the spectral distortions for neutrinos and antineutrinos with non-zero $\xi_{\nu}$ differ and lie on opposite sides of the spectral distortions for $\xi_{\nu}=0$.
+Antineutrinos with non-zero $\xi_{\nu}$ exhibit larger spectral distortions compared to those for $\xi_{\nu}=0$, while neutrinos display smaller distortions.
+Moreover, the larger the $\xi_{\nu}$, the greater the discrepancy between the spectral distortions of neutrinos and antineutrinos.
+These properties hold in both the flavor and mass bases and arise from the differing Pauli-blocking effects for neutrinos and antineutrinos in the presence of primordial neutrino asymmetries.
+Finally, the above conclusions regarding neutrino density matrices also apply to negative $\xi_{\nu}$, with the density matrices for neutrinos and antineutrinos swapped compared to those for positive $\xi_{\nu}$.
+In other words, the density distributions and spectral distortions of antineutrinos in Figs.
+3 and 4 also apply to neutrinos with opposite $\xi_{\nu}$.
+In summary, we have presented the most accurate results of neutrino decoupling in the presence of primordial neutrino asymmetries and discussed the corresponding $N_{\rm eff}$ and spectral distortions of neutrinos and antineutrinos.
+In the following sections, we will explore the effects of primordial neutrino asymmetries on BBN, CMB, and structure formation, focusing on the corrections induced by the spectral distortions of neutrinos and antineutrinos.
 4 Impacts of the Primordial Neutrino Asymmetry on BBN
-Soon after neutrino decoupling, there is an epoch of primordial nucleosynthesis or the so-called Big Bang Nucleosynthesis (BBN), where the light elements are produced after protons and neutrons decouple from the SM thermal bath.
-Since the decoupling of protons depends on electron neutrinos, the resulting abundances of light elements after BBN depends on the properties of electron neutrinos and antineutrinos.
-BBN is affected by primordial neutrino asymmetries in three ways as follow.
+Soon after neutrino decoupling, primordial nucleosynthesis, or Big Bang Nucleosynthesis (BBN), occurs, leading to the production of light elements as protons and neutrons decouple from the Standard Model (SM) thermal bath.
+The decoupling of protons is influenced by electron neutrinos, causing the resulting abundances of light elements after BBN to depend on the properties of both electron neutrinos and antineutrinos.
+BBN is impacted by primordial neutrino asymmetries in three ways (for reviews, see, $e.g.,$ Pitrou:2018cgg ):
   1. 1.
-    The leading effect is the change in the weak interaction rate at the onset of BBN due to the chemical potential of the electron neutrinos, since the conversion between proton and neutron is related to the electron neutrino through the weak interaction, so that the difference of the chemical potentials of proton and neutron is decided by the chemical potential of electron neutrino, $i.e.,$ $\mu_{n}-\mu_{p}=\mu_{e}-\mu_{\nu_{e}}\simeq-\mu_{\nu_{e}}$.
-Therefore, the resulting neutron-to-proton ratio and the subsequent BBN processes are modified according to the absolute value of the electron neutrino chemical potential and its sign.
+    The primary effect is the alteration of the weak interaction rate at the onset of BBN, which is influenced by the chemical potential of electron neutrinos.
+The conversion between protons and neutrons is mediated by the weak interaction, leading to a difference in the chemical potentials of protons and neutrons determined by the chemical potential of electron neutrinos, i.e., $\mu_{n}-\mu_{p}=\mu_{e}-\mu_{\nu_{e}}\simeq-\mu_{\nu_{e}}$.
+Consequently, the resulting neutron-to-proton ratio and subsequent BBN processes are modified according to the absolute value and sign of the electron neutrino chemical potential.
   2. 2.
-    In addition, the expansion rate of the universe during BBN is changed by the additional neutrino energy density contributed by the chemical potentials.
-As we have seen in the last section, the neutrino energy density after neutrino decoupling is usually represented by $N_{\rm eff}$, and we have $N_{\rm eff}\simeq 3.044+3\left(\frac{30}{7\pi^{2}}\xi_{\nu}^{2}+\frac{15}{7\pi^{4}}\xi_{\nu}^{4}\right)$.
-The corresponding correction in BBN therefore depends on the absolute asymmetries of all three neutrino flavors.
+    Additionally, the expansion rate of the universe during BBN is affected by the increased neutrino energy density contributed by the chemical potentials.
+As discussed in the previous section, the neutrino energy density following neutrino decoupling is typically represented by $N_{\rm eff}$, which can be expressed as $N_{\rm eff}\simeq 3.044+3\left(\frac{30}{7\pi^{2}}\xi_{\nu}^{2}+\frac{15}{7\pi^{4}}\xi_{\nu}^{4}\right)$.
+Thus, the corresponding corrections in BBN depend on the absolute asymmetries of all three neutrino flavors.
   3. 3.
-    Beyond the approximation of the thermal Fermi-Dirac distributions for neutrinos and antineutrinos, the spectral distortions resulting from the non-instantaneous decoupling of the neutrinos will also induce a correction of the weak interaction rate and thus affect the BBN.
-Such corrections are usually neglected in previous studies, especially in the presence of primordial neutrino asymmetries.
-In the following, we first review the general treatment of the weak interaction rate and introduce the new correction induced by the neutrino spectral distortions, then we discuss the resulting light element abundances from the BBN.
+    Beyond the approximation of thermal Fermi-Dirac distributions for neutrinos and antineutrinos, spectral distortions arising from the non-instantaneous decoupling of neutrinos will also induce corrections to the weak interaction rate, thereby affecting BBN.
+Such corrections have often been overlooked in previous studies, particularly in the context of primordial neutrino asymmetries.
+In the following sections, we will first review the general treatment of the weak interaction rate and introduce the new corrections induced by neutrino spectral distortions, after which we will discuss the resulting light element abundances from BBN.
 4.1 The Weak Correction from Neutrino Spectral Distortions
-At the onset of the BBN, the conversion between neutron ($n$) and proton ($p$) is dominated by following interaction reactions,
+At the onset of BBN, the conversion between neutron ($n$) and proton ($p$) is dominated by the following interaction reactions:
 $n+\nu_{e}$ $\leftrightarrow p+e^{-},$ (66)
@@
 $n+e^{+}$ $\leftrightarrow p+\bar{\nu}_{e}.$ (68)
-In the remainder of this section, we will omit the subscript for the electron-neutrino and the superscript for the electron/positron as long as there is no ambiguity.
-Due to the forceful coupling between photons and electrons and positrons, it is appropriate to assume that all particles expect for neutrinos are in local thermal equilibrium.
-Therefore, the neutron and proton number densities obey the Boltzmann equations
+In the remainder of this section, we will omit the subscript for the electron neutrino and the superscript for the electron/positron whenever there is no ambiguity.
+Due to the strong coupling between photons, electrons, and positrons, it is appropriate to assume that all particles except for neutrinos are in local thermal equilibrium.
+Consequently, the number densities of neutrons and protons obey the Boltzmann equations:
 $\dot{n}_{n}+3Hn_{n}$ $=-n_{n}\Gamma_{n\to p}+n_{p}\Gamma_{p\to n},$ (69)
-$\dot{n}_{p}+3Hn_{p}$ $=-n_{p}\Gamma_{p\to n}+n_{n}\Gamma_{n\to p},$ (69)
-where the reaction rates $n_{n}\Gamma_{n\to p}$ and $n_{p}\Gamma_{p\to n}$ incorporate all relevant collision terms
+$\dot{n}_{p}+3Hn_{p}$ $=-n_{p}\Gamma_{p\to n}+n_{n}\Gamma_{n\to p}.$ (70)
+Here, the reaction rates $n_{n}\Gamma_{n\to p}$ and $n_{p}\Gamma_{p\to n}$ incorporate all relevant collision terms:
 $\Gamma_{n\to p}$ $=$ $\Gamma_{n+\nu\to p+e}+\Gamma_{n\to p+e+\bar{\nu}}+\Gamma_{n+e\to p+\bar{\nu}},$ (71)
 $\Gamma_{p\to n}$ $=$ $\Gamma_{p+e\to n+\nu}+\Gamma_{p+e+\bar{\nu}\to n}+\Gamma_{p+\bar{\nu}\to n+e}.$ (72)
-And the reaction rate for a specific interaction reaction with the nucleon $a$ in the initial state and nucleon $b$ in the final state takes the form as
+The reaction rate for a specific interaction involving nucleon $a$ in the initial state and nucleon $b$ in the final state is given by:
 $n_{a}\Gamma_{a\to b}=$ $\int\mathrm{d}\pi_{a}\mathrm{d}\pi_{b}\mathrm{d}\pi_{e}\mathrm{d}\pi_{\nu}\delta^{(4)}\left({p}_{a}-p_{b}+\alpha_{e}p_{e}+\alpha_{\nu}p_{\nu}\right)\left|M\right|^{2}_{a\to b}{}$ (73)
 ×$\times f_{a}(E_{a})f_{b}(-E_{b})f_{e}(\alpha_{e}{E}_{e})\tilde{f}_{\nu}(E_{\nu}),$
-where $\mathrm{d}\pi_{i}=\frac{1}{(2\pi)^{3}}\frac{\mathrm{d}^{3}\mathbf{p}_{i}}{2E_{i}}$, $\alpha_{e}(\alpha_{\nu})=1$ if the electron (neutrino) is in the initial state and $\alpha_{e}(\alpha_{\nu})=-1$ if the electron (neutrino) is in the final state.
-In addition, $f_{i}(E)\equiv\frac{1}{e^{E/T_{\gamma}}+1}$ for $i=a,b,e$ is the Fermi-Dirac distribution with vanishing chemical potential, and we have used the relation $1-f_{i}(E)=f_{i}(-E)$ to incorporate the corresponding Pauli blocking factors.
-Lastly, $\tilde{f}_{\nu}(E_{\nu})$ equals $\varrho_{\nu}(E_{\nu})\,(\varrho_{\bar{\nu}}(E_{\nu}))$ for the neutrino (antineutrino) in the initial state and equals $1-\varrho_{\nu}(E_{\nu})\,(1-\varrho_{\bar{\nu}}(E_{\nu}))$ for the neutrino (antineutrino) in the initial state.
+where $\mathrm{d}\pi_{i}=\frac{1}{(2\pi)^{3}}\frac{\mathrm{d}^{3}\mathbf{p}_{i}}{2E_{i}}$, $\alpha_{e}(\alpha_{\nu})=1$ if the electron (neutrino) is in the initial state and $\alpha_{e}(\alpha_{\nu})=-1$ if it is in the final state.
+Additionally, $f_{i}(E)\equiv\frac{1}{e^{E/T_{\gamma}}+1}$ for $i=a,b,e$ is the Fermi-Dirac distribution with a vanishing chemical potential, and we have used the relation $1-f_{i}(E)=f_{i}(-E)$ to account for the corresponding Pauli blocking factors.
+Lastly, $\tilde{f}_{\nu}(E_{\nu})$ equals $\varrho_{\nu}(E_{\nu})$ (or $\varrho_{\bar{\nu}}(E_{\nu})$) for the neutrino (or antineutrino) in the initial state, and equals $1-\varrho_{\nu}(E_{\nu})$ (or $1-\varrho_{\bar{\nu}}(E_{\nu})$) for those in the final state.
 […]
@@
 $\bar{\Gamma}_{p\to n}$ $=K\int_{0}^{\infty}p^{2}\mathrm{d}p[\chi_{-}(E)+\chi_{-}(-E)]\,,$ (75)
-with $E=\sqrt{p^{2}+m_{e}^{2}}$ is the energy of the electron and $K=1/(\tau_{n}\lambda_{0}m_{e}^{5})$, where $\tau_{n}$ is the life time of the neutron and $\lambda_{0}$ the Born coefficient.
-The $\chi^{\pm}$ function is defined as
-$\chi_{\pm}(E)\equiv(E_{\nu}^{\mp})^{2}f_{e}(-E)f_{\nu}^{\pm}(E^{\mp}_{\nu})\,,\,\,E^{\mp}_{\nu}\equiv E\mp\Delta,$ (75)
-where $\Delta\equiv m_{n}-m_{p}$ is the energy gap between neutron and proton.
-The first term in Eq.
-(73) corresponds to the $n\to p+e$ processes Eq.
-(65) or Eq.
-(66) depending on whether $E-\Delta$ is positive or negative, and the second term corresponds to the $n+e\to p$ process Eq.
-(67) and similarly for the terms in Eq.
-(74).
-Note that here $E^{\mp}_{\nu}$ is equal to the energy of the neutrino $E_{\nu}$ for the processes where the neutrino is in the initial state, or its opposite $-E_{\nu}$ for the processes where the neutrino is in the final state.
-And the term $f_{\nu}^{\pm}(E^{\mp}_{\nu})$ in the function $\chi^{\pm}$ gives the corresponding distribution or the Pauli-blocking factor by the relation $1-f_{\nu}^{+}(E_{\nu})=f_{\nu}^{-}(-E_{\nu})$.
-Now we introduce the corrections induced by the spectral distortions of neutrinos and antineutrinos. In the infinite nucleon mass limit, the corrections to the Born rates can be written as
-$\delta\bar{\Gamma}_{n\to p}^{\rm SD}$ $=K\int_{0}^{\infty}p^{2}\mathrm{d}p[\delta\chi_{+}(E)+\delta\chi_{+}(-E)]\,,$ (76)
-$\delta\bar{\Gamma}_{p\to n}^{\rm SD}$ $=K\int_{0}^{\infty}p^{2}\mathrm{d}p[\delta\chi_{-}(E)+\delta\chi_{-}(-E)]\,,$ (77)
-with
-$\delta\chi_{\pm}(E)\equiv$ $(E_{\nu}^{\mp})^{2}f_{e}(-E){\rm sgn}(E^{\mp}_{\nu}){}$ (78)
-×$\times\left(\frac{1+\Theta(\pm E^{\mp}_{\nu})\delta\varrho_{\nu}+\Theta(\mp E^{\mp}_{\nu})\delta\varrho_{\bar{\nu}}}{e^{\left|E^{\mp}_{\nu}\right|/T_{\rm CM}\mp{\rm sgn}(E^{\mp}_{\nu})\xi_{\nu}}+1}-\frac{1}{e^{\left|E^{\mp}_{\nu}\right|/T_{\nu}\mp{\rm sgn}(E^{\mp}_{\nu})\xi_{\nu}}+1}\right)\,,$
-where $\Theta(x)$ is the Heaviside step function, $\delta\varrho_{\nu}$ and $\delta\varrho_{\bar{\nu}}$ are the spectral distortions of electron neutrinos and antineutrinos.
-For a particular process, the various sgn functions and the $\Theta$ functions together represent the nature and the position of the neutrino particle: Whether it is a neutrino or an antineutrino, and whether it is in the initial state or the final state.
-We also note that $T_{\rm CM}$ and $T_{\nu}$ has slight difference after neutrino decoupling.
-If we neglect such difference and set $\xi_{\nu}=0$, so that $\delta\varrho_{\nu}=\delta\varrho_{\bar{\nu}}$, then Eq.
-(78) coincides with Eq.
+where $E=\sqrt{p^{2}+m_{e}^{2}}$ represents the energy of the electron and $K=1/(\tau_{n}\lambda_{0}m_{e}^{5})$, with $\tau_{n}$ being the lifetime of the neutron and $\lambda_{0}$ the Born coefficient.
+The $\chi^{\pm}$ functions are defined as
+$\chi_{\pm}(E)\equiv(E_{\nu}^{\mp})^{2}f_{e}(-E)f_{\nu}^{\pm}(E^{\mp}_{\nu})\,,\quad E^{\mp}_{\nu}\equiv E\mp\Delta,$ (76)
+where $\Delta\equiv m_{n}-m_{p}$ is the energy gap between the neutron and proton.
+The first term in (74) corresponds to the $n\to p+e$ processes (66) or (67), depending on whether $E-\Delta$ is positive or negative, while the second term corresponds to the $n+e\to p$ process (68).
+Similarly, the terms in (75) can be interpreted.
+Here, $E^{\mp}_{\nu}$ is equal to the energy of the neutrino $E_{\nu}$ when the neutrino is in the initial state, and $-E_{\nu}$ when it is in the final state.
+The term $f_{\nu}^{\pm}(E^{\mp}_{\nu})$ in the function $\chi^{\pm}$ provides the corresponding distribution or Pauli-blocking factor, following the relation $1-f_{\nu}^{+}(E_{\nu})=f_{\nu}^{-}(-E_{\nu})$.
+Now, we introduce the corrections induced by the spectral distortions of neutrinos and antineutrinos. In the infinite nucleon mass limit, the corrections to the Born rates can be expressed as follows:
+$\delta\bar{\Gamma}_{n\to p}^{\rm SD}$ $=K\int_{0}^{\infty}p^{2}\mathrm{d}p\left[\delta\chi_{+}(E)+\delta\chi_{+}(-E)\right]\,,$ (77)
+$\delta\bar{\Gamma}_{p\to n}^{\rm SD}$ $=K\int_{0}^{\infty}p^{2}\mathrm{d}p\left[\delta\chi_{-}(E)+\delta\chi_{-}(-E)\right]\,,$ (78)
+where
+$\delta\chi_{\pm}(E)\equiv$ $(E_{\nu}^{\mp})^{2}f_{e}(-E)\,{\rm sgn}(E^{\mp}_{\nu})$
+×$\times\left(\frac{1+\Theta(\pm E^{\mp}_{\nu})\delta\varrho_{\nu}+\Theta(\mp E^{\mp}_{\nu})\delta\varrho_{\bar{\nu}}}{e^{\left|E^{\mp}_{\nu}\right|/T_{\rm CM}\mp{\rm sgn}(E^{\mp}_{\nu})\xi_{\nu}}+1}-\frac{1}{e^{\left|E^{\mp}_{\nu}\right|/T_{\nu}\mp{\rm sgn}(E^{\mp}_{\nu})\xi_{\nu}}+1}\right)\,,$ (79)
+where $\Theta(x)$ is the Heaviside step function, and $\delta\varrho_{\nu}$ and $\delta\varrho_{\bar{\nu}}$ represent the spectral distortions of electron neutrinos and antineutrinos, respectively.
+For a given process, the various sgn and $\Theta$ functions together characterize the nature and state of the neutrino particle, indicating whether it is a neutrino or an antineutrino and whether it is in the initial or final state.
+Additionally, we note that $T_{\rm CM}$ and $T_{\nu}$ exhibit slight differences after neutrino decoupling.
+If we ignore this difference and set $\xi_{\nu}=0$, resulting in $\delta\varrho_{\nu}=\delta\varrho_{\bar{\nu}}$, then Eq.
+(4.1) coincides with Eq.
 (29) in Ref.
-Froustey:2019owm , where they considered the corrections to the weak rates from neutrino spectral distortions in the SM.
-Finally, we briefly review the other corrections to the weak interaction rates that we have considered in this work.
-First are the zero temperature radiative corrections, including the so-called pure radiative corrections corresponding to the emission and absorption of a virtual photon in the interaction and the so-called bremsstrahlung correction due to the emission of a real photon.
-Second are the finite temperature radiative corrections due to the finite temperature corrections to the proton-electron interactions and the electron energy.
-Third and fourth are the corrections from the finite nucleon mass and the weak magnetism.
-Combining all these corrections, the total weak reaction rates are given by
+Froustey:2019owm, where corrections to the weak rates from neutrino spectral distortions in the SM were considered.
+Finally, we provide a brief overview of other corrections to the weak interaction rates considered in this work.
+First, we address the zero-temperature radiative corrections, which include pure radiative corrections corresponding to the emission and absorption of a virtual photon during the interaction, as well as bremsstrahlung corrections arising from the emission of a real photon.
+Second, we consider finite-temperature radiative corrections related to the finite temperature effects on proton-electron interactions and electron energy.
+Third and fourth, we include corrections due to the finite nucleon mass and weak magnetism.
+The total weak reaction rates, combining all these corrections, are expressed as
 $\Gamma_{n\to p}=\Gamma^{\rm RC0}_{n\to p}+\delta\Gamma^{\rm SD,RC0}_{n\to p}+\Gamma^{T}_{n\to p}+\delta\Gamma^{\rm BS}_{n\to p}+\delta\Gamma^{\rm FM}_{n\to p}\,,$ (80)
-where the superscript ’RC0’, ’SD’, ’T’, ’BS’ and ’FM’ represent the zero temperature radiative correction, the spectral distortion, the finite temperature radiative correction, the bremsstrahlung and the finite nucleon mass correction, respectively.
-And the form for the $p\to n$ rates is similar.
-Note that the effects of different corrections may be coupled.
-For example, the effects of the zero temperature radiative correction on the corrections induced by the spectral distortions are also considered in this work.
-Specifically, the Born rates with the zero temperature radiative correction take the form of
-${\Gamma}^{{\rm RC}0}_{n\to p}$ $=K\int_{0}^{\infty}p^{2}\mathrm{d}p[\mathcal{F}_{+}(E)R(E,|\Delta-E|)\chi_{+}(E)+\mathcal{F}_{+}(-E)R(E,\Delta+E)\chi_{+}(-E)]\,,$ (80)
-${\Gamma}^{{\rm RC}0}_{p\to n}$ $=K\int_{0}^{\infty}p^{2}\mathrm{d}p[\mathcal{F}_{-}(E)R(E,\Delta+E)\chi_{-}(E)+\mathcal{F}_{-}(-E)R(E,|E-\Delta|)\chi_{-}(-E)]\,.$ (81)
-The explicit expressions for the Fermi function $F_{\pm}$ and the radiative correction factor $R(E,k_{\rm max})$ can be found in $e.g.,$   Ref.
-Pitrou:2018cgg and are not repeated here for brevity.
-We see that the zero temperature radiative correction only modifies the coefficients of the $\chi_{\pm}$ function.
-The $\delta\Gamma^{\rm RC0,SD}_{n\to p}$ and $\delta\Gamma^{\rm RC0,SD}_{p\to n}$ can also be obtained simply by the replacement of $\chi_{\pm}\to\delta\chi_{\pm}$.
-In principle, the other corrections are also affected by the spectral distortions of neutrinos and antineutrinos.
-However, such effects must be negligible since the corrections themselves are already sub-leading.
-Therefore, we neglect these interferences in this work.
+where the superscripts ’RC0’, ’SD’, ’T’, ’BS’, and ’FM’ denote the zero-temperature radiative correction, spectral distortion, finite-temperature radiative correction, bremsstrahlung, and finite nucleon mass correction, respectively.
+The form for the $p\to n$ rates is analogous.
+It is important to note that the effects of different corrections may be coupled.
+For instance, the impact of zero-temperature radiative corrections on those induced by spectral distortions is also addressed in this work.
+Specifically, the Born rates with zero-temperature radiative correction are given by
+${\Gamma}^{{\rm RC}0}_{n\to p}$ $=K\int_{0}^{\infty}p^{2}\mathrm{d}p\left[\mathcal{F}_{+}(E)R(E,|\Delta-E|)\chi_{+}(E)+\mathcal{F}_{+}(-E)R(E,\Delta+E)\chi_{+}(-E)\right]\,,$ (81)
+${\Gamma}^{{\rm RC}0}_{p\to n}$ $=K\int_{0}^{\infty}p^{2}\mathrm{d}p\left[\mathcal{F}_{-}(E)R(E,\Delta+E)\chi_{-}(E)+\mathcal{F}_{-}(-E)R(E,|E-\Delta|)\chi_{-}(-E)\right]\,.$ (82)
+The explicit expressions for the Fermi function $F_{\pm}$ and the radiative correction factor $R(E,k_{\rm max})$ can be found in, e.g., Ref.
+Pitrou:2018cgg, and are not repeated here for brevity.
+It is evident that the zero-temperature radiative correction modifies only the coefficients of the $\chi_{\pm}$ functions.
+The terms $\delta\Gamma^{\rm RC0,SD}_{n\to p}$ and $\delta\Gamma^{\rm RC0,SD}_{p\to n}$ can be derived simply by replacing $\chi_{\pm}$ with $\delta\chi_{\pm}$.
+In principle, other corrections are also influenced by the spectral distortions of neutrinos and antineutrinos.
+However, these effects are expected to be negligible, as the corrections themselves are already sub-leading.
+Thus, we disregard these interferences in this work.
 4.2 Results of BBN Abundances
-In this subsection we present the resulting light element abundances incorporating all the corrections discussed in the last subsection.
-To do this, we have added the aforementioned correction induced by the neutrino spectral distortions in the public BBN code PRIMATPitrou:2018cgg and solved the abundances of the light elements.
-For the computation we have used the latest values of other physical and cosmological parameters, $e.g.,$ the neutron lifetime is taken as $\tau_{n}=879.4s$, the axial coupling of the nucleons is $g_{A}=1.2756$, and the baryon density is $\omega_{b}h^{2}=0.024$.
-For convenience, we define the mass fraction of isotope $i$ as $X_{i}\equiv A_{i}\frac{n_{i}}{n_{b}}$, where $A_{i}$ and $n_{i}$ are the baryon number and the number density of isotope $i$ and $n_{b}$ is the total number density of baryon.
-Customarily, we also denote $Y_{P}\equiv X_{{}^{4}{\rm He}}$ and $i/\rm H\equiv n_{i}/n_{{}^{1}\rm H}$.
+In this subsection, we present the resulting light element abundances, incorporating all corrections discussed previously.
+Specifically, we have implemented the corrections induced by neutrino spectral distortions into the public BBN code PRIMAT Pitrou:2018cgg and computed the abundances of the light elements.
+For this computation, we utilized the latest values of relevant physical and cosmological parameters; for instance, the neutron lifetime is taken to be $\tau_{n}=879.4\,\text{s}$, the axial coupling of the nucleons is $g_{A}=1.2756$, and the baryon density is $\omega_{b}h^{2}=0.024$.
+For convenience, we define the mass fraction of isotope $i$ as $X_{i}\equiv A_{i}\frac{n_{i}}{n_{b}}$, where $A_{i}$ and $n_{i}$ are the baryon number and number density of isotope $i$, respectively, and $n_{b}$ is the total baryon number density.
+We also customarily denote $Y_{P}\equiv X_{{}^{4}{\rm He}}$ and $i/\rm H\equiv n_{i}/n_{{}^{1}\rm H}$.
 Figure 5: Light element abundances as a function of the neutrino degeneracy parameter $\xi_{\nu}$, normalized to the results for $\xi_{\nu}=0$.
-The ${}^{3}{\rm He}$ stands for ${}^{3}{\rm He}+\rm T$ and ${}^{7}{\rm Li}$ stands for ${}^{7}{\rm Li}+\rm^{7}Be$ due to the slow radioactive decays.
-All the optional weak rate corrections in PRIMAT (see Ref.
-Pitrou:2018cgg ) and the new correction from neutrino spectral distortions are included in the calculation.
-The $68\%$ constraints on the helium abundance $Y\rm_{P}$ from the CMB (${Y\rm_{P}}=0.242\pm 0.012$) and the recent EMPRESS experiment (${Y\rm_{P}}=0.2370^{+0.0034}_{-0.0033}$) are indicated by the light and dark blue shaded regions, respectively.
-We show in Fig.
-5 the resulting light element abundances, including all the aforementioned weak rate corrections in (79) and the QED corrections up to $\mathcal{O}(e^{3})$ for plasma thermodynamics and certain nuclear rates, as a function of the neutrino degeneracy parameter $\xi_{\nu}$.
-The abundance of ${}^{3}{\rm He}$ stands for ${}^{3}{\rm He}+\rm T$ and ${}^{7}{\rm Li}$ for ${}^{7}{\rm Li}+\rm^{7}Be$ to account for the slow radioactive decays.
-Since the abundances for different elements can vary by several orders of magnitude, we normalize the element abundances by their corresponding abundance with $\xi_{\nu}=0$.
-Therefore all curves in Fig.
+The ${}^{3}{\rm He}$ notation refers to ${}^{3}{\rm He}+\rm T$, and ${}^{7}{\rm Li}$ denotes ${}^{7}{\rm Li}+\rm^{7}Be$, accounting for slow radioactive decays.
+All optional weak rate corrections in PRIMAT (see Ref.
+Pitrou:2018cgg) and the new corrections from neutrino spectral distortions are included in the calculations, see text for more details.
+The $68\%$ constraints on the helium abundance $Y_{P}$ from the CMB ($Y_{P}=0.242\pm 0.012$) and the recent EMPRESS experiment ($Y_{P}=0.2370^{+0.0034}_{-0.0033}$) are indicated by the light and dark blue shaded regions, respectively.
+Figure 5 illustrates the resulting light element abundances, incorporating all weak rate corrections from (80) and the QED corrections up to $\mathcal{O}(e^{3})$ for plasma thermodynamics and certain nuclear rates, as a function of the neutrino degeneracy parameter $\xi_{\nu}$.
+The abundance of ${}^{3}{\rm He}$ refers to ${}^{3}{\rm He}+\rm T$, while ${}^{7}{\rm Li}$ corresponds to ${}^{7}{\rm Li}+\rm^{7}Be$, accounting for slow radioactive decays.
+Since the abundances of different elements can vary by several orders of magnitude, we normalize the element abundances to their respective values at $\xi_{\nu}=0$.
+Consequently, all curves in Fig.
 5 intersect at $\xi_{\nu}=0$.
-We see that the abundances of both $\rm D,\,^{3}{\rm He},\,^{4}{\rm He}$ and ${}^{7}{\rm Li}$ decrease with increasing $\xi_{\nu}$ due to the fact that the larger the $\xi_{\nu}$, the smaller the number density ratio between neutron and proton.
-In addition, the variation of the ${}^{4}{\rm He}$ abundance is the most sensitive to the variation of $\xi_{\nu}$, while the ${}^{3}{\rm He}$ abundance is the least sensitive.
-We also show in Fig.
-5 the constraints on the helium abundance from obtained by CMB and EMPRESS experiments,
-${Y\rm_{P}}|_{\rm CMB}=0.242\pm 0.012,\,\,{Y\rm_{P}}|_{\rm EMPRESS}=0.2370^{+0.0034}_{-0.0033},$ (82)
-by the light and dark blue shaded regions.
-From our results, the constraints on ${Y\rm_{P}}$ within the $68\%$ confidence interval can be easily translated into constraints on the neutrino degeneracy parameter $\xi_{\nu}$ as
-${\rm CMB}:\,\,-0.026\leq$ $\,\xi_{\nu}\leq 0.069\,,$ (83)
-EMPRESS:  ${\rm EMPRESS}:\,\,0.027\leq$ $\,\xi_{\nu}\leq 0.054\,.$ (84)
-Clearly, the EMPRESS experiment prefer a nonzero positive $\xi_{\nu}$ while the CMB experiment is still compatible with the SM case $\xi_{\nu}=0$.
-Figure 6: The relative difference of the light element abundances compared to the results in the neutrino instantaneous decoupling limit as a function of the neutrino degeneracy parameter $\xi_{\nu}$.
-The solid and dotted curves correspond to the results with and without the weak corrections induced by the neutrino spectral distortions, respectively.
-The meaning of the color for the curves is the same as in the plot Fig.
-5, except that the negative results of ${}^{7}{\rm Li}$ are shown here for simplicity.
+We observe that the abundances of $\rm D,\,^{3}{\rm He},\,^{4}{\rm He}$, and ${}^{7}{\rm Li}$ decrease with increasing $\xi_{\nu}$ due to the corresponding reduction in the neutron-to-proton number density ratio.
+Furthermore, the variation in ${}^{4}{\rm He}$ abundance is most sensitive to changes in $\xi_{\nu}$, while the ${}^{3}{\rm He}$ abundance is the least sensitive.
+Additionally, Fig.
+5 shows the constraints on helium abundance derived from the CMB and EMPRESS experiments:
+$Y_{P}|_{\rm CMB}=0.242\pm 0.012,\quad Y_{P}|_{\rm EMPRESS}=0.2370^{+0.0034}_{-0.0033},$ (83)
+as represented by the light and dark blue shaded regions.
+Our results indicate that the constraints on $Y_{P}$ within the $68\%$ confidence interval can be translated into constraints on the neutrino degeneracy parameter $\xi_{\nu}$:
+${\rm CMB}:\quad-0.026\leq$ $\,\xi_{\nu}\leq 0.069\,,$ (84)
+${\rm EMPRESS}:\quad 0.027\leq$ $\,\xi_{\nu}\leq 0.054\,.$ (85)
+Clearly, the EMPRESS experiment favors a non-zero positive $\xi_{\nu}$, while the CMB experiment remains compatible with the Standard Model case of $\xi_{\nu}=0$.
+Figure 6: The relative difference in light element abundances compared to the results in the neutrino instantaneous decoupling limit as a function of the neutrino degeneracy parameter $\xi_{\nu}$.
+The solid and dotted curves correspond to results with and without weak corrections induced by neutrino spectral distortions, respectively.
+The color scheme for the curves is consistent with that in Fig.
+5, except that negative results for ${}^{7}{\rm Li}$ are presented here for simplicity.
 BBN Framework ${Y\rm_{P}}$ ${\rm D}/{\rm H}\times 10^{5}$ ${{}^{3}{\rm He}}/{\rm H}\times 10^{5}$ ${{}^{7}{\rm Li}}/{\rm H}\times 10^{10}$
@@
 No-SD, $\xi_{\nu}=-0.1$ 0.27375 2.59246 1.06239 5.85555
-No-SD, $\xi_{\nu}=0$ 0.24712 2.43725 1.03926 5.5077
-No-SD, $\xi_{\nu}=0.1$ 0.2225 2.31162 1.02011 5.12186
+No-SD, $\xi_{\nu}=0$ 0.24712 2.43725 1.03926 5.50770
+No-SD, $\xi_{\nu}=0.1$ 0.22250 2.31162 1.02011 5.12186
 Actual-SD, $\xi_{\nu}=-0.1$ 0.27375 2.59247 1.06239 5.85558
@@
 Actual-SD, $\xi_{\nu}=0.1$ 0.22266 2.31244 1.02024 5.12449
-Table 1: Light element abundances including all weak rate corrections in (79), for various implementations of neutrino decoupling and degeneracy parameters.
-${}^{3}{\rm He}$ stands for $(^{3}{\rm He}+\mathrm{T})$ and ${}^{7}{\rm Li}$ stands for $(^{7}{\rm Li}+\rm Be)$ to account for slow radioactive decays.
+Table 1: Light element abundances, including all weak rate corrections in (80), for various implementations of neutrino decoupling and degeneracy parameters.
+${}^{3}{\rm He}$ denotes $(^{3}{\rm He}+\mathrm{T})$ and ${}^{7}{\rm Li}$ denotes $(^{7}{\rm Li}+\rm Be)$ to account for slow radioactive decays.
 BBN Framework $\delta{Y\rm_{P}}\,(\%)$ $\delta{\rm D}/{\rm H}\,(\%)$ $\delta{{}^{3}{\rm He}}/{\rm H}\,(\%)$ $\delta{{}^{7}{\rm Li}}/{\rm H}\,(\%)$
 No-SD, $\xi_{\nu}=-0.1$ 0.058 0.4 0.137 -0.411
-No-SD, $\xi_{\nu}=0$ 0.007 0.367 0.122 -0.43
-No-SD, $\xi_{\nu}=0.1$0.048 0.343 0.112 -0.46
+No-SD, $\xi_{\nu}=0$ 0.007 0.367 0.122 -0.430
+No-SD, $\xi_{\nu}=0.1$0.048 0.343 0.112 -0.460
 Actual-SD, $\xi_{\nu}=-0.1$ 0.059 0.4 0.137 -0.411
@@
 Actual-SD, $\xi_{\nu}=0.1$ 0.026 0.379 0.125 -0.408
-Table 2: Relative difference of the light element abundances for the case with and without the corrections induced by the neutrino spectral distortions compared to the results in the neutrino instantaneous decoupling limit.
-${}^{3}{\rm He}$ stands for $(^{3}{\rm He}+\mathrm{T})$ and ${}^{7}{\rm Li}$ stands for $(^{7}{\rm Li}+\rm Be)$ to account for slow radioactive decays.
-It is also interesting to discuss the effects of the corrections induced by the neutrino spectral distortions on BBN.
-Focusing on the corrections induced by different treatments of the neutrino decoupling process, we consider the following three implementations: (1) The neutrino instantaneous decoupling limit, $i.e.,$ the $N_{\rm eff}^{\rm inst}$ as discussed in the last section and the thermal Fermi-Dirac distributions for neutrinos and antineutrinos are used in the calculation of the BBN; (2) The no-SD case, where the actual $N_{\rm eff}$ and neutrino temperatures are used, including the non-instantaneous decoupling effect, but considering only the thermal Fermi-Dirac distributions for neutrinos and antineutrinos.
-This is the most common treatment for BBN in the studies to date; (3) The actual SD case, using the actual $N_{\rm eff}$ with the non-instantaneous decoupling effect and the actual density distributions with the spectral distortions for neutrinos and antineutrinos, corresponding to the results shown in Fig.
+Table 2: Relative differences in light element abundances, comparing cases with and without corrections induced by neutrino spectral distortions to results in the neutrino instantaneous decoupling limit.
+${}^{3}{\rm He}$ denotes $(^{3}{\rm He}+\mathrm{T})$ and ${}^{7}{\rm Li}$ denotes $(^{7}{\rm Li}+\rm Be)$ to account for slow radioactive decays.
+It is also interesting to discuss the effects of corrections induced by neutrino spectral distortions on BBN.
+We examine the impact of different treatments of the neutrino decoupling process, considering three implementations: (1) the neutrino instantaneous decoupling limit, $i.e.,$ , $N_{\rm eff}^{\rm inst}$, as discussed in the previous section, where thermal Fermi-Dirac distributions for neutrinos and antineutrinos are used in the BBN calculations; (2) the no-SD case, which employs the actual $N_{\rm eff}$ and neutrino temperatures, incorporating the non-instantaneous decoupling effect while only using thermal Fermi-Dirac distributions for neutrinos and antineutrinos.
+This treatment is the most common in current BBN studies; (3) the actual SD case, which utilizes the actual $N_{\rm eff}$, accounting for the non-instantaneous decoupling effect and incorporating actual density distributions with spectral distortions for neutrinos and antineutrinos.
+This case corresponds to the results shown in Fig.
 5.
-Using the BBN results in the instantaneous decoupling limit as the benchmark, we show the relative differences of the results for the no-SD and the actual-SD cases with the dotted and solid curves in Fig.
+Using the BBN results in the instantaneous decoupling limit as a benchmark, we present the relative differences for the no-SD and actual-SD cases, indicated by dotted and solid curves in Fig.
 6, respectively.
-Note that we have plotted the opposite of the relative differences for ${}^{7}{\rm Li}$ in Fig.
-6 as it is negative due to the particular nuclear reaction rates used in PRIMAT.
-We also report the explicit values of the abundances of helium-4, deuterium, helium-3 and lithium-7 for the degeneracy parameters $\xi_{\nu}=0,\pm 1$ in table 1, and the associated relative variations compared to the baseline abundance in table 2.
-We also note that our results with $\xi_{\nu}=0$ are consistent with the results reported in Froustey:2020mcq , where they also considered the effects of neutrino spectral distortions to BBN within the SM case.
-As we can see in Fig.
-6 and the tables, the actual-SD case and the no-SD case yield almost the same light element abundances for $\xi_{\nu}=-0.1$.
-And the variation of the BBN abundances from the results in the instantaneous decoupling limit for the actual-SD case show little dependence on $\xi_{\nu}$, while the traditional treatment without the spectral distortions predicts a stronger dependence for both elements.
-In particular, the no-SD case predicts negative variations for the helium abundance with positive $\xi_{\nu}$ compared to the result in the instantaneous decoupling limit.
-However, the more precise results with the spectral distortions show that the variations of the helium abundance are positive, showing that the inclusion of the neutrino spectral distortions is necessary for a precise and accurate prediction of the BBN abundances.
-Finally, we also provide a set of fitting functions for the light element abundances as
+Note that we have plotted the negative of the relative differences for ${}^{7}{\rm Li}$ in Fig.
+6 due to the specific nuclear reaction rates utilized in PRIMAT.
+We report the explicit abundances of helium-4, deuterium, helium-3, and lithium-7 for the degeneracy parameters $\xi_{\nu}=0,\pm 1$ in table 1, along with the associated relative variations compared to the baseline abundance in table 2.
+Our results with $\xi_{\nu}=0$ are consistent with those reported in Froustey:2020mcq , which also considered the effects of neutrino spectral distortions on BBN within the standard model (SM).
+As shown in Fig.
+6 and the accompanying tables, the actual-SD and no-SD cases yield nearly identical light element abundances for $\xi_{\nu}=-0.1$.
+The BBN abundances in the actual-SD case show minimal dependence on $\xi_{\nu}$, while the traditional no-SD treatment predicts a stronger dependence for both elements.
+Notably, the no-SD case predicts negative variations for the helium abundance with positive $\xi_{\nu}$ compared to the instantaneous decoupling limit.
+In contrast, the more precise results with spectral distortions indicate positive variations in helium abundance, demonstrating that incorporating neutrino spectral distortions is crucial for accurate predictions of BBN abundances.
+Finally, we provide fitting functions for the light element abundances as follows:
 $Y_{\mathrm{P}}\left(\xi_{\nu}\right)$ $\simeq\left.Y_{\mathrm{P}}\right|_{\xi_{\nu}=0}\times e^{-1.033\xi-0.131\xi^{2}},\,\,$ (86)
@@
 ${}^{7}\mathrm{Li/H}\left(\xi_{\nu}\right)$ $\simeq\left.{}^{7}\mathrm{Li/H}\right|_{\xi_{\nu}=0}\times e^{-0.63\,\xi_{\nu}(1+\xi_{\nu})},$ (89)
-where the subscript $\xi_{\nu}=0$ labels the results obtained in the SM case $\xi_{\nu}=0$ using the actual density matrices, see the explicit values in Table 1.
-Such fitting functions benefit for further studies, and the fractional errors are smaller than $0.1\%$ with the range $-0.1\leq\xi_{\nu}\leq 0.1$.
-[^3] The coefficients of our fitting functions are slightly different from the results in Pitrou:2018cgg ; Froustey:2024mgf , which we think is due to the different corrections to the weak interaction rates we considered.
-5 Impacts of the Primordial Neutrino Asymmetry on CMB and LSS
-In this section we focus on the impacts of primordial neutrino asymmetries on the evolution of the late universe, especially the cosmic microwave background (CMB) and the large scale structure (LSS).
-In general, the primordial neutrino asymmetry influence CMB and LSS in three ways.
+where the subscript $\xi_{\nu}=0$ indicates results obtained in the SM case using actual density matrices, as detailed in Table 1.
+These fitting functions facilitate further studies, yielding fractional errors smaller than $0.01\%$ within the range $-0.1\leq\xi_{\nu}\leq 0.1$.
+[^3] The coefficients of our fitting functions differ slightly from those in Pitrou:2018cgg ; Froustey:2024mgf , which is mainly due to the different neutron lifetimes we have chosen.
+In addition, our fitting function is more precise.
+5 Impacts of Primordial Neutrino Asymmetry on the CMB and LSS
+In this section, we examine the effects of primordial neutrino asymmetries on the evolution of the late universe, particularly focusing on the cosmic microwave background (CMB) and large-scale structure (LSS).
+Primordial neutrino asymmetry affects the CMB and LSS in three primary ways:
   1. 1.
-    The expansion rate of the Universe is changed by the increased $N_{\rm eff}$ due to the primordial neutrino asymmetries.
-Particularly, the time of matter-radiation equality is delayed by the additional radiation density, leading to an enhanced early integrated Sachs-Wolfe effect for CMB angular power spectrum.
+    The expansion rate of the universe is altered by the increased $N_{\rm eff}$ resulting from primordial neutrino asymmetries.
+Specifically, the additional radiation density delays the time of matter-radiation equality, enhancing the early integrated Sachs-Wolfe effect in the CMB angular power spectrum.
   2. 2.
-    As we have seen in Fig.
-5, the abundance of helium is significantly modified by the electron neutrino asymmetry, which will in turn alter the evolution of the late Universe, such as the tail of the CMB angular power spectrum by the amount of diffusion damping and the baryon acoustic oscillation (BAO).
+    As shown in Fig.
+5, the electron neutrino asymmetry significantly modifies the helium abundance, which subsequently affects the late-time evolution of the universe.
+This impacts features such as the tail of the CMB angular power spectrum due to diffusion damping and the baryon acoustic oscillations (BAO).
   3. 3.
-    In addition, in order to incorporate the effect of neutrino oscillation in the process of neutrino decoupling, the mass of neutrinos must be considered.
-The neutrino masses will alter the evolution of the late universe by such as, affect the matter-radiation equality, even in the SM case with the thermal Fermi-Dirac distributions.
-For the SM case, the experiment on CMB and BAO can put a relative stringent bound on the sum of neutrino masses $\sum m_{\nu}<0.07\,{\rm eV}$ DESI:2024mwx , we will return to the constraint on neutrino masses in the next section.
-Here we emphasize that the neutrino Boltzmann hierarchy for massive neutrinos will be furthermore modified by the neutrino degeneracy parameter and the spectral distortions.
-In the following we first review the impact of massive neutrinos on the universe and present the Boltzmann hierarchies of electrons, photons and neutrinos.
-Then we discuss the effects of the primordial neutrino asymmetries on the CMB temperature anisotropies and the linear matter power spectrum computed with the Einstein-Boltzmann code CLASS Lesgourgues:2011re ; Lesgourgues:2011rh .
-5.1 Effects of the primordial neutrino asymmetries in structure formation
-To describe the flavor oscillations in the neutrino decoupling process, the mass of the neutrinos must be taken into account, which in turn has significant implications for the process of structure formation, since massive neutrinos are the only species in the SM that scale as radiation in the early universe and as matter in the late universe.
-In the early universe, neutrinos and antineutrinos remain relativistic after decoupling as long as their effective temperatures are much larger than their masses.
-In this case, the neutrino and antineutrino energy densities contribute to the total radiation energy density, parameterised by the aforementioned $N_{\rm eff}$.
-Since the neutrino effective temperatures decrease as $a^{-1}$ with the expansion of the universe, neutrinos and antineutrinos become non-relativistic when their effective temperatures fall below their masses, and contribute to matter after that.
-The non-relativistic transitions and the free-steaming effects of massive neutrinos then leave significant imprints on the CMB anisotropies and the LSS.
-As mentioned above, the situation is modified in the presence of the primordial neutrino asymmetries.
-In addition to the indirect change induced by the modified Hubble rate and the helium abundance, we highlight several other modifications due to the primordial neutrino asymmetries in the following.
+    Additionally, to account for neutrino oscillations during neutrino decoupling, neutrino masses must be considered.
+These masses influence the late-time evolution of the universe, particularly affecting matter-radiation equality, even in the standard model (SM) scenario with thermal Fermi-Dirac distributions.
+For the SM case, CMB and BAO observations can place stringent constraints on the sum of neutrino masses, yielding $\sum m_{\nu}<0.07\,{\rm eV}$ DESI:2024mwx .
+We will revisit these constraints in the next section.
+Furthermore, the neutrino Boltzmann hierarchy for massive neutrinos is modified by the neutrino degeneracy parameter and spectral distortions.
+In the following, we first review the impact of massive neutrinos on the universe and present the Boltzmann hierarchies for electrons, photons, and neutrinos.
+We then discuss the effects of primordial neutrino asymmetries on CMB temperature anisotropies and the linear matter power spectrum, computed using the Einstein-Boltzmann code CLASS Lesgourgues:2011re ; Lesgourgues:2011rh .
+5.1 Effects of Primordial Neutrino Asymmetries in Structure Formation
+To accurately describe flavor oscillations during neutrino decoupling, neutrino mass must be considered, which has significant implications for structure formation.
+Massive neutrinos are the only species in the SM that behave as radiation in the early universe and as matter in the late universe.
+In the early universe, neutrinos and antineutrinos remain relativistic after decoupling, as long as their effective temperatures are much higher than their masses.
+In this regime, their energy densities contribute to the total radiation energy density, parameterized by $N_{\rm eff}$.
+As the universe expands and the effective temperatures decrease as $a^{-1}$, neutrinos and antineutrinos become non-relativistic when their effective temperatures fall below their masses, at which point they contribute to the matter content.
+The transition to the non-relativistic phase and the free-streaming effects of massive neutrinos leave noticeable imprints on CMB anisotropies and LSS.
+As discussed, primordial neutrino asymmetries modify this scenario.
+In addition to indirect changes from the altered Hubble rate and helium abundance, we emphasize several direct effects of primordial neutrino asymmetries.
 5.1.1 Effects on Neutrino Abundance
-First, the total relic abundance of neutrinos and antineutrinos are directly increased due to the primordial neutrino asymmetries. At present, the total neutrino relic abundance can be written as
+First, the total relic abundance of neutrinos and antineutrinos increases directly due to primordial neutrino asymmetries. At present, the total neutrino relic abundance is given by
 $\Omega_{\nu}h^{2}$ $=\frac{\sum\limits_{i}T_{\rm CM,0}^{4}\int_{0}^{\infty}\mathrm{d}y\,y^{2}\sqrt{y^{2}+m_{i}}\left(\varrho_{i}(y)+\bar{\varrho}_{i}(y)\right)/(2\pi^{2})}{\rho_{\rm crit}}$ (90)
 $\overset{\delta\varrho=0}{\simeq}\frac{\sum\limits_{i}m_{i}}{93.12\,{\rm eV}}\frac{{\rm Li}_{3}(-\exp(-\xi_{\nu}))+{\rm Li}_{3}(-\exp(\xi_{\nu}))}{3\zeta(3)/2},$
-where $y=p/T_{\rm CM}$ is the comoving momentum defined in section 3, $T_{\rm CM,0}$ is the current comoving temperature, $\rho_{\rm crit}$ is the critical energy density, $\varrho_{i}(y)$ and $\bar{\varrho}_{i}(y)$ are the respective comoving density distributions for neutrinos and antineutrinos with $i=1,2,3$, corresponding to the three mass eigenstates with mass $m_{i}$.
-Obviously, the present neutrino abundance is also modified by the spectral distortions of neutrinos and antineutrinos, and the integrals over the neutrino distributions cannot be carried out analytically in this case.
-The effect of the primordial neutrino asymmetries will be clearer by neglecting the spectral distortions and performing the integral analytically, as seen in the second row of (89).
-The first factor in the second row $\frac{\sum\limits_{i}m_{i}}{93.12\,{\rm eV}}$ is the usual neutrino relic abundance in the SM case, while the second factor represents the effect of neutrino asymmetries with ${\rm Li}_{j}(x)$ the polylogarithms of order $j$ and $\zeta(x)$ the Riemann zeta function.
-For the SM case one finds ${\rm Li}_{3}(-1)=3\zeta(3)/4$, so the second factor becomes 1.
-For non-zero $\xi_{\nu}$, the second factor depends only on the absolute value of $\xi_{\nu}$, and the larger the $\xi_{\nu}$, the larger the second factor and the larger the neutrino abundance.
-However, we note that such corrections from primordial neutrino asymmetries are quite small.
-For example, the resulting relative difference of $\Omega_{\nu}h^{2}$ for $\xi_{\nu}=0.2$ is only $1.54\%$.
-Similarly, the corrections to the neutrino abundance induced by the spectral distortions of neutrinos and antineutrinos are also negligible.
+where $y=p/T_{\rm CM}$ is the comoving momentum defined in section 3, $T_{\rm CM,0}$ is the current comoving temperature, $\rho_{\rm crit}$ is the critical energy density, and $\varrho_{i}(y)$ and $\bar{\varrho}_{i}(y)$ represent the comoving density distributions of neutrinos and antineutrinos, respectively, for $i=1,2,3$, corresponding to the three mass eigenstates with mass $m_{i}$.
+The present neutrino abundance is further modified by spectral distortions of neutrinos and antineutrinos, though the integrals over their distributions cannot be carried out analytically in this case.
+The effect of primordial neutrino asymmetries becomes clearer when spectral distortions are neglected, allowing an analytical evaluation of the integrals, as shown in the second row of (90).
+The first factor, $\frac{\sum\limits_{i}m_{i}}{93.12\,{\rm eV}}$, represents the usual neutrino relic abundance in the SM, while the second factor accounts for the effects of neutrino asymmetries, with ${\rm Li}_{j}(x)$ being the polylogarithms of order $j$ and $\zeta(x)$ the Riemann zeta function.
+In the SM case, ${\rm Li}_{3}(-1)=3\zeta(3)/4$, so the second factor equals 1.
+For non-zero $\xi_{\nu}$, the second factor depends on the absolute value of $\xi_{\nu}$, with larger $\xi_{\nu}$ leading to a larger factor and, consequently, a larger neutrino abundance.
+However, the corrections from primordial neutrino asymmetries are small.
+For instance, the relative difference in $\Omega_{\nu}h^{2}$ for $\xi_{\nu}=0.2$ is only $1.54\%$.
+Similarly, the corrections to the neutrino abundance from spectral distortions of neutrinos and antineutrinos are negligible.
 5.1.2 Effects on the Neutrino Boltzmann Hierarchies
-In order to obtain the accurate CMB spectral power spectrum and the matter power spectrum, we need to solve a coupled set of the Boltzmann-Einstein equations for the perturbations of the metric, photon, baryon, cold dark matter and massive neutrinos.
-Here we briefly review how neutrino asymmetries and spectral distortions affect the Boltzmann-Einstein equations.
+To obtain accurate CMB spectral power and matter power spectra, we must solve the coupled Boltzmann-Einstein equations for perturbations in the metric, photons, baryons, cold dark matter, and massive neutrinos.
+Here, we briefly review how neutrino asymmetries and spectral distortions affect the Boltzmann-Einstein equations.
 Following Ref.
@@
 […]
-$ds^{2}=a^{2}(\tau)\{-d\tau^{2}+(\delta_{ij}+h_{ij})dx^{i}dx^{j}\}.$ (91)
-The perturbation of the metric $h_{ij}$ can be decomposed as $h_{ij}=h\delta_{ij}+h_{ij}^{\parallel}+h_{ij}^{\perp}+h_{ij}^{T}$, where $h\equiv h_{ii}$ is the trace part of $h_{ij}$ , and $h_{ij}^{\parallel}$, $h_{ij}^{\perp}$ and $h_{ij}^{T}$ are the traceless parts, satisfying
+In this work, we adopt the synchronous gauge, where the line element is written as
+$ds^{2}=a^{2}(\tau)\left\{-d\tau^{2}+(\delta_{ij}+h_{ij})dx^{i}dx^{j}\right\},$ (92)
+with the metric perturbation $h_{ij}$ decomposed as $h_{ij}=h\delta_{ij}+h_{ij}^{\parallel}+h_{ij}^{\perp}+h_{ij}^{T}$, where $h\equiv h_{ii}$ is the trace of $h_{ij}$, and $h_{ij}^{\parallel}$, $h_{ij}^{\perp}$, and $h_{ij}^{T}$ are the traceless components, satisfying
 $\epsilon_{ijk}\partial_{j}\partial_{l}h_{lk}^{\parallel}=0,\quad\partial_{i}\partial_{j}h_{ij}^{\perp}=0,\quad\partial_{i}h_{ij}^{T}=0.$ (93)
-By definition, $h_{ij}^{\parallel}$ and $h_{ij}^{\perp}$ can be written in terms of a scalar filed $\mu$ and a divergenceless vector $\vec{A}$,
-$h_{ij}^{\parallel}$ $=\left(\partial_{i}\partial_{j}-\frac{1}{3}\delta_{ij}\nabla^{2}\right)\mu$ (93)
+By definition, $h_{ij}^{\parallel}$ and $h_{ij}^{\perp}$ can be written in terms of a scalar field $\mu$ and a divergenceless vector $\vec{A}$ as
+$h_{ij}^{\parallel}$ $=\left(\partial_{i}\partial_{j}-\frac{1}{3}\delta_{ij}\nabla^{2}\right)\mu,$ (94)
 $h_{ij}^{\perp}$ $=\partial_{i}A_{j}+\partial_{j}A_{i},\quad\partial_{i}A_{i}=0.$
-Therefore, the scalar modes of the metric perturbations are characterised by the two scalar fields $h$ and $\mu$, while the vector and the tensor modes are characterised by $\vec{A}$ and $h_{ij}^{T}$, respectively.
-In this work we will only consider the scalar modes of $h_{ij}$, which can be Fourier transformed as
+Thus, the scalar modes of the metric perturbations are characterized by the two scalar fields $h$ and $\mu$, while the vector and tensor modes are characterized by $\vec{A}$ and $h_{ij}^{T}$, respectively.
+In this work, we focus solely on the scalar modes of $h_{ij}$, which can be Fourier transformed as
 $h_{ij}(\vec{x},\tau)=\int d^{3}k~{}e^{i\vec{k}\cdot\vec{x}}\left\{\frac{k_{i}k_{j}}{k^{2}}h(\vec{k},\tau)+6\eta(\vec{k},\tau)\left(\frac{k_{i}k_{j}}{k^{2}}-\frac{1}{3}\delta_{ij}\right)\right\},$ (95)
-where $h(\vec{k},\tau)$ and $\eta(\vec{k},\tau)$ are the trace part and the traceless part of the scalar modes in the Fourier space which are related to the scalar fields $h$ and $\mu$ in the real space.
-Since neutrinos are collisionless after decoupling, the evolution of $\Psi_{i}$ obeys the first-order collsionless Boltzmann equation in the Fourier-space
+where $h(\vec{k},\tau)$ and $\eta(\vec{k},\tau)$ represent the trace and traceless parts of the scalar modes in Fourier space, related to the scalar fields $h$ and $\mu$ in real space.
+Since neutrinos are collisionless after decoupling, the evolution of $\Psi_{i}$ follows the first-order collisionless Boltzmann equation in Fourier space:
 $\frac{\partial\Psi_{i}}{\partial\tau}+i\frac{q}{\epsilon}(\vec{k}\cdot\hat{n})\Psi_{i}+\frac{d\ln\varrho_{i}^{0}}{d\ln q}\left[\dot{\eta}-\frac{\dot{h}+6\dot{\eta}}{2}(\hat{k}\cdot\hat{n})^{2}\right]=0,$ (96)
-where $\epsilon\equiv aE=\sqrt{q^{2}+a^{2}m^{2}_{\nu_{i}}}$ is the comoving energy. Since Eq. (95) is independent of the azimuth angle, we can expand $\Psi$ in a Legendre series
+where $\epsilon\equiv aE=\sqrt{q^{2}+a^{2}m^{2}_{\nu_{i}}}$ is the comoving energy. Since Eq. (96) is independent of the azimuthal angle, we can expand $\Psi_{i}$ in a Legendre series:
 $\Psi_{i}(\vec{k},\hat{n},q,\tau)=\sum_{l=0}^{\infty}(-i)^{l}(2l+1)\Psi_{i,l}(k,q,\tau)P_{l}(\hat{k}\cdot\hat{n}).$ (97)
-Inserting Eq. (96) into Eq. (95) and using the orthonormality of the Legendre polynomials, we can obtain the Boltzmann hierarchies for $\Psi_{i,l}$:
+Substituting Eq. (97) into Eq. (96) and using the orthonormality of Legendre polynomials, we obtain the Boltzmann hierarchies for $\Psi_{i,l}$:
 $\dot{\Psi}_{i,0}$ $=-\frac{qk}{\epsilon}\Psi_{1}+\frac{\dot{h}}{6}\frac{d\ln\varrho_{i}^{0}}{d\ln q},$ (98)
@@
 $\dot{\Psi}_{i,l\geq 3}$ $=\frac{qk}{(2l+1)\epsilon}\left[l\Psi_{l-1}-(l+1)\Psi_{l+1}\right].$
-Note that the effect of neutrino asymmetries and spectral distortions is manifested by the neutrino density distribution term $\frac{{d}\ln\varrho_{i}^{0}}{{d}\ln q}$.
-Since the density matrices for neutrinos and antineutrinos with different flavors are different in the presence of primordial neutrino asymmetries, the Boltzmann hierarchies for the corresponding perturbations $\Psi_{i}$ and $\bar{\Psi}_{i}$ are in general different and should be treated separately.
-With the moments $\Psi_{i,l}$, the perturbations in physical quantities, such as the energy density, pressure, energy flux and shear stress, can be computed straightforwardly by integrating different moments.
-For our purpose, the fractional density perturbation for the neutrinos with family number $i$ can be expressed as
+The effect of neutrino asymmetries and spectral distortions manifests through the neutrino density distribution term $\frac{d\ln\varrho_{i}^{0}}{d\ln q}$.
+Due to primordial neutrino asymmetries, the density matrices for neutrinos and antineutrinos of different flavors differ, leading to distinct Boltzmann hierarchies for the perturbations $\Psi_{i}$ and $\bar{\Psi}_{i}$, which must be treated separately.
+With the moments $\Psi_{i,l}$, perturbations in physical quantities such as energy density, pressure, energy flux, and shear stress can be computed straightforwardly by integrating different moments.
+Specifically, the fractional density perturbation for neutrinos of family $i$ is given by
 […]
-where $\rho_{i,0}(t)$ is the zeroth-order homogeneous background energy density for neutrinos.
-Eventually, the effects of neutrino masses and asymmetries at different length scales can be reflected in the evolution of the Fourier modes $\delta_{i,k}$, which we obtain via the CLASS code.
+where $\rho_{i,0}(t)$ represents the zeroth-order homogeneous background energy density for neutrinos.
+The effects of neutrino masses and asymmetries across different length scales are reflected in the evolution of the Fourier modes $\delta_{i,k}$, which we obtain via the CLASS code.
 5.1.3 Effects on the BAO
-Following Ma:1995ey , the evolution of the perturbations for photon obey the Boltzmann hierarchy,
-$\dot{\delta}_{\gamma}$ $=$ $-{4\over 3}\theta_{\gamma}-{2\over 3}\dot{h}\,,$
+Following Ref. Ma:1995ey , the evolution of photon perturbations is governed by the Boltzmann hierarchy:
+$\dot{\delta}_{\gamma}$ $=$ $-\frac{4}{3}\theta_{\gamma}-\frac{2}{3}\dot{h},$
 $\dot{\theta}_{\gamma}$ $=$ $k^{2}\left(\frac{1}{4}\delta_{\gamma}-\sigma_{\gamma}\right)+an_{e}\sigma_{T}(\theta_{b}-\theta_{\gamma}),$
 […]
-$\dot{F}_{\gamma\,l}$ $=$ ${k\over 2l+1}\left[lF_{\gamma\,(l-1)}-(l+1)F_{\gamma\,(l+1)}\right]-an_{e}\sigma_{T}F_{\gamma\,l}\,,\quad l\geq 3\,,$
+$\dot{F}_{\gamma\,l}$ $=$ $\frac{k}{2l+1}\left[lF_{\gamma\,(l-1)}-(l+1)F_{\gamma\,(l+1)}\right]-an_{e}\sigma_{T}F_{\gamma\,l},\quad l\geq 3,$
 […]
-where $\delta_{\gamma}$ is the density fluctuation of photon, $\theta_{\gamma}$ is the divergence of photon velocity, and ${F}_{\gamma\,l}$ is the density perturbation of photon.
-In addition, $n_{e}$ is the electron number density and $\sigma_{T}$ is the Thomson scattering cross section.
-The Boltzmann equations for baryon are
-$\dot{\delta}_{b}$ $=$ $-\theta_{b}-{1\over 2}\dot{h}\,,$
-$\dot{\theta}_{b}$ $=$ $-{\dot{a}\over a}\theta_{b}+c_{s}^{2}k^{2}\delta_{b}+{4\bar{\rho}_{\gamma}\over 3\bar{\rho}_{b}}an_{e}\sigma_{T}(\theta_{\gamma}-\theta_{b})\,,$ (100)
-where $\delta_{b}$ and $\theta_{b}$ is the baryon density fluctuation and the divergence of baryon fluid velocity, and $c_{s}^{2}$ is the square of the baryon sound speed.
-Due to the electromagnetic interactions between photons and electrons, they evolve differently before and after recombination.
-Before recombination, photons and baryons are tightly coupled so that they can be treated as a single photon-baryon fluid, called as the "tight-coupling approximation".
-During such "tight-coupling epoch", the perturbations of baryon and photon undergo harmonic motion with very slowly decaying amplitude (Peebles:1970ag, ; Sunyaev:1970bma, ).
-After recombination, photons release their hold on the baryons, so that the baryon acoustic oscillations (BAO) become frozen-in, and the the baryon perturbations begin to grow as $\propto a$.
-The BAO feature then gets transferred to the gravitational potential and the matter fluctuations, leaving imprints in the galaxy power spectrum that depends on the sound horizon at decoupling and the angular diameter distance to the surface of photon-electron last-scattering.
-In the presence of primordial neutrino asymmetries, the matter-radiation equality will be directly modified by the additional contribution to $N_{\rm eff}$, leading to the modification of the BAO properties.
-In addition, the effects of primordial neutrino asymmetries on the evolution of neutrino perturbations are also transported to photons and baryons through metric perturbations encoded in $h$ and $\eta$, even though neutrinos are collisionless.
-As a consequence, the primordial neutrino asymmetries have significant implications for both the CMB angular power spectrum and the BAO in the matter power spectrum, as we will see later.
+Here, $\delta_{\gamma}$ is the photon density fluctuation, $\theta_{\gamma}$ is the divergence of photon velocity, and $F_{\gamma\,l}$ represents the photon density perturbation.
+Additionally, $n_{e}$ is the electron number density, and $\sigma_{T}$ is the Thomson scattering cross section.
+The Boltzmann equations for baryons are
+$\dot{\delta}_{b}$ $=$ $-\theta_{b}-\frac{1}{2}\dot{h},$
+$\dot{\theta}_{b}$ $=$ $-\frac{\dot{a}}{a}\theta_{b}+c_{s}^{2}k^{2}\delta_{b}+\frac{4\bar{\rho}_{\gamma}}{3\bar{\rho}_{b}}an_{e}\sigma_{T}(\theta_{\gamma}-\theta_{b}),$ (101)
+where $\delta_{b}$ and $\theta_{b}$ are the baryon density fluctuation and the divergence of baryon fluid velocity, respectively, and $c_{s}^{2}$ is the square of the baryon sound speed.
+Due to electromagnetic interactions between photons and electrons, the evolution of photons and baryons differs before and after recombination.
+Prior to recombination, photons and baryons are tightly coupled and can be treated as a single photon-baryon fluid, known as the "tight-coupling approximation." During this "tight-coupling epoch," the perturbations of baryons and photons undergo harmonic motion with slowly decaying amplitude (Peebles:1970ag, ; Sunyaev:1970bma, ).
+After recombination, photons decouple from baryons, allowing the baryon acoustic oscillations (BAO) to freeze, and baryon perturbations grow as $\propto a$.
+The BAO feature is then imprinted on the gravitational potential and matter fluctuations, influencing the galaxy power spectrum, which depends on the sound horizon at decoupling and the angular diameter distance to the last-scattering surface.
+Primordial neutrino asymmetries modify the matter-radiation equality via their contribution to $N_{\rm eff}$, altering the BAO properties.
+Moreover, neutrino asymmetries affect the evolution of neutrino perturbations, which in turn influence photon and baryon evolution through metric perturbations encoded in $h$ and $\eta$, despite the collisionless nature of neutrinos.
+As a result, primordial neutrino asymmetries significantly impact both the CMB angular power spectrum and the BAO in the matter power spectrum, as we will explore later.
 5.2 Results of the CMB Anisotropies and the Matter Power Spectrum
-Figure 7: Left panel: The CMB angular power spectrum $C_{l}^{\rm TT}$ within the SM case for different total neutrino masses, compared to the results with massless neutrinos.
-Right panel: The linear matter power spectrum $P(k)$ within the SM case for different total neutrino masses, compared to the results with massless neutrinos.
-We present in this subsection the numerical results of the CMB and LSS obtained via the CLASS code.
-Since both observable are affected by the neutrino masses and asymmetries at the same time, we first discuss the effect of neutrino masses on the CMB and LSS within the SM.
-We present the ratio of the CMB angular power spectrum between the results with massive neutrinos and massless neutrinos $C_{l}^{\rm TT}/C_{l}^{\rm TT}(\sum m_{\nu}=0)$ and the similar ratio for the matter power spectrum $P_{k}/P_{k}(\sum m_{\nu}=0)$ for several choices of total neutrino masses in the left and right panels of Fig.
-7, respectively.
-Note that the normal order of the neutrino mass hierarchy is considered in this work, so we adjust the $\mu_{\nu_{1}}$ for each choice of total mass, while $\mu_{\nu_{2}}$ and $\mu_{\nu_{3}}$ are obtained by the fixed mass differences between neutrinos.
-Focusing on the left panel, the ratio of the $C_{l}^{\rm TT}$ between massive and massless neutrinos can be divided into three stages.
-At large angular scales (lower multipoles ${\ell\lesssim 20}$), massive neutrinos induce a small suppression which comes from the late time integrated Sachs-Wolfe effect.
-On the intermediate scales (${20\lesssim\ell\lesssim 500}$), dominated by the imprints of the acoustic oscillations in the baryon-photon fluid, the first peak of $C_{l}^{\rm TT}$ is affected by the modified evolution of the gravitational potential due to the correlation between massive neutrinos and metric perturbations.
-As we can see, the height of the first peak given by the early integrated Sachs-Wolfe effect is increased for larger neutrino masses compared to the case with massless neutrinos.
-The acoustic peaks at small scales (${\ell\gtrsim 500}$) are also affected by the massive neutrinos.
-Similar with the first peak, the height of the acoustic peaks is increased for larger neutrno masses.
-Similarly, the resulting matter power spectrum with massive neutrinos can also be divided to three stages.
-At large scales corresponding to $k\lesssim 10^{-3}\,h/{\rm Mpc}$, neutrinos do not enter the horizon, and the small suppression of $P(k)$ is induced by the modification of the matter-radiation equality from the non-relativistic transition of massive neutrinos.
-At intermediate scales corresponding to $10^{-3}\,h/{\rm Mpc}\lesssim k\lesssim 1\,h/{\rm Mpc}$, neutrinos gradually enter the horizon, so that the free streaming of these neutrinos results in an additional suppression of the matter power spectrum.
-Therefore, $P_{k}/P_{k}(\sum m_{\nu}=0)$ decreases with increasing wavenumber $k$.
-And at small scales ($k\gtrsim 1\,h/{\rm Mpc}$), the majority of neutrinos stream freely, leading to an almost constant suppression of the power spectrum.
-The larger the total neutrino mass, the greater the suppression of the matter power spectrum.
-It is also worth noting that the matter power spectrum is more sensitive to the mass of the neutrinos, as can be seen from the plots.
-Therefore, it is more efficient to constrain the neutrino masses with the LSS observations, as we will see in the next section.
-Figure 8: Left panel: The CMB angular power spectrum $C_{l}^{\rm TT}$ for different neutrino degeneracy parameters $\xi_{\nu}$, compared to the results with $\xi_{\nu}=0$.
-Right panel: The linear matter power spectrum $P(k)$ for different neutrino degeneracy parameters $\xi_{\nu}$, compared to the results with $\xi_{\nu}=0$.
-The total neutrino mass is fixed to $\sum m_{\nu}=0.1{\rm eV}$ for both panels.
-Figure 9: The sound horizon at the baryon drag epoch of the early universe $r_{d}$ as a function of the primordial neutrino asymmetry $\xi_{\nu}$.
+Figure 7: Left panel: The CMB angular power spectrum $C_{l}^{\rm TT}$ in the SM for different total neutrino masses, compared to massless neutrinos.
+Right panel: The linear matter power spectrum $P(k)$ in the SM for different total neutrino masses, compared to massless neutrinos.
+This subsection presents the numerical results for the CMB and LSS obtained using the CLASS code.
+Since both observables are influenced by neutrino masses and asymmetries, we first analyze the effect of neutrino masses on the CMB and LSS within the SM.
+In Fig.
+7, we show the ratio of the CMB angular power spectrum, $C_{l}^{\rm TT}/C_{l}^{\rm TT}(\sum m_{\nu}=0)$, and the matter power spectrum, $P_{k}/P_{k}(\sum m_{\nu}=0)$, for several choices of total neutrino mass.
+The left and right panels depict the results for the CMB and matter power spectra, respectively.
+This analysis assumes the normal hierarchy for neutrino masses, where $\mu_{\nu_{1}}$ is adjusted for each total mass, while $\mu_{\nu_{2}}$ and $\mu_{\nu_{3}}$ are derived from the fixed mass differences between neutrinos.
+Focusing on the left panel, the ratio of $C_{l}^{\rm TT}$ between massive and massless neutrinos can be categorized into three stages.
+At large angular scales (lower multipoles, $\ell\lesssim 20$), massive neutrinos cause a small suppression due to the late-time integrated Sachs-Wolfe effect.
+At intermediate scales ($20\lesssim\ell\lesssim 500$), dominated by acoustic oscillations in the baryon-photon fluid, the first peak of $C_{l}^{\rm TT}$ is influenced by the modified gravitational potential evolution, resulting from the interaction between massive neutrinos and metric perturbations.
+Notably, the height of the first peak, due to the early integrated Sachs-Wolfe effect, is enhanced for larger neutrino masses compared to massless neutrinos.
+At smaller scales ($\ell\gtrsim 500$), the acoustic peaks are similarly affected, with their height increasing for larger neutrino masses.
+The matter power spectrum with massive neutrinos can also be divided into three stages.
+At large scales ($k\lesssim 10^{-3}\,h/{\rm Mpc}$), neutrinos remain outside the horizon, and the suppression of $P(k)$ results from the shift in the matter-radiation equality, caused by the transition of neutrinos to non-relativistic speeds.
+At intermediate scales ($10^{-3}\,h/{\rm Mpc}\lesssim k\lesssim 1\,h/{\rm Mpc}$), neutrinos gradually enter the horizon, and their free-streaming produces additional suppression in the matter power spectrum, with $P_{k}/P_{k}(\sum m_{\nu}=0)$ decreasing as $k$ increases.
+Finally, at small scales ($k\gtrsim 1\,h/{\rm Mpc}$), most neutrinos freely stream, leading to an almost constant suppression of the power spectrum.
+The larger the total neutrino mass, the greater the suppression.
+It is worth noting that the matter power spectrum is more sensitive to neutrino masses than the CMB, as shown in the plots.
+Thus, LSS observations are more effective in constraining neutrino masses, a point that will be further discussed in the next section.
+Figure 8: Left panel: The CMB angular power spectrum $C_{l}^{\rm TT}$ for different neutrino degeneracy parameters $\xi_{\nu}$, compared to the result for $\xi_{\nu}=0$.
+Right panel: The linear matter power spectrum $P(k)$ for different neutrino degeneracy parameters $\xi_{\nu}$, compared to the result for $\xi_{\nu}=0$.
+The total neutrino mass is fixed at $\sum m_{\nu}=0.1\,\rm{eV}$ in both panels.
+Figure 9: The sound horizon at the baryon drag epoch, $r_{d}$, as a function of the primordial neutrino asymmetry $\xi_{\nu}$.
 We now present the results for non-zero $\xi_{\nu}$.
-To isolate the effects of neutrino asymmetries from the total neutrino mass, the left panel of Fig.
-8 shows the ratio of the CMB angular power spectrum for $\xi_{\nu}\neq 0$ to $\xi_{\nu}=0$, with $\sum m_{\nu_{i}}=0.1\,\text{eV}$.
-A similar ratio for the matter power spectrum is shown in the right panel.
-As seen in the left panel, the effects of neutrino asymmetries on the CMB angular power spectrum manifest primarily at small scales ($\ell\gtrsim 400$).
-At $\ell\lesssim 400$, $C_{l}^{\rm TT}$ for non-zero $\xi_{\nu}$ matches the results for $\xi_{\nu}=0$.
+To isolate the effects of neutrino asymmetry from the total neutrino mass, the left panel of Fig.
+8 shows the ratio of the CMB angular power spectrum for $\xi_{\nu}\neq 0$ to that for $\xi_{\nu}=0$, with $\sum m_{\nu_{i}}=0.1\,\text{eV}$.
+The corresponding ratio for the matter power spectrum is shown in the right panel.
+As seen in the left panel, the effects of neutrino asymmetry on the CMB angular power spectrum are primarily observed at small scales ($\ell\gtrsim 400$).
+For $\ell\lesssim 400$, $C_{l}^{\rm TT}$ remains nearly identical for both $\xi_{\nu}\neq 0$ and $\xi_{\nu}=0$.
 However, as $\ell$ increases, deviations due to $\xi_{\nu}$ become apparent.
 These deviations depend on both the magnitude and sign of $\xi_{\nu}$: positive $\xi_{\nu}$ enhances the small-scale power, while negative $\xi_{\nu}$ leads to suppression.
-This behavior is largely driven by changes in helium abundance, which alters the CMB spectrum via diffusion damping.
-In addition to overall enhancement or suppression, the acoustic peaks are also affected by $\xi_{\nu}$, as larger $\xi_{\nu}$ values modify the oscillations at the last scattering surface.
-The amplitude of these modifications grows with the absolute value of $\xi_{\nu}$, with negative $\xi_{\nu}$ producing more pronounced effects.
-For the linear matter power spectrum in the right panel of Fig.
-8, we observe that at large scales, the result for non-zero $\xi_{\nu}$ is identical to that for $\xi_{\nu}=0$.
-At small scales, however, the free-streaming of neutrinos is enhanced with larger $\xi_{\nu}$, resulting in a slightly stronger suppression of the matter power spectrum, although this suppression is insensitive to the sign of $\xi_{\nu}$.
-The more intriguing effect of non-zero $\xi_{\nu}$ is the oscillations of the matter power spectrum on the scales $10^{-1}\,h/{\rm Mpc}\gtrsim k\gtrsim 10^{0}\,h/{\rm Mpc}$, which represents the effect of the BAO.
-Similar to the acoustic oscillations in the CMB angular power spectrum, we find that the amplitudes of the BAO in the matter power spectrum are also larger with a larger absolute value of $\xi_{\nu}$, while the negative $\xi_{\nu}$ leads to a larger effect than that for the positive $\xi_{\nu}$.
-In addition to the imprints of BAO on the matter power spectrum, we show in Fig.
-9 the sound horizon at the baryon drag epoch of the early universe $r_{d}$ as a function of the primordial neutrino asymmetry $\xi_{\nu}$.
-Due to the additional contribution to $N_{\rm eff}$ for non-zero $\xi_{\nu}$ according to Eq.
-63, the larger the $\left|\xi_{\nu}\right|$, the smaller the $r_{d}$, resulting in the significant imprints on the matter power spectrum.
-These results imply that BAO experiments can also provide powerful constraints on the primordial neutrino asymmetry, as we will in Section 6, which is often overlooked in previous studies.
-Note also that these conclusions are drawn with a fixed total neutrino mass $\sum m_{\nu_{i}}=0.1\,\text{eV}$.
-However, they remain valid for different total neutrino masses, although the resulting ratios $C_{l}^{\rm TT}/C_{l}^{\rm TT,\,\xi_{\nu}=0}$ and $P(k)/P^{\xi_{\nu}=0}(k)$ may vary slightly at intermediate scales due to the non-relativistic transition of massive neutrinos.
+This behavior is primarily driven by changes in helium abundance, which affect the CMB spectrum through diffusion damping.
+In addition to the overall enhancement or suppression, the acoustic peaks are also influenced by $\xi_{\nu}$, as larger values modify the oscillations at the last scattering surface.
+The amplitude of these modifications increases with the absolute value of $\xi_{\nu}$, with negative $\xi_{\nu}$ producing more pronounced effects.
+For the linear matter power spectrum, shown in the right panel of Fig.
+8, we observe that at large scales, the results for non-zero $\xi_{\nu}$ are identical to those for $\xi_{\nu}=0$.
+At smaller scales, however, the free-streaming of neutrinos is enhanced with increasing $\xi_{\nu}$, resulting in a slightly stronger suppression of the matter power spectrum, while this suppression is largely insensitive to the sign of $\xi_{\nu}$.
+A more notable effect of non-zero $\xi_{\nu}$ is the oscillations in the matter power spectrum on scales $10^{-1}\,h/{\rm Mpc}\gtrsim k\gtrsim 10^{0}\,h/{\rm Mpc}$, which reflect the influence of the BAO.
+Similar to the acoustic oscillations in the CMB angular power spectrum, the amplitude of BAO in the matter power spectrum increases with the absolute value of $\xi_{\nu}$, with negative $\xi_{\nu}$ leading to a larger effect than positive $\xi_{\nu}$.
+In addition to BAO features in the matter power spectrum, Fig.
+9 shows the sound horizon at the baryon drag epoch, $r_{d}$, as a function of the primordial neutrino asymmetry $\xi_{\nu}$.
+Due to the additional contribution to $N_{\rm eff}$ for non-zero $\xi_{\nu}$, as per Eq.
+63, larger $\left|\xi_{\nu}\right|$ corresponds to a smaller $r_{d}$, resulting in significant imprints on the matter power spectrum.
+These results imply that BAO experiments can provide stringent constraints on the primordial neutrino asymmetry, as will be discussed in Section 6, a consideration often overlooked in previous studies.
+Note that these conclusions are drawn with a fixed total neutrino mass of $\sum m_{\nu_{i}}=0.1\,\text{eV}$; however, they remain valid for different total neutrino masses, with only slight variations in the ratios $C_{l}^{\rm TT}/C_{l}^{\rm TT,\,\xi_{\nu}=0}$ and $P(k)/P^{\xi_{\nu}=0}(k)$ at intermediate scales due to the non-relativistic transition of massive neutrinos.
 Figure 10: The CMB angular power spectrum $C_{l}^{\rm TT}$ (left panel) and the linear matter power spectrum $P(k)$ (right panel) for different neutrino degeneracy parameters $\xi_{\nu}$, compared to the corresponding results with $\xi_{\nu}=0$.
-The total neutrino mass is fixed to $\sum m_{\nu}=0.1{\rm eV}$ for both panels.
-The solid curves are the same with Fig.
-5, corresponding to the results including the spectral distortions of the neutrinos and antineutrinos, while the dashed curves correspond to the results without the spectral distortions.
-Finally, we briefly discuss the effects of spectral distortions of the neutrinos and antineutrinos on the CMB and LSS.
-Similar to Fig.
-8, we present the resulting CMB angular power spectrum $C_{l}^{\rm TT}$ and linear matter power spectrum $P(k)$ with non-zero $\xi_{\nu}$, neglecting the neutrino spectral distortions, as the dashed curves in the left and right panels of Fig.
-10.
-The solid curves in Fig.
-10 are the same as those in Fig.
-8, corresponding to the results including the spectral distortions of the neutrinos and antineutrinos.
-Comparing the resulting CMB angular power spectrum $C_{l}^{\rm TT}$ with and without the neutrino spectral distortions, we see that the primary effect of including the spectral distortions is a slight suppression of the first three acoustic peaks, with an almost negligible effect on the remaining peaks in the small scales.
-For the linear power spectrum, we find that the effect of spectral distortions on the BAO is similar with that on the CMB angular power spectrum.
-Here the first two acoustic peaks are slightly enhanced while the effects on the remaining peaks are negligible.
-In addition, we find that the calculations without the neutrino spectral distortions cannot capture the right behavior of the matter power spectrum at small scales, where the results without the spectral distortions predict a near constant suppression at the scales $k\gtrsim 1\,h/{\rm Mpc}$, while the preciser results with the spectral distortions show that the matter power spectrum receives larger suppression at these scales.
-However, the discussion here refers only to the linear matter power spectrum, and we expect that a comprehensive study of the effects of neutrino spectral distortions on the LSS in the nonlinear regime can be done in the future with the N-body simulations.
+The total neutrino mass is fixed to $\sum m_{\nu}=0.1\,\text{eV}$ for both panels.
+The solid curves are identical to those in Fig.
+5, representing the results including spectral distortions of the neutrinos and antineutrinos, while the dashed curves represent results without spectral distortions.
+Finally, we briefly discuss the effects of spectral distortions of neutrinos and antineutrinos on the CMB and LSS.
+As in Fig.
+8, we present in the left and right panels of Fig.
+10 the CMB angular power spectrum $C_{l}^{\rm TT}$ and the linear matter power spectrum $P(k)$ for non-zero $\xi_{\nu}$, compared to the corresponding results with $\xi_{\nu}=0$.
+The solid curves correspond to results that include spectral distortions, as shown in Fig.
+8, while the dashed curves correspond to results without the neutrino spectral distortions.
+Comparing $C_{l}^{\rm TT}$ with and without spectral distortions, we see that the primary effect of including these distortions is a slight suppression of the first three acoustic peaks, with negligible impact on the smaller-scale peaks.
+For the linear matter power spectrum, the effect on the BAO is similar to that on the CMB spectrum, with the first two acoustic peaks slightly enhanced while the effects on subsequent peaks remain minimal.
+Furthermore, we find that calculations that exclude spectral distortions fail to accurately capture the behavior of the matter power spectrum at small scales.
+Without distortions, the results predict a near-constant suppression at $k\gtrsim 1\,h/{\rm Mpc}$, whereas more precise calculations including distortions show greater suppression at these scales.
+However, this analysis is limited to the linear matter power spectrum.
+A comprehensive study of the impact of neutrino spectral distortions on LSS in the nonlinear regime would require future N-body simulations, which is beyond the scope of this paper and left for future work.
 6 Combined Analysis of Primordial Neutrino Asymmetries
-In previous sections, we have shown that the primordial neutrino asymmetry significantly influences the neutrino decoupling process, leaving significant observable imprints on the Universe, such as the abundance of light elements produced during the BBN, the CMB anisotropies, and the BAO imprints in the LSS.
-These cosmological observations can then be used to constrain the primordial neutrino asymmetries.
-In this section, we consolidate the results of the previous sections and discuss the constraints on the primordial neutrino asymmetry derived from contemporary observations of the BBN, CMB and BAO.
-In addition, we briefly explore the corresponding constraints on UV model parameters using a specific model that generates primordial neutrino asymmetries in the early universe.
+In previous sections, we have demonstrated that primordial neutrino asymmetry significantly influences the neutrino decoupling process, leaving observable imprints on the Universe, such as the abundance of light elements produced during BBN, the anisotropies in the CMB, and the BAO features in the LSS.
+These cosmological observations can be used to constrain primordial neutrino asymmetries.
+In this section, we consolidate the results from previous sections and discuss the constraints on primordial neutrino asymmetries derived from present BBN, CMB, and BAO observations.
+Furthermore, we briefly explore the associated constraints on UV model parameters by using a specific model that generates primordial neutrino asymmetries in the early universe.
 6.1 Observational Constraints on Primordial Neutrino Asymmetry
@@
 $\mathbf{\Lambda}$CDM + $\mathbf{\xi_{\nu}}$ + $\mathbf{m_{\nu_{1}}}$ $0.024\pm 0.012$ $2.200\pm 0.011$ $0.1205\pm 0.0009$ 0.131
-Table 3: Summary of constraints on the primordial lepton asymmetry $\xi_{\nu}$, the baryon abundance $\omega_{b}$, the cold dark matter abundance $\omega_{cdm}$ and the total neutrino mass $\sum m_{\nu}$ from various datasets and cosmological models.
-In this work we utilize the following datasets for BBN, CMB, and BAO in our analysis:
-BBN
-The EMPRESS measurement of helium abundance, ${Y\rm_{P}}=0.2370^{+0.0034}_{-0.0033}$ empress , and the deuterium abundance adopted by the PDG group, $10^{5}\times{\rm D/H}=2.547\pm 0.025$ pdg , which is based on the results of Cooke:2017cwo .
-CMB
+Table 3: Summary of constraints on the primordial lepton asymmetry $\xi_{\nu}$, the baryon density $\omega_{b}$, the cold dark matter density $\omega_{cdm}$, and the total neutrino mass $\sum m_{\nu}$ in normal hierarchy from various datasets and cosmological models.
+The datasets used for BBN, CMB, and BAO analyses in this work are as follows:
+BBN:
+The EMPRESS measurement of helium abundance, ${Y\rm_{P}}=0.2370^{+0.0034}_{-0.0033}$ empress , and the deuterium abundance from the PDG, $10^{5}\times{\rm D/H}=2.547\pm 0.025$ pdg , based on Cooke:2017cwo .
+CMB:
 The Planck 2018 low-$l$ and high-$l$ TT, TE, and EE spectra, along with the reconstructed CMB lensing spectrum Planck:2018vyg ; Planck:2018lbu ; Planck:2019nip .
-BAO
+BAO:
 The BOSS DR12 BOSS:2016wmc with its full covariance matrix for BAO, as well as the 6dFGS Beutler:2011hx and MGS of SDSS Ross:2014qpa .
-We note that the helium abundance from the BBN data set primarily constrains the primordial neutrino asymmetries, while the deuterium abundance is more sensitive to the baryon density $\omega_{b}$.
-The BAO dataset is mainly used to constrain the total neutrino mass $\sum m_{\nu}$, but as we will discuss, it also places additional constraints on primordial neutrino asymmetries.
-To extract the constraints from these datasets, we conduct MCMC analyses using the Montepython engine Audren:2012wb ; Brinckmann:2018cvx .
-For analysis, the $N_{\rm eff}$ is adopted as Eq.
-(63), derived in Section 3.
-The helium and deuterium abundances predicted by BBN are computed using our fitting formulas in Equation (85).
+It should be noted that the helium abundance from the BBN dataset primarily constrains primordial neutrino asymmetries, while the deuterium abundance is more sensitive to the baryon density $\omega_{b}$.
+The BAO dataset primarily constrains the total neutrino mass $\sum m_{\nu}$, but also provides additional constraints on primordial neutrino asymmetries.
+To extract constraints from these datasets, we perform MCMC analyses using the Montepython engine Audren:2012wb ; Brinckmann:2018cvx .
+For the analysis, $N_{\rm eff}$ is derived from Eq.
+(63) as discussed in Section 3.
+The predicted helium and deuterium abundances from BBN are calculated using our fitting formulas in Eq.
+(86).
 As shown in Fig.
-10, the effects of neutrino spectral distortions on the CMB angular power spectrum and the BAO in the matter power spectrum are small.
-Therefore, for simplicity, the (anti)neutrino spectral distortions are neglected in the MCMC analysis.
-The parameter set for the CMB and BAO datasets includes the six independent parameters of the $\Lambda$CDM model, along with the neutrino degeneracy parameter $\xi_{\nu}$ and the mass of the lightest neutrino for the normal hierarchy $m_{\nu_{1}}$, i.e., $\{\omega_{b},\omega_{cdm},\theta_{s},A_{s},n_{s},\tau_{\mathrm{reio}},\xi_{\nu},m_{\nu_{1}}\}$, where $\theta_{s}$ represents the sound horizon, $A_{s}$ is the amplitude of the primordial curvature perturbations at the pivot scale $k=0.05\,\mathrm{Mpc}^{-1}$, $n_{s}$ is the scalar spectral index, and $\tau_{\mathrm{reio}}$ is the optical depth to reionization.
-On the other hand, we only consider $\{\omega_{b},\xi_{\nu}\}$ for the BBN dataset, as the predicted light element abundances from BBN are largely independent of other cosmological parameters within the relevant range.
-Figure 11: 2D and 1D marginalised posterior distributions for the baryon abundance $\omega_{b}$, the neutrino degeneracy parameter $\xi_{\nu}$, the cold dark matter abundance $\omega_{cdm}$, the mass of the lightest neutrino for the normal hierarchy $m_{\nu_{1}}$, and scalar spectral index $n_{s}$.
-The red, blue, and green regions represent constraints from BBN, CMB + BAO, and their combination, respectively.
-Figure 12: 1D marginalised distributions for different parameter sets derived from the BBN + CMB + BAO datasets.
-With the aforementioned choices of parameters and datasets, we modified the MontePython sampler and CLASS codes to perform the MCMC analysis.
-The Gelman-Rubin criterion for all chains converged to $R-1<0.01$.
-The resulting $1\,\sigma$ cosmological constraints are summarized in Table 3, where the $95\%$ confidence limits on the total neutrino mas are also presented from the constraints on the mass of the lightest neutrino for the normal hierarchy $m_{\nu_{1}}$.
-For the same model $\Lambda$CDM + ${\xi_{\nu}}$ + ${m_{\nu_{1}}}$, we present the 1D and 2D marginalized posterior distributions of key parameters that show significant variations across different datasets in Fig.
+10, the impact of neutrino spectral distortions on the CMB angular power spectrum and BAO in the matter power spectrum is minor.
+Therefore, for simplicity, (anti)neutrino spectral distortions are neglected in the MCMC analysis.
+For the CMB and BAO datasets, the parameter set includes the six independent parameters of the $\Lambda$CDM model, along with the neutrino degeneracy parameter $\xi_{\nu}$ and the mass of the lightest neutrino in the normal hierarchy $m_{\nu_{1}}$, i.e., $\{\omega_{b},\omega_{cdm},\theta_{s},A_{s},n_{s},\tau_{\mathrm{reio}},\xi_{\nu},m_{\nu_{1}}\}$.
+Here, $\theta_{s}$ is the sound horizon, $A_{s}$ is the amplitude of primordial curvature perturbations at the pivot scale $k=0.05\,\mathrm{Mpc}^{-1}$, $n_{s}$ is the scalar spectral index, and $\tau_{\mathrm{reio}}$ is the optical depth to reionization.
+For the BBN dataset, we consider only $\{\omega_{b},\xi_{\nu}\}$, as the predicted light element abundances from BBN are largely independent of other cosmological parameters within the relevant range.
+Figure 11: 2D and 1D marginalized posterior distributions for the baryon abundance $\omega_{b}$, the neutrino degeneracy parameter $\xi_{\nu}$, the cold dark matter abundance $\omega_{cdm}$, the scalar spectral index $n_{s}$, and the mass of the lightest neutrino for the normal hierarchy $m_{\nu_{1}}$.
+The green, blue and red regions represent constraints from BBN, CMB + BAO, and their combination, respectively.
+Figure 12: 1D marginalized distributions for different parameter sets derived from the BBN + CMB + BAO datasets.
+Using the specified parameters and datasets, we modified the MontePython sampler and CLASS codes to perform the MCMC analysis.
+All chains converged with a Gelman-Rubin criterion of $R-1<0.01$.
+The resulting $1\,\sigma$ cosmological constraints are summarized in Table 3, where the $95\%$ confidence limits on the total neutrino mass are also presented, derived from constraints on the mass of the lightest neutrino for the normal hierarchy, $m_{\nu_{1}}$.
+For the model $\Lambda$CDM + ${\xi_{\nu}}$ + ${m_{\nu_{1}}}$, we present the 1D and 2D marginalized posterior distributions of key parameters that show significant variations across different datasets in Fig.
 11, including the baryon abundance $\omega_{b}$, the neutrino degeneracy parameter $\xi_{\nu}$, the cold dark matter abundance $\omega_{cdm}$, the sound horizon $\theta_{s}$, and the mass of the lightest neutrino for the normal hierarchy $m_{\nu_{1}}$.
-Concerning the constraints on the primordial neutrino asymmetries, we see that the recent EMPRESS data in the BBN dataset points to a positive primordial neutrino asymmetry $\xi_{\nu}=0.039\pm 0.013$, which is different from zero with a $3.0\,\sigma$ significance.
-Combining the BBN and CMB data, the resulting constraints on $\xi_{\nu}$ are significantly reduced to $\xi_{\nu}=0.028\pm 0.012$, different from zero with a $2.3\,\sigma$ significance.
-Compared to the constraints $\xi_{\nu}=0.034\pm 0.014$ reported in previous study with the same BBN and CMB datasets Escudero:2022okz , our result indicate a $\sim 18\%$ smaller central value and a slight reduction in the uncertainty.
-Incorporating BOSS BAO observations in the MCMC analysis, the primordial neutrino asymmetry constraint further reduce to $\xi_{\nu}=0.024\pm 0.013$, indicating a $\sim 29\%$ reduction in the central value compared to previous results, while the significance of a positive neutrino asymmetry is reduced to $2.0\,\sigma$.
-Our study demonstrate the additional constraining power of the BAO experiments on the primordial neutrino asymmetry, which are not aware in previous studies.
-We find that the constraints on other parameters also vary significantly with different datasets, especially for $\omega_{b}$, $\omega_{cmb}$, $n_{s}$ and $m_{\nu_{1}}$, as shown in Fig.
+Regarding constraints on primordial neutrino asymmetries, the EMPRESS data in the BBN dataset indicates a positive primordial neutrino asymmetry, $\xi_{\nu}=0.039\pm 0.013$, which differs from zero at $3.0\,\sigma$ significance.
+Combining BBN and CMB data significantly reduces the constraint to $\xi_{\nu}=0.028\pm 0.012$, which differs from zero at $2.3\,\sigma$ significance.
+Compared to previous results ($\xi_{\nu}=0.034\pm 0.014$) using the same BBN and CMB datasets Escudero:2022okz , our results show an 18% reduction in the central value and a slight decrease in uncertainty.
+Incorporating the BOSS BAO observations further reduces the constraint on primordial neutrino asymmetry to $\xi_{\nu}=0.024\pm 0.012$, representing a 29% reduction in the central value relative to previous findings, while the significance of a positive neutrino asymmetry decreases to $2.0\,\sigma$.
+Our analysis highlights the additional constraining power of BAO experiments on primordial neutrino asymmetries, which was not fully explored in previous studies.
+We find that constraints on other parameters, such as $\omega_{b}$, $\omega_{cdm}$, $n_{s}$, and $m_{\nu_{1}}$, also vary significantly depending on the dataset used, as shown in Fig.
 11.
-For the constraints on $\omega_{b}$, the difference is largely due to the fact that we chose the nuclear reaction rates in PRIMAT in the process of BBN, which predicted a deuterium abundance lower than the current constraints from PDG, thus leading to significantly lower constraints on $\omega_{b}$ Pitrou:2018cgg ; Burns:2022hkq ; Escudero:2022okz .
-While the constraints on the primordial neutrino asymmetry $\xi_{\nu}$ do not strongly depend on the choice of nuclear reaction rates, as discussed in  Escudero:2022okz , a detailed study with different nuclear reaction rates will be performed in our companion paper.
-The inclusion of the BOSS BAO data significantly changes the corresponding constraints on $\omega_{cmb}$, $n_{s}$ and $m_{\nu_{1}}$ due to its sensitivity to the evolution of the late universe.
-Especially, the resulting $95\%$ upper limit for the total mass of neutrinos are greatly tightened from $0.414$ eV by BBN + CMB to $0.131$ eV by BBN + CMB + BAO, since the BAO data help to break the CMB degeneracy.
-Recently, the BAO observations of DESI have put an even stronger bound on neutrino masses $\sum m_{\nu_{i}}\lesssim 0.07\,eV$ DESI:2024mwx .
-The relevant studies on primordial neutrino asymmetry and neutrino masses with DESI BAO data will also be discussed in our companion paper.
-Finally we briefly discuss the effects of different parameter sets on the resulting constraints, especially the interplay between the primordial neutrino asymmetry and the total mass of neutrinos.
-Using the data set of BBN + CMB + BAO, we present in Fig.
-12 the resulting 1d marginalised distributions for three choices of the parameter set, first is the set of $\Lambda CDM+m_{\nu_{1}}$, second is the $\Lambda CDM+\xi_{\nu}$, while the third is $\Lambda CDM+m_{\nu_{1}}+\xi_{\nu}$.
-As we can see, while the different choices of the parameter set lead to quite modifications to the constraints on the six parameters of $\Lambda CDM$, especially for $\omega_{b}$ and $n_{s}$, the resulting constraints on the additional parameters are almost the same for different parameter sets.
-The constraints on the total mass of neutrinos are only slightly tightened while including primordial neutrino asymmetry $\xi_{\nu}$ as a free parameter, while the constraints on $\xi_{\nu}$ are almost the mass as the inclusion of free $\sum m_{\nu}$.
-Therefore, we conclude that the constraints on neutrino masses and the primordial neutrino asymmetries can indeed be analysed separately if one do not concern the constraints on the other parameters in $\Lambda CDM$, otherwise the parameter sets with both the neutrino masses and the primordial neutrino asymmetries must be considered.
+For $\omega_{b}$, the discrepancy is mainly attributed to the choice of nuclear reaction rates in the PRIMAT code during the BBN process, which predicts a deuterium abundance lower than current constraints from PDG, leading to a significantly lower $\omega_{b}$ Pitrou:2018cgg ; Burns:2022hkq ; Escudero:2022okz .
+While constraints on the primordial neutrino asymmetry $\xi_{\nu}$ do not strongly depend on the choice of nuclear reaction rates, as discussed in Escudero:2022okz , a detailed study using different nuclear reaction rates will be presented in our companion paper.
+The inclusion of BOSS BAO data significantly modifies the constraints on $\omega_{cdm}$, $n_{s}$, and $m_{\nu_{1}}$ due to its sensitivity to the evolution of the late Universe.
+In particular, the $95\%$ upper limit on the total neutrino mass tightens from $0.414$ eV (BBN + CMB) to $0.131$ eV (BBN + CMB + BAO), as BAO data help break the CMB degeneracy.
+Recently, DESI BAO observations have placed even stronger constraints on neutrino masses, $\sum m_{\nu_{i}}\lesssim 0.07\,\mathrm{eV}$ DESI:2024mwx .
+Relevant studies on primordial neutrino asymmetry and neutrino masses using DESI BAO data will be addressed in our companion paper.
+Finally, we briefly discuss the impact of different parameter sets on the resulting constraints, particularly the interplay between primordial neutrino asymmetry and total neutrino mass.
+Using the BBN + CMB + BAO dataset, we present in Fig.
+12 the 1D marginalized distributions for three parameter sets: $\Lambda$CDM + $m_{\nu_{1}}$, $\Lambda$CDM + $\xi_{\nu}$, and $\Lambda$CDM + $m_{\nu_{1}}$ + $\xi_{\nu}$.
+As shown, while different parameter sets lead to significant modifications in constraints on the six $\Lambda$CDM parameters, especially $\omega_{b}$ and $n_{s}$, the resulting constraints on additional parameters are consistent across different parameter sets.
+The corresponding constraints from different parameter sets are also summarized in Table 3.
+We find that the constraints on the total neutrino mass are only slightly tightened when $\xi_{\nu}$ is included as a free parameter, while constraints on $\xi_{\nu}$ remain nearly unchanged when $\sum m_{\nu}$ is a free parameter.
+Therefore, we conclude that constraints on neutrino masses and primordial neutrino asymmetries can be analyzed independently if one is not concerned with constraints on other $\Lambda$CDM parameters.
+However, if constraints on other $\Lambda$CDM parameters are of interest, a parameter set including both of neutrino mass and primordial neutrino asymmetry must be considered.
 6.2 Constraints on UV Model Parameters
-With the constraints on the primordial neutrino asymmetries, we discuss in this subsection the corresponding constraints on the UV model parameters with a specific model to generate the primordial neutrino asymmetries.
-Here we consider a Q-ball decay model Kawasaki:2022hvx , where Q-balls are produced in the Affleck-Dine mechanism Kusenko:1997si ; Enqvist:1997si ; Kasuya:1999wu , which utilizes flat directions in the minimal supersymmetric standard model (MSSM).
-When some flat direction with lepton number has a large field value during inflation, it generates lepton number by field dynamics after inflation.
-At the same time, Q-balls are produced and the generated lepton number is all confined inside the Q-balls.
-Thus, such Q-balls have a large lepton number and are called L-balls.
-Since the sphaleron effect is inactive inside the L-balls, the lepton asymmetry produced by the Affleck-Dine mechanism are kept without being converted to a baryon asymmetry Kawasaki:2002hq ; Gelmini:2020ekg .
-The L-balls decay into neutrinos below the electroweak scale and realize a large lepton asymmetry in the universe.
-Following the results of Kawasaki:2022hvx , the primordial lepton asymmetry $\eta_{L}$ is defined as the ratio between the difference of the number density of neutrinos over antineutrinos and the photon number density,
+In this subsection, we discuss the constraints on UV model parameters based on the constraints on primordial neutrino asymmetries.
+Specifically, we consider a Q-ball decay model Kawasaki:2022hvx , where Q-balls are produced through the Affleck-Dine mechanism Kusenko:1997si ; Enqvist:1997si ; Kasuya:1999wu , which operates along flat directions in the minimal supersymmetric standard model (MSSM).
+When a flat direction associated with lepton number reaches a large field value during inflation, it generates a lepton number via field dynamics after inflation.
+At the same time, Q-balls form, and the generated lepton number is confined within them.
+These Q-balls, referred to as L-balls, possess a large lepton number.
+Since the sphaleron process is inactive inside L-balls, the lepton asymmetry generated by the Affleck-Dine mechanism is preserved without conversion to a baryon asymmetry Kawasaki:2002hq ; Gelmini:2020ekg .
+After the electroweak scale, the L-balls decay into neutrinos, resulting in a significant lepton asymmetry in the universe.
+Following Kawasaki:2022hvx , the primordial lepton asymmetry $\eta_{L}$ is defined as the ratio of the difference between the number densities of neutrinos and antineutrinos to the photon number density:
 $\eta_{L}\equiv\sum_{\alpha}\frac{n_{\nu_{\alpha}}-n_{\bar{\nu}_{\alpha}}}{s}\simeq 0.105\,\xi_{\nu},$ (102)
-where $s=\frac{2\pi^{2}}{45}g_{\star}T_{\gamma}^{3}$ is the entropy density and $g_{\star}=10.75$ is the relativistic degree of freedom at the neutrino decoupling epoch.
-In addition, the resulting $\eta_{L}$ in the Q-ball decay model can be expressed as
+where $s=\frac{2\pi^{2}}{45}g_{\star}T_{\gamma}^{3}$ is the entropy density, and $g_{\star}=10.75$ is the relativistic degree of freedom at neutrino decoupling.
+In the Q-ball decay model, $\eta_{L}$ can also be expressed as:
 $\eta_{L}=\frac{3T_{\mathrm{D}}}{4m_{3/2}},$ (103)
-where $m_{3/2}$ is the the gravitino mass, $T_{\mathrm{D}}$ is the decay temperature of the L-balls and can be expressed with the other UV model parameters with
+where $m_{3/2}$ is the gravitino mass and $T_{\mathrm{D}}$ is the L-balls’ decay temperature, which is a function of other UV model parameters:
 $T_{\mathrm{D}}$ $\simeq 2.69~{}\mathrm{MeV}\,$
 ×$\hskip 15.0pt\times\left(\frac{g_{*}}{10.75}\right)^{-1/4}\left(\frac{\beta}{6\times 10^{-4}}\right)^{-5/8}\left(\frac{m_{3/2}}{0.5~{}\mathrm{GeV}}\right)^{5/2}\left(\frac{M_{F}}{5\times 10^{6}~{}\mathrm{GeV}}\right)^{-2}\left(\frac{N_{\ell}}{3}\right)^{1/2}\left(\frac{\zeta}{2.5}\right)^{1/2},$ (104)
-where $g_{*}(T_{\mathrm{D}})$ is the relativistic number of degrees of freedom at $T_{\mathrm{D}}$, $\beta\simeq 6\times 10^{-4}$ is a dimensionless constant, $M_{F}$ is the SUSY breaking scale, $N_{\ell}$ is the number of decay channels.
-Therefore, once we have obtained the constraints on the lepton asymmetry from cosmological observations, they can be directly translated into constraints on the model parameters of the UV model.
-According to our study, the constraints on $\eta_{L}$ from the BBN + CMB + BAO datasets are $\eta_{L}=0.0025\pm 0.0013$. Therefore, the constraint on the L-balls decay temperature is
+where $g_{*}(T_{\mathrm{D}})$ is the relativistic degrees of freedom at $T_{\mathrm{D}}$, $\beta\simeq 6\times 10^{-4}$ is a dimensionless constant, $M_{F}$ is the SUSY breaking scale, and $N_{\ell}$ represents the number of decay channels.
+Hence, the cosmological constraints on lepton asymmetry can be translated into constraints on the UV model parameters.
+Our analysis of the BBN + CMB + BAO datasets yields a constraint on the lepton asymmetry of $\eta_{L}=0.0025\pm 0.0013$. The corresponding constraint on the L-balls’ decay temperature is:
 $T_{\mathrm{D}}$ $=\frac{4}{3}\eta_{L}m_{3/2}\simeq 3.3~{}\mathrm{MeV}\,\left(\frac{m_{3/2}}{0.5~{}\mathrm{GeV}}\right)\left(\frac{\eta_{L}}{5\times 10^{-3}}\right)$
 […]
-and conversely,
+and, conversely,
 $m_{3/2}$ $=\frac{3T_{\mathrm{D}}}{4T\eta_{L}}\simeq 0.5~{}\mathrm{GeV}\,\left(\frac{T_{\mathrm{D}}}{3.3~{}\mathrm{MeV}}\right)\left(\frac{5\times 10^{-3}}{\eta_{L}}\right)$
 […]
-These constraints can also be converted into the constraints on other UV model parameters, such as the SUSY breaking scale, as seen in the preliminary study of Kawasaki:2022hvx .
+These constraints can be further translated into constraints on other UV model parameters, such as the SUSY breaking scale, as shown in Kawasaki:2022hvx .
 7 Conclusion
-The presence of primordial neutrino asymmetries before the neutrino decoupling epoch has significant implications for both the neutrino decoupling process and subsequent cosmological phenomena, and can therefore be constrained by observations of BBN, CMB and LSS.
-However, the effects of primordial neutrino asymmetries on neutrino decoupling—particularly concerning the resulting $N_{\rm eff}$ and the spectral distortions of (anti)neutrinos, including the impacts of flavor oscillations and FTQED corrections—remain ambiguous.
-These ambiguities affect the interpretation of subsequent processes.
-Consequently, studies of primordial neutrino asymmetries in relation to BBN and the CMB often employ approximations, such as using thermal Fermi-Dirac distributions for (anti)neutrinos, without thorough scrutiny.
-In this work, we aim to address these ambiguities by numerically solving the neutrino decoupling process and analyzing the implications of the resulting (anti)neutrino density matrices for BBN, CMB, and LSS.
-We begin by re-deriving the quantum kinetic equations (QKEs) for neutrinos and antineutrinos using the Closed Time Path (CTP) formalism for two-point functions.
-Focusing on primordial neutrino asymmetries with complete flavor equilibration, where $\xi_{\nu_{e}}=\xi_{\nu_{\mu}}=\xi_{\nu_{\tau}}=\xi_{\nu}$, we numerically solve the QKEs in the early universe using the public code FortEPiaNO Gariazzo:2019gyi ; Bennett:2020zkv , including the effects of neutrino oscillations, matter effects, FTQED corrections up to $\mathcal{O}(3)$, and comprehensive neutrino-electron and neutrino-neutrino collision terms.
-The final $N_{\rm eff}$ is then evaluated based on the resulting density matrices, and can be expressed as $N_{\rm eff}\simeq N_{\rm eff}^{\rm SM}+3\left(\frac{30}{7\pi^{2}}\xi_{\nu}^{2}+\frac{15}{7\pi^{4}}\xi_{\nu}^{4}\right)+0.0102\xi_{\nu}^{2}$, where $N_{\rm eff}^{\rm SM}=3.0440\pm 0.0002$, while the second and last term represent the corrections from nonzero $\xi_{\nu}$ in the instantaneous decoupling limit and from the non-instantaneous effect, respectively.
-We also confirm that the spectral distortions of neutrinos and antineutrinos with non-zero $\xi_{\nu}$ differ due to distinct Pauli-blocking effects, in agreement with the priliminary study in Grohs:2016cuu , where neutrino spectral distortions were evaluated while neglecting flavor oscillations and FTQED corrections.
-Using the density matrices for neutrinos and antineutrinos, we further investigate the implications of primordial neutrino asymmetries for BBN, CMB and LSS.
-For BBN, we introduce a novel correction to the weak interaction rates arising from the distinct spectral distortions of neutrinos and antineutrinos.
-The light element abundances are then computed with the public BBN code PRIMAT Pitrou:2018cgg , incorporating the other additional corrections on weak rates.
-According to the the EMPRESS measurement on the helium abundance $Y_{P}|_{\rm EMPRESS}=0.2370^{+0.0034}_{-0.0033}$, our results immediately indicate a positive primordial neutrino asymmetry $0.032\leq\xi_{\nu}\leq 0.052$ with fixed baryon abundance $\omega_{b}$.
-The effects of the new proposed correction are also examined, and we find that the spectral distortions for neutrinos and antineutrinos are important for precise predictions for BBN, especially for the abundance of Helium-4.
-Compared with the results in the instantaneous limit, the traditional treatment with Fermi-Dirac distributions predict a negative relative difference for a positive primordial neutrino asymmetry, while our precise treatment with spectral distortions show that the relative difference is still positive.
-Additionally, we have provided semi-analytic expressions for these abundances as functions of $\xi_{\nu}$, which will be useful for future studies.
-For the CMB and LSS, a novel finding of our study is that primordial neutrino asymmetries also affect the amplitudes of the acoustic oscillations, both for the CMB and for baryon acoustic oscillations (BAO).
-This shows that BAO observations could serve as a tool for constraining primordial neutrino asymmetries.
-Based on our systematic understanding of primordial asymmetries, we perform a joint MCMC analysis including the neutrino degeneracy parameter $\xi_{\nu}$ and the total neutrino mass $\sum m_{\nu}$, using data from EMPRESS BBN, Planck CMB, and BOSS BAO observations.
-Combining the EMPRESS BBN and Planck CMB datasets, the primordial neutrino asymmetry constraint yields $\xi_{\nu}=0.028\pm 0.013$, indicating a $\sim 18\%$ smaller central value and a slight reduction in the uncertainty compared to previous results $\xi_{\nu}=0.034\pm 0.014$ obtained in Escudero:2022okz .
-Incorporating BOSS BAO observations in the MCMC analysis, the primordial neutrino asymmetry constraint further reduce to $\xi_{\nu}=0.024\pm 0.013$, indicating a $\sim 29\%$ reduction in the central value compared to previous results and a $2.0\,\sigma$ significance of a positive neutrino asymmetry.
-We also discuss the impacts of treating the sum of neutrino mass $\sum m_{\nu}$ and the primordial neutrino asymmetry $\xi_{\nu}$ as free parameters on the resulting constraints for each of them.
-Finally, we also show how the constraints on the neutrino asymmetries can be translated into constraints on the UV model parameters with a specific model in which the large neutrino asymmetries are generated by the decay of Q-balls Kawasaki:2002hq ; Gelmini:2020ekg .
-Our work provides the framework to obtain the implications of new physics beyond the SM and $\Lambda$CDM model for neutrino decoupling and subsequent processes, and the corresponding cosmological constraints from present and future BBN, CMB and LSS observations.
-Thus, our work facilitates a more comprehensive analysis for primordial neutrino asymmetry and other models with new physics, such as precise consideration of the neutrino synchronous oscillations Froustey:2021azz ; Froustey:2024mgf or neutrino non-standard interactions Du:2021idh ; Du:2023upj , based on more and more precise cosmological observations in the future.
+The existence of primordial neutrino asymmetries prior to the neutrino decoupling epoch has significant implications for the neutrino decoupling process and subsequent cosmological phenomena.
+These asymmetries can therefore be constrained through observations of BBN, CMB, and LSS.
+However, the effects of primordial neutrino asymmetries on neutrino decoupling, particularly with respect to the resulting $N_{\rm eff}$ and spectral distortions of (anti)neutrinos—including the impacts of flavor oscillations and FTQED corrections-remain uncertain.
+These uncertainties make it difficult to interpret subsequent cosmological processes accurately.
+Consequently, studies of primordial neutrino asymmetries often rely on approximations, such as the use of thermal Fermi-Dirac distributions for (anti)neutrinos, without rigorous verification.
+In this work, we address these ambiguities by numerically solving the neutrino decoupling process and analyzing the resulting (anti)neutrino density matrices in the context of BBN, CMB, and LSS.
+We rederive the quantum kinetic equations (QKEs) for neutrinos and antineutrinos using the Closed Time Path (CTP) formalism for two-point functions.
+Assuming complete flavor equilibration before neutrino decoupling, where $\xi_{\nu_{e}}=\xi_{\nu_{\mu}}=\xi_{\nu_{\tau}}=\xi_{\nu}$, we numerically solve the QKEs in the early universe using an extended version of the public code FortEPiaNO Gariazzo:2019gyi ; Bennett:2020zkv .
+Our calculations include the effects of neutrino oscillations, matter interactions, FTQED corrections up to $\mathcal{O}(3)$, and comprehensive neutrino-electron and neutrino-neutrino collision terms.
+The final $N_{\rm eff}$ is then computed based on the resulting density matrices and can be expressed as $N_{\rm eff}\simeq N_{\rm eff}^{\rm SM}+3\left(\frac{30}{7\pi^{2}}\xi_{\nu}^{2}+\frac{15}{7\pi^{4}}\xi_{\nu}^{4}\right)+0.0102\xi_{\nu}^{2}$, where $N_{\rm eff}^{\rm SM}=3.0440\pm 0.0002$.
+The second and third terms represent corrections from nonzero $\xi_{\nu}$ in the instantaneous decoupling limit and non-instantaneous effects, respectively.
+We also confirm that spectral distortions of neutrinos and antineutrinos with nonzero $\xi_{\nu}$ differ due to distinct Pauli-blocking effects, in agreement with preliminary studies Grohs:2016cuu .
+Using the neutrino and antineutrino density matrices, we further investigate the effects of primordial neutrino asymmetries on BBN, CMB, and LSS.
+For BBN, we introduce a novel correction to the weak interaction rates arising from distinct neutrino and antineutrino spectral distortions.
+We calculate the light element abundances using the public BBN code PRIMAT Pitrou:2018cgg , incorporating additional corrections to weak rates.
+According to the EMPRESS measurement of the helium abundance $Y_{P}|_{\rm EMPRESS}=0.2370^{+0.0034}_{-0.0033}$, our results suggest a positive primordial neutrino asymmetry in the range $0.032\leq\xi_{\nu}\leq 0.052$, assuming a fixed baryon abundance $\omega_{b}=0.024$.
+Our analysis also highlights the importance of spectral distortions for precise BBN predictions, particularly for Helium-4 abundance.
+Compared to results in the instantaneous limit, the traditional treatment using Fermi-Dirac distributions predicts a negative relative difference for positive primordial neutrino asymmetry, whereas our precise treatment with spectral distortions shows a positive relative difference.
+Additionally, we provide semi-analytic expressions for these abundances as functions of $\xi_{\nu}$, which will be valuable for future studies.
+For CMB and LSS, our study reveals that primordial neutrino asymmetries influence the amplitudes of acoustic oscillations in both the CMB and baryon acoustic oscillations (BAO), suggesting that BAO observations can serve as a tool for constraining these asymmetries.
+To further constrain primordial asymmetries, we perform a joint MCMC analysis including the neutrino degeneracy parameter $\xi_{\nu}$ and the total neutrino mass $\sum m_{\nu}$, using data from EMPRESS BBN, Planck CMB, and BOSS BAO observations.
+Combining the EMPRESS BBN and Planck CMB datasets, we constrain the primordial neutrino asymmetry to $\xi_{\nu}=0.028\pm 0.012$, a reduction of $\sim 18\%$ in the central value and a slight reduction in the uncertainty compared to previous results of $\xi_{\nu}=0.034\pm 0.014$ Escudero:2022okz .
+Incorporating BOSS BAO observations further refines the constraint to $\xi_{\nu}=0.024\pm 0.012$, yielding a $\sim 29\%$ reduction in the central value and a $2.0\sigma$ significance for a positive neutrino asymmetry.
+We also discuss the interplay between $\sum m_{\nu}$ and $\xi_{\nu}$ when both are treated as free parameters, and how these constraints translate into bounds on UV model parameters, using a specific model where large neutrino asymmetries are generated by Q-ball decay Kawasaki:2002hq ; Gelmini:2020ekg .
+Our work establishes a framework for exploring the implications of physics beyond the Standard Model and $\Lambda$CDM for neutrino decoupling and related cosmological processes, providing corresponding constraints from current and future BBN, CMB, and LSS observations.
+It paves the way for more precise studies of primordial neutrino asymmetries and other new physics scenarios, such as synchronous neutrino oscillations Froustey:2021azz ; Froustey:2024mgf and non-standard neutrino interactions Du:2021idh ; Du:2023upj , driven by increasingly precise cosmological observations.
 Acknowledgements.
+We thank J.
+Froustey and C.
+Pitrou for helpful communications about their relevant work.
 This work is supported by the National Science Foundation of China under Grants No.

### v3

@@
-Chinese Academy of Sciences, Beijing 100190, China
+Primordial lepton asymmetries: neutrino transport, spectral distortions and cosmological constraints
 Abstract
-The primordial neutrino asymmetry leave profound imprints on the evolution history of the universe, which can be constrained by cosmological observations, including Big Bang Nucleosynthesis (BBN), Cosmic Microwave Background (CMB), and Large-Scale Structure (LSS).
-We present comprehensive analysis on implications and constraints of the primordial neutrino asymmetry $\xi_{\nu}$, based on a precise treatment of neutrino decoupling by solving the complete (anti)neutrino quantum kinetic equations in the Closed-Time-Path formalism.
-The effective number of neutrinos $N_{\rm eff}$ and (anti)neutrino spectral distortions are calculated, and we find that the non-instantaneous decoupling correction is given by $\delta N_{\rm eff}=0.0440+0.0102\,\xi_{\nu}^{2}$.
+The primordial neutrino asymmetry leaves profound imprints on the evolution history of the universe, which can be constrained by cosmological observations, including Big Bang Nucleosynthesis (BBN), Cosmic Microwave Background (CMB), and Large-Scale Structure (LSS).
+We present comprehensive analysis on implications and constraints of the primordial neutrino asymmetry, based on a precise treatment of neutrino decoupling by solving the complete (anti)neutrino quantum kinetic equations in the Closed-Time-Path formalism.
+Assuming the same primordial asymmetry $\xi_{\nu}$ for all neutrino flavors, the effective number of neutrinos $N_{\rm eff}$ and (anti)neutrino spectral distortions are calculated, and we find that the non-instantaneous decoupling correction is given by $\delta N_{\rm eff}=0.0440+0.0102\,\xi_{\nu}^{2}$.
 Then we perform the state-of-the-art calculation for the abundance of light elements including (anti)neutrino spectral distortions, which indicate a positive asymmetry $0.032\leq\xi_{\nu}\leq 0.052$ from EMPRESS data.
@@
 1 Introduction
-[…]
+The presence of neutrinos, as evidenced by various cosmological observations, is crucial for understanding the evolution of the universe Lesgourgues:2013sjj .
+In the early universe, neutrinos and antineutrinos remain in thermal equilibrium with the Standard Model (SM) thermal bath as long as the interaction rates between (anti)neutrinos and other SM particles exceed the Hubble expansion rate.
+This equilibrium breaks down when the temperature falls below approximately $1\,\text{MeV}$, leading to the decoupling of neutrinos from the SM thermal bath.
+Shortly after decoupling, the annihilation of electrons and positrons reheats the SM thermal bath, causing the photon temperature to exceed that of the neutrinos, with $T_{\gamma}/T_{\nu}\simeq(11/4)^{1/3}$.
+Subsequent processes in the universe, in particular Big Bang Nucleosynthesis (BBN) and the formation of the Cosmic Microwave Background (CMB) anisotropies and Large Scale Structure (LSS), are both affected by the relic neutrinos, allowing constraints on neutrino properties to be obtained from relevant observations, see $e.g.,$ Pitrou:2018cgg ; Planck:2018vyg ; empress ; DESI:2024mwx .
 A precise understanding of the neutrino decoupling process is essential for accurate constraints on neutrino properties.
 Notably, the partial reheating of neutrinos and antineutrinos due to the small overlap between neutrino decoupling and $e^{\pm}$ annihilations results in slightly non-thermal spectra and an increased total neutrino energy density, typically described by the effective number of relativistic species, $N_{\rm eff}$.
-The values of $N_{\rm eff}$ and the spectral distortions of neutrinos and antineutrinos are used as initial conditions to evaluate the impact of relic neutrinos in subsequent processes.
+The values of $N_{\rm eff}$ and the spectral distortions of neutrinos and antineutrinos are then used as initial conditions to evaluate the impact of relic neutrinos in subsequent processes.
 For the Standard Model case, where asymmetries between neutrinos and antineutrinos are assumed to be negligible, significant progress has been made in detailing the neutrino decoupling process over recent decades Hannestad:1995rs ; Dolgov:1997mb ; Mangano:2001iu ; Mangano:2005cc ; Grohs:2015tfy ; deSalas:2016ztq ; Akita:2020szl ; Froustey:2019owm ; EscuderoAbenza:2020cmq ; Froustey:2020mcq ; Bennett:2020zkv .
-Including flavor oscillations, matter effects, Finite-Temperature QED (FTQED) corrections up to $\mathcal{O}(3)$, and full neutrino-electron and neutrino-neutrino collision terms, the current standard prediction yields $N_{\rm eff}^{\rm SM}=3.0440\pm 0.0002$ Froustey:2020mcq ; Bennett:2020zkv ; Akita:2020szl .
+Including flavor oscillations, matter effects, Finite-Temperature QED (FTQED) corrections up to $\mathcal{O}(e^{3})$, where $e$ is the elementary electric charge, and full neutrino-electron and neutrino-neutrino collision terms, the current standard prediction yields $N_{\rm eff}^{\rm SM}=3.0440\pm 0.0002$ Froustey:2020mcq ; Bennett:2020zkv ; Akita:2020szl .
 On the other hand, inspired by the recent EMPRESS survey on primordial helium abundance empress and its potential to alleviate cosmological tensions Barenboim:2016lxv ; Yeung:2020zde ; Seto:2021tad ; Kumar:2022vee ; Yeung:2024krv , the scenario where asymmetries between neutrinos and antineutrinos are not negligible has recently attracted significant attention.
-Due to the sphaleron processes in the early universe Kuzmin:1985mm ; Khlebnikov:1988sr ; Harvey:1990qw ; Dreiner:1992vm , one might naively expect neutrino asymmetries, or lepton asymmetries, to be of the same order as the baryon asymmetry of the universe, which is strongly constrained by BBN and CMB observations: $\eta_{B}\equiv(n_{B}-n_{\bar{B}})/n_{\gamma}=(6.14\pm 0.04)\times 10^{-10}$ Planck:2018vyg , where $n_{B}$, $n_{\bar{B}}$, and $n_{\gamma}$ denote the number densities of baryons, antibaryons and photons, respectively.
+Typically, such neutrino asymmetries can be parameterised with the degeneracy parameters $\xi_{\nu_{\alpha}}\equiv\mu_{\nu_{\alpha}}/T_{\nu_{\alpha}}$, defined as the chemical potential for the flavour $\alpha$ neutrino normalised to its temperature, so that
+$\eta_{\nu}\equiv\frac{1}{n_{\gamma}}\sum_{\alpha=e,\mu,\tau}(n_{\nu_{\alpha}}-n_{\bar{\nu}_{\alpha}})=\frac{\pi^{2}}{33\zeta(3)}\sum_{\alpha=e,\mu,\tau}\left(\xi_{\nu_{\alpha}}+\frac{\xi_{\nu_{\alpha}}^{3}}{\pi^{2}}\right)\ ,$ (1)
+where $n_{\gamma}$ is the photon number density and $n_{\nu_{\alpha}}(n_{\bar{\nu}_{\alpha}})$ is the (anti)neutrino number density with flavour $\alpha$.
+For the second equality, the SM value for the neutrino-photon temperature ratio $T_{\nu_{i}}/T_{\gamma}=(4/11)^{1/3}$ is assumed.
+Beyond thermal equilibrium, an accurate $\eta_{\nu}$ must be obtained by a precise treatment of the neutrino decoupling process for given $\xi_{\nu_{\alpha}}$ as initial conditions.
+Due to the sphaleron processes in the early universe Kuzmin:1985mm ; Khlebnikov:1988sr ; Harvey:1990qw ; Dreiner:1992vm , one might naively expect neutrino asymmetries, or lepton asymmetries, to be of the same order as the baryon asymmetry of the universe, which is strongly constrained by BBN and CMB observations: $\eta_{B}\equiv(n_{B}-n_{\bar{B}})/n_{\gamma}=(6.14\pm 0.04)\times 10^{-10}$ Planck:2018vyg , where $n_{B}$ and $n_{\bar{B}}$ denote the number densities of baryons and antibaryons, respectively.
 However, substantial lepton asymmetries compared to the baryon asymmetry before the neutrino decoupling epoch can still be generated by various models, see $e.g.,$ Casas:1997gx ; Dolgov:1989us ; Bajc:1997ky ; Asaka:2005pn ; Asaka:2005an ; Pilaftsis:2003gt ; Borah:2022uos ; Kawasaki:2002hq ; Kawasaki:2022hvx .
@@
 Consequently, observations of BBN, CMB, and LSS can constrain primordial neutrino asymmetries Simha:2008mt ; Oldengott:2017tzj ; Pitrou:2018cgg ; Matsumoto:2022tlr ; Burns:2022hkq ; Escudero:2022okz .
-Notably, recent findings from the EMPRESS survey suggest a preference for a positive asymmetry in electron neutrinos Matsumoto:2022tlr ; Burns:2022hkq ; Escudero:2022okzx.
-Specifically, combining BBN and CMB observations yields $\xi_{\nu_{e}}=0.034\pm 0.014$ Escudero:2022okz , where $\xi_{\nu_{e}}=\mu_{\nu_{e}}/T_{\nu}$ is the degeneracy parameter for electron neutrinos.
+Notably, recent findings from the EMPRESS survey suggest a preference for a positive asymmetry in electron neutrinos Matsumoto:2022tlr ; Burns:2022hkq ; Escudero:2022okz .
+Specifically, combining BBN and CMB observations yields $\xi_{\nu_{e}}=0.034\pm 0.014$ Escudero:2022okz .
 However, the current situation is less satisfactory compared to the Standard Model (SM) scenario due to persistent uncertainties concerning neutrino decoupling.
@@
 In this work, we focus on evaluating the neutrino decoupling process and its subsequent effects on BBN, CMB, and LSS in the presence of primordial neutrino asymmetries.
+Since the synchronous oscillations before neutrino decoupling will redistribute the neutrino asymmetries between different flavors, the resulting neutrino asymmetries of different flavors are usually of the same order, so that $\xi_{\nu_{e}}\simeq\xi_{\nu_{\mu}}\simeq\xi_{\nu_{\tau}}$ Dolgov:2002ab ; Pastor:2008ti ; Mangano:2010ei ; Mangano:2011ip ; Castorina:2012md ; Barenboim:2016shh ; Johns:2016enc ; Froustey:2021azz .
+Therefore, a complete flavor equilibration with $\xi_{\nu_{e}}=\xi_{\nu_{\mu}}=\xi_{\nu_{\tau}}=\xi_{\nu}$ is usually assumed in the study of the cosmological constraints on neutrino asymmetries Simha:2008mt ; Oldengott:2017tzj ; Pitrou:2018cgg ; Matsumoto:2022tlr ; Burns:2022hkq ; Escudero:2022okz .
+The framework we have developed in this work can also be straightforwardly applied to studies of the general flavour equilibration process.
+However, a detailed illustration is beyond the scope of this paper.
+Therefore, we will also assume complete flavor equilibration in this paper and leave further studies of general flavor equilibration to future work.
+Relevant previous discussions of the exact flavor equilibration process and its cosmological implications can also be found in, $e.g.,$ Pastor:2008ti ; Mangano:2011ip ; Froustey:2024mgf .
 To accurately describe the evolution of neutrino density matrices during decoupling, including flavor oscillations and full collision effects, we employ the complete set of neutrino quantum kinetic equations (QKEs).
@@
   • We re-derive the most general form of the QKEs for neutrinos and antineutrinos using the CTP formalism, incorporating primordial neutrino asymmetry.
-  • We numerically solve the QKEs for the neutrino decoupling process, accounting for flavor oscillations, matter interactions, FTQED corrections up to $\mathcal{O}(3)$, and complete neutrino-electron and neutrino-neutrino collision terms, matching the accuracy of state-of-the-art Standard Model (SM) predictions Froustey:2020mcq ; Bennett:2020zkv  [^1] Additional QED corrections, such as those to interaction vertices Cielo:2023bqp ; Jackson:2023zkl ; Drewes:2024wbw , remain under debate for the SM case and are not included in this study..
+  • We numerically solve the QKEs for the neutrino decoupling process, accounting for flavor oscillations, matter interactions, FTQED corrections up to $\mathcal{O}(e^{3})$, and complete neutrino-electron and neutrino-neutrino collision terms, matching the accuracy of state-of-the-art Standard Model (SM) predictions Froustey:2020mcq ; Bennett:2020zkv  [^1] Additional QED corrections, such as those to interaction vertices Cielo:2023bqp ; Jackson:2023zkl ; Drewes:2024wbw , remain under debate for the SM case and are not included in this study..
   • We use the resulting density matrices to study the impacts of primordial neutrino asymmetry on BBN, CMB, and LSS, with a particular emphasis on (anti)neutrino spectral distortions by comparing these results with those obtained using the Fermi-Dirac distribution for neutrinos.
@@
 Additionally, $\tilde{\Sigma}(x,y)$ is the neutrino proper self-energy, which is generally a functional of the Green functions and can be decomposed into a local singular term and a matrix in CTP space:
-$\tilde{\Sigma}(x,y)=-\mathbb{1}_{2\times 2}i\Sigma(x)\delta^{4}(x-y)+\tilde{\Pi}\,,\quad\tilde{\Pi}=\begin{pmatrix}\left(\Pi_{ab}\right)^{T}&\left(\Pi_{ab}\right)^{<}\\ \left(\Pi_{ab}\right)^{>}&-\left(\Pi_{ab}\right)^{\bar{T}}\end{pmatrix}\,.$ (6)
+$\tilde{\Sigma}(x,y)=-\mathbb{1}_{2\times 2}i\Sigma(x)\delta^{4}(x-y)+\tilde{\Pi}\,,\quad\tilde{\Pi}=\begin{pmatrix}\left(\Pi_{ab}\right)^{T}&\left(\Pi_{ab}\right)^{<}\\ \left(\Pi_{ab}\right)^{>}&-\left(\Pi_{ab}\right)^{\bar{T}}\end{pmatrix}\,$.
+(7)
 In principle, $\Sigma(x)$ arises from the forward scattering potentials contributed by one-loop diagrams, while $\Pi^{\pm}(x,y)$ represents contributions from non-forward scatterings and annihilations that occur in at least two loops.
@@
 In conclusion, our power counting reads
-$\frac{\partial_{X},m,\Sigma,S^{H}}{E_{\nu}}=O(\epsilon)\,,\quad\frac{\tilde{\Pi}^{\pm}}{E_{\nu}}=O(\epsilon^{2})\,,$ (15)
+$\frac{\partial_{X},m,\Sigma,S^{H}}{E_{\nu}}=\mathcal{O}(\epsilon)\,,\quad\frac{\tilde{\Pi}^{\pm}}{E_{\nu}}=\mathcal{O}(\epsilon^{2})\,$, (16)
 where $E_{\nu}$ is the neutrino energy.
@@
     Beyond $\mathcal{O}(\epsilon^{0})$, all spinor components of $S^{F}$ receive additional corrections due to neutrino masses and interactions.
-Thus, solve the kinetic equation (13b) at $\mathcal{O}(\epsilon^{1})$ to find the constraint relations among these small components.
+Thus, solve the kinetic equation (14b) at $\mathcal{O}(\epsilon)$ to find the constraint relations among these small components.
   3. 3.
-    Expand the kinetic equation (13b) to $\mathcal{O}(\epsilon^{2})$ and reorganize it using the constraint relations obtained at $\mathcal{O}(\epsilon^{1})$.
+    Expand the kinetic equation (14b) to $\mathcal{O}(\epsilon^{2})$ and reorganize it using the constraint relations obtained at $\mathcal{O}(\epsilon)$.
   4. 4.
@@
 $\left(F_{T}^{R}\right)_{\mu\nu}(k,x)$ $=e^{i\varphi(k)}(\hat{\kappa}(k)\wedge\hat{x}^{+}(k))_{\mu\nu}\,\Phi^{\dagger}(k,x)\,.$ (18c)
-Here, we have introduced a set of basis vectors used to express the Lorentz tensors and components of the two-point functions.
-The basis consists of two light-like vectors $\hat{\kappa}^{\mu}(k)=({\rm sign}(k^{0}),\hat{k})$ and $\hat{\kappa}^{\prime\mu}(k)=({\rm sign}(k^{0}),-\hat{k})$, and two transverse four-vectors $\hat{x}_{1,2}(k)$ which satisfy the relations $\hat{\kappa}\cdot\hat{\kappa}=\hat{\kappa}^{\prime}\cdot\hat{\kappa}^{\prime}=0$, $\hat{\kappa}\cdot\hat{\kappa}^{\prime}=2$, $\hat{\kappa}\cdot\hat{x}_{i}=\hat{\kappa}^{\prime}\cdot\hat{x}_{i}=0$, and $\hat{x}_{i}\cdot\hat{x}_{j}=-\delta_{ij}$.
-Additionally, we define $\hat{x}^{\pm}\equiv\hat{x}_{1}\pm i\hat{x}_{2}$, such that $\hat{x}^{+}\cdot\hat{x}^{-}=-2$.
+[…]
 In addition, $F_{L,R}(k,x)$ and $\Phi(k,x)$ can be organized into a $2n_{f}\times 2n_{f}$ matrix:
@@
 $F(\vec{p},x)=\left(\begin{array}[]{cc}\varrho_{LL}&\varrho_{LR}\\ \varrho_{RL}&\varrho_{RR}\end{array}\right)\,;\qquad\qquad\bar{F}(\vec{p},x)=\left(\begin{array}[]{cc}\bar{\varrho}_{RR}&\bar{\varrho}_{RL}\\ \bar{\varrho}_{LR}&\bar{\varrho}_{LL}\end{array}\right)\,$, (21)
-where $n_{f}$ is the total number of neutrino flavors, and each block is a square $n_{f}\times n_{f}$ matrix with $\varrho_{hh^{\prime}}=\varrho_{hh^{\prime}}^{ij}$, where $i,\,j$ denote the flavors and $h,\,h^{\prime}\in\{L,R\}$ denote the chirality.
+where $n_{f}$ is the total number of neutrino flavors, and each block is a square $n_{f}\times n_{f}$ matrix with $\varrho_{hh^{\prime}}=\varrho_{hh^{\prime}}^{ij}$, where i,ji,\,j denote the flavors and $h,\,h^{\prime}\in\{L,R\}$ denote the chirality.
 Among the various components, $\varrho_{LL}$ and $\bar{\varrho}_{RR}$ represent the density matrices for active left-handed neutrinos and right-handed antineutrinos, respectively.
@@
 ${\cal F}=\left(\begin{array}[]{cc}\varrho&\phi\\ \phi^{\dagger}&\bar{\varrho}^{T}\end{array}\right)~,$ (22)
-where $\varrho$ and $\bar{\varrho}$ denote the density matrices for active neutrinos and antineutrinos, respectively, and $\phi$ describes the quantum spin coherence between neutrinos and antineutrinos.
-Similar to the Dirac case, the density matrix for Majorana neutrinos, ${\cal F}$, can be extracted from the two-point functions via
-$-2\int_{0}^{\infty}\frac{dk^{0}}{2\pi}\hat{F}(k,x)=\mathcal{F}(\vec{k},x)-\frac{1}{2}\mathbb{1}\,.$ (22)
+[…]
+$2\int_{0}^{\infty}\frac{dk^{0}}{2\pi}\hat{F}(k,x)=\mathcal{F}(\vec{k},x)-\frac{1}{2}\mathbb{1}\,.$ (23)
 Beyond $\mathcal{O}(\epsilon^{0})$, the equations (20) and (23) serve as the definitions of density matrices for Dirac and Majorana neutrinos, respectively, and are used to extract the QKEs from the evolution equation of $\hat{F}$.
 Additionally, to obtain the QKEs at $\mathcal{O}(\epsilon^{2})$, only the $\mathcal{O}(\epsilon^{0})$ expression for the vector component of the spectral function is required, which is given by $\hat{\rho}(k)=2i\pi|\vec{k}|\delta(k^{2})\text{sgn}(k^{0})$.
-2.2.2 Kinetic equation to $\mathcal{O}(\epsilon^{1})$
+2.2.2 Kinetic equation to $\mathcal{O}(\epsilon)$
 Beyond $\mathcal{O}(\epsilon^{0})$, we still consider $F_{L,R}(k,x)$ and $\Phi(k,x)$ as the four independent spinor components of $S^{F}(k,x)$, which can be isolated using the projections:
@@
 where the upper (lower) signs and indices refer to $\Phi$ ($\Phi^{\dagger}$).
-The dispersion relations for the functions $F_{L,R}(k,x)$ and $\Phi(k,x)$, as well as for the other spinor components of $S^{F}(k,x)$, are modified by neutrino masses and interactions at $\mathcal{O}(\epsilon^{1})$.
+The dispersion relations for the functions $F_{L,R}(k,x)$ and $\Phi(k,x)$, as well as for the other spinor components of $S^{F}(k,x)$, are modified by neutrino masses and interactions at $\mathcal{O}(\epsilon)$.
 Therefore, the general statistical function $S^{F}(k,x)$ can be expressed as
@@
 where $\Delta$ represents the $\mathcal{O}(\epsilon)$ small components.
-Expand the kinetic equation for $S^{F}(k,x)$ to $O(\epsilon)$, we obtain
+Expand the kinetic equation for $S^{F}(k,x)$ to $\mathcal{O}(\epsilon)$, we obtain
 […]
-Substituting the general statistical function $S^{F}(k,x)$, including the $O(\epsilon)$ corrections, into the kinetic equation and decomposing into scalar, vector, and tensor components, we derive the dispersion relations for $F_{L,R}(k,x)$ and $\Phi(k,x)$ at $O(\epsilon)$:
-$\left(k\cdot\hat{\kappa}\right)F_{R}-\frac{1}{2}\left\{\Sigma_{L}^{\kappa},F_{R}\right\}$ $=O\left(\epsilon^{2}\right),$ (26)
-$\left(k\cdot\hat{\kappa}\right)F_{L}-\frac{1}{2}\left\{\Sigma_{R}^{\kappa},F_{L}\right\}$ $=O\left(\epsilon^{2}\right),$ (27)
-$\left(k\cdot\hat{\kappa}\right)P_{+}^{ij}F_{T}^{j}-\frac{1}{2}\left(\Sigma_{R}^{\kappa}P_{+}^{ij}F_{T}^{j}+P_{+}^{ij}F_{T}^{j}\Sigma_{L}^{\kappa}\right)$ $=O\left(\epsilon^{2}\right),$ (28)
-$\left(k\cdot\hat{\kappa}\right)P_{-}^{ij}F_{T}^{j}-\frac{1}{2}\left(\Sigma_{L}^{\kappa}P_{-}^{ij}F_{T}^{j}+P_{-}^{ij}F_{T}^{j}\Sigma_{R}^{\kappa}\right)$ $=O\left(\epsilon^{2}\right),$ (29)
+Substituting the general statistical function $S^{F}(k,x)$, including the $\mathcal{O}(\epsilon)$ corrections, into the kinetic equation and decomposing into scalar, vector, and tensor components, we derive the dispersion relations for $F_{L,R}(k,x)$ and $\Phi(k,x)$ at $\mathcal{O}(\epsilon)$:
+$\left(k\cdot\hat{\kappa}\right)F_{R}-\frac{1}{2}\left\{\Sigma_{L}^{\kappa},F_{R}\right\}$ $=\mathcal{O}\left(\epsilon^{2}\right),$ (27)
+$\left(k\cdot\hat{\kappa}\right)F_{L}-\frac{1}{2}\left\{\Sigma_{R}^{\kappa},F_{L}\right\}$ $=\mathcal{O}\left(\epsilon^{2}\right),$ (28)
+$\left(k\cdot\hat{\kappa}\right)P_{+}^{ij}F_{T}^{j}-\frac{1}{2}\left(\Sigma_{R}^{\kappa}P_{+}^{ij}F_{T}^{j}+P_{+}^{ij}F_{T}^{j}\Sigma_{L}^{\kappa}\right)$ $=\mathcal{O}\left(\epsilon^{2}\right),$ (29)
+$\left(k\cdot\hat{\kappa}\right)P_{-}^{ij}F_{T}^{j}-\frac{1}{2}\left(\Sigma_{L}^{\kappa}P_{-}^{ij}F_{T}^{j}+P_{-}^{ij}F_{T}^{j}\Sigma_{R}^{\kappa}\right)$ $=\mathcal{O}\left(\epsilon^{2}\right),$ (30)
 and the expressions for the small components
-$\Delta_{L/R}^{\kappa}$ $=O\left(\epsilon^{2}\right)\ \ \ \ \,,\ \ \ \ \ \Delta_{T}^{i}=O\left(\epsilon^{2}\right),$ (30)
+$\Delta_{L/R}^{\kappa}$ $=\mathcal{O}\left(\epsilon^{2}\right)\ \ \ \ \,,\ \ \ \ \ \Delta_{T}^{i}=\mathcal{O}\left(\epsilon^{2}\right),$ (31)
 $\Delta_{S}$ $=\frac{1}{2\left|\vec{k}\right|}\left(m^{\dagger}F_{R}+F_{L}m^{\dagger}\right)+\frac{P_{+}^{ij}}{2\left|\vec{k}\right|}\left(i\partial^{i}F_{T}^{j}-\left(\Sigma_{R}^{i}F_{T}^{j}-F_{T}^{j}\Sigma_{L}^{i}\right)\right)$ (32)
@@
 2.2.3 Kinetic equation to $\mathcal{O}(\epsilon^{2})$
-At $O(\epsilon^{2})$, the kinetic equation $S^{F}(k,x)$ becomes
+At $\mathcal{O}(\epsilon^{2})$, the kinetic equation $S^{F}(k,x)$ becomes
 […]
@@
 with
-[…]
+$\hat{S}^{\pm}=-i\hat{\rho}\,\mathbb{1}\pm\hat{F}~$, (41)
 and the $2n_{f}\times 2n_{f}$ self energy matrices $\hat{\Pi}^{\pm}(k)$ are given by
@@
 where
-[…]
-[…]
+$\Pi^{\pm}(\vec{k})$ $=\int_{0}^{\infty}dk^{0}\,\hat{\Pi}^{\pm}(k^{0},\vec{k})\,\delta(k^{0}-\left|\vec{k}\right|)\,$, (48)
+$\bar{\Pi}^{\pm}(\vec{k})$ $=-\int_{-\infty}^{0}dk^{0}\,\hat{\Pi}^{\mp}(k^{0},-\vec{k})\,\delta(k^{0}+\left|\vec{k}\right|)\,.$ (49)
 On the other hand, the QKEs for Majorana neutrinos can be derived by integrating Eq. (37a) over positive frequencies using Eq. (23), yielding:
@@
 However, since the temperature of neutrino decoupling is near that of electron-positron annihilation, neutrinos are also heated through weak interactions with electrons and positrons, leading to a slightly increased $N_{\rm eff}$.
-The most precise calculation for $N_{\rm eff}$ in the SM, including neutrino oscillations, matter effects, FTQED corrections up to $\mathcal{O}(3)$, and neutrino-electron and neutrino-neutrino collision terms, gives $N_{\rm eff}^{\rm SM}=3.0440\pm 0.0002$ Froustey:2020mcq ; Bennett:2020zkv .
+The most precise calculation for $N_{\rm eff}$ in the SM, including neutrino oscillations, matter effects, FTQED corrections up to $\mathcal{O}(e^{3})$, and neutrino-electron and neutrino-neutrino collision terms, gives $N_{\rm eff}^{\rm SM}=3.0440\pm 0.0002$ Froustey:2020mcq ; Bennett:2020zkv .
 Figure 2: Left panel: Neutrino non-instantaneous decoupling correction $\delta N_{\rm eff}^{\rm dec}=N_{\rm eff}-N_{\rm eff}^{\rm inst}$ as a function of the neutrino degeneracy parameter $\xi_{\nu}$.
@@
 $N_{\rm eff}^{\rm inst}\simeq 3+\sum_{\alpha}\left(\frac{30}{7\pi^{2}}\xi_{\nu_{\alpha}}^{2}+\frac{15}{7\pi^{4}}\xi_{\nu_{\alpha}}^{4}\right)=3+3\left(\frac{30}{7\pi^{2}}\xi_{\nu}^{2}+\frac{15}{7\pi^{4}}\xi_{\nu}^{4}\right),$ (64)
-[…]
+where the second equality uses the condition $\xi_{\nu_{e}}=\xi_{\nu_{\mu}}=\xi_{\nu_{\tau}}=\xi_{\nu}$.
+Beyond the instantaneous decoupling limit, we explore the non-instantaneous decoupling corrections, $\delta N_{\rm eff}^{\rm dec}=N_{\rm eff}-N_{\rm eff}^{\rm inst}$, in the presence of the primordial neutrino asymmetry $\xi_{\nu}$.
+In the left panel of Fig.
+2, we show $\delta N_{\rm eff}^{\rm dec}$ as a function of $\xi_{\nu}$.
+For the SM case with $\xi_{\nu}=0$, our calculation yields $\delta N_{\rm eff}^{\rm dec,SM}=0.0441$, consistent with the literature result, $\delta N_{\rm eff}^{\rm dec,SM}=0.0440\pm 0.0002$.
+[^2] The $\sim 0.0001$ difference between our result and the central value in previous literature arises mainly from different values of $N_{y}$ used for discretizing the neutrino momentum grid.
+Our result agrees with previous results using $N_{y}=40$; see Table 2 in Bennett:2020zkv .
+For non-zero $\left|\xi_{\nu}\right|$, $\delta N_{\rm eff}^{\rm dec}$ increases beyond $\delta N_{\rm eff}^{\rm dec,SM}$, as expected since a higher $\left|\xi_{\nu}\right|$ results in more energy transfer to neutrinos from electrons and positrons.
+We find that $\delta N_{\rm eff}^{\rm dec}$ as a function of $\xi_{\nu}$ can be well fitted by $\delta N_{\rm eff}^{\rm dec}=\delta N_{\rm eff}^{\rm dec,SM}+0.0102\,\xi_{\nu}^{2}$, as shown in the left panel of Fig.
+2.
+In conclusion, the final $N_{\rm eff}$ after neutrino decoupling, in the presence of primordial neutrino asymmetry $\xi_{\nu}$, is given by
 $N_{\rm eff}(\xi_{\nu})=N_{\rm eff}^{\rm SM}+3\left(\frac{30}{7\pi^{2}}\xi_{\nu}^{2}+\frac{15}{7\pi^{4}}\xi_{\nu}^{4}\right)+0.0102\,\xi_{\nu}^{2},$ (65)
@@
 $n_{a}\Gamma_{a\to b}=$ $\int\mathrm{d}\pi_{a}\mathrm{d}\pi_{b}\mathrm{d}\pi_{e}\mathrm{d}\pi_{\nu}\delta^{(4)}\left({p}_{a}-p_{b}+\alpha_{e}p_{e}+\alpha_{\nu}p_{\nu}\right)\left|M\right|^{2}_{a\to b}{}$ (74)
-×$\times f_{a}(E_{a})f_{b}(-E_{b})f_{e}(\alpha_{e}{E}_{e})\tilde{f}_{\nu}(E_{\nu}),$
+$\times f_{a}(E_{a})f_{b}(-E_{b})f_{e}(\alpha_{e}{E}_{e})\tilde{f}_{\nu}(E_{\nu}),$
 where $\mathrm{d}\pi_{i}=\frac{1}{(2\pi)^{3}}\frac{\mathrm{d}^{3}\mathbf{p}_{i}}{2E_{i}}$, $\alpha_{e}(\alpha_{\nu})=1$ if the electron (neutrino) is in the initial state and $\alpha_{e}(\alpha_{\nu})=-1$ if it is in the final state.
@@
 Lastly, $\tilde{f}_{\nu}(E_{\nu})$ equals $\varrho_{\nu}(E_{\nu})$ (or $\varrho_{\bar{\nu}}(E_{\nu})$) for the neutrino (or antineutrino) in the initial state, and equals $1-\varrho_{\nu}(E_{\nu})$ (or $1-\varrho_{\bar{\nu}}(E_{\nu})$) for those in the final state.
-[…]
+To leading order, two additional approximations are typically made.
+First, we neglect neutrino spectral distortions and use the Fermi-Dirac distribution with the corresponding chemical potentials $f_{\nu}^{\pm}(E)\equiv\frac{1}{e^{E/T_{\nu}\mp\xi_{\nu}}+1}$ for neutrinos and antineutrinos, where $T_{\nu}$ is the effective temperature of the neutrino, defined as the temperature that satisfies $3\frac{7}{8}\frac{\pi^{2}}{15}T_{\nu}^{4}=\rho_{\nu}$.
+Furthermore, since the temperature during the BBN epoch (approximately $0.1\,{\rm MeV}$) is much smaller than the masses of the neutron and proton (approximately $1\,{\rm GeV}$), we can work within the infinite nucleon mass limit, also known as the Born approximation.
+Under these two approximations and following the notation of Pitrou:2018cgg , the reaction rates are significantly simplified as follows:
 $\bar{\Gamma}_{n\to p}$ $=K\int_{0}^{\infty}p^{2}\mathrm{d}p[\chi_{+}(E)+\chi_{+}(-E)]\,$, (75)
@@
 Similarly, the terms in (76) can be interpreted.
-Here, $E^{\mp}_{\nu}$ is equal to the energy of the neutrino $E_{\nu}$ when the neutrino is in the initial state, and $-E_{\nu}$ when it is in the final state.
+Here, $E^{\mp}_{\nu}$ is equal to the energy of the neutrino $E_{\nu}$ when the neutrino is in the initial state, and $E_{\nu}$ when it is in the final state.
 The term $f_{\nu}^{\pm}(E^{\mp}_{\nu})$ in the function $\chi^{\pm}$ provides the corresponding distribution or Pauli-blocking factor, following the relation $1-f_{\nu}^{+}(E_{\nu})=f_{\nu}^{-}(-E_{\nu})$.
@@
 $\delta\chi_{\pm}(E)\equiv$ $(E_{\nu}^{\mp})^{2}f_{e}(-E)\,{\rm sgn}(E^{\mp}_{\nu})$
-×$\times\left(\frac{1+\Theta(\pm E^{\mp}_{\nu})\delta\varrho_{\nu}+\Theta(\mp E^{\mp}_{\nu})\delta\varrho_{\bar{\nu}}}{e^{\left|E^{\mp}_{\nu}\right|/T_{\rm CM}\mp{\rm sgn}(E^{\mp}_{\nu})\xi_{\nu}}+1}-\frac{1}{e^{\left|E^{\mp}_{\nu}\right|/T_{\nu}\mp{\rm sgn}(E^{\mp}_{\nu})\xi_{\nu}}+1}\right)\,,$ (79)
+$\times\left(\frac{1+\Theta(\pm E^{\mp}_{\nu})\delta\varrho_{\nu}+\Theta(\mp E^{\mp}_{\nu})\delta\varrho_{\bar{\nu}}}{e^{\left|E^{\mp}_{\nu}\right|/T_{\rm CM}\mp{\rm sgn}(E^{\mp}_{\nu})\xi_{\nu}}+1}-\frac{1}{e^{\left|E^{\mp}_{\nu}\right|/T_{\nu}\mp{\rm sgn}(E^{\mp}_{\nu})\xi_{\nu}}+1}\right)\,$, (80)
 where $\Theta(x)$ is the Heaviside step function, and $\delta\varrho_{\nu}$ and $\delta\varrho_{\bar{\nu}}$ represent the spectral distortions of electron neutrinos and antineutrinos, respectively.
@@
 (4.1) coincides with Eq.
-(29) in Ref.
-Froustey:2019owm, where corrections to the weak rates from neutrino spectral distortions in the SM were considered.
+(29) in Froustey:2019owm , where corrections to the weak rates from neutrino spectral distortions in the SM were considered.
 Finally, we provide a brief overview of other corrections to the weak interaction rates considered in this work.
@@
 ${\Gamma}^{{\rm RC}0}_{p\to n}$ $=K\int_{0}^{\infty}p^{2}\mathrm{d}p\left[\mathcal{F}_{-}(E)R(E,\Delta+E)\chi_{-}(E)+\mathcal{F}_{-}(-E)R(E,|E-\Delta|)\chi_{-}(-E)\right]\,.$ (83)
-The explicit expressions for the Fermi function $F_{\pm}$ and the radiative correction factor $R(E,k_{\rm max})$ can be found in, e.g., Ref.
-Pitrou:2018cgg, and are not repeated here for brevity.
+The explicit expressions for the Fermi function $F_{\pm}$ and the radiative correction factor $R(E,k_{\rm max})$ can be found in, e.g., Pitrou:2018cgg , and are not repeated here for brevity.
 It is evident that the zero-temperature radiative correction modifies only the coefficients of the $\chi_{\pm}$ functions.
@@
 Specifically, we have implemented the corrections induced by neutrino spectral distortions into the public BBN code PRIMAT Pitrou:2018cgg and computed the abundances of the light elements.
-For this computation, we utilized the latest values of relevant physical and cosmological parameters; for instance, the neutron lifetime is taken to be $\tau_{n}=879.4\,\text{s}$, the axial coupling of the nucleons is $g_{A}=1.2756$, and the baryon density is $\omega_{b}h^{2}=0.024$.
+For this computation, we utilized the latest values of relevant physical and cosmological parameters; for instance, the neutron lifetime is taken to be $\tau_{n}=879.4\,\text{s}$, the axial coupling of the nucleons is $g_{A}=1.2756$, and the baryon density is $\omega_{b}=0.0224$.
 For convenience, we define the mass fraction of isotope $i$ as $X_{i}\equiv A_{i}\frac{n_{i}}{n_{b}}$, where $A_{i}$ and $n_{i}$ are the baryon number and number density of isotope $i$, respectively, and $n_{b}$ is the total baryon number density.
 We also customarily denote $Y_{P}\equiv X_{{}^{4}{\rm He}}$ and $i/\rm H\equiv n_{i}/n_{{}^{1}\rm H}$.
-Figure 5: Light element abundances as a function of the neutrino degeneracy parameter $\xi_{\nu}$, normalized to the results for $\xi_{\nu}=0$.
+Figure 5: Light element abundances as a function of the neutrino degeneracy parameter $\xi_{\nu}$, normalized with the results for $\xi_{\nu}=0$.
+The baryon density is fixed at $\omega_{b}=0.0224$, and the nuclear rates of the BBN are driven by the PRIMAT approach.
 The $^{3}{\rm He}$ notation refers to $^{3}{\rm He}+\rm T$, and $^{7}{\rm Li}$ denotes $^{7}{\rm Li}+\rm^{7}Be$, accounting for slow radioactive decays.
-All optional weak rate corrections in PRIMAT (see Ref.
-Pitrou:2018cgg) and the new corrections from neutrino spectral distortions are included in the calculations, see text for more details.
-The $68\%$ constraints on the helium abundance $Y_{P}$ from the CMB ($Y_{P}=0.242\pm 0.012$) and the recent EMPRESS experiment ($Y_{P}=0.2370^{+0.0034}_{-0.0033}$) are indicated by the light and dark blue shaded regions, respectively.
-Figure 5 illustrates the resulting light element abundances, incorporating all weak rate corrections from (80) and the QED corrections up to $\mathcal{O}(e^{3})$ for plasma thermodynamics and certain nuclear rates, as a function of the neutrino degeneracy parameter $\xi_{\nu}$.
+All optional weak rate corrections in PRIMAT (see Pitrou:2018cgg ) and the new corrections from neutrino spectral distortions are included in the calculations, see text for more details.
+The $68\%$ constraints on the helium abundance $Y_{P}$ from the CMB ($Y_{P}=0.242\pm 0.012$), the PDG recommended value ($Y_{P}=0.245\pm 0.003$) and the recent EMPRESS experiment ($Y_{P}=0.2370^{+0.0034}_{-0.0033}$) are indicated by the light and dark blue shaded regions, respectively.
+Fig.
+5 illustrates the resulting light element abundances, incorporating all weak rate corrections from (81) and the QED corrections up to $\mathcal{O}(e^{3})$ for plasma thermodynamics and certain nuclear rates, as a function of the neutrino degeneracy parameter $\xi_{\nu}$.
 The abundance of $^{3}{\rm He}$ refers to $^{3}{\rm He}+\rm T$, while $^{7}{\rm Li}$ corresponds to $^{7}{\rm Li}+\rm^{7}Be$, accounting for slow radioactive decays.
@@
 Additionally, Fig.
-5 shows the constraints on helium abundance derived from the CMB and EMPRESS experiments:
-$Y_{P}|_{\rm CMB}=0.242\pm 0.012,\quad Y_{P}|_{\rm EMPRESS}=0.2370^{+0.0034}_{-0.0033},$ (83)
-as represented by the light and dark blue shaded regions.
-Our results indicate that the constraints on $Y_{P}$ within the $68\%$ confidence interval can be translated into constraints on the neutrino degeneracy parameter $\xi_{\nu}$:
+5 shows the constraints on helium abundance derived from the CMB observation, the EMPRESS observation and the PDG recommended value:
+$Y_{P}|_{\rm CMB}=0.242\pm 0.012,$ (84)
+$Y_{P}|_{\rm EMPRESS}=0.2370^{+0.0034}_{-0.0033},$ (85)
+$Y_{P}|_{\rm PDG}=0.245\pm 0.003$ (86)
+as represented by the blue shaded regions.
+Through our numerical results with the fixed baryon density $\omega_{b}=0.0224$, the constraints on $Y_{P}$ within the $68\%$ confidence interval can be easily translated into constraints on the neutrino degeneracy parameter $\xi_{\nu}$:
 ${\rm CMB}:\quad-0.026\leq$ $\,\xi_{\nu}\leq 0.069\,$, (87)
-${\rm EMPRESS}:\quad 0.027\leq$ $\,\xi_{\nu}\leq 0.054\,.$ (85)
-Clearly, the EMPRESS experiment favors a non-zero positive $\xi_{\nu}$, while the CMB experiment remains compatible with the Standard Model case of $\xi_{\nu}=0$.
+${\rm EMPRESS}:\quad 0.027\leq$ $\,\xi_{\nu}\leq 0.054\,$, (88)
+${\rm PDG}:\quad-0.003\leq$ $\,\xi_{\nu}\leq 0.021\,.$ (89)
+Clearly, while the CMB observation yields a significantly weaker constraint on the helium abundance $Y_{P}$, the constraint from the EMPRESS observation has similar uncertainties but a much smaller central value with respect to the PDG recommended observation.
+Consequently, the EMPRESS observation favours a non-zero positive $\xi_{\nu}$, while the CMB and PDG recommended observations remain compatible with the Standard Model case of $\xi_{\nu}=0$.
+Key Nuclear Reactions
+[…]
+[…]
+3He+D $\to$ 4He+p 3He+4He $\to$ $\gamma 7Be$+n $\to$ 7Li+p 7Li+p $\to$ 4He+4He
+Table 1: The key nuclear reactions that possessing different reaction rates in the PRIMAT and NACRE II approaches.
+Figure 6: Light element abundances as a function of the neutrino degeneracy parameter $\xi_{\nu}$ that driven by the PRIMAT approach (solid curves) and the NACRE II approach (dash-dotted curves), normalized with the results for $\xi_{\nu}=0$ in the PRIMAT approach.
+The $^{3}{\rm He}$ notation refers to $^{3}{\rm He}+\rm T$, and $^{7}{\rm Li}$ denotes $^{7}{\rm Li}+\rm^{7}Be$, accounting for slow radioactive decays.
+All optional weak rate corrections in PRIMAT (see Pitrou:2018cgg ) and the new corrections from neutrino spectral distortions are included in the calculations, see text for more details.
+The $68\%$ constraints on the deuterium abundance $D/H$ of the PDG recommended value ($D/H\times 10^{5}=2.547\pm 0.025$) are indicated by the light red shaded region.
+It is also interesting to consider the deuterium abundance constraint.
+For this purpose, we first briefly discuss the two most widely used databases for thermonuclear rates relevant to BBN, as these are very relevant to the deuterium abundance predicted by BBN.
+The first database is tabulated by the public BBN code PRIMAT with an extensive catalogue of more than 400 reactionsPitrou:2018cgg , implemented according to the refined statistical analyses of Refs.
+Descouvemont:2004cw ; Iliadis:2016vkw ; InestaGomez:2017eya .
+We will refer to the predicted BBN abundances obtained with this database as the PRIMAT approach, and use it to obtain most of the numerical results in this paper.
+On the other hand, the NACRE II database Xu:2013fha gives an evaluation of 12 key nuclear reaction rates with mass number $A<16$, based on the nuclear cross sections described by the so-called potential model Angulo:1999zz , as shown in Tab.
+1.
+With the same nuclear rates of the other reactions with the PRIMAT approach, we will denote it as the NACRE II approach.
+We present the resulting BBN abundances as a function of the neutrino degeneracy parameter $\xi_{\nu}$ in these two approaches in Fig.
+6, where the solid curves refer to the PRIMAT approach, which is identical to the results in Fig.
+5, and the dashed curves refer to the NACRE II approach.
+All curves are normalized using the resulting for $\xi_{\nu}=0$ in the PRIMAT approach, and the $68\%$ constraints on the deuterium abundance $D/H$ of the PDG recommended value ($D/H\times 10^{5}=2.547\pm 0.025$) are indicated by the light red shaded region.
+Comparing the light element abundances obtained by the two approaches, we find that they give exactly the same predictions for $Y_{P}$ and similar predictions for $^{3}{\rm He}/H$.
+However, the NACRE II approach predicts a a significantly larger $D/H$ and smaller $^{7}{\rm Li}/H$ than the PRIMAT approach.
+Consequently, the constraints on the neutrino degeneracy parameter $\xi_{\nu}$ from the PDG constraint $D/H\times 10^{5}=2.547\pm 0.025$ with the fixed baryon density $\omega_{b}=0.0224$ are also significantly different in the two approaches:
+${\rm PRIMAT}:\quad-0.088\leq$ $\,\xi_{\nu}\leq-0.057\,$, (90)
+${\rm NACRE\,II}:\quad-0.042\leq$ $\,\xi_{\nu}\leq-0.009\,.$ (91)
+We see that, while the PRIMAT approach points to a much smaller $\xi_{\nu}$, both approaches favor a negative negative neutrino asymmetry $\xi_{\nu}$.
+Such results indicate a mild tension between current observations of helium and deuterium abundances.
+In addition, we note that despite the mild tension between $Y_{P}$ and $D/H$ observations, the preference for positive $\xi_{\nu}$ from the EMPRESS observation on $Y_{P}$ is more reliable, since $Y_{P}$ is more sensitive to the neutrino asymmetry $\xi_{\nu}$, while $D/H$ is more sensitive to the baryon density $\omega_{b}$, which is fixed in our upper analysis.
+Therefore, the preference for a negative $\xi_{\nu}$ from the $D/H$ observation can be greatly mitigated with a slightly different $\omega_{b}$.
+Consequently, an overall MCMC analysis combining helium and deuterium observations still indicates a preference for positive $\xi_{\nu}$, as we will see in Sec.
+6.
+For the sake of simplicity, we will also only consider the PRIMAT approach in the following, and leave a detailed comparison between different nuclear reaction rates to our subsequent companion paper Li:2025rjr .
 Figure 7: The relative difference in light element abundances compared to the results in the neutrino instantaneous decoupling limit as a function of the neutrino degeneracy parameter $\xi_{\nu}$.
@@
 No-SD, $\xi_{\nu}=0$ 0.007 0.367 0.122 -0.430
-No-SD, $\xi_{\nu}=0.1$0.048 0.343 0.112 -0.460
+No-SD, $\xi_{\nu}=0.1$ -0.048 0.343 0.112 -0.460
 Actual-SD, $\xi_{\nu}=-0.1$ 0.059 0.4 0.137 -0.411
@@
 It is also interesting to discuss the effects of corrections induced by neutrino spectral distortions on BBN.
-We examine the impact of different treatments of the neutrino decoupling process, considering three implementations: (1) the neutrino instantaneous decoupling limit, $i.e.,$ , $N_{\rm eff}^{\rm inst}$, as discussed in the previous section, where thermal Fermi-Dirac distributions for neutrinos and antineutrinos are used in the BBN calculations; (2) the no-SD case, which employs the actual $N_{\rm eff}$ and neutrino temperatures, incorporating the non-instantaneous decoupling effect while only using thermal Fermi-Dirac distributions for neutrinos and antineutrinos.
+We examine the impact of different treatments of the neutrino decoupling process, considering three implementations: (1) the neutrino instantaneous decoupling limit, $i.e.,$ $N_{\rm eff}^{\rm inst}$, as discussed in the previous section, where thermal Fermi-Dirac distributions for neutrinos and antineutrinos are used in the BBN calculations; (2) the no-SD case, which employs the actual $N_{\rm eff}$ and neutrino temperatures, incorporating the non-instantaneous decoupling effect while only using thermal Fermi-Dirac distributions for neutrinos and antineutrinos.
 This treatment is the most common in current BBN studies; (3) the actual SD case, which utilizes the actual $N_{\rm eff}$, accounting for the non-instantaneous decoupling effect and incorporating actual density distributions with spectral distortions for neutrinos and antineutrinos.
@@
 Using the BBN results in the instantaneous decoupling limit as a benchmark, we present the relative differences for the no-SD and actual-SD cases, indicated by dotted and solid curves in Fig.
-6, respectively.
+7, respectively.
 Note that we have plotted the negative of the relative differences for $^{7}{\rm Li}$ in Fig.
 7 due to the specific nuclear reaction rates utilized in PRIMAT.
-We report the explicit abundances of helium-4, deuterium, helium-3, and lithium-7 for the degeneracy parameters $\xi_{\nu}=0,\pm 1$ in table 1, along with the associated relative variations compared to the baseline abundance in table 2.
+We report the explicit abundances of helium-4, deuterium, helium-3, and lithium-7 for the degeneracy parameters $\xi_{\nu}=0,\pm 0.1$ in table 2, along with the associated relative variations compared to the baseline abundance in table 3.
 Our results with $\xi_{\nu}=0$ are consistent with those reported in Froustey:2020mcq , which also considered the effects of neutrino spectral distortions on BBN within the standard model (SM).
@@
 where the subscript $\xi_{\nu}=0$ indicates results obtained in the SM case using actual density matrices, as detailed in Table 2.
-These fitting functions facilitate further studies, yielding fractional errors smaller than $0.01\%$ within the range $-0.1\leq\xi_{\nu}\leq 0.1$.
+These fitting functions facilitate further studies, yielding fractional errors smaller than $0.01\%$ within the range $0.1\leq\xi_{\nu}\leq 0.1$.
 [^3] The coefficients of our fitting functions differ slightly from those in Pitrou:2018cgg ; Froustey:2024mgf , which is mainly due to the different neutron lifetimes we have chosen.
@@
 $\Omega_{\nu}h^{2}$ $=\frac{\sum\limits_{i}T_{\rm CM,0}^{4}\int_{0}^{\infty}\mathrm{d}y\,y^{2}\sqrt{y^{2}+m_{i}}\left(\varrho_{i}(y)+\bar{\varrho}_{i}(y)\right)/(2\pi^{2})}{\rho_{\rm crit}}$ (96)
-$\overset{\delta\varrho=0}{\simeq}\frac{\sum\limits_{i}m_{i}}{93.12\,{\rm eV}}\frac{{\rm Li}_{3}(-\exp(-\xi_{\nu}))+{\rm Li}_{3}(-\exp(\xi_{\nu}))}{3\zeta(3)/2},$
+$\overset{\delta\varrho=0}{\simeq}\frac{\sum\limits_{i}m_{i}}{93.12\,{\rm eV}}\left(-\frac{{\rm Li}_{3}(-\exp(-\xi_{\nu}))+{\rm Li}_{3}(-\exp(\xi_{\nu}))}{3\zeta(3)/2}\right),$
 where $y=p/T_{\rm CM}$ is the comoving momentum defined in section 3, $T_{\rm CM,0}$ is the current comoving temperature, $\rho_{\rm crit}$ is the critical energy density, and $\varrho_{i}(y)$ and $\bar{\varrho}_{i}(y)$ represent the comoving density distributions of neutrinos and antineutrinos, respectively, for $i=1,2,3$, corresponding to the three mass eigenstates with mass $m_{i}$.
@@
 The first factor, $\frac{\sum\limits_{i}m_{i}}{93.12\,{\rm eV}}$, represents the usual neutrino relic abundance in the SM, while the second factor accounts for the effects of neutrino asymmetries, with ${\rm Li}_{j}(x)$ being the polylogarithms of order $j$ and $\zeta(x)$ the Riemann zeta function.
-In the SM case, ${\rm Li}_{3}(-1)=3\zeta(3)/4$, so the second factor equals 1.
+In the SM case, ${\rm Li}_{3}(-1)=-3\zeta(3)/4$, so the second factor equals 1.
 For non-zero $\xi_{\nu}$, the second factor depends on the absolute value of $\xi_{\nu}$, with larger $\xi_{\nu}$ leading to a larger factor and, consequently, a larger neutrino abundance.
@@
 $\varrho_{i}\left(\vec{x},\vec{p},\tau\right)=\varrho_{i}^{0}(y)\left[1+\Psi_{i}\left(\vec{x},y,\hat{n},\tau\right)\right],$ (97)
-[…]
+where $\tau\equiv\int dt/a(t)$ is the conformal time, $y\equiv p/T_{\rm CM}$ is the comoving momentum, $\hat{n}\equiv\vec{p}/p$ is the direction of the momentum, and $\varrho_{i}^{0}(y)$ represents the unperturbed zeroth-order neutrino density distributions obtained in Section 3.
 In this work, we adopt the synchronous gauge, where the line element is written as
@@
 where $\epsilon\equiv aE=\sqrt{q^{2}+a^{2}m^{2}_{\nu_{i}}}$ is the comoving energy. Since Eq. (102) is independent of the azimuthal angle, we can expand $\Psi_{i}$ in a Legendre series:
-$\Psi_{i}(\vec{k},\hat{n},q,\tau)=\sum_{l=0}^{\infty}(-i)^{l}(2l+1)\Psi_{i,l}(k,q,\tau)P_{l}(\hat{k}\cdot\hat{n}).$ (97)
+[…]
 Substituting Eq. (103) into Eq. (102) and using the orthonormality of Legendre polynomials, we obtain the Boltzmann hierarchies for $\Psi_{i,l}$:
@@
 Specifically, the fractional density perturbation for neutrinos of family $i$ is given by
-[…]
+$\delta_{i}(\vec{x},t)\equiv\frac{\rho_{i}(\vec{x},t)-\rho_{i,0}(t)}{\rho_{i,0}(t)}=\frac{4\pi T_{\rm CM}^{4}\int dy~y^{2}\varrho_{i}^{0}(y)\Psi_{i,0}\epsilon}{\rho_{i,0}(t)},$ (105)
 where $\rho_{i,0}(t)$ represents the zeroth-order homogeneous background energy density for neutrinos.
@@
 Following Ref. Ma:1995ey , the evolution of photon perturbations is governed by the Boltzmann hierarchy:
-$\dot{\delta}_{\gamma}$ $=$ $-\frac{4}{3}\theta_{\gamma}-\frac{2}{3}\dot{h},$
+$\dot{\delta}_{\gamma}$ $=-\frac{4}{3}\theta_{\gamma}-\frac{2}{3}\dot{h},$ (106)
 $\dot{\theta}_{\gamma}$ $=k^{2}\left(\frac{1}{4}\delta_{\gamma}-\sigma_{\gamma}\right)+an_{e}\sigma_{T}(\theta_{b}-\theta_{\gamma}),$
-[…]
+$\dot{F}_{\gamma\,2}$ $=2\dot{\sigma}_{\gamma}=\frac{8}{15}\theta_{\gamma}-\frac{3}{5}kF_{\gamma\,3}+\frac{4}{15}\dot{h}+\frac{8}{5}\dot{\eta}-\frac{9}{5}an_{e}\sigma_{T}\sigma_{\gamma}+\frac{1}{10}an_{e}\sigma_{T}(G_{\gamma\,0}+G_{\gamma\,2}),$
 $\dot{F}_{\gamma\,l}$ $=\frac{k}{2l+1}\left[lF_{\gamma\,(l-1)}-(l+1)F_{\gamma\,(l+1)}\right]-an_{e}\sigma_{T}F_{\gamma\,l},\quad l\geq 3,$
-[…]
+$\dot{G}_{\gamma\,l}$ $=\frac{k}{2l+1}\left[lG_{\gamma\,(l-1)}-(l+1)G_{\gamma\,(l+1)}\right]+an_{e}\sigma_{T}\left[-G_{\gamma\,l}+\frac{1}{2}\left(F_{\gamma\,2}+G_{\gamma\,0}+G_{\gamma\,2}\right)\left(\delta_{l0}+\frac{\delta_{l2}}{5}\right)\right].$
 Here, $\delta_{\gamma}$ is the photon density fluctuation, $\theta_{\gamma}$ is the divergence of photon velocity, and $F_{\gamma\,l}$ represents the photon density perturbation.
@@
 In Fig.
-7, we show the ratio of the CMB angular power spectrum, $C_{l}^{\rm TT}/C_{l}^{\rm TT}(\sum m_{\nu}=0)$, and the matter power spectrum, $P_{k}/P_{k}(\sum m_{\nu}=0)$, for several choices of total neutrino mass.
+8, we show the ratio of the CMB angular power spectrum, $C_{l}^{\rm TT}/C_{l}^{\rm TT}(\sum m_{\nu}=0)$, and the matter power spectrum, $P_{k}/P_{k}(\sum m_{\nu}=0)$, for several choices of total neutrino mass.
 The left and right panels depict the results for the CMB and matter power spectra, respectively.
@@
 For the linear matter power spectrum, shown in the right panel of Fig.
-8, we observe that at large scales, the results for non-zero $\xi_{\nu}$ are identical to those for $\xi_{\nu}=0$.
+9, we observe that at large scales, the results for non-zero $\xi_{\nu}$ are identical to those for $\xi_{\nu}=0$.
 At smaller scales, however, the free-streaming of neutrinos is enhanced with increasing $\xi_{\nu}$, resulting in a slightly stronger suppression of the matter power spectrum, while this suppression is largely insensitive to the sign of $\xi_{\nu}$.
@@
 Due to the additional contribution to $N_{\rm eff}$ for non-zero $\xi_{\nu}$, as per Eq.
-63, larger $\left|\xi_{\nu}\right|$ corresponds to a smaller $r_{d}$, resulting in significant imprints on the matter power spectrum.
+64, larger $\left|\xi_{\nu}\right|$ corresponds to a smaller $r_{d}$, resulting in significant imprints on the matter power spectrum.
 These results imply that BAO experiments can provide stringent constraints on the primordial neutrino asymmetry, as will be discussed in Section 6, a consideration often overlooked in previous studies.
@@
 As in Fig.
-8, we present in the left and right panels of Fig.
+9, we present in the left and right panels of Fig.
 11 the CMB angular power spectrum $C_{l}^{\rm TT}$ and the linear matter power spectrum $P(k)$ for non-zero $\xi_{\nu}$, compared to the corresponding results with $\xi_{\nu}=0$.
 The solid curves correspond to results that include spectral distortions, as shown in Fig.
-8, while the dashed curves correspond to results without the neutrino spectral distortions.
+9, while the dashed curves correspond to results without the neutrino spectral distortions.
 Comparing $C_{l}^{\rm TT}$ with and without spectral distortions, we see that the primary effect of including these distortions is a slight suppression of the first three acoustic peaks, with negligible impact on the smaller-scale peaks.
@@
 BBN + CMB $0.028\pm 0.012$ $2.189\pm 0.012$ $0.1225\pm 0.0013$ 0.414
-BBN + CMB + BAO $0.024\pm 0.012$ $2.200\pm 0.011$ $0.1205\pm 0.0009$ 0.131
+[…]
 Bounds from the datasets BBN + CMB + BAO
@@
 $\mathbf{\Lambda}$CDM + $\mathbf{m_{\nu_{1}}}$ $2.211\pm 0.010$ $0.1203\pm 0.0009$ 0.136
-$\mathbf{\Lambda}$CDM + $\mathbf{\xi_{\nu}}$ $0.023\pm 0.012$ $2.200\pm 0.011$ $0.1207\pm 0.0009$
-$\mathbf{\Lambda}$CDM + $\mathbf{\xi_{\nu}}$ + $\mathbf{m_{\nu_{1}}}$ $0.024\pm 0.012$ $2.200\pm 0.011$ $0.1205\pm 0.0009$ 0.131
+[…]
+[…]
 Table 4: Summary of constraints on the primordial lepton asymmetry $\xi_{\nu}$, the baryon density $\omega_{b}$, the cold dark matter density $\omega_{cdm}$, and the total neutrino mass $\sum m_{\nu}$ in normal hierarchy from various datasets and cosmological models.
@@
 The BAO dataset primarily constrains the total neutrino mass $\sum m_{\nu}$, but also provides additional constraints on primordial neutrino asymmetries.
-To extract constraints from these datasets, we perform MCMC analyses using the Montepython engine Audren:2012wb ; Brinckmann:2018cvx .
-For the analysis, $N_{\rm eff}$ is derived from Eq.
-(63) as discussed in Section 3.
-The predicted helium and deuterium abundances from BBN are calculated using our fitting formulas in Eq.
-(86).
-As shown in Fig.
-10, the impact of neutrino spectral distortions on the CMB angular power spectrum and BAO in the matter power spectrum is minor.
-Therefore, for simplicity, (anti)neutrino spectral distortions are neglected in the MCMC analysis.
-For the CMB and BAO datasets, the parameter set includes the six independent parameters of the $\Lambda$CDM model, along with the neutrino degeneracy parameter $\xi_{\nu}$ and the mass of the lightest neutrino in the normal hierarchy $m_{\nu_{1}}$, i.e., $\{\omega_{b},\omega_{cdm},\theta_{s},A_{s},n_{s},\tau_{\mathrm{reio}},\xi_{\nu},m_{\nu_{1}}\}$.
-Here, $\theta_{s}$ is the sound horizon, $A_{s}$ is the amplitude of primordial curvature perturbations at the pivot scale $k=0.05\,\mathrm{Mpc}^{-1}$, $n_{s}$ is the scalar spectral index, and $\tau_{\mathrm{reio}}$ is the optical depth to reionization.
-For the BBN dataset, we consider only $\{\omega_{b},\xi_{\nu}\}$, as the predicted light element abundances from BBN are largely independent of other cosmological parameters within the relevant range.
+[…]
 Figure 12: 2D and 1D marginalized posterior distributions for the baryon abundance $\omega_{b}$, the neutrino degeneracy parameter $\xi_{\nu}$, the cold dark matter abundance $\omega_{cdm}$, the scalar spectral index $n_{s}$, and the mass of the lightest neutrino for the normal hierarchy $m_{\nu_{1}}$.
@@
 Figure 13: 1D marginalized distributions for different parameter sets derived from the BBN + CMB + BAO datasets.
-Using the specified parameters and datasets, we modified the MontePython sampler and CLASS codes to perform the MCMC analysis.
-All chains converged with a Gelman-Rubin criterion of $R-1<0.01$.
-The resulting $1\,\sigma$ cosmological constraints are summarized in Table 3, where the $95\%$ confidence limits on the total neutrino mass are also presented, derived from constraints on the mass of the lightest neutrino for the normal hierarchy, $m_{\nu_{1}}$.
-For the model $\Lambda$CDM + ${\xi_{\nu}}$ + ${m_{\nu_{1}}}$, we present the 1D and 2D marginalized posterior distributions of key parameters that show significant variations across different datasets in Fig.
-11, including the baryon abundance $\omega_{b}$, the neutrino degeneracy parameter $\xi_{\nu}$, the cold dark matter abundance $\omega_{cdm}$, the sound horizon $\theta_{s}$, and the mass of the lightest neutrino for the normal hierarchy $m_{\nu_{1}}$.
+[…]
 Regarding constraints on primordial neutrino asymmetries, the EMPRESS data in the BBN dataset indicates a positive primordial neutrino asymmetry, $\xi_{\nu}=0.039\pm 0.013$, which differs from zero at $3.0\,\sigma$ significance.
@@
 We find that constraints on other parameters, such as $\omega_{b}$, $\omega_{cdm}$, $n_{s}$, and $m_{\nu_{1}}$, also vary significantly depending on the dataset used, as shown in Fig.
-11.
+12.
 For $\omega_{b}$, the discrepancy is mainly attributed to the choice of nuclear reaction rates in the PRIMAT code during the BBN process, which predicts a deuterium abundance lower than current constraints from PDG, leading to a significantly lower $\omega_{b}$ Pitrou:2018cgg ; Burns:2022hkq ; Escudero:2022okz .
-While constraints on the primordial neutrino asymmetry $\xi_{\nu}$ do not strongly depend on the choice of nuclear reaction rates, as discussed in Escudero:2022okz , a detailed study using different nuclear reaction rates will be presented in our companion paper.
+While constraints on the primordial neutrino asymmetry $\xi_{\nu}$ do not strongly depend on the choice of nuclear reaction rates, as discussed in Escudero:2022okz , a detailed study using different nuclear reaction rates will be presented in our companion paper Li:2025rjr .
 The inclusion of BOSS BAO data significantly modifies the constraints on $\omega_{cdm}$, $n_{s}$, and $m_{\nu_{1}}$ due to its sensitivity to the evolution of the late Universe.
@@
 Recently, DESI BAO observations have placed even stronger constraints on neutrino masses, $\sum m_{\nu_{i}}\lesssim 0.07\,\mathrm{eV}$ DESI:2024mwx .
-Relevant studies on primordial neutrino asymmetry and neutrino masses using DESI BAO data will be addressed in our companion paper.
-Finally, we briefly discuss the impact of different parameter sets on the resulting constraints, particularly the interplay between primordial neutrino asymmetry and total neutrino mass.
-Using the BBN + CMB + BAO dataset, we present in Fig.
-12 the 1D marginalized distributions for three parameter sets: $\Lambda$CDM + $m_{\nu_{1}}$, $\Lambda$CDM + $\xi_{\nu}$, and $\Lambda$CDM + $m_{\nu_{1}}$ + $\xi_{\nu}$.
-As shown, while different parameter sets lead to significant modifications in constraints on the six $\Lambda$CDM parameters, especially $\omega_{b}$ and $n_{s}$, the resulting constraints on additional parameters are consistent across different parameter sets.
-The corresponding constraints from different parameter sets are also summarized in Table 3.
-We find that the constraints on the total neutrino mass are only slightly tightened when $\xi_{\nu}$ is included as a free parameter, while constraints on $\xi_{\nu}$ remain nearly unchanged when $\sum m_{\nu}$ is a free parameter.
-Therefore, we conclude that constraints on neutrino masses and primordial neutrino asymmetries can be analyzed independently if one is not concerned with constraints on other $\Lambda$CDM parameters.
-However, if constraints on other $\Lambda$CDM parameters are of interest, a parameter set including both of neutrino mass and primordial neutrino asymmetry must be considered.
+Relevant studies on primordial neutrino asymmetry and neutrino masses using DESI BAO data will be addressed in our companion paper Li:2025rjr .
+[…]
 6.2 Constraints on UV Model Parameters
@@
 After the electroweak scale, the L-balls decay into neutrinos, resulting in a significant lepton asymmetry in the universe.
-Following Kawasaki:2022hvx , the primordial lepton asymmetry $\eta_{L}$ is defined as the ratio of the difference between the number densities of neutrinos and antineutrinos to the photon number density:
-$\eta_{L}\equiv\sum_{\alpha}\frac{n_{\nu_{\alpha}}-n_{\bar{\nu}_{\alpha}}}{s}\simeq 0.105\,\xi_{\nu},$ (102)
-where $s=\frac{2\pi^{2}}{45}g_{\star}T_{\gamma}^{3}$ is the entropy density, and $g_{\star}=10.75$ is the relativistic degree of freedom at neutrino decoupling.
-In the Q-ball decay model, $\eta_{L}$ can also be expressed as:
-$\eta_{L}=\frac{3T_{\mathrm{D}}}{4m_{3/2}},$ (103)
-where $m_{3/2}$ is the gravitino mass and $T_{\mathrm{D}}$ is the L-balls’ decay temperature, which is a function of other UV model parameters:
+Following Kawasaki:2022hvx , the primordial lepton asymmetry $\eta_{L}$ generated by the decay of Q-balls can be written as [^4] The difference between the expression here and the expression in  Kawasaki:2022hvx is due to the different definition of the primordial lepton asymmetry $\eta_{L}$.
+$\eta_{L}=\frac{3\pi^{4}g_{\star}T_{\mathrm{D}}}{495\zeta(3)m_{3/2}},$ (108)
+where $g_{\star}=10.75$ is the relativistic degree of freedom at neutrino decoupling, $m_{3/2}$ is the gravitino mass and $T_{\mathrm{D}}$ is the L-balls’ decay temperature, which is a function of other UV model parameters:
 $T_{\mathrm{D}}$ $\simeq 2.69~\mathrm{MeV}\,$
-×$\hskip 15.0pt\times\left(\frac{g_{*}}{10.75}\right)^{-1/4}\left(\frac{\beta}{6\times 10^{-4}}\right)^{-5/8}\left(\frac{m_{3/2}}{0.5~{}\mathrm{GeV}}\right)^{5/2}\left(\frac{M_{F}}{5\times 10^{6}~{}\mathrm{GeV}}\right)^{-2}\left(\frac{N_{\ell}}{3}\right)^{1/2}\left(\frac{\zeta}{2.5}\right)^{1/2},$ (104)
+$\hskip 15.0pt\times\left(\frac{g_{*}}{10.75}\right)^{-1/4}\left(\frac{\beta}{6\times 10^{-4}}\right)^{-5/8}\left(\frac{m_{3/2}}{0.5~\mathrm{GeV}}\right)^{5/2}\left(\frac{M_{F}}{5\times 10^{6}~\mathrm{GeV}}\right)^{-2}\left(\frac{N_{\ell}}{3}\right)^{1/2}\left(\frac{\zeta}{2.5}\right)^{1/2},$ (109)
 where $g_{*}(T_{\mathrm{D}})$ is the relativistic degrees of freedom at $T_{\mathrm{D}}$, $\beta\simeq 6\times 10^{-4}$ is a dimensionless constant, $M_{F}$ is the SUSY breaking scale, and $N_{\ell}$ represents the number of decay channels.
@@
 $T_{\mathrm{D}}$ $=\frac{4}{3}\eta_{L}m_{3/2}\simeq 3.3~\mathrm{MeV}\,\left(\frac{m_{3/2}}{0.5~\mathrm{GeV}}\right)\left(\frac{\eta_{L}}{5\times 10^{-3}}\right)$
-[…]
+$\simeq\left(\frac{m_{3/2}}{0.5~\mathrm{GeV}}\right)\,(1.65\pm 0.86)~\mathrm{MeV}\quad\,[\textbf{BBN+CMB+BAO}],$ (110)
 and, conversely,
 $m_{3/2}$ $=\frac{3T_{\mathrm{D}}}{4T\eta_{L}}\simeq 0.5~\mathrm{GeV}\,\left(\frac{T_{\mathrm{D}}}{3.3~\mathrm{MeV}}\right)\left(\frac{5\times 10^{-3}}{\eta_{L}}\right)$
-[…]
+$\simeq\left(\frac{T_{\mathrm{D}}}{3.3~\mathrm{MeV}}\right)(1.00^{+1.08}_{-0.34})~\mathrm{GeV}\quad\,[\textbf{BBN+CMB+BAO}].$ (111)
 These constraints can be further translated into constraints on other UV model parameters, such as the SUSY breaking scale, as shown in Kawasaki:2022hvx .
@@
 We rederive the quantum kinetic equations (QKEs) for neutrinos and antineutrinos using the Closed Time Path (CTP) formalism for two-point functions.
-Assuming complete flavor equilibration before neutrino decoupling, where $\xi_{\nu_{e}}=\xi_{\nu_{\mu}}=\xi_{\nu_{\tau}}=\xi_{\nu}$, we numerically solve the QKEs in the early universe using an extended version of the public code FortEPiaNO Gariazzo:2019gyi ; Bennett:2020zkv .
-Our calculations include the effects of neutrino oscillations, matter interactions, FTQED corrections up to $\mathcal{O}(3)$, and comprehensive neutrino-electron and neutrino-neutrino collision terms.
+Focusing on the complete flavor equilibration cases with $\xi_{\nu_{e}}=\xi_{\nu_{\mu}}=\xi_{\nu_{\tau}}=\xi_{\nu}$ , we numerically solve the QKEs in the early universe using an extended version of the public code FortEPiaNO Gariazzo:2019gyi ; Bennett:2020zkv .
+Our calculations include the effects of neutrino oscillations, matter interactions, FTQED corrections up to $\mathcal{O}(e^{3})$, and comprehensive neutrino-electron and neutrino-neutrino collision terms.
 The final $N_{\rm eff}$ is then computed based on the resulting density matrices and can be expressed as $N_{\rm eff}\simeq N_{\rm eff}^{\rm SM}+3\left(\frac{30}{7\pi^{2}}\xi_{\nu}^{2}+\frac{15}{7\pi^{4}}\xi_{\nu}^{4}\right)+0.0102\xi_{\nu}^{2}$, where $N_{\rm eff}^{\rm SM}=3.0440\pm 0.0002$.
@@
 We also confirm that spectral distortions of neutrinos and antineutrinos with nonzero $\xi_{\nu}$ differ due to distinct Pauli-blocking effects, in agreement with preliminary studies Grohs:2016cuu .
+The formalism we develop can also be straightforwardly applied to the study of flavor equilibration processes, see Pastor:2008ti ; Mangano:2011ip ; Froustey:2024mgf for previous relevant discussions.
 Using the neutrino and antineutrino density matrices, we further investigate the effects of primordial neutrino asymmetries on BBN, CMB, and LSS.
@@
 We also discuss the interplay between $\sum m_{\nu}$ and $\xi_{\nu}$ when both are treated as free parameters, and how these constraints translate into bounds on UV model parameters, using a specific model where large neutrino asymmetries are generated by Q-ball decay Kawasaki:2002hq ; Gelmini:2020ekg .
-Our work establishes a framework for exploring the implications of physics beyond the Standard Model and $\Lambda$CDM for neutrino decoupling and related cosmological processes, providing corresponding constraints from current and future BBN, CMB, and LSS observations.
-It paves the way for more precise studies of primordial neutrino asymmetries and other new physics scenarios, such as synchronous neutrino oscillations Froustey:2021azz ; Froustey:2024mgf and non-standard neutrino interactions Du:2021idh ; Du:2023upj , driven by increasingly precise cosmological observations.
+[…]
 Acknowledgements.
@@
 Pitrou for helpful communications about their relevant work.
-This work is supported by the National Science Foundation of China under Grants No.
+This work is supported by the NSFC under Grants No.
 12347105, No.
 12375099 and No.
-12047503, and the National Key Research and Development Program of China Grant No.
+12447101, and the National Key Research and Development Program of China Grant No.
 2020YFC2201501, No.
@@
 where, as before, $\varrho^{(i)},\bar{\varrho}^{(i)}$ are the corresponding density matrices for neutrinos and antineutrinos with momentum $p_{i}$, and $\mathrm{Tr}(\cdots)$ means the trace of the term in front of it.
-Finally, the collision terms for the antineutrinos ¯¯$\bar{\mathcal{I}}\equiv\bar{\mathcal{I}}_{\rm sc}+\bar{\mathcal{I}}_{\rm ann}+\bar{\mathcal{I}}_{\nu\nu}$ can be obtained simply by the transformations $\varrho\leftrightarrow\bar{\varrho}$ and $L\leftrightarrow R$.
+Finally, the collision terms for the antineutrinos $\bar{\mathcal{I}}\equiv\bar{\mathcal{I}}_{\rm sc}+\bar{\mathcal{I}}_{\rm ann}+\bar{\mathcal{I}}_{\nu\nu}$ can be obtained simply by the transformations $\varrho\leftrightarrow\bar{\varrho}$ and $L\leftrightarrow R$.
 A.2 Reduced expressions for the comoving collision terms
@@
 $\tilde{\mathcal{I}}$ $=$ $\frac{G_{F}^{2}}{(2\pi)^{3}y^{2}}\left(\tilde{\mathcal{I}}_{\rm sc}+\tilde{\mathcal{I}}_{\rm ann}+\tilde{\mathcal{I}}_{\nu\nu}\right)\,$, (122)
-~$\tilde{\mathcal{I}}_{\rm sc}$ $=$ $\int{\rm d}y_{2}{\rm d}y_{3}\frac{y_{2}}{E_{2}}$
+$\tilde{\mathcal{I}}_{\rm sc}$ $=$ $\int{\rm d}y_{2}{\rm d}y_{3}\frac{y_{2}}{E_{2}}$
 $\left\{\left(\Pi_{2}^{s}(y,y_{4})+\Pi_{2}^{s}(y,y_{2})\right)\left[F_{\mathrm{sc}}^{LL}(\nu^{(1)},e^{(2)},\nu^{(3)},e^{(4)})+F_{\mathrm{sc}}^{RR}(\nu^{(1)},\bar{e}^{(2)},\nu^{(3)},\bar{e}^{(4)})\right]\right$.
 $\left.-2(x^{2}+\delta m_{e}^{2})\Pi_{1}^{s}(y,y_{3})\left[F_{\mathrm{sc}}^{LR}(\nu^{(1)},e^{(2)},\nu^{(3)},e^{(4)})+F_{\mathrm{sc}}^{LR}(\nu^{(1)},\bar{e}^{(2)},\nu^{(3)},\bar{e}^{(4)})\right]\right\}\,$,
-~$\tilde{\mathcal{I}}_{\rm ann}$ $=$ $\int{\rm d}y_{2}{\rm d}y_{3}\frac{y_{3}}{E_{3}}$
+$\tilde{\mathcal{I}}_{\rm ann}$ $=$ $\int{\rm d}y_{2}{\rm d}y_{3}\frac{y_{3}}{E_{3}}$
 $\left\{\Pi_{2}^{a}(y,y_{4})F_{\rm ann}^{LL}(\nu^{(1)},\bar{\nu}^{(2)},e^{(3)},\bar{e}^{(4)})+\Pi_{2}^{a}(y,y_{3})F_{\rm ann}^{RR}(\nu^{(1)},\bar{\nu}^{(2)},e^{(3)},\bar{e}^{(4)})\right$.
 $\left.+(x^{2}+\delta m_{e}^{2})\Pi_{1}^{a}(y,y_{2})\left[F_{\rm ann}^{RL}(\nu^{(1)},\bar{\nu}^{(2)},e^{(3)},\bar{e}^{(4)})+F_{\rm ann}^{LR}(\nu^{(1)},\bar{\nu}^{(2)},e^{(3)},\bar{e}^{(4)})\right]\right\}\,$,
-~$\tilde{\mathcal{I}}_{\nu\nu}$ $=$ $\frac{1}{4}\int{\rm d}y_{2}{\rm d}y_{3}\,$
+$\tilde{\mathcal{I}}_{\nu\nu}$ $=$ $\frac{1}{4}\int{\rm d}y_{2}{\rm d}y_{3}\,$
 $\left\{\Pi_{2}^{\nu}(y,y_{2})F_{\nu\nu}(\nu^{(1)},\bar{\nu}^{(2)},\nu^{(3)},\bar{\nu}^{(4)})+\Pi_{2}^{\nu}(y,y_{4})F_{\nu\bar{\nu}}(\nu^{(1)},\bar{\nu}^{(2)},\nu^{(3)},\bar{\nu}^{(4)})\right\}\,$,
@@
 $D_{3}(a,b,c,d)$ $=$ $\frac{16}{\pi}\int_{0}^{\infty}\frac{\mathrm{d}\lambda}{\lambda^{6}}\prod_{i=a,b,c,d}\Big{[}\lambda i\cos(\lambda i)-\sin(\lambda i)\Big{]}\,.$ (136)
-Again, the comoving collision terms for the antineutrinos ¯~¯~$\tilde{\bar{\mathcal{I}}}\equiv\tilde{\bar{\mathcal{I}}}_{\rm sc}+\tilde{\bar{\mathcal{I}}}_{\rm ann}+\tilde{\bar{\mathcal{I}}}_{\nu\nu}$ can be obtained simply by the transformations $\varrho\leftrightarrow\bar{\varrho}$ and $L\leftrightarrow R$.
+Again, the comoving collision terms for the antineutrinos $\tilde{\bar{\mathcal{I}}}\equiv\tilde{\bar{\mathcal{I}}}_{\rm sc}+\tilde{\bar{\mathcal{I}}}_{\rm ann}+\tilde{\bar{\mathcal{I}}}_{\nu\nu}$ can be obtained simply by the transformations $\varrho\leftrightarrow\bar{\varrho}$ and $L\leftrightarrow R$.
 Appendix B Finite Temperature QED Corrections